A Complete Bibliography of Publications in North American Journal of Fisheries Management: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.04

Title word cross-reference

+ [VGM00]. ^{18} [MFW^{+}09]. _{2} [SH07]. _{n} [HN05]. _{r} [HN05]. α [LKS00]. δ [MFW^{+}09]. F_{1} [RJB07]. F_{MSY} [JCW02]. × [Box02a, GWBH02, GAR^{+}09, WAB^{+}02]. W_{r} [BMB03]. W_{s} [BWB^{+}00, DBM04].

-ATPase [VGM00]. -challenged [JMP07]. -Methyltestosterone [LKS00].

1.5-Length [BT03]. 12-mm [CRK^{+}05, Tat09]. 13 [Pit02]. 13-mm [JB00]. 15-Inch [FHG^{+}01]. 16-Inch [NH00]. 16-mm [JB00].

20-Year [Ano06, BB06]. 21st [BI02]. 222 [SSS07]. 23-cm [BWKS02]. 23-mm [BG06]. 24 [Kol08]. 254-mm [Box02b, HJ02].

3-Trifluoromethyl-4-Nitrophenol [SSS08].

40-Inch [MA00a].

5.87-cm [GAKB04].

6.03-cm [GAKB04].

Abalone [LRBJ08, RPD09]. Abbott [You09]. Ability [LB09, PHK03]. aboard [PGS01]. Absorbable [SVP^{+}04]. Abundance [ACBS05, ATM03, AHJ^{+}06, AIHRdSD09, BM00, Bar04, BMK^{+}00, BAC05, BWC08, BRSA07, BHM^{+}07, BTM07, CKC^{+}09, CPW^{+}06, CM05, DS05a, DP03b, DB09b, Dew08, Dew09, EYI08, GSWS04, HB05a, HDPS07, HNTM00, HP04, HMSM08, HBHB00, How06, HLH05, JJ07, LF05, Mad08, MSES03, MEL09, MDKM06, MHW00, NBHH00, NBHH01, PPS03, PJBT09, QGBH06, RG01, Rad06, RHGW05, RHB05, RF00, RD05b, RD06, SB05a, SM00a, SMBM03, SM00b, SAC05, TAM^{+}03, TTRF05, TPG06, TFW00, TCSS07, Und00, WMSG09, WF04, WHB03, Wil06, ZB00]. Abundances [MZ02]. Acceptance [PWG02]. Access [HL09, Sta01, SMR08, VPR06]. Acclimated [GCM^{+}04]. Acclimation [KLS01, SLS^{+}04]. Accounting [BPB07, HBB07, MTL09, NL06]. Accumulation [LD04]. Accuracy [HB05b, HSH04, JEB03, LL06, MC04a, MNB09, NGWK08, RR05, RCK05, SH04c, SMR08]. Accurate [GCM^{+}03b, STSS05]. Acid [CRD05, SJS00]. Acidified [HDB00]. Acoustic [AAF^{+}06, AGB^{+}04, BSSC04, CJKS05, SL09, SPR05c, Yul00, YASG07]. Acoustic-Tracking [BSSC04]. Acquired [GBLC02]. across [AHJ^{+}06, HB05a, MH05]. Actinospore [NTSJ03]. Actions [Aus01, BPRG08, PH05]. Active [CWP03, DK04, PH03a, Sul03a, WB03]. Actively [Rad03]. Activities [GJB09, WC01]. Activity [MMVL05, MS04b, SHMR03, VGM00]. Actual [WG08]. Acute [LEJ06, Mea04]. Addition [PRSA06]. Additions [HS08]. Additive [WMSG09]. Adfluvial [DHMHJ06, HW07]. Adherence [GGZH08]. Adipose [VBHT05]. Adirondack [BKJ06, JKS01]. Adjacent [FSH06, She04]. Adjoining [BKD^{+}09]. Adjusting [CGM^{+}06]. Adult [AHJ^{+}06, BE02, Bai03, BPK08, BG00, BHM^{+}07, BJ06, CH00a, DB06, EYI08, EFS04, GLA00, GTZM09, Har03, HZG^{+}09, KB04b, KPRB04, KPJ^{+}04, KBPC08, KMB^{+}09, KJS^{+}09, KHJ^{+}08, LPL04, LKS00, LH04b, MS03a, MDH00, MPR08, MTM09, MC01, NBHH01, NCK^{+}06, NJP^{+}09, NFQ07, PSB^{+}07, PHC^{+}09, PP00, RW03, RB03, RVM^{+}08, SVW^{+}08b, SJP02, UBK04, UGM^{+}07, VW07, YI04, YI06]. Adults [GCM^{+}04]. Advanced [MCS^{+}08]. Advanced-Fingerling [MCS^{+}08]. Advances [GHM^{+}07]. Advantages [Vok03]. Adverse [MOB^{+}08]. Advice [Dor02, OJL08]. Aerial [CH00b, LR05, RR05, VPR06]. Affect [GHA^{+}02, LH04b, MHRR07]. Affected [KSB02, RTBF09]. Affecting [CM05, DPW05, FSSR08, GSV^{+}09, JZMS05, MC06, NGWK08, PR04a, SM09c, YASG08]. Affects [DCB09, HV03, Tat09]. after [BKJ06, BG06, BK09b, LF05, Mea04, MC08, PM08, RD05a, Sch03, SRGC05, SHPCQ00, WF04, WME05]. against [SM07b]. Age [BG06, BSEM05, BM01b, BE01, BP01, BM04b, CPWM06, CHJ07, DJSG09, EC00, ESPW03, GM04, GCQVMAVP09, GGCF08, GM00, GL02b, HNE04, HWKN00, Hor01, HAW01, HLH05, HSH04, HKL08, IMSW03, IK05a, Ise07a, JJ07, KD00, KQ07, KQH09, KWP06, LL06, MHO01, MMH09, MSR08, MZ02, MEHJ08, MSP^{+}08, MC01, NB05, NN08, NMJ01, OV01, OGS04, ON05, PS02, PEW07, PF02, PF03, QSGS04, QJBH07, RCK05, SI08, SDSQ09, SB04, SH08, Stu06, SN02, SB06b, TAM^{+}03, VBP08, WJKW07, WESW09, WMO08, WTW^{+}04, YS04]. Age-0 [BG06, BSEM05, BP01, BM04b, EC00, GM04, GM00, GL02b, HNE04, JJ07, KD00, MSR08, MZ02, OV01, OGS04, ON05, PS02, PF02, PF03, SI08, SB04, SN02, TAM^{+}03, WJKW07, WESW09]. Age-Specific [NMJ01]. Age-Structured [CPWM06, YS04]. Agents [SMB07]. Ages [BIBP02, BH03b, DB06, KC00, Log07, LF01]. Aggregate [PCIB08, WP03]. Aggregation [JZMS05]. Aggregations [SPR05c]. Aging [MB05b, MSP^{+}08]. Agricultural [MFP^{+}08, SMSM01, TPG02]. Air [CJPB09, DO01, DO02, DP04, SRGC05, TCD^{+}08]. Air-Exposed [TCD^{+}08]. Alabama [EYI08, GM00, HMM09, HM08, IDK03, KD00, LM02, Mac03, MHM09, MMH09, SIJH06, SM09b, SM09d]. ALaboratory [DO02]. Alaska [AP00a, BCM^{+}05, Bry00, BCW04, BLM^{+}09, Dew08, Dew09, HWB^{+}03, HGHG04, HB06, KBF^{+}00, Mar01, MTN09, RBM00, SSR07, Und00, UBK04, UGM^{+}07, WT09]. Alaskan [Mek04, SH04a, WHC09]. Alberta [PvPR^{+}06, Sul02, Sul03a, Sul03b, WFS07]. Alder [EM06]. Alewife [Bau02, CKC^{+}09, EV09, PPE03, SM00b]. Alewives [HP09, LL06, MOB^{+}08, PM08]. Algal [MHRR07]. Alive [Dia08]. Alizarin [COG^{+}07, COS^{+}09, vdWF03]. Allacustrine [DBG07]. Alleles [WCR^{+}05]. Allocation [CGM^{+}06, CWDH03, Goo07, PW09]. Allometric [DTMS06]. Allow [PCIB08]. Allowing [JPB^{+}07]. Alluvial [CH00b]. along [HS03, RM01, TCSS07, WGSW07]. Alter [BSB03]. Alternative [BM00, BMB03, GHAS04, MHW00, PALdlN05, PMPS03]. Alternatives [GD02, HPM02, MGAC00]. Amarillo [GCQVMAVP09]. ambigua [COS^{+}09]. America [Goe02, HB00, QGSS03]. American [BHO02, CSI^{+}04, HHAK02, HLO08, MHO01, MHO02, MHOH06, MH08b, McC01, MMLM06, MDH00, RH04, SLG02, SOM09, Spr05b]. americana [SEGG05]. among [BW06, BF06, CLB04, CLB^{+}07, CWDH03, ICBM04, KBPC08, KGM01, LB09, ML04, OGK07, PR05b, QBH^{+}09, RW03, RBB08, ZHBB09]. Amount [CK09]. Amphibians [WP08]. Amplification [BK09a]. Amplified [LZW^{+}08]. Anadara [FPRRHQ09]. Anadromous [DS07, Har03, HHWL00, HLO08, MH08b, SMF^{+}09]. Anadromy [BC05]. Analyses [GJB09, Ian02, VTK^{+}07]. Analysis [ARC^{+}07, BHO^{+}08, BKD^{+}09, CRC^{+}08, CPWM06, Dor02, GMHF03, GDH04, GD02, HCSCT06, IMW02, Jac05, KHJC06, KRB^{+}04, Kha06, KQ07, KBF^{+}00, MPPB02, MTN09, MPBM06, MHRR07, MSP^{+}08, MBHK02, RPW^{+}08, RDBD08, Ryp08, SDSQ09, SRN00, TWP03, You09]. Analytical [PvPR^{+}06]. Analyzing [BWM08, CPWM06]. Anatomical [DH09, JSMV07, LSK^{+}04]. Anchor [BI00, HJ06, PR05a, VW01, WW04]. Anesthetic [PP00]. Anesthetics [SH07]. Anesthetized [SSS07]. Angled [BTI05, OSC06, TCD^{+}08]. Angler [AHH00, ATM03, BCC03, BHL03, Box02b, BS01, CRBD06, DMB09, EWD^{+}06, FS06, FPK07a, MC04a, MGH03a, Mek04, MWN08, PSWD07, SLEM07, SR02b, SZW02, VFJ^{+}05, VJK^{+}07, VPR06, WP04a, ZS03]. Angler-Caught [BHL03]. Angler-Reported [MC04a]. Anglers [Arl06, CBK00, DCSM01, HB07, IWLB05, KG05, PGR^{+}04, PPS03, RB07, SGG03, Sul03b, TWPP06, WG08]. Angling [ASD04, AKKC08, ARPP06, BRCC03, CPW^{+}05, CPSM00, CTB07, DHNH05, DHNH07b, DD04, FHG^{+}01, HBH00, HBH05, KSM^{+}03, LPO01, MMKC05, MD00, NHNE00, PC00, SL02a, SRGC05, Sch04b, Sul02, SKM^{+}03, UHRP01, WFS07, YH04]. Annapolis [GM02]. Annual [DB06, DA05, HVN06, LPL05, VR05]. Annulus [TWP03, WTW^{+}04]. Antenna [CB01, IYJS02]. Antennas [BGH04]. Antibiotic [IYJS02]. Antimycin [DM07, HMW^{+}09, RBK09, SL02a]. antipodarum [ROS04, VB08]. Any [JCW02]. Apache [ARC06, PN03]. Apalachicola [EYI08]. Appalachian [GAG^{+}01, HWKN00, KM00]. Apparatus [PGS01, VH00]. Apparently [Arl06]. Application [AAF^{+}06, AWWG08, BWW^{+}08, BWJ^{+}08, DV05, GBLC02, HLO08, MHO01, MTL09, PMPS03, PvPR^{+}06, PAAK07, SVW08a, Zho03]. Applications [BCM^{+}05, CHD03, GB03, GHM^{+}07, Ker08, MMS08, MBHK02, PM01]. Applied [HN05, VPR06]. Applying [GBLC02, SVW^{+}08b]. Approach [AHH00, BWC08, DMM02, GHAS04, JRNS09, LMA^{+}03, NMP07, OJL08, PLP01, Pea04, PvPR^{+}06, PPSC03, TFW00]. Approaches [BPRG08]. Appropriateness [UGM^{+}07]. Approximation [WG03]. Aquaculture [WHC09]. Aquatic [BTK04, BBD02, CB04, DM07, HHB05, KS02, KSB02, LWHP07, MSW^{+}08, MHB07, PSW06, RDN00, ZB00]. AQUI [SH07]. AQUI-S [SH07]. Arapaima [Cas04]. Archimedes [WBH02]. Archipelago [Orr07]. Archived [RDBD08]. Arctic [DB06, FTWS06, KB04a, Rik00]. Area [AWWG08, FPK07b, PW09, PBI03, Sha04, WFW02, WP08, WP04b, BNV02]. Area-under-the-Curve [PBI03, Sha04]. Areal [SH04b]. Areas [DJC^{+}06, EM06, GAM^{+}08, HB05a, MP01, SVS^{+}09, TAME03]. Arizona [Cla04, BG00, DM07, MRSR01, MSRP01, PR04a]. Ark [FPRRHQ09]. Arkansas [BH05, CRL08, DP03b, HRL06, HBJ05, HBKJ08, HBB04, RL02]. Arm [NSH^{+}08]. Array [CJKS05]. Arrays [KH01]. Arrival [NFQ07]. Arrowtooth [WT09]. Artifacts [Dew08, Dew09]. Artificial [BSSS02, BMK^{+}01, CRBD06, FWEP08, GCR08, HKFM08, OGE01, Zho03]. Artisanal [AMVA09]. Asian [JT07, RSF06, SN07]. Aspects [MP04b]. Assemblage [CBP07, HBJE00, PSP^{+}01, RPW09a, RPW09b, RUKR07, VWDK00]. Assemblages [BW08, FSLS09, LPSL08a, LZS^{+}09, OMH05, PCG^{+}01].

Assess [BSSC04, HTA^{+}09, JEB03, LCM06b, RBM00, YASG07]. Assessing [AK05, CSW05, DMB09, Fie04, HLV^{+}07, KFG^{+}01, KLW00, OS05, ON05, RHK^{+}03, SP07, TC01]. Assessment [AKARF08, BBO03, BBGVM05, BHO02, BABC08, BS01, BM04b, BMCB09, Cas04, CB01, DJSG09, Ger07, GM05, HJC09, HBKJ08, Ian02, JCW02, JZMS05, KP07, LZW^{+}08, LM02, MHM09, MH08b, MRSR01, OJL08, PVMR01, PMPS03, PAAK07, RJY08, SH07, Sch04a, SH04c, SBR^{+}05, SLB^{+}06, VEK09, WME05, WHQ^{+}08, YS04]. Assessments [PCG^{+}01, SW09]. Assignment [IK05a, Ise07a, SDSQ09]. Associated [ARMB04, BH03a, BHDA05, BKNB04, BJ09, BKZA08, Cla04, GCN^{+}08, IWLB05, IC08, KS02, MRSR01, MMKC05, MRK08, OSC06, PNSB01, SB06a, SMWS02, SW04, TP05, ZS03]. Association [LB09]. Associations [ML06, SAC05]. Assumptions [DHNH05, MTN09]. Atimycin [RSF06]. Atimycin-A [RSF06]. Atlantic [CJM^{+}02, GMH09, HJC09, BTK04, BRGO03, CR04, DC06, DV00, ERO^{+}09, FSK04, FF04, GJB09, GD02, GTZM09, GL02a, GL02b, HCSCT06, HVN06, HR06, HEB09, HHM05, HLV^{+}07, HZG^{+}09, HOFM06, Joh04, KRK^{+}02, KWP06, LT01, MFT08, MGAC00, MJ05, MSES08, MFRS09, MMF02, Moh03, MRW05, OGS04, Orr07, OJL08, PBH00, RPJB05, RD09, RBB08, SPR^{+}05a, SHD^{+}05, SJ03, SFS04, SVS^{+}09, SEJCC04, SLB^{+}06, SMM^{+}07, VC02, WCR^{+}05, WGSW07]. ATPase [VGM00]. Attempt [RJB07]. Attempts [Rad03]. Attitudes [HB07]. Attraction [Spr05b, YMR03]. Attributed [AHC^{+}09]. Attributes [BRSA07]. Australia [Stu06]. Australian [DGAG08, FRW09]. Automated [MRS00]. Autumn [CLB04, MDHK08]. Autumn-to-Spring [CLB04]. Availability [DHG02, GMHX08, PSP^{+}01, VFJ08]. Available [CB04, FHSM08]. Averaged [WG03]. Avoidance [AGB^{+}04, MC04b, RCB07]. away [PJC00].

B.A.S.S. [WSP02]. Back [BE01, MNB09]. Back-Calculate [BE01]. Back-Calculated [MNB09]. Backcountry [YH04]. Backpack [DSW01, SSG05, VWDK00]. Backwater [SGB07]. Backwaters [CRL08]. Bacterial [JMP07]. Bag [BCC03, CKC^{+}09, FS06, VC02]. Bahía [FPRRHQ09]. Bait [AKKC08, AB02, DP07, FS07, Jen03, SNO02, WMB^{+}00]. Baited [JSL05]. Baitfish [OSC06]. Baitfishing [DK04]. Baiting [Sto05]. Baja [FPRRHQ09, RPD09]. Bank [GHASLC09, ML06, OJL08, GDH05, GHDM09, MDM^{+}04]. Bar [BBZW04, KH01, PR05a]. Baraboo [CBP07]. Barbed [DP07, SH02a]. Barbless [DP07, SH02a]. Barndoor [GHDM09]. Barotrauma [GC08, SGD09]. Barramundi [DGAG08, GG09a]. Barrier [BBO03]. Barriers [ARC06, Cla04, JKS01]. Bars [SBR^{+}05]. Based [AHH00, BRS02, BWJ^{+}08, BPB07, BMW^{+}02, BLK^{+}08, CPWM06, CGM^{+}06, CWP03, DeV07, DB09b, DP02, DV05, FDM^{+}06, LRBJ08, MB05a, McH07, Mir02, MSP^{+}08, NMP07, PPSC03, RMBZ08, VBP08, WVT09, WG03, WTW^{+}04, Wil02, AIHRdSD09, MTN09]. Bases [Bau02]. Basin [BRB09, HMW^{+}09, McC01, NBGK09, RPM08, APS05, BE02, CHR^{+}08, DJL09, GCM^{+}04, KPJ^{+}04, KRB^{+}04, KRDC08, MB05b, MPR08, MPM09a, PPBM04, QGBH06, QJBH07, QBH^{+}09, RFLN01, SM07a, SCD^{+}08, SSWC05, TFF^{+}07, WPV^{+}04, ZB00]. Basins [FB07, JGW^{+}09, ZC06]. Basinwide [SLB^{+}06]. Basis [BCW^{+}09]. Bass [ATM03, ARMB04, AWM08, BBK09, Bau02, BRGO03, BP02, BTI05, BAC05, BABC08, BSG05, BM04b, BRSA07, Buc01, BSB03, BH03b, BBS05, BTHE04, BM02, CHD03, CK09, Col02a, Col02b, CRL08, CH00a, CPSM00, CBM00, CB01, CPDS01, CM05, DF07, DMB09, DD02, DW09, DSM07, DCSM01, ENJO04a, ENJO04b, FS00, GAR^{+}09, GC08, GM00, GSC02a, GSC02b, HJ04, HJ06, HRL06, HB05c, HA02, HM08, IYJS02, JH01, JPB^{+}07, KB04b, LB08, LLCBF07, LF01, LF05, LM02, MB01, Mac04, MMS08, Mau08, MCPM02, MC00, MCS^{+}08, MNS01, MMKC05, Mue03, MTAB08, NB05, NBN04, NLC01, NN02, NN08, NH00, OS05, OCW04, ON05, PPE03, PR04b, QGBS02, RMD02, RAP09, RWZ09, RP09, RWJD04, RHHFB00, RBB08, SB06a, SH05a, SWG^{+}06, SD06, SGR02, SR02a, SR02b, Sch04b, SJS00]. Bass [SM09b, SWPO06, SWD08, SOR05, SM09d, SKM^{+}03, SRN00, SN02, TFF^{+}07, TAME03, TAM^{+}03, TWRH07, TCD^{+}08, VBR04, VW01, WN04, WMB^{+}00, WSP02, WPB02, WCR^{+}05, YI04, YI06, YDJM09]. Basses [AIHRdSD09]. Batch [MB08, MMLL03]. Bath [MC08]. Bay [AP00a, AIHRdSD09, Dew08, Dew09, Fie02, GBFJA07, HGHG04, KBF^{+}00, RW06, BPAA03, BSSC04, CH01, NMJ01, NTH08, RHHFB00, SSB01, SBR^{+}05]. Bayesian [Dor02, MPBM06, MBHK02]. Bays [BBJ^{+}00, CKSG09, KSB^{+}08, SMM^{+}07]. Be [She04, GSV^{+}09]. Beach [PCG^{+}01]. Bear [CKH05]. Bearing [DeV07, MB05a, McH07]. Beaver [HK04, PPBM04]. Bed [KVMM07, MP04a]. Beds [MSW^{+}08]. Been [MFW^{+}09]. before [WF04]. Behavior [BTS01, BWC08, BSSC04, BTSB07, BJ06, CSB03a, GC08, JLSL06, KBPC08, KJS^{+}09, MMS08, NRJ05, PK01, Pea08, PPAR06, RCB07, RTBF09, SSR07, SR02b, TBL04, TWPP06, TCD^{+}08, TRW00, TCSS07, VRK00, WESW09, WG08]. Behavioral [DP04, GFPJ09, HBC08, YH04]. below [MRSR01, MSRP01]. Benefit [MPBM06]. Benefits [CBS^{+}05]. Benthic [FSLS09, HBS^{+}05, PGS01, Ron03, WLG03]. Benthos [PNSB01]. Bering [JML02]. Best [MR02]. Better [HPM02]. Between [RP09, ACBS05, Ano06, Arl06, BB06, BABC08, BPH08, BP01, BMB03, CH00b, DTMS06, FHE05, FM08a, Goo07, HNE04, HMS^{+}03, HRB^{+}04, KLH^{+}08, LB08, LGA02, MMH09, ML06, MTF^{+}08, PP01, PTN06, PH03b, RHB05, RH03, SLG02, SIB02, SPM06, SWA^{+}08, SJS00, SGPK09, SSC07, SM00b, SN02, VFJ08, VS05, VR05, WP04b]. Bias [BSRS04, BJ09, BR06, Che04, GASH05, HB06, LF01, YS04]. Biased [GCN^{+}08]. Biases [GHAS04, WCE^{+}09, ZS03]. Big [CHAO08, SHNE05, WP03]. Bighead [RBK09]. Bight [GMH09]. Billfish [GD02]. Bioassays [AFM^{+}06]. Bioassessment [MSES08, UF06]. Biochemical [MWC02]. Bioenergetics [JMHB08, MBS^{+}08, NSH^{+}08]. Biological [BPS00, DMM02, GP03, MPBM06, NBH^{+}03, OJL08, RHB08, WC01]. Biologically [BLK^{+}08]. Biologist [RAP09]. Biology [WPB02]. Biomanipulation [PRSA06]. Biomass [ICB04, LI09, Mac02, SYH^{+}07, YASG07]. Biomonitoring [Dra07]. Biota [RDN00]. Biotic [BPT08, DP03a, DP02, DV05, FM08a, Gol09, MOH07, RHK^{+}03]. Biphenyl [BRS02]. Bird [RFLN01]. Bivalve [FPRRHQ09]. Black [BWKS02, BPH08, Buc01, DAJB09, ENJO04a, ENJO04b, GWBH02, GBPT08, IBSC02, IC08, JB00, LF05, MB01, MCPM02, MC05, MC06, Mue03, MAC00, NDH04, PR03, PR05a, PR05b, PLW^{+}09, Pin00, RWZ09, RP09, RBB08, SIB02, SB05b, SH02b, WSP02, FPRRHQ09]. Black-Nosed [IBSC02]. Blackfoot [Sch00, Sch01, SPL02]. Blacktail [RR08]. Blackwater [PRP00]. Bleach [HSB^{+}00]. Bleach-Dipped [HSB^{+}00]. Block [PBT05, SSG05]. Blom [Ger09]. Bloodworms [SABW08]. Blue [AHC^{+}09, BS09, CH01, Dia08, GAKB04, HBS^{+}08, HBB04, HMM09, KNL02, NHS08, RT09, RW06, SEC01, WKAP04]. Blue-Ribbon [KNL02]. Bluefin [BDS^{+}02, MGAC00]. Bluefish [FSSR08]. Bluegill [BM04b, CSB^{+}03b, CBS^{+}05, CM05, EWD^{+}06, JB00, Jac05, PWG02, PWK02, PH05, SPM06, SL02a, SH05b, TP05, WEC07]. Bluegills [CMN08, CA06, HAW01, HW09, NB05, PR05b, PWB04, SM09a, UHRP01, WDB06]. Bluehead [QBH^{+}09]. Boat [Cun04, DS05c, LCM06b, Mir05, OS05, RUKR07, Sta01]. Boating [LR05]. Bocaccio [MGGL04]. Bodied [HBS^{+}05]. Body [CTP08, HH01a, KMB^{+}09, MFW^{+}09, MDHK08, PNSB01, SHMR03, SJ08]. Bog [PTD03]. Boise [MDHK08]. Bonneville [BTM07, CKH05, GAM^{+}08, HK04, PJC00, RB03, Wer07]. Boreal [Ste03, WTH08]. borealis [LI09]. Bottlenecks [HBFL08, MMLM06]. Bottom [Han03, MS07, MD04a, SYH^{+}07, YASG08]. Boulder [RVM^{+}08]. Bow [PvPR^{+}06]. Bowfin [KQH09]. Box [CB04, RO06]. Brazos [DW08]. Breaching [SDPB00]. Breeding [BTS01]. Bridge [BKZA08]. Bridgelip [Ric07]. Bristol [AP00a, Dew08, Dew09, HGHG04, KBF^{+}00]. British [BCM^{+}05, BWJ^{+}08, BK09b, CTB07, HS03, MPBM06, MHRR07, MRK08, NWC^{+}03, Orr07, PSB^{+}07, PHC^{+}09, SPR05c, BPB07, BPS00, CW02, HMM^{+}05, PH09, PRP00, RH03, SSP08]. Broad [LMA^{+}03]. Broad-Scale [LMA^{+}03]. Broodstock [SBCH07]. Brook [BKJ06, CBM03, DMMW08, DCB09, DH09, DP07, FM08a, GAG^{+}01, HMSF08, HK04, HOFM06, HBFL08, HDB00, HTGS08, HH01b, JKS01, JRWK08, KWA00, KPL09, LKYM09, LH04b, MLS06, MMF02, MS04b, PFWC08, PBM03, RJB07, RD05a, SRGC05, Sch08, SCD^{+}08, She04, SH08, SHM^{+}08, WAB^{+}02, WSM^{+}08, YMR03]. Broughton [Orr07]. Brown [BKJ06, BFR04, BB04, Car01, DH09, FM08a, HV03, JEB07, OV01, OGE01, PN05, QK00, RPW^{+}08, SJP05, SSC07, SWWR04, TWN^{+}07, Wil06, WTH08, YH04]. Brownlee [SMB03]. Bubble [BE02, BERH02]. Buffalo [WW01]. Buffer [BNV02, SM07b]. Build [CRBD06]. Bull [ACBS05, ACB07, BPT08, BC05, DHMHJ06, DBG07, DRD01, DRC03, GG09b, HMSM08, JGW^{+}09, JMP07, MP01, MDHK08, MGH^{+}03b, MM05, PSWD07, PMPS03, RA01, Sch03, TPG06, ZJ02, ZM06]. Bullhead [SJP05]. Bundles [WSH09]. Burbot [PH09, SM07b]. Bycatch [CH01, DO01, DO02, DAJB09, GCM^{+}03b, GMH09, HPB05, MD04a, Pow05, PR04b, SFS04, TKW00, WMJBC04, ZK00, dSCT01]. Bypass [APS05, JAAW05, MSWS01, SMSM01, Wer07].

calcarifer [DGAG08, GG09a]. Calcein [COS^{+}09, HOFM06, HKFM08, MMF02, Moh03, MB08, NT04]. Calcein-Marked [MMF02]. Calcified [DTMS06]. Calculate [BE01]. Calculated [MNB09]. Calibration [GBLC02]. California [BG00, FPRRHQ09, RPD09, WM05, ACPHT09, AMVA09, Bar05, FSH06, GG05, JR03, KLH^{+}08, LRBJ08, MS04a, MSW^{+}08, MW07, OM04, Ris00, SMSM01, SAC05, WHHR01, WP03]. californicus [CPE09]. Callinectes [CH01, GAKB04, RT09]. Camera [CB04]. Campeche [GHASLC09]. Can [GSV^{+}09, MMW08, PPS03, SBCH07]. Canada [BPB07, FF04, HS03, HL09, MFT08, Orr07, PH09, PvPR^{+}06, SPR^{+}05a, SEJCC04]. Canadian [FTWS06, PPS03, STSS05, VEK09]. Canal [CBM00]. Canals [GZC08, PvPR^{+}06]. Canister [AB03]. Canyon [MRSR01, MSRP01, PR04a]. Capacity [KFG^{+}01]. Cape [HRK03b, NMJ01]. Capital [MTF^{+}08]. Caps [MP08, Mir02, QSGS04]. Captive [CCKF^{+}09, MMF02]. Captively [BTS01, BBTC07]. Captivity [FRW09]. Capture [AMVA09, BPK08, BTS02, BUH07, BS09, CPWM06, CH01, DF07, DD04, DK04, FRW09, GFPJ09, GOG09, HPM08, HB06, HLH05, JSL05, Jon05, LKYM09, Mek04, OSC06, Rul07, UBK04]. Captured [ASD04, GSWS04, RH04, RD05b, RD06, VW07]. Capturing [HHM05]. Carcass [BK09a, EM06, MPM09b, SSR09]. Carcasses [PSB^{+}07]. Cardiac [CPDS01]. Carlin [DC06]. Carolina [JDB^{+}04, OZR08, BWB05, BPBS08, BKNB04, BDS^{+}02, BTHE04, DIC04, GAG^{+}01, LPWH01, LB08, MRC09, PFS^{+}06, RDBD08, RT09]. Carp [BBD02, JT07, PSW06, PASP07, PEW07, RBK09, SSB01, SWMH06, WHCL02, KGM01]. Casco [BBJ^{+}00]. Case [Arl06, GJB09, HGHG04, HI01, Ise07b, KM05, PTN06, SDPB00, SJP02, SL02b, SVS^{+}09, UWBL05, WVT09]. Catch [ASD04, ATM03, AWWG08, Arl06, ARPP06, BKJ06, BP01, CDPA09, CH04, CGM^{+}06, CPSM00, Dau05, DMB09, DB09b, Dia08, EO03, FSSR08, Goo07, HBH00, HNE04, IWLB05, JB00, JSMV07, Jen03, JPB^{+}07, KPL09, KRK^{+}02, KG05, Mac04, MHOH06, MGH03a, MC00, Mek04, MWN08, MMKC05, MTF^{+}08, NLC01, PMBM08, PHB^{+}09, RAP09, RBB08, RT09, SLG02, SL02a, SGG03, SRGC05, SGR02, SR02b, SH05b, SL01, SCHK03, Sto05, SM06, TWH01, TMS05, VFJ^{+}05, VJK^{+}07, VW07, VPR06, WG08, WW08]. Catch-and-Release [ASD04, ARPP06, CPSM00, FSSR08, JSMV07, KRK^{+}02, Mek04, MMKC05, NLC01, RAP09, SL02a, SRGC05, SM06, TWH01, VW07, WG08, WW08].

Catch-Curve [CDPA09]. Catchability [HPB05, HBH00, HBH05, MC06, NHNE00, PSP^{+}06, RHB03, SH05a, SWWR04, VFJ^{+}05, YH04]. Catchable [BSD09]. Catchable-Sized [BSD09]. Catches [ALA07, AB02, BJ09, SSG05, Sul03b, YASG08]. Catching [WP04a]. Categories [BWB^{+}00]. Catenary [RGP^{+}07]. Catfish [AWWG08, AB02, BMP07b, BI00, BIBP02, BS09, CPG^{+}08, Cun04, DS05a, DS05b, DS05c, DB09a, DPW05, FS07, HP04, HMM09, KWP06, MP08, MHM09, MMH09, MS02, Mic05, MWN08, MNB09, Mic09, SJP05, SIJH06, SI08, SJ05, SMB03, SJ08, SMPM02, TC01, VR01, VR05, WKAP04, WW01]. Catostomids [QJBH07]. Catskill [BWEM08]. Cattle [GG09b]. Caught [BMP07b, BHL03, BFF^{+}02, DD04, DK04, DP07, DCSM01, ENJO04a, FCM^{+}09, GSWS05, GC08, KG08, LSK^{+}04, MFW^{+}09, NRJ05, PRHS03, RB07, SWG^{+}06, STSS05]. Cause [Mic05]. Caused [SKM^{+}03, UBK04]. Causes [MJ04, PGR^{+}04]. Cautionary [BM01b, CMN08, WTW^{+}04]. Cayuga [KCR^{+}05]. Cedar [TBL04]. Ceded [SHB^{+}04]. Central [WP03, Cla04, DP02, DV05, Dra07, JR03, KG05, LLCBF07, MRK08, MC01, NDH04, NWC^{+}03, Sch04a, SAC05, WM05]. Centrarchid [Bar04, CCBW01, CRC^{+}08, OMH05, RRS07]. Century [BI02]. Ceratomyxa [MVH^{+}09]. cerebralis [BLA^{+}07, GGZH08, NTSJ03, SLS^{+}04, WAB^{+}02, ZB00]. CFD [Kha06]. Chain [WH09]. challenged [JMP07]. Challenges [BI02, PH03a]. Chambers [HNTM00]. Change [BTS02, WHNC09]. Changes [BK09b, BHL03, BM02, CWK09, CLB04, CP04, DHG02, GLA00, HF00, Han03, HBHB00, MB01, MFP^{+}08, MHRR07, MC01, PC00, PNSB01, PM08, Rad06, SHM^{+}08, SKM^{+}03, TGR02, TGR06]. Changing [BB08, FD01, MH08a, SHB^{+}04]. Channel [AWWG08, AB02, BWEM08, BMP07b, BM01a, BFH^{+}08, BI00, BIBP02, BS09, CPG^{+}08, DPW05, FS07, GLA00, HP04, HMM09, ML04, MRJL05, MS02, Mic05, MWN08, MNB09, Mic09, NGWK08, RRCJ08, SI08, SJ05, SMB03, SMPM02, VR01]. Channelized [JJ07]. Char [Rik00, SWA^{+}08]. Characteristics [Bai03, BRGO03, CJPB09, Col02b, DS05a, GHA^{+}02, Gol09, GM00, HW07, HMM09, HTGS08, KDSF07, MEL09, Muh02, NTH08, PW03, PLW^{+}09, RH03, RDME02, SIB02, Sch00, SVW^{+}08b, SB05b, TWP03, YS04, YGG04]. Characterization [CCKF^{+}09]. Characterizing [LZS^{+}09]. Characters [HEB09]. Charlotte [BWJ^{+}08]. Charter [MTF^{+}08]. Chase [Cun04, DS05c]. Chassahowitzka [SMP04]. Chautauqua [MM02]. Chehalis [FSC03]. Chemical [MHB07, MC08]. Chemistry [BH05, CAM06, CLB^{+}07, EM06]. Chesapeake [RHHFB00, SSB01, SBR^{+}05]. Chickamauga [HB05c]. Chicot [RL02]. Chilean [WHC09]. Chilko [BPS00, MPBM06]. Chinook [ARC^{+}07, AGB^{+}04, BHO^{+}08, BWW^{+}08, BWJ^{+}08, BM01a, BTS01, BJ06, CKC^{+}09, CBWB02, CB03, CBYB03, CSB03a, CP04, CSA^{+}04, CHJ07, CTB07, Dar03, DHGP03, GM04, GG05, GCM^{+}04, GTRC02, GHA^{+}02, GAM^{+}08, GMHX08, GCR08, HTA^{+}09, HLMB01, HI01, HMS^{+}03, HCFK08, JR03, KPRB04, KPJ^{+}04, KJR^{+}09, KLH^{+}08, KJS^{+}09, LSK^{+}04, MC04b, MS03a, MVH^{+}09, MRJL05, MMVL05, MS04a, MSW^{+}08, MSW^{+}01, MPM09a, MPM09b, NJP^{+}09, Neg03, NT04, PCIB08, PJBT09, PB00, PB03, PM01, RFL04, SP07, SMWS02, SMH^{+}03, STSS05, SHN05, TFF^{+}07, TRW00, TGR02, TRS04, TGR06, TCGR06, TTM^{+}05, VBHT05, VGM00, VRK00, WBH02, WF04, WME05, WM05, ZSH03]. Chiwawa [MPM09b]. Chlorophyll [BHVS06]. Choice [AHH00, ERO^{+}09, GH08, HBB07]. Chub [CPW^{+}06, HRK03a, HKL08, Sto05]. Chum [AGM^{+}09, BCW^{+}09, BUH07, CP04, GHA^{+}02, GAM^{+}08, HDPS07, HPSW08, SHPCQ00, SH04c, UBK04, UGM^{+}07]. Cichlids [BM04a]. Circle [ASD04, CBS^{+}05, Dia08, GH08, Jon05]. Cisco [SEB^{+}09]. Ciscoes [KMG09]. City [TCSS07]. Class [BPBS08, BM00, CDPA09, DD02, IBSC02, JDB^{+}04, MHW00, ON05, PP01, Qui07, RL02, SFUS07, SB04]. Classes [MCS^{+}08, RHBS04]. Classifications [Gol09, PWH02]. Classify [CJPB09, PBM03]. Classifying [MMW08]. Claytor [PMH05]. Clearwater [CBWB02]. Climate [CH02, WHNC09]. Clinch [BB04]. Clipping [PF02, VBHT05, WESW09]. Clips [BSB09, DC06, THE05]. Closed [HBJ05]. Closed-System [HBJ05]. Clove [PP00, SSS07]. Club [ENJO04a, ENJO04b]. Cluster [MTN09]. Cluster-based [MTN09]. cm [BWKS02, GAKB04]. CO [SH07]. Coast [GBFJA07, HS03, MRK08, MP02, NWC^{+}03, BRGO03, Bro02, Dor02, GCQVMAVP09, Han03, HPB05, HPM02, Ian02, Mac02, ML04, OZR08, Ral02, WB03, WCR^{+}05, WGSW07]. Coastal [AMVA09, BGB09, BTHE04, EV09, Mar01, MHS08, RMF^{+}02, RH03, RW06, SSR07, SW04, TBB^{+}07, WSH09]. Coaster [DMMW08, HMSF08, HBFL08, Sch08, SCD^{+}08]. Coasts [VTK^{+}07]. Coastwide [HLV^{+}07]. Cobble [DeV07, MB05a, McH07, RM01]. Cod [FF04, MFRS09, NMJ01, OGS04, RD09, ZK00]. Coded [ARC^{+}07, BT03, BLB^{+}05, BTSB07, Buc01, BSB09, Fri01, HCS09, IF00, LH04a, MPB07, MMLL03, NH01, PCIB08, VSB02, VBHT05]. Coho [BMK^{+}00, BCD02, BT03, BND^{+}07, BFF^{+}02, EH01, FSC03, GG05, HCS09, HOFM06, MSK05, MW07, PPBM04, RFL04, RVM^{+}08, RRCJ08, SBNN03, TTM^{+}05, VTK^{+}07, WMSG09, WHHR01, WHC09]. Coldwater [MFP^{+}08]. Collapsed [RJB07]. Collected [GCR08]. Collecting [IHW02]. Collection [BM01a, FM08b]. Collections [ICBM04, PBB03]. Colonies [RFLN01]. Color [Cur06]. Colorado [BG00, CPW^{+}06, MW04, QGBH06, QJBH07, QBH^{+}09, RC01, SM07a, ZB00, BW06, CHR^{+}08, JMHB08, MRSR01, MWC02, NTSJ03, PR04a, RC01, SWWR04]. Columbia [BERH02, BLA^{+}07, BPB07, GAM^{+}08, HS03, HMS^{+}03, MSW^{+}01, PH09, PJC00, RDN00, WSM01, APS05, BE02, BCM^{+}05, BWJ^{+}08, BM01a, BK09b, BS00, BPS00, CTB07, CW02, DHGP03, EWK^{+}08, GTRC02, GHA^{+}02, GMHX08, HTA^{+}09, HMM^{+}05, JGW^{+}09, KP07, KPRB04, KPJ^{+}04, KBPC08, KRB^{+}04, MS03a, MPBM06, MRJL05, MHRR07, MB05b, MRK08, MPR08, MBM01a, MBM01b, NCK^{+}06, NWC^{+}03, Orr07, PSB^{+}07, PHC^{+}09, PRP00, RB03, RH03, RFLN01, SSP08, SPR05c, TRW00, TRS04, TGR06, TCGR06, WPV^{+}04]. Combinations [SCHK03]. Combine [PCIB08].

Combining [DMMW08, VPR06]. Come [CRBD06]. Comment [BM00, BB08, DeV07, Ger09, Ise07a, MH08a, McH07, MHW00, New06]. Comments [CBS^{+}05, PM01]. Commercial [BSD09, DB09b, DAJB09, GMHF03, Goo07, GMH09, HGHG04, KG08, KQP^{+}09, PW09, RBB08, SL01, SVW^{+}08b, TC01, WP03]. Commercially [CB04, FHSM08]. Common [BLB^{+}05, PSW06, PASP07, PBT05, PEW07, SWMH06, TWH01, TWPP06]. Communities [BG08, CAM06, CJL07, Dra07, FSH06, LCM06b, MOB^{+}08, MFP^{+}08, TPG02, WFW02]. Community [AK05, HS08, LPL04, MB01, MSES08, ML03, NSH^{+}08, PTN06, SB00, SA04, SVPP08]. Comparative [BMP07b, HMW^{+}09, Joh04, SGR02, Wil06]. Compare [KHJC06]. Compared [BI00, KG05]. Comparing [BE01, MPPB02, MHOH06, VPR06]. Comparison [ACBS05, ARK09, BMP07a, BM08, BGH04, BG08, BP02, BABC08, BW06, CSW05, CJL07, CR04, Col02a, CHJ07, DB06, DeV07, DJC^{+}06, EH01, ESPW03, Fay03, FJ02, GW04, GTRC02, GCR08, HMS^{+}03, IHW02, JJ07, KPL09, KC00, LCM06a, LR05, MBS05, MS03a, MB05a, McH07, MSES08, MSES03, Mue03, MEHJ08, MP02, NB05, NN08, NH07, NMJ01, PHB^{+}09, Pin00, PHC^{+}09, RA04, RF00, RUKR07, SIJH06, SH05b, SCHK03, SBSL06, SB06b, TAME03, THE05, WG08, WFW02, WMO08, WCE^{+}09, WDB06, WLG03, You09, Yul00, ZAF05]. Comparisons [MMQ01, PMBM08, PC00, QGSS03, SLG02]. Compensatory [Box02b]. Competition [BM04b, SSC07]. Complete [KPL09]. Complex [BRB09, BG08, HPB05, OJL08, OZR08, PCIB08, RD09]. Complexone [vdWF03]. Component [HHPH03, JDB^{+}04, PHHC02]. Components [BP02]. Composition [AK05, BSRS04, BSB03, CHJ07, CRD05, GGCF08, GMH09, HLO08, MFW^{+}09, MHRR07, SCHK03, SHMR03, Und00]. Compound [DS07]. Computational [Ger07, Kha06]. Compute [GASH05]. Computer [IK05a, Ise07a, MMW08]. Computer-Driven [MMW08]. Concentrations [BHVS06, DHM01, EFS04, SSS07, vdWF03]. Concurrently [OBG^{+}00]. Condition [CTP08, GL08, GSC02b, HN05, Hor01, KC01b, MOH07, MR06, PR04a, PHC^{+}09, PPE03, PNSB01, RV08]. Conditional [CH04, GOG09]. Conditioning [WH04]. Conditions [BLB^{+}05, HH01a, KRB^{+}04, SWG^{+}06, SD06, SMH^{+}03, SBR^{+}05, TPG02]. Configurations [CB01]. Confinement [CCG09]. Confounding [Ach01]. Conjoint [AHH00, GD02]. Conjoint-Based [AHH00]. Connected [HP09]. Connecticut [ENJO04a, ENJO04b, PN05]. Connectivity [SGB07]. Consequences [RG01, SGD09, WHNC09]. Conservation [BK09a, BC05, CBS^{+}05, HWB^{+}03, HGF06, MRC09, RA01, SMU05, WSM^{+}08]. Conserve [MKP05]. Conserving [HK00]. Considerations [FF04, TTRF05]. Consistency [Mir05]. Conspecifics [LKYM09, YMR03]. Constant [CH04]. Constructed [ARC06, BHDA05, BBO03, MPM09a, WSH09]. Construction [CB04]. Consuming [PB03]. Consumption [KC01b, LPL04, LPL05, LS08, LI09, MBS^{+}08, PM08]. Containing [MP01].

Content [HP09, PALdlN05]. Contents [HH01a]. Context [Aus01, HBB07]. Continental [LGA02, LS08, SFS04]. Continuing [BI02]. Continuous [JJ07]. Contrasting [WTH08]. Contrasts [vPP05]. Contribution [CRL08, HRL06, HB05c, IBSC02, Luc02, PGR^{+}04, PF03, RL02, VBJR07, VB03]. Control [BMP07a, BLA^{+}07, HBHB00, LPSL08a, LPSL08b, RSF06, RJB07, ROS04, RW06, SVW08a, SGB07]. Conventional [RPW^{+}08]. Cooling [PFS^{+}06, YDJM09]. Coolwater [LZS^{+}09]. Copper [SSR07]. Core [NTSJ03]. Cormorant [SZL00]. Correcting [HN05]. Correction [Che04]. corrections [Dew09]. Cortisol [SSS07, WBH02]. Cost [LCL05, MPBM06, SMR08]. Cost-Effective [SMR08]. Cost-Effectiveness [LCL05]. Costs [Win04]. Cottids [TBL04]. Count [Cas04, LR05, PBI03]. Counter [SH04c]. Counting [SH04c]. Counts [ACBS05, DRD01, HB05b, HW07, JGW^{+}09, TPG06]. County [IDK03, KD00, WH09]. Cover [CH00b, KSB02, KCA06, PSW06, SHM^{+}08]. Coverage [PH03b, TAM^{+}03]. Crab [AHC^{+}09, AHYC09, CH01, Dew08, Dew09, GMH09, GAKB04, HBS^{+}08, KBF^{+}00, RW06, ZK00]. Crabbers [SEC01]. Crabs [BSSC04, RT09, SEC01]. Crappie [BI02, Box02a, BHVS06, DMM02, DAJB09, GWBH02, IMW02, ISBC02, JB00, Mac03, MC05, MMR08, MD00, MA00b, NDH04, PLW^{+}09, RP09, SH02b, SB04]. Crappies [BWKS02, BPH08, Box02b, BSG05, DMH02, HJ02, IBSC02, IC08, Mac04, MC06, MM02, MAC00, PR05a, PR05b, Pin00, RL02, RCK05, SBIC02, SIB02]. Crassostrea [GBPT08, SBR^{+}05]. Crayfish [IS01, LDMW08]. Crayfishes [WSH09]. Created [MMW^{+}05]. Creek [Bai03, BHO^{+}08, BM02, DHMHJ06, HB06, HK04, MMQ01, VBR04, WP03]. Creel [BP02, DHNH07a, DST00, IWBL07, MC04a, SBSL06]. Crested [SZL00]. Criteria [BLK^{+}08, LL08, RA01, YGG04]. Critical [BTM07]. Critically [MKP05]. Crooked [SHNE05]. Crop [HGHG04]. Cryopreservation [WLHT08]. Cryptic [GFPJ09]. Cucumber [CPE09]. Cull [RT09]. Culture [BSD09]. Cultured [BMP07b, SJS00]. Culverts [BMCB09]. Cunner [NMJ01]. Cured [SLW09]. Current [DMM02, LB09]. Currituck [PR04b]. Curve [CDPA09, PBI03, Sha04]. Curves [BJ04, Chi09, Mac04]. Cutthroat [ACPHT09, BH03a, BGB09, BK09b, BLM^{+}09, BTM07, CKH05, DSR03, DSW01, GZC08, HFY00, HKT00, HK00, Hil02, HK04, KHR00, LH04b, MSR08, MP01, MSES03, NTW^{+}04, PFWC08, RBM00, RMF^{+}02, SSR07, SH07, Sch00, Sch01, Sch03, SGPK09, SSC07, She04, SMU05, SSWC05, WHNC09, YGG04]. Cycle [MD04b, VR05]. Cyprinid [DW08]. Cyprinids [QJBH07].

Dace [RJY08]. Daily [BCC03, BKZA08, CJPB09, DW08, Jac05, RWBG04, SI08]. Dakota [BWKS02, GWBH02, HWB08, ICB04, IWLB05, IWBL07, JEB07, Luc02, SL02b, WPB02]. Dam [AB03, BJ06, CBP07, EYI08, GAM^{+}08, HHWL00, JAAW05, MMW^{+}05, MRSR01, MSRP01, MMVL05, PJC00, PHC^{+}09, SM00a, Sch03, SBT^{+}09, SDPB00, Wer07, BM01a, DIC04, Har03, PKW01, RB03].

Damage [AHYC09, AHC^{+}09]. Dams [APS05, HZG^{+}09, McC01, MDH00, MSWS01, SGP05, SVPP08, NCK^{+}06]. Daphnia [ICB04]. Darkhouse [PC00]. Darroch [UGM^{+}07]. Darter [PF09]. Darters [RA04]. Data [BM01b, BWC08, BWM08, BCW04, CPWM06, DB09b, FDM^{+}06, GG05, GGCF08, GBLC02, GMH09, GCR08, HI01, HLO08, HHB05, Ian02, KFG^{+}01, Loo06, MHO02, MPM09b, MBHK02, PW09, RDBD08, TC01, Vok03, Win04, YS04, YI04]. Davis [Ris00]. Day [How06, SYH^{+}07, TPG06, TGR06, ZS03]. Daylight [MMVL05]. Days [DSW01]. DC [Car01, KWA00]. Debris [BKNB04, BPK^{+}00, KSB02, MFT08, Mar01, Ron03, RH03, SLN07]. Debris-Laden [SLN07]. Decades [SM07a, SBNN03]. Decision [MPPB02, MSP^{+}08, PLP01]. Decisions [SIB02]. Deck [PRHS03]. Decline [MR04]. Declines [SHM^{+}08, YS05]. Declining [DHG02, PBH00]. Decompression [PSKP09]. Decreased [FS06]. Deep [Bai03, BCP03]. Deeply [DP07]. Deepwater [HPB05, SHJ00, ZK09]. Defaunated [RD05a]. Defining [Aus01, LZS^{+}09]. Deflectors [MSWS01]. Degradation [GM05]. Delaware [BPAA03, BSSC04, CK09, CH01, HHAK02, NTH08, VPR06]. Delay [VRK00]. Delayed [BTB^{+}02, DO01, GSWS05, SP07]. Delta [SSR07]. DeLury [GDH04]. Demersal [MFRS09]. Demise [MWR^{+}01]. Demographic [GHM^{+}07, NL06]. Demographics [CPG^{+}08, DHMHJ06, EWD^{+}06]. Demography [HP04]. Denil [SPL02]. Denmark [SVS^{+}09]. Densities [JA07, Yul00]. Density [Box02a, BR06, BHVS06, CP04, CRZ05, FM08a, GW04, GGCF08, HNE04, HBH05, HL08, LDMW08, MC00, Mic09, PTM03, PPF07, RTBF09, SHB^{+}04, SHNE05, SWWR04, Vok03, ZAF05]. Density-Dependent [Box02a, GGCF08, HL08, Mic09]. Dependence [HBH05, PTM03, SHB^{+}04, SHNE05]. Dependent [Box02a, GGCF08, HL08, Mic09, MHF^{+}03, TGR02, VS05]. Depletion [BM00, BPL02, MHW00]. Deposited [HWN09]. Deposition [SI08, ST00]. Depth [AB03, BCD02, FH02, HPM08, ML04, RGP^{+}07]. Derelict [HBS^{+}08, KP07]. Derived [NGWK08, SZW02, SMR08, CRZ05]. Deschutes [NJP^{+}09]. Description [BG08, GCM^{+}03a]. Desert [DBM04, SLEM07, ZC06]. Desiccation [SEGG05]. Design [AAF^{+}06, BWEM08, DS07, DV00, DS04, HWH05, JGW^{+}09, LR08, RO06]. Designatable [VEK09]. Designed [GCM^{+}03a, SVW^{+}08b]. Designing [JR03, PH03b]. Designs [BW08, BABC08, MPPB02, SSP08, SMR08, VPR06]. Detect [LRD^{+}04, Mac04, WVT09]. Detectability [YASG07]. Detecting [AP00b, BW08, BCP03, CJKS05, CJMP08, HSB^{+}00, VSB02]. Detection [APS05, BT03, BPL02, Clo00, CJ02, FH02, GHAS04, NT04, RFLN01, SL09, SLW09, SMW06]. Detectors [CRK^{+}05]. Deter [FCM^{+}09, WFS07]. Determinants [AKKC08]. Determination [BBK09, Cur06, DJSG09, Neg03]. Determinations [Ian02, SH08]. Determine [AFM^{+}06, BHO^{+}08, CWG04, DMMW08, GHM^{+}07, GMH09, JGW^{+}09, Mac02, MFW^{+}09, SVS^{+}09]. Determined [MC05]. Determining [BJ06, DD02, FHN05, PWPC07, Stu06]. Developing [DB09b, HS03, MJ07]. Development [BLK^{+}08, COS^{+}09, DP02, FSLS09, GWSS02, GSC02b, KCA06, MVH^{+}09, PMPS03, PM01, RG01, RP09, RV08, Ric07, SLB^{+}09, TRW00, UF06, WCSJ04]. Developmental [LT01]. Device [FM08b, MRS00, O’M08, WMJBC04]. Devices [CH01, GCN^{+}08, GG09a, SH04b]. Dewatered [MRJL05]. Diadromous [KRDC08]. Diameter [APS05, Han03, SVP^{+}04]. Diamondback [CH01]. Diary [BS01]. DIDSON [BWC08, MS07]. Diel [BPT08, BRB09, BM01a, HPB05, MGH^{+}03b, PCG^{+}01]. Diet [CLB04, GRM07, HK04, NTW^{+}04, PNSB01]. Dietary [BSD09, HOFM06]. Diets [WKAP04]. Differences [Bau02, BMB03, Gol09, JMP07, LB08, MCPM02, QBH^{+}09, SPM06]. Different [BW08, BSS06, BSB09, DGAG08, FWEP08, IMSW03, Pin00, RHB05, SCHK03, SEGG05, TWPP06, TMS05, WPV^{+}04]. Differential [PSKP09]. Differentiation [ICBM04, SJS00, WG03]. Differing [JM00]. Difficult [YS05]. Dike [NH09]. Dilemmas [Sul03a]. Dimensional [Kha06, TGR02]. Dimensions [HMSF08]. Dimorphic [PN03]. Diploid [DST00, NN08, TSMD03]. Dipped [HSB^{+}00]. Direct [DB06, MBHK02]. Directions [WSM^{+}08]. Discard [PRHS03, TKW00]. Discarded [CK09, DP04, MFRS09, O’M08]. Discards [Wil02]. Discharge [BLM^{+}09, CBM00, DHG02, MSRP01, PHC^{+}09, SOR05]. Disconnect [Arl06]. Discrete [MP08]. Discriminates [CLB^{+}07, HCSCT06]. Discriminating [FGSJ05]. Discrimination [GG05, HEB09]. Disease [GGZH08, JMP07, ZB00]. Disinfectants [SVW08a]. Disinfection [MC08]. Dispersal [BB02, DHMHJ06, HM09, HB05c, HM08, MMF^{+}03, PB00, RPM08, YI06]. Displaced [HM08]. Displacement [She04]. Disposition [CK09]. Dissatisfaction [Box02b]. Dissemination [GGZH08]. Dissolved [RDN00]. Distance [FH02, HKH^{+}02, PPF07]. Distances [SL09]. Distinct [HB05a, PMH05]. Distinguish [WCSJ04]. Distinguishing [RRS07, TWP03]. Distribution [AMK^{+}05, BWM08, BRSA07, CP04, DMB09, Dew08, Dew09, DRC03, GCM^{+}04, HWB08, HMSM08, HCFK08, How06, HTGS08, KBPC08, LWHP07, LF05, MMQ01, MEL09, MTM09, MPM09b, NTSJ03, PJBT09, PRP00, RPW^{+}08, SB05a, SCHK03, SHW00, SSWC05, SW04, SYH^{+}07, TCSS07, VTK^{+}07, WF04, WHHR01, ZB00, ZAF05]. Distributions [BM01a, CBP07, DHM01, HRB^{+}04, HH01c, ML06, RAP09, SH05b, WLG03]. District [SMSM01]. Disturbance [CPDS01]. Diver [HB05b]. Divergence [AKARF08, BF06, NPE^{+}06]. Diversion [GM02, ZJ02]. Diversions [GSV^{+}09]. Diversity [HWB^{+}03, HMW^{+}09, LCM06a, LZW^{+}08, NBGK09, SM00a]. Diving [Mue03]. DNA [BK09a, BCM^{+}05, BW06, GMHF03, SDSQ09, TWN^{+}07, WCSJ04, WCR^{+}05].

DNA-Level [BCM^{+}05]. Do [MSK05, Sch02, WP03]. Documentation [PM01]. Does [GH08, Mic05, SNO02, Sch02]. Dogfish [Rul07]. Dolly [BLM^{+}09]. Domestic [Wil06]. Dominance [MJ04]. Dominated [OM04]. Dorsal [BE01, Log07]. Double [BT03, SZL00, Zho02]. Double-Crested [SZL00]. Double-Index-Tagging [Zho02]. Double-Length [BT03]. Down [Box02a]. Downrigger [BCP03]. Downstream [CJMP08, GAM^{+}08, GGG08, GLA00, MDHK08, SMBM03, SAC05, WSM01]. Drainage [BMCB09, DBG07, MBM01b, PRP00, RRB^{+}07, Sch01]. Drainwater [SMSM01]. Drawdown [RWZ09, TGR06]. Dredge [GDH04, GDH05, PAAK07]. Dredging [BPAA03, WC01]. Dreissenid [PM08]. Dried [LL08]. Drift [BFH^{+}08]. Drifting [GOG09]. Driftless [BNV02]. Driven [MMW08]. Drop [BG08]. Dropping [FD01]. Drowned [HP09]. Drum [BPBS08, BSR07, CKSG09, DS08, GBPT08, JDB^{+}02, JDB^{+}04, KSB^{+}08, LPWH01, MC01, RDBD08, SMP04, VC02, VW07]. Drummond [MED08]. Dry [PMBM07]. Dual [BWC08, GM04, MS07, TRS04]. Dual-Frequency [BWC08, MS07, TRS04]. Dual-Marking [GM04]. Duration [AP00b, BSG05, GGG01, Mek04, PRHS03]. during [BSD09, BMCB09, CK09, CHSS08, DHG02, ENJO04a, GHASLC09, HH01a, Hor01, KPRB04, KGM01, MC05, MP02, NLC01, RGP^{+}07, RD05a, SWG^{+}06, SBCH07, WBJ02, WBE^{+}05]. Duty [MD04b]. Dwelling [BPT08, CBM03, LDMW08, PR01, PBT05]. Dynamics [BFH^{+}08, CJM^{+}02, CHSS08, CM05, CBM03, FSH06, FPK07a, GHDM09, HHL^{+}08, HR06, Kha06, KC01a, LPL04, LHSS07, MED08, MTF^{+}08, NL06, PSP^{+}01, PTD03, PTN06, RHBS04, SB00, SMB03, WBE^{+}05, vPP05].

Earlier [BTB^{+}02, EV09]. Early [BBZW04, Fie00, KMCM01, PT07, VTK^{+}07]. East [MRK08]. East-Central [MRK08]. Eastern [ACB07, GMHF03, GBPT08, HP09, HTGS08, KF09, LHSS07, PAAK07, PF03, SBR^{+}05, WPB02]. ebena [LH04a]. Ebonyshell [LH04a]. Ecological [LGA02, PW03, PR04b, RFL04, WP03]. Ecologically [HB05a]. Ecology [GSC02a, HBFL08, QGBS02]. Economic [BDS^{+}02, CHD03, CPW^{+}05, KNL02, Loo06, SL01, SPL03]. Economically [WW08]. Economics [JM00]. Ecoregion [BNV02, GWSS02]. Ecosystem [OJL08, ZK09, LS08]. Ecosystems [PSW06]. Eel [MJ04]. Eels [CSI^{+}04, McC01, MMW08, RSF06, RH04, SOM09, SN07]. Effect [AB02, BPK08, BTS02, BT03, CHCD02, CH01, DP03a, Dia08, DD04, DK04, DCSM01, DSW01, FWEP08, HPM08, HBH05, HWN09, HKFM08, HL09, KLS01, KJR^{+}09, MGAC00, MHB07, OGE01, PHK03, PR05a, PWG02, Pea04, PASP07, PF02, RT09, SHH^{+}04, SSG05, SHMR03, SH02b, SW04, Sto05, TBL04, TWPP06, TMS05, ZM06]. Effective [RA01, SSS07, SMR08]. Effectiveness [ARC06, BK09a, BBZW04, Bin04, CQ05, Cla04, GZC08, GM02, JBG05, LCL05, Mit02, OM04, PSW06, PBM03, RWZ09, RSF06, RHB08, SPM06, VSB02, WFS07]. Effects [ASD04, AP00b, AGB^{+}04, BW08, BBTC07, BBD02, BSG05, BUH07, BSS06, Car01, CCBW01, CAM06, CBP07, CDPA09, Chi09, CCG09, CPE09, CKC^{+}09, CH02, CRC^{+}08, CK09, CFL^{+}03, CGM^{+}06, CH00a, CHJ07, CTB07, CBM03, DP04, DGAG08, DH09, DM07, EWD^{+}06, ENJO04b, FTWS06, FS06, FPK07b, GL08, GMHX08, GTZM09, HMW^{+}09, HTA^{+}09, HBS^{+}08, HLV^{+}07, How06, HA02, IYJS02, IK05b, IWBL07, JSMV07, JED^{+}03, KMB^{+}09, KD00, KJS^{+}09, KQP^{+}09, KCA06, LT01, Loo06, MP04a, MOB^{+}08, MP08, MA00a, MC04b, MFP^{+}08, MTN09, MC00, MHS08, MWN08, Mic09, MMR08, NJP^{+}09, OCW04, PSW06, PWK02, PTD03, Pit02, PPAR06, PH05, RMD02, RD09, Rik00, RFL04, RA04, RRCJ08, RDN00, SMSM01, SIB02, SGP05, SHB^{+}04, SOM09, SWG^{+}06, SMBM03, SM09b, SHD^{+}05, SZL00, SHW00, SMH02, SVPP08, SWWR04]. Effects [SLB^{+}09, SVP^{+}04, TC01, UHRP01, VBHT05, WESW09, WEC07, WH04, WC01, WMB^{+}00, Wil02, WHC09, GH08]. Efficacy [BSB03, Cun04, FSLS09, HBS^{+}05, Mau08, SPL02]. Efficiencies [OSC06]. Efficiency [ARC^{+}07, AK05, BSB03, BBS05, BS09, CBS^{+}05, DS05c, DHNH07a, DMH02, DD04, DK04, GDH04, GDH05, Jon05, LDMW08, MRS00, MDH00, PR01, PAAK07, RWJD04, SMH02, SABW08, WMJBC04]. Efficient [SLN07]. Effort [BG05, BPAA03, BCC03, CGM^{+}06, CW02, CWP03, DP03a, DP03b, DHNH07b, DB09b, Dra07, FS06, GCM^{+}03a, GCM^{+}03b, Han03, LCM06b, LR05, MOH07, MC00, PSWD07, PHRG00, PC00, RHK^{+}03, SMR08, VPR06]. Egg [BBTC07, BRGO03, CRD05, Joh04, LSK^{+}05, LRBJ08, MSK05, Orr07, RWJD04, SM09c]. Egg-per-Recruit [LRBJ08]. Eggs [BSSS02, DV00, FTWS06, FSF06, RWJD04]. Eight [BBJ^{+}00, SLEM07]. Eighteen [DST00]. Elasmobranchs [GHM^{+}07, LGA02]. Elastomer [BLB^{+}05, BSR07, Clo00, CJ02, Cur06, FSK04, HJ06, JRWK08, OV01, PF09, RB09, SLW09, SCPC06, SMW06, THE05, WW04]. Electric [FCM^{+}09, WFW02]. Electrical [Cla04, DMH02, GGG01]. Electrified [FSLS09]. Electrodes [SH04b]. Electrofisher [MK09, MS05]. Electrofishers [PVMR01, WFW02]. Electrofishing [BM00, BPH08, BP01, Car01, Col02a, Cun04, DS05c, DP03a, DP03b, DF07, DMH02, DM04, GFPJ09, HNE04, HKH^{+}02, JJ07, KM00, LCM06b, MOH07, MK09, MC00, MSES08, MLS06, MD04b, Mir05, Mue03, MHW00, OS05, PFWC08, PCG^{+}01, PBB03, RJY08, RHK^{+}03, RH04, RHB03, RD05b, RD06, RUKR07, SSG05, SH05a, SH05b, SH04b, SWPO06, SWWR04, Sta01, TAME03, VWDK00, WWB04]. Electromyogram [BTSB07, GBLC02]. Electronic [BT03, GCM^{+}03a, GCM^{+}03b]. Electroshock [DSW01, HR02, HRK03a, HRK03b, SE00, VGM00, WEC07]. Electroshock-Induced [HR02, HRK03a, HRK03b, SE00]. Electroshocking [GL08, KWA00, SHD^{+}05]. Elevated [DO01]. Elevation [MDH00]. Eleven [Yul00, BH05, HBKJ08]. Eliminate [MB09]. Eliminating [KM00, TC01]. Elkhorn [BM02, VBR04]. Embeddedness [DeV07, MB05a, McH07]. Embryo [CR04, CRD05]. Embryonic [KWA00]. Embryos [KCA06]. Emergence [BBGVM05, MRJL05]. Emergent [CME02, HGF06, RG01, Rad06]. Emergent-Fry [CME02]. Emigrating [NC01, RO06]. Emigration [JKS01, KG00, MMW^{+}05, PFS^{+}06]. Emphasis [IMSW03, Mun02, RAP09, SCPC06]. Empirical [JZMS05, LPL05, RR08]. Encroachment [MSW^{+}08]. Endangered [AFM^{+}06, BW06, HEB09, HRK03b, PH08, SM07a, Sto05, WLHT08]. Endemic [BTM07, LB08]. Endothall [MMS08]. Energetic [CLB04, TTM^{+}05]. Energetics [BTSB07, Kha06]. Energy [HP09, PALdlN05]. Enforcement [WFS07]. Engineered [FSH06]. England [MD04a]. Enhance [MPPB02]. Enhanced [BSD09]. Enhancement [ATM03, LCL05, MS04a, MSW^{+}08]. Enhancing [SRN00]. Enough [HK00]. Entrained [You09, YDJM09]. Entrainment [GZC08, GSV^{+}09, HLV^{+}07, JZMS05, JBG05, PvPR^{+}06]. Entry [JED^{+}03]. Environment [KMB^{+}09, KLW00, PWK02, RTBF09, VH00, WH04]. Environmental [BM02, DRC03, JPB^{+}07, MC00, MEL09, MDHK08, NL06, Pit02, SGG03, SWWR04]. Equation [NHS08, OW09, RR08]. Equations [BWB^{+}00, DBM04, GHAS04, GASH05, HH01b, Ric07]. Equilibrium [GJB09, MTN09]. Equipment [Kol08, Mir05, MB09]. Eradication [PFWC08]. Erie [GMHF03, IK05b, SM07b, TJKB06, VBP08, WVT09]. Errata [Ano02, Ano03a, Ano03b, Ano04]. Erratum [Ano06, Ano08a, Ano08b, RD06]. Error [Cur06, DRD01, HBH05, HW07, NL06, PGR^{+}04]. Errors [KS02]. Escanaba [NHNE00]. Escape [GAKB04]. Escapement [BM08, BPB07, FWEP08, GG05, HI01, NJP^{+}09, PBI03, Sha01, Sha04, Zho03]. Escaping [HPM08]. Essential [SVW^{+}08b]. Establish [QSGS04]. Established [MMQ01, vPP05]. Establishing [JT07, Mir02, ZK09]. Estimate [BG05, BB08, DKM07, GDH04, Ger09, HHB05, HKH^{+}02, KQH09, Loo06, Mac07, MH08a, MOH07, MBS^{+}08, PHRG00, QJBH07, RWJD04, RGP^{+}07, Sha01, Sha04, SBT^{+}09, TFW00, YS05]. Estimated [CHJ07, GGCF08, LF01, MRJL05, Orr07]. Estimates [BM08, BSRS04, BPB07, BE01, Dew08, Dew09, GCN^{+}08, GW04, GDH05, GBFJA07, HNTM00, HSH04, IK05b, LPWH01, LS08, LI09, MC05, MNB09, MMR08, MBD02, MSW^{+}01, NJP^{+}09, NGWK08, RHJ^{+}06, RD05b, RD06, STSS05, SMR08, SB06b, TP06, TMS05, Vok03, VPR06, WMO08, WP08, WSM01, Wil02, YS04, YI04, Yul00]. Estimating [ARC^{+}07, ACBS05, BM00, BMK^{+}00, BM01b, BWC08, Bry00, BSS06, BIBP02, BH03b, CTP08, CHD03, DHNH05, DHNH07a, Dra07, GG05, GCM^{+}03b, HWKN00, HLO08, JCW02, JBHP07, KPL09, KC00, LDMW08, LPL05, LSK^{+}04, LR05, Log07, MR02, MHO01, MZ02, MDKM06, MHW00, NC01, New06, PSB^{+}07, PEW07, PPF07, RHB05, RF00, SLG02, SDSQ09, TTRF05, TPG06, VC02, You09, Zho02, Zho04]. Estimation [Ger07, HHPH03, HAW01, IMSW03, KS02, LL06, MEHJ08, MSP^{+}08, PBI03, PHHC01, PHHC02, RCK05, Sha04, TKW00, VBP08, WTW^{+}04, WPS03]. Estimations [HGP04]. Estimator [UGM^{+}07]. Estuaries [BJ09, KSB^{+}08, WC01]. Estuary [GSV^{+}09, JDB^{+}04, SEC01]. Ethanol [CGP00, PMBM07]. Ethnic [HD02]. Eulachon [SPR05c]. Eurasian [UHRP01]. Evacuation [NH01, PB03]. Evaluate [Bon02, DS04, GM05]. Evaluating [ACPHT09, BND^{+}07, BR06, BCW04, CQ05, DHNH05, DHNH07a, GP03, Ise07b, Jen03, Luc02, MK09, MPM09b, PR01, RWZ09, SLB^{+}09, VRS01, WVT09, WLG03, ZC06, LGA02]. Evaluation [AP00b, ALA07, BPAA03, BLA^{+}07, BRCC03, BCD02, BJ04, BSEM05, BWKS02, BS01, BLB^{+}05, BDLR02, BTSB07, BI00, BSR07, CCBW01, CJPB09, CWP03, COS^{+}09, DH09, DC06, DMM02, DR07, DIC04, FHG^{+}01, FS06, FS07, FCM^{+}09, GG05, GCM^{+}03b, GAKB04, HPSW05, HPSW08, HJ04, HPB05, HWB08, HJ02, HMM^{+}05, ISBC02, IMSW03, JKHE05, JRNS09, JAAW05, KMCM01, LL06, MC05, MS04a, MLS06, MD04a, MRC09, MSP^{+}08, MMLL03, NSH^{+}08, NC01, NH00, OV01, OGS04, PH08, PF09, PJC00, PHHC01, Qui07, RAP09, RJY08, RV08, Ric07, SB06a, SH04a, SSP08, SCPC06, SMF^{+}09, SH08, SLB^{+}06, TP06, VWDK00, WCE^{+}09, Wan06, WKAP04, Wer07, WPV^{+}04]. Evaluations [LM02, O’M08, RMBZ08]. Event [Dau05, Mea04]. Event-Triggered [Dau05]. Events [How06]. Evidence [BMW^{+}02, BTB^{+}02, DB03, FSH06, JPB^{+}07, PBH00, PN03, SWPO06, SM07b, UBK04, WGSW07, WP03]. Ex [TKT02]. Exaggeration [Sul03b]. Examine [NTSJ03, RPW^{+}08]. Example [BCW04, FHN05, HLV^{+}07]. Exclosures [BL08]. Excluder [GCN^{+}08]. Exclusion [BBZW04]. Exemptions [ENJO04b]. Exercise [KRK^{+}02, WH04]. Exhaustive [KRK^{+}02]. Exotic [MHB07, MB09]. Experience [BTB^{+}02, Mek04, SVW^{+}08b]. Experienced [DCSM01]. Experiment [MPBM06]. Experimental [BM04b, Jac05, MPPB02, MJ07, Pea08, SPR^{+}05a, Ste03]. Explain [CRZ05]. Explicit [GDH04, KF09, TGR02]. Explicitly [MTL09]. Exploitation [CW02, DHNH05, FPK07a, FPK07b, GP03, GBFJA07, HBJ05, HPM02, IK05b, IWLB05, JZMS05, MWN08, MBD02, SLEM07, SR02a, SR02b, WT09]. Exploited [MHM09]. Exploring [LRBJ08]. Explosives [FTWS06]. Exposed [DMYC02, TCD^{+}08]. Exposure [AGM^{+}09, DO01, DO02, DP04, FHS06, SRGC05]. Expression [JMP07]. Expulsion [PASP07]. Extending [HGHG04]. Extensively [OBG^{+}00]. Extent [BSG05, BC05, FDM^{+}06, Pow05]. External [CB01, COG^{+}07, HEB09, HKFM08, NT04, SVP^{+}04, WDB06]. Extraction [BK09a, Mic05]. Extreme [KC01b]. extremes [MC08]. Eyeballs [JT07].

Face [AKARF08, Ise07b]. Facility [SOM09]. Factor [MR06]. Factors [BH03a, DPW05, ENJO04a, FSSR08, FD01, GSV^{+}09, HB05a, HFY00, HAW01, HW09, JZMS05, Mac03, Mar01, MC06, MM02, MEL09, MDHK08, MMF^{+}03, NGWK08, PR04a, PRP08, Pit02, QGSS03, RB07, She04, SMWS02, SM09c, SJ03, TP05, YASG08]. Failure [You09]. Fall [BUH07, CBWB02, CB03, CBYB03, CSB03a, CSA^{+}04, DHGP03, GCM^{+}04, GTRC02, GHA^{+}02, GAM^{+}08, GMHX08, HTA^{+}09, HH01a, JR03, JKS01, KLH^{+}08, KHJ^{+}08, MVH^{+}09, MC05, MRJL05, MMVL05, MHF^{+}03, MBM01a, NJP^{+}09, SMH^{+}03, TRW00, TGR02, TRS04, TGR06, TCGR06, VRK00, WM05, GSC02b, PM01]. Fall-Run [KLH^{+}08, MVH^{+}09, WM05]. Fallback [NCK^{+}06, RB03]. Fallen [ACPHT09]. Fallowing [MRW05]. Farmed [Orr07]. Farmington [PN05]. Farms [MRW05, MRK08]. Fast [ON05]. Fast-Growing [ON05]. Fate [BB02, GC08]. Fathead [BBZW04]. Fathom [MDM^{+}04]. Fatty [CRD05, SJS00]. Faucet [MC08]. Fear [HRK03b]. Feasibility [BND^{+}07, GM04, PR03]. Features [CH00b, CRZ05, NBH^{+}03]. Fecundity [LSK^{+}05, NMJ01, SHNE05]. Fed [VB08]. Federal [HGHG04, KRB^{+}04]. Federally [PF09]. Feeding [GRM07, WEC07]. Female [HSB^{+}00, JSL05, JLSL06, KGM01, MPM09a]. Feral [SSB01]. Fermented [BSD09]. Fertile [KG00, KGM01]. Fertilization [BPS00, BMK^{+}01, MPBM06, MHRR07]. Fertilized [BSSS02]. Fidelity [BHM^{+}07, JH01, NN02, PWB04, SJP02, WN04]. Field [BLB^{+}05, FJ02, GBLC02, HHB05, LaV02, PP00, RMBZ08]. Field-Based [RMBZ08]. Fill [SHPCQ00]. Film [ČSB05]. Fin [BE01, BSB09, DC06, HSH04, KHJC06, KQ07, Moh03, PHK03, PF02, SH08, SB06b, THE05, VBHT05, WESW09, WMO08, WTW^{+}04, ZM06]. Fin-Clipping [WESW09]. Fine [DeV07, FWEP08, HWN09, MB05a, McH07]. Fine-Sediment [DeV07, FWEP08, MB05a, McH07]. Fines [LD04]. Fingerling [Buc01, BSB03, BBS05, Clo00, CJ02, Mau08, MCS^{+}08, Mit02, MAC00]. Fingerlings [COG^{+}07, JA07, KH09, Luc02, OBG^{+}00, ZHBB09]. Fingerprinting [SVS^{+}09]. First [SWD08, TWP03, TPG06]. First- [TPG06]. First-Year [SWD08]. Fish [AB03, AP00b, AMVA09, Ano06, AWWG08, AK05, ARPP06, APS05, BW08, BWEM08, BKNB04, BBO03, BL08, BM01a, BG08, BBZW04, BWB^{+}00, Bon02, BB06, BWC08, BR06, BUH07, BBJ^{+}00, BTSB07, Bry00, ČSB05, CTP08, CAM06, CBP07, CDPA09, Che04, CWK09, CSW05, CJL07, Cla04, CJKS05, CWDH03, CRZ05, CRK^{+}05, DMYC02, Dau05, DP03b, DKM07, DP04, DGAG08, DH09, DA05, DP02, DV05, Dra07, FHE05, FSH06, FM08b, FDM^{+}06, FSLS09, FHSM08, GZC08, GW04, GM02, GM05, GFPJ09, GSWS04, GSV^{+}09, HNTM00, HBJE00, HWB08, HLV^{+}07, HKE^{+}01, How06, HS08, HKH^{+}02, IK05a, Ise07a, JEB03, JJ07, JBHP07, KVMM07, Kha06, KMCM01, KRDC08, LCM06a, LCM06b, LPL04, LPL05, LR08, LPSL08a, LPSL08b, LZS^{+}09, MFW^{+}09, MOB^{+}08, MHO01, MKP05, MFP^{+}08, Mar01, Mea04, Mek04, MSES08, MRS00]. Fish [Mir05, MHB07, ML03, MB08, MRK08, NSH^{+}08, NL06, NH07, NH09, OM04, OZR08, PGR^{+}04, PFS^{+}06, PR01, PSP^{+}01, PCG^{+}01, PRP00, PvPR^{+}06, PBB03, RB09, RTBF09, Ris00, RMBZ08, RPW09a, RPW09b, RUKR07, RW06, Ryp08, SB00, SGP05, Sch04a, SPL02, SA04, SGD09, SZL00, SVW^{+}08b, SVPP08, Ste03, SLB^{+}09, SYH^{+}07, SEB^{+}09, Stu06, TPG02, TFW00, TCSS07, UBK04, UF06, VWDK00, WFW02, WP08, WC01, WP04a, Win04, YASG07, CRBD06, UHRP01]. Fish-Based [DP02, DV05]. Fished [PW09]. Fisher [RPD09, SVW^{+}08b, LR05]. Fisheries [ARC^{+}07, ARMB04, Aus01, BM08, BB08, CBS^{+}05, CW02, FPRRHQ09, FCM^{+}09, FHSM08, GP03, GMHF03, Ger07, HBH00, HBH05, HGHG04, HKE^{+}01, ISBC02, JML02, KFG^{+}01, LMA^{+}03, MH08a, MA00b, MB09, MTF^{+}08, New06, PW09, PLP01, PH03a, PH03b, PHHC02, PMPS03, PPSC03, RHJ^{+}06, RR05, SPL03, SS09, STSS05, Sul03a, TKW00, YWM^{+}00, Zho02, Zho04, ZS03, vPP05]. Fishermen [Goo07]. Fishers [Cas04, WP03]. Fishery [AHH00, AWM08, AMVA09, AHC^{+}09, BPAA03, BWJ^{+}08, BRCC03, BP02, BDS^{+}02, BS01, CHD03, CH04, CK09, Col02b, DHNH07a, DB09b, EO03, GCM^{+}03a, GCM^{+}03b, GGCF08, GDH04, GD02, GMH09, GBFJA07, HS03, Han03, HPB05, HHPH03, HBB04, HMM09, JRNS09, KNL02, KRK^{+}02, KBF^{+}00, LRBJ08, LSK^{+}04, Loo06, MB01, MHM09, Mek04, MMLM06, MD04a, MFRS09, MP02, NRJ05, NFQ07, OFM^{+}09, PGR^{+}04, PRSA06, PR04b, Rad03, RBB08, RT09, SNO02, SLG02, SH02a, SL01, SMPM02, SCB00, VBJR07, WVT09, WMJBC04, WB03, dSCT01, vPP05]. Fishery-Independent [WVT09]. Fishes [AAF^{+}06, AFM^{+}06, BG05, BRS02, CCBW01, CHR^{+}08, DBM04, DM04, DS04, DW08, FM09, GL08, HBS^{+}05, HLO08, JMHB08, KD00, LB09, PWPC07, QGBH06, RPJB05, RCB07, RC01, Ron03, SMSM01, SB05a, SM00a, SMBM03, SGB07, SW04, TTRF05, TBB^{+}07, Vok03, WH04, WSH09, WP04b, WLG03, YASG08, ZK09]. Fishing [Arl06, BTS02, CK09, CGM^{+}06, CWP03, CA06, CBM03, DCB09, DAJB09, FD01, GHM^{+}07, Han03, HBJ05, HD02, HBB07, HL09, HB07, Jen03, JBHP07, JPB^{+}07, KPL09, KG05, KM05, LPWH01, MHO02, McC01, MSF^{+}09, MTF^{+}08, MC01, New06, OMH05, RGP^{+}07, SGG03, SCMF02, SMR08, SD05, TWRH07, TF06, VW07, WG08, WW08, WSP02, WPS03, Zho02, Zho04]. FishLength [DO02]. Fishway [Kha06, PHC^{+}09, RB03]. Fishways [DKM07]. Five [KG05, PEW07, RHJ^{+}06]. Fixed [BMP07a, QGBH06]. Fixed-Site [QGBH06]. Flaming [HB05a]. Flannelmouth [MW04, PR04a, QBH^{+}09]. Flat [RMBZ08]. Flat-Plate [RMBZ08]. Flatfish [HPB05, MD04a]. Flathead [AB02, Cun04, DS05a, DS05b, DS05c, DB09a, KWP06, MP08, MHM09, MGH^{+}03b, MM05, MBS^{+}08, SIJH06, SJ08, VR05, WKAP04, WW01]. Fleet [Dia08, MK09]. Float [AHYC09]. Floating [PBB03, RG01, Rad06]. Floating-Leaf [RG01, Rad06]. Flood [BSG05, FSH06]. Flooding [SMBM03]. Floodplain [CJL07, HGF06, RRCJ08, SHN05]. Floodplains [FSH06]. Florida [GBFJA07, ATM03, Ano06, BAC05, BB06, BABC08, BPH08, CAM06, CWK09, CA06, DAJB09, EYI08, MH08b, MCS^{+}08, MP02, MC01, MAC00, NDH04, NN02, SMP04, TAME03, TAM^{+}03, TWH01, TWPP06, WMJBC04]. Flounder [PALdlN05, SVW^{+}08b, WT09]. Flow [AKARF08, BMCB09, CBYB03, CP04, DHM01, GB03, GMHX08, HB00, JR03, JED^{+}03, JAAW05, KVMM07, MMVL05, MJ07, MSWS01, MS04b, SPL02, SHW00, TGR02, Wer07]. Flow-Dependent [TGR02]. Flows [CAM06, KLW00, MSRP01]. Floy [BBO03, Rik00]. Fluctuation [CRC^{+}08]. Fluctuations [BKZA08, GMHX08]. Fluid [Kha06, MFW^{+}09]. Fluorescent [CCBW01, COG^{+}07, Moh03, RA04]. Flushing [SHW00]. Fluvial [ACB07, CKH05, HB05b, Sch00, Sch01]. Flux [MDKM06]. Following [CPSM00, GLA00, ATM03, Box02a, Dew09, FHS06, GL02b, IDK03, KRK^{+}02, MSRP01, PB00, VGM00]. Food [HK04, MHRR07]. Foothills [LH04b]. Footrope [Han03]. Forage [SZL00, TTRF05]. Foraging [SVS^{+}09, WESW09]. Forebay [VRK00]. Forebays [AB03]. Forecasting [CSB00, HPSW05, HDPS07, HPSW08, HVN06, HCS09, Mad08, WMSG09, Zho03]. Forest [SSR09]. Forested [FB07]. Forestry [HL09, SLB^{+}09]. Forests [GWSS02]. Fork [CKH05, HKT00, How06, SGPK09]. Forks [MMQ01]. Form [WLG03]. Formalin [CGP00, PMBM07]. Formation [DW08, RWBG04, RH03, SFUS07, WTW^{+}04]. Formula [You09]. Foster [PFWC08]. Fountain [PF09]. Four [AP00a, BAC05, CKSG09, Col02b, CSB^{+}03b, DW09, DBM04, FHE05, Fri01, HH01c, LM02, OM04, PB00, Sch04a, Sch00, ZJ02]. Fourier [WG03]. Fox [SGP05]. Fragment [LZW^{+}08]. Fragmentation [CHR^{+}08]. Fragmented [Sch03]. Frame [BMP07a]. Framework [MPPB02]. Francis [SJP02, SL02b, Ded07]. franciscanus [SSP08]. Fraser [SPR05c, BPB07, PSB^{+}07]. Free [Kha06]. Frequencies [BM00, MHW00]. Frequency [Bon02, BWC08, Car01, FH02, GM02, MS07, PJC00, RAP09, RCB07, TRS04]. Freshwater [BTI05, CRBD06, GSV^{+}09, HGF06, HD02, LB09, LPL05, MFW^{+}09, PBB03, SGG03, Stu06, WW08, WMB^{+}00]. Fry [BKZA08, CME02, CR04, FWEP08, HMM^{+}05, Joh04, KB04a, LT01, LPB04, Luc02, Mit02, MMF02, RTBF09, SJ03, TBL04, UWBL05, ZJ02, vdWF03]. Ft. [Box02b]. Fullness [PWPC07]. Fully [BSD09]. Fulton [MR06]. Fusconaia [LH04a]. Future [AP00a, BHM^{+}07, MED08]. Fyke [BR06, CJL07, RHB03, RUKR07].

Gap [NTSJ03]. Gape [SJ08]. Gas [AGM^{+}09, BE02, BERH02, RDN00]. Gaseous [SH07]. Gastric [BDLR02, FM08b, HJ04, SMF^{+}09, Wan06, WKAP04]. Gatewell [BM01a]. Gatewells [AB03]. Gear [BR06, CPG^{+}08, CH00a, FSF06, GDH04, HB06, KG08, MP04a, PHB^{+}09, SIB02, AHC^{+}09]. Gear-Specific [CPG^{+}08]. Gears [AWWG08, BS09, MBS05, NH07, RHB05, SH04a, YASG07]. Gene [AKARF08, WM05]. Genera [RA04, RRS07]. General [BPRG08, DS04, Gol09]. Generalized [WMSG09]. Generating [GM02, MS07]. Generation [SPR^{+}05a]. Genetic [AKARF08, BK09a, BF06, BSB03, BTHE04, CKSG09, CCKF^{+}09, DMMW08, DR07, ICBM04, KSB^{+}08, LZW^{+}08, MGGL04, MRC09, NPE^{+}06, NBGK09, PCIB08, RA01, RDBD08, TWN^{+}07, VTK^{+}07, WME05, WHQ^{+}08, WPV^{+}04, WGSW07]. Genetics [MS03b, MWC02, WSM^{+}08]. Genotypes [FGSJ05]. Genotypic [RDBD08]. Genotyping [SDSQ09]. Geographic [ICBM04, JEB03, Mar01, TFW00]. Geographical [FDM^{+}06, NGWK08, RPW^{+}08]. Geographically [ST00]. Geometry [RGP^{+}07]. Geomorphic [BW09, FB07, Mar01]. Georges [ML06, OJL08, GDH05, GHDM09]. Georgia [BMK^{+}00, BJ04, BJ09, MB01, SIJH06, SM09a, Sch04a, SBNN03, WCSJ04]. German [Arl06]. Gertrude [HB06]. Ghost [KP07]. Giant [CPE09, Ron03]. gigas [Cas04]. Gill [AK05, BM08, BFF^{+}02, CK09, Col02a, DAJB09, FCM^{+}09, FRW09, GSWS04, KP07, PR04b, Rul07, VGM00, WVT09]. Gill-Net [BFF^{+}02, CK09, DAJB09, FCM^{+}09, FRW09, PR04b, Rul07]. GIS [CRZ05]. GIS-Derived [CRZ05]. Gizzard [Bau02, IDK03, KD00, SGR02, ZHBB09]. Glacial [ICB04, WPB02]. Glen [MRSR01, MSRP01]. Global [RHB08]. Goals [Ach01, BPB07, CKC^{+}09, PH08, WHB03]. Goby [DJC^{+}06, DTMS06]. Golden [COG^{+}07, COS^{+}09, Stu06]. Gonad [GSC02b]. Gorge [HB05a]. GPP [MK09, MS05, PVMR01]. Grade [LPSL08a, LPSL08b]. Gradient [FCM^{+}09]. Gradients [CAM06, GAM^{+}08]. Grand [PR04a]. Grande [ARK09, CCG09, SSC07]. Granite [Har03, JAAW05, NBN04]. Graphical [JRNS09]. Grass [BBD02, SSB01, WHCL02]. Grassbed [RWZ09]. Grassland [SMSM01]. Grating [MPR08]. Gravel [MSW^{+}08, SHPCQ00]. Grayling [DB06, KB04a]. Grayson [BMK^{+}01]. Grazing [BL08]. Great [HBB04, MOB^{+}08, SEB^{+}09, WHQ^{+}08, ZK09, CBK00, DJL09, DJSG09, HMW^{+}09, KM05, NC01, OGK07, PS02, PHRG00, PWB04, TBB^{+}07, WCE^{+}09]. Greater [KNL02]. Green [ICBM04]. Greenback [HFY00, YGG04]. Gripping [GG09a]. Groundfish [Bro02, JML02, LGA02, Mac02, ML06, Ral02, TKW00]. Grounds [Fie00]. Grouper [GHASLC09, OZR08]. Groups [BMB03, HD02]. Growing [ON05, WP04b]. Growth [ASD04, AW04, BKJ06, BG06, BT03, BAC05, Box02a, Chi09, CH02, CRL08, CB03, DCB09, DB06, DW09, DPW05, DW08, DSW01, EC00, ESPW03, FPRRHQ09, GCQVMAVP09, GL08, GGCF08, GHDM09, GRM07, GSC02b, HB05a, HWN09, HL08, HRL06, HB05c, HLH05, HKL08, HMM^{+}05, IS01, JDB^{+}04, Joh04, KH09, KMG09, KJS^{+}09, KQ07, KWP06, LT01, MC04b, MMH09, MNB09, MMR08, NB05, NN02, NN08, OBG^{+}00, PR05a, PTM03, PH05, PF02, QGSS03, Rik00, RC01, RBM00, SIJH06, SM09a, SHB^{+}04, SL02a, SBCH07, SHD^{+}05, SWD08, SSWC05, SH02b, SEJCC04, Tat09, TSMD03, TP05, VFJ08, VBHT05, VB08, WJKW07, WESW09, Wil06, ZM06]. Grunt [MP02]. Guadalupe [LLCBF07]. Guidance [KH01]. Guides [KG05]. Guiding [PJC00]. Gulf [AKARF08, KG08, dSCT01, AMVA09, Dia08, GCM^{+}03a, GCM^{+}03b, GGCF08, GBFJA07, HVN06, HJC09, LI09, ML06, MMLM06, MP02, OJL08, RD09, RRB^{+}07, WCR^{+}05].

Habitat [ATM03, Ano06, Aus01, BPT08, BHDA05, BCAK07, BKNB04, BFR04, BHO02, Bin04, BB06, BG00, BSSC04, BBJ^{+}00, BCW04, CQ05, CHAO08, CSW05, CRZ05, CM05, DHG02, DS05b, ERO^{+}09, FM08a, FSH06, FPK07b, GMHX08, GGG08, GWSS02, GM05, HB00, HTA^{+}09, HBJE00, HFT08, HK04, HGP04, HA02, JH01, JZMS05, JEB07, KB04b, KS02, KSB02, KG00, KRB^{+}04, KRDC08, KHJ^{+}08, LH04b, Mar01, MS04a, MBM01a, MBM01b, MGH^{+}03b, MM05, NTW^{+}04, OM04, PKW01, PLW^{+}09, Pie04, PPF07, QK00, RPJB05, RWZ09, RJY08, RHB08, RMF^{+}02, RPW09a, RPW09b, SGP05, SW09, SVW^{+}08b, SBR^{+}05, SHN05, ST00, SAC05, SM09d, SMM^{+}07, TPG02, TGR02, TGR06, WW01, ZDK^{+}03, ZC06]. Habitat-Modified [BKNB04]. Habitat-Specific [PPF07]. Habitats [Bar04, BG08, BTK04, DHGP03, GTRC02, MP04a, RRCJ08, TCGR06]. Habituation [SB06a]. Hake [MD04a]. Half [HS08]. Half-Log [HS08]. Halibut [CH02, CH04, DO01, TKW00, WB03]. Haliotidae [RPD09]. Hammerheads [HJC09]. Hand [BJ09]. Hand-Retrieved [BJ09]. Handheld [VSB02]. Handline [SPL03]. Handling [AKKC08, BPK08, BUH07, CCG09, CPDS01, Har03, UBK04, VGM00]. Hanford [TCGR06]. Hard [KRK^{+}02, KQH09, Mau08, QJBH07]. Hardwood [OM04]. Hardwood-Dominated [OM04]. Hardy [MTN09]. Harm [Mir05]. Harvest [AHH00, BMP07a, BSD09, BKD^{+}09, Bro02, CH01, DHNH07b, Dor02, Fie04, HW09, IWLB05, JPB^{+}07, KPL09, KG05, KQP^{+}09, LF05, MC04a, MBS05, MGH03a, Orr07, PC00, PMPS03, PM01, QGP^{+}02, Rad03, Ral02, SPM06, SMR08, Sul02, SABW08, TC01, WFS07, Wil02]. Harvesting [Li00, MP04a, MGAC00, NL00, UHRP01]. Harvests [HVN06]. Hatch [ZHBB09]. Hatched [LPB04]. Hatchery [BMPU03, BSD09, BSRS04, BHO^{+}08, BW06, BG00, BSB09, CKSG09, CLB^{+}07, CSA^{+}04, CWDH03, FSC03, GCM^{+}04, HHAK02, HM09, HCFK08, HMM^{+}05, KSB^{+}08, KLS01, KJS^{+}09, MMQ01, MHS08, MP01, MHF^{+}03, MRC09, MTM09, NRJ05, PJBT09, PB00, PPAR06, PN05, RFL04, RPM08, SE00, SB06a, SSC07, SBNN03, TWP03, TWN^{+}07, WW04, WWB04, WF04, ZSH03]. Hatchery- [HCFK08]. Hatchery-Origin [BHO^{+}08, NRJ05]. Hatchery-Produced [BSRS04, CWDH03]. Hatchery-Raised [RPM08, SSC07]. Hatchery-Reared [BG00, CLB^{+}07, HHAK02, MHF^{+}03, PB00, RFL04, SB06a, SBNN03, WWB04]. Hatchery-Released [KSB^{+}08]. Hatchery-Supplemented [CKSG09]. Hatching [BBGVM05]. Hatteras [BDS^{+}02]. Haul [Win04]. Haverstraw [SMM^{+}07]. Hawaii [PLP01, SL01, SPL03]. Hawaiian [O’M08]. Head [IV05, SGP05, SVPP08]. Headboat [MP02]. Headwater [BGB09, CHR^{+}08, Muh02, QGBH06]. Healing [SE00]. Health [HKE^{+}01, MRSR01, Sch04a, SCB00]. Heated [MC08, SH05b]. Held [BTI05]. Hematological [BPK08]. Hemorrhages [SE00]. Henderson [SCMF02]. Herbicides [SLB^{+}09]. Heron [HBB04]. Herring [Fie00, HKE^{+}01, MRK08, OJL08, SMF^{+}09, TKT02]. Hickory [MH08b]. Hidrostal [HLMB01, WBH02]. Hierarchical [RBB^{+}02]. High [BH03a, BBZW04, Car01, DM07, GM02, JED^{+}03, MC04b, PHHC01, PHHC02, RCB07]. High-Flow [JED^{+}03]. High-Frequency [Car01, RCB07]. High-Frequency-Sound [GM02]. High-Mountain [BH03a]. High-Reward [PHHC01, PHHC02]. High-Velocity [BBZW04]. Highland [DP03b]. Highlands [RPJB05]. Hillsborough [CAM06]. Historic [BW06, WHB03]. Historical [BABC08, GCN^{+}08, MHO02, MHOH06, MSES03, PC00, Rad06]. Histories [Ise07b]. History [CBWB02, GHASLC09, HBFL08, KM05, MCPM02, MEL09, NL00, PTN06, SDSQ09, SDPB00]. Hokkaido [AMK^{+}05, MSF^{+}09]. Holding [PHK03, WH04]. Holdover [PN05]. Home [DS05b, PWB04, SM09d, Vok03, WW01]. Homing [BT03, HHAK02, KLS01]. Homogenization [WM05]. Hood [TWP03]. Hook [AB02, CSB^{+}03b, CBS^{+}05, DD04, DK04, DP07, GH08, HB06, Mek04, OSC06, RGP^{+}07, VW07]. Hook-and-Line [HB06]. Hooked [DP07]. Hooking [AKKC08, BTI05, BVH00, CBS^{+}05, DO01, DCSM01, HW09, JSMV07, LSK^{+}04, RB07, WMB^{+}00]. Hooks [ASD04, Dia08, DP07, Jon05, SH02a]. Hoop [DBG00, FS07, LKYM09, MS02, Sto05]. Hoop-Net [FS07]. Horizontal [HB05b, RMBZ08, Yul00]. Horseshoe [BSSC04, GMH09]. Host [ZB00]. Household [SVW08a]. Hudson [SMM^{+}07, MMKC05, PBH00, YDJM09]. Hugo [BSG05]. Human [HMSF08, RG01]. Humpback [CPW^{+}06, Sto05]. Huron [DB09b, ARC^{+}07, DB03, Fie02, MDM^{+}04, MED08, MEHJ08, PM08, RAH01, WBJ02]. Hybrid [IYJS02, SJS00]. Hybridization [HKT00, KLH^{+}08, MMR08, SWA^{+}08, SGPK09]. Hybrids [GAR^{+}09, QBH^{+}09, WAB^{+}02]. Hydraulic [GAM^{+}08, KLW00]. Hydrilla [TAM^{+}03]. Hydroacoustic [BM08, GW04, HNTM00, TTRF05]. Hydrodynamics [Kha06]. Hydroelectric [AB03, APS05, DHGP03, GM02, HZG^{+}09, MMVL05, SM00a, SBT^{+}09]. Hydrogen [DHM01]. Hydrologic [How06, Mac03]. Hydrology [SB00, SBIC02, SGR02]. Hydrometeorological [SEJCC04]. Hydropeaking [SPR^{+}05a]. Hydropower [HTA^{+}09, MS03a, WSM01]. Hydrosystem [BTB^{+}02, SP07]. Hyporheic [AGM^{+}09, GHA^{+}02]. Hypothetical [DMM02]. Hypoxia [SEC01, SW04]. Hypural [HF00].

Ice [BHDA05, CWK09, LH04b, MB09, PMBM07]. Ice-Water [MB09]. Iconic [DGAG08]. Idaho [CKH05, PH09, SGPK09, CLB04, CHJ07, DST00, DHMHJ06, DBG07, HHL^{+}08, HKT00, HMSM08, HM09, HK04, HW07, KB04b, MOH07, MSES03, MEL09, MDHK08, NBGK09, PH08, PB00, SLEM07, TSMD03, VS05, ZAF05, ZC06]. Identification [BCM^{+}05, BWW^{+}08, BWJ^{+}08, BCW^{+}09, BH05, BWC08, FI01, FHSM08, KSB^{+}08, MS07, TWN^{+}07, TRS04, WPV^{+}04]. Identify [EFS04, MMW08]. Identifying [BPRG08, GAR^{+}09, NMP07, SVW^{+}08b]. If [CRBD06]. II [KSB02]. Illegal [Sul02, WFS07]. Illinois [BHHW02, EWD^{+}06, HAW01, PLW^{+}09, SGP05]. Image [HHWL00, Ryp08]. Images [MMW08]. Imaging [CWG04, EFS04, TRS04]. Immature [MHO01]. Immersible [MB09]. Immersion [BM04a, vdWF03]. Immigration [SGB07]. Immobilization [DMH02, GGG01]. Immobilized [DM04]. Imnaha [HCFK08]. Impact [DSM07, FS00, KP07, McC01, MR06, PPE03]. Impacts [ARMB04, BCC03, BDS^{+}02, CHD03, DHGP03, DC06, DAJB09, ENJO04b, HMM09, JPB^{+}07, MMS08, PT07, SL01, WC01]. Impairment [CPSM00, DP04]. Impediment [CL04]. Imperiled [MKP05]. Impingement [HLV^{+}07, Pea04]. Implant [BI00, BSR07, Clo00, CJ02, Cur06, FSK04, HJ06, IS01, JRWK08, OV01, OGS04, PF09, RB09, SLW09, SCPC06, SMW06, WW04]. Implantation [BG06, GL02b, Tat09]. Implanted [CJKS05, CFL^{+}03, CB01, NH01, SJP05, Tat09, WJKW07]. Implanting [ARK09, SH07]. Implications [AWM08, BC05, BHM^{+}07, BBS05, CSW05, CKC^{+}09, CBM00, CL04, CW02, CA06, DO01, EV09, FHE05, Fie00, GGZH08, GGG08, HKT00, HGF06, HW07, HBHB00, HW09, HH01c, KRK^{+}02, LS08, LI09, MED08, MPBM06, MJ04, MJ05, MS03b, MSF^{+}09, MJ07, NWC^{+}03, NFQ07, NTH08, PSP^{+}01, Rei07, RHGW05, SPM06, SGB07, SYH^{+}07, SEB^{+}09, SEJCC04, SMM^{+}07, VBP08, WHB03, WT09, YASG08, ZHBB09]. Importance [DD02, FSH06, PPBM04, RMF^{+}02, SGG03]. Important [WW08]. Impounded [CSI^{+}04]. Impoundment [BWKS02, BM04b, MMW^{+}05, PLW^{+}09, PPAR06, SMBM03, SH05b, VRK00]. Impoundments [FS07, MS02, MWN08]. Improve [MSK05, Mir05, YWM^{+}00]. Improved [Box02a, HDPS07, LaV02]. Improvement [VC02]. Improvements [Loo06]. Improving [AAF^{+}06, BBS05, MDH00, OM04, SMR08]. In-Channel [ML04]. In-Season [HCS09]. In-Transit [BPL02]. Inch [FHG^{+}01, MA00a, NH00]. Incidence [BE02, BERH02, SGD09]. Incidental [AMVA09]. Incidentally [BFF^{+}02]. Inclined [BBZW04, DS07]. Incomplete [KPL09]. Inconnu [Und00]. Incorporating [IK05a, Ise07a, PBI03]. Incorporation [RAP09]. Increase [EWK^{+}08, GHM^{+}07, GBPT08, SL02a]. Increased [MSRP01]. Increasing [CKC^{+}09, DS05c, HHL^{+}08]. Increment [DW08, RWBG04]. Increments [DB06, MNB09]. Incubating [AGM^{+}09, FTWS06]. Incubator [BBGVM05, DV00]. Incubators [BSSS02, KB04a]. Independent [WVT09]. Index [BMB03, BWM08, BBJ^{+}00, DP03a, DP02, DV05, GDH05, GWSS02, Gol09, HN05, OFM^{+}09, Qui07, Sch04a, Zho02]. Index-Removal [GDH05]. Indiana [KDSF07, LSK^{+}05, ML03]. Indicate [PMH05]. Indicated [CSI^{+}04]. Indicates [MGGL04]. Indicators [CDPA09, GP03, TTM^{+}05]. Indices [DB09b, PBM03, RV08, SCB00, WCE^{+}09]. Individual [CCG09, HA02, IK05a, Ise07a]. Induced [CSB^{+}03b, DO01, HR02, HRK03a, HRK03b, JDB^{+}02, JLSL06, KWA00, KMCM01, Rul07, SE00, WCE^{+}09]. Inducements [TMS05]. Induction [COG^{+}07, COS^{+}09, Moh03]. Industry [Ach01]. Inequality [MWN08]. Inexpensive [CME02, CB04, Ryp08, VH00]. Infected [SD06]. Infection [WAB^{+}02]. Inference [JRNS09]. Inferring [MBHK02]. Infestation [MRW05, MRK08]. Inflatable [MMW^{+}05]. Inflated [AWWG08]. Inflow [DD02]. Influence [AKKC08, AGM^{+}09, Box02b, CBYB03, CPDS01, CBS^{+}05, DHM01, DMH02, DSR03, EM06, EO03, FH02, GWBH02, GAM^{+}08, GM00, HK04, HGP04, Jon05, KCR^{+}05, KC01b, MFT08, Mac03, Mar01, Mek04, MMB06, MH05, PSB^{+}07, PPS03, PCG^{+}01, QGP^{+}02, Rei07, RB03, SNO02, SP07, SR02b, SWD08, SSWC05, SMH^{+}03, VJK^{+}07]. Influenced [RCB07]. Influences [BPT08, CHR^{+}08, DRC03, GC08, MDHK08, SGR02]. Influencing [BRSA07, HFY00, HAW01, HW09, MM02, MMF^{+}03, PRP08, QGSS03, RB07, RDME02, She04, SJ03, TCGR06]. Information [FDM^{+}06, JEB03, NGWK08, RPW^{+}08, TFW00]. Infrared [HHWL00, SH04c]. Ingestion [DCB09]. Inhabiting [SMSM01]. Initial [AKKC08, BBS05, IBSC02, IC08, Rad03, SB06a, WSP02]. Injectable [CCBW01, RA04]. Injected [SMB07]. Injuries [DGAG08, GG09a, KWA00]. Injury [AKKC08, ČSB05, CH00a, CPDS01, CSB^{+}03b, CBS^{+}05, DMH02, DM04, DCSM01, DSW01, HLMB01, HR02, HRK03a, Jon05, Mek04, OSC06, RH04, WWB04]. Inland [CGP00, CRD05, DP02, HK00]. Innovations [Jen03]. Insertion [Buc01, Pie04, SMF^{+}09]. Inshore [MTL09, SNO02]. Insight [VTK^{+}07]. Instream [BHDA05, CQ05, DH09, DV00, FB07, GB03, KLW00, OM04, TPG02]. Insurance [HGHG04]. Intake [Pea04]. Integrated [ARK09, BG06, BGB09, BJ06, Dar03, DB09a, DH09, DMM02, FHSM08, GL02b, IC08, KJS^{+}09, LRD^{+}04, MBHM00, MMB06, MSW^{+}01, PB03, RFLN01, Tat09, WJKW07]. Integrity [DP03a, DP02, DV05, FM08a, Gol09, RHK^{+}03]. Intended [WG08]. Intensity [MP04a, TBL04]. Intensively [OBG^{+}00]. Interactions [FHE05, KGM01, KD00, LGA02, OGK07, PTD03, PTN06]. Interactive [SHD^{+}05]. Intercalibration [TJKB06]. Interconnected [PR05b]. Interdam [Spr05b]. Intergeneric [SWA^{+}08]. Intergrade [NN02]. Interior [BPS00]. Interjurisdictional [FHSM08]. Intermediate [Sta01]. Intermediate-Size [Sta01]. Internal [CB01, FI01, VW01]. Interpolation [SW09]. Interpretation [BS00, CMN08]. Interpreting [Gol09, MHO02]. Interrelationships [SPL03]. Interrogation [CJMP08]. Interval [HGP04, LR05]. Intragastric [KPRB04]. Intraperitoneally [CFL^{+}03]. Intrapopulation [CRD05]. Intraspecific [MC00, SWWR04]. Intrinsic [FSH06, GHM^{+}07]. Introduced [KWP06, NPE^{+}06, SIJH06, SGPK09]. Introducing [RJB07]. Introduction [GWBH02, LF05, PPE03, Sch08, WME05]. Introgressed [LLCBF07]. Introgression [BK09b, PN03, WCR^{+}05]. Inundation [BSG05, SEGG05]. Invasion [She04]. Invasions [PM08]. Invasive [PP00, RSF06]. Invertebrate [BBJ^{+}00, Mea04]. Invertebrates [DM07]. Iodide [HSB^{+}00]. Iowa [FS07, LPL04, LPSL08a, LPSL08b, Mit02, PSP^{+}01, RPW09a, RPW09b]. Iran [EO03]. Irrigation [GZC08, PvPR^{+}06, SDPB00, UWBL05]. Island [HS03, MRK08, MED08]. Islands [BWJ^{+}08]. Isle [LPO01]. Isolation [TKK^{+}01]. Isotope [KHJC06, MFW^{+}09]. issue [Dew09]. Issues [CHD03, WSM^{+}08]. Iteroparity [EWK^{+}08].

J [Dia08, ASD04]. J-Type [ASD04]. J. [SB04]. Jackson [FRW09]. Japan [AMK^{+}05]. Japanese [SWA^{+}08]. Jaw [IK05b]. Jersey [BPAA03]. Jet [JED^{+}03]. Jim [EYI08]. Joaquin [SMSM01]. Jocassee [BFR04]. John [How06, TGR06]. Joseph [DS05a, DS05b]. journal [Dew09]. Junction [CB04]. Justicia [SEGG05]. Juvenile [ASD04, AMVA09, AGB^{+}04, APS05, BPBS08, BERH02, BM04a, BCAK07, BMK^{+}00, BMP07b, BCD02, BM01a, BTK04, BS00, BM04b, BLB^{+}05, BSR07, CSB00, CBWB02, Ded07, DC06, DHMHJ06, DW08, Fri01, GGCF08, GGG08, GRM07, GL02a, HHM05, HGF06, HR02, IF00, JDB^{+}04, JED^{+}03, JAAW05, JBG05, LRD^{+}04, LKS00, MC04b, ML04, MJ05, MHRR07, MRS00, MSF^{+}09, MR06, MRK08, MJ07, MMB06, MBS^{+}08, MSWS01, NBN04, Neg03, OSC06, PMBM07, PMBM08, Pea04, PB03, PSP^{+}06, PJC00, PPAR06, RWZ09, Rik00, RFL04, RC01, RRCJ08, RFLN01, SPR^{+}05a, SBCH07, SMP04, SHD^{+}05, SHW00, SHMR03, SLS^{+}04, SHN05, SAC05, SVP^{+}04, SMM^{+}07, TFF^{+}07, Tat09, TTM^{+}05, VW01, VTK^{+}07, VGM00, VRK00, Wan06, WBH02, WF04, WHHR01, WBJ02, WSM01].

Kansas [GSC02b, QGBS02, QGSS03, QSGS04, SSG05, SGR02, SR02a, SR02b, Sch04b]. Kelp [HKE^{+}01]. Kelt [Wer07]. Kelts [EWK^{+}08]. Kennebec [McC01]. Kentucky [BMK^{+}01, BM02, VBR04]. Kernel [Vok03]. Kettle [BMW^{+}02]. Keys [IK05a, Ise07a]. Kidney [JMP07]. King [Dew08, Dew09, KBF^{+}00]. Kirikuchi [SWA^{+}08]. Kissimmee [ATM03]. Klamath [Bar05, MVH^{+}09, PM01]. Klickitat [NPE^{+}06]. Knowledge [PR04b]. Known [Mac02]. Known-Biomass [Mac02]. Kokanee [BM08, CLB04, FHS06]. Kokanees [BMPU03]. Kootenai [MBM01b, PK01, PKW01, PH08, PH09]. Korea [HKL08].

Laboratory [BLB^{+}05, DCB09, GP05, GBLC02, GG09a, LaV02, RTBF09, VH00]. Laboratory-Acquired [GBLC02]. Lacs [PGR^{+}04, RB07]. Ladders [SPL02]. Laden [SLN07]. Lagoon [AMVA09]. Lahontan [ACPHT09, DSR03]. Lake [ACPHT09, AW04, Ano06, BPK08, BRB09, BSEM05, BB06, BMW^{+}02, BKD^{+}09, BPS00, BSSS02, BHM^{+}07, BMK^{+}01, CHAO08, CME02, CHSS08, CBM03, DB09b, DB03, DMM02, DR07, EC00, FTWS06, FS06, Fie00, Fie02, FPK07a, FPK07b, GGG01, GWBH02, GP03, GMHF03, HHL^{+}08, HEB09, HFT08, HL08, HB05c, HKFM08, HP09, Hor01, HLH05, HV03, IDK03, JM00, KHS05, KF09, KGM01, KD00, LL06, LPL04, LHSS07, LF05, MDM^{+}04, MED08, MR04, MPBM06, MTHH04, MM02, MHRR07, MNS01, Mic09, Mit02, MEHJ08, MSP^{+}08, MAC00, NDH04, NBHH00, NSH^{+}08, NHNE00, NH00, PMH05, PWB04, PSP^{+}01, PTD03, PNSB01, PM08, RAH01, RV08, RHBS04, RM01, RW06, RDME02, SHH^{+}02, SHH^{+}04, SHNE05, SM09b, SM00b, SB05b, SCMF02, SM07b, SN02, SVP^{+}04, UHRP01, VFJ^{+}05, VBJR07, VJK^{+}07, VFJ08]. Lake [VWDK00, VEK09, VS05, WAB^{+}02, WW01, WHQ^{+}08, WBJ02, WHB03, WBE^{+}05, WP04b, WHCL02, AAF^{+}06, ARC^{+}07, ATM03, AW04, BBK09, BMPU03, BSRS04, BSEM05, BKD^{+}09, BPS00, BSSS02, BHM^{+}07, CKC^{+}09, CA06, DB03, EH01, FS00, GTRC02, GRM07, HHL^{+}08, HMSF08, HMM09, HRB^{+}04, HP09, HBHB00, HM08, IK05b, KHS05, KG00, KCR^{+}05, KJR^{+}09, LSK^{+}05, LHSS07, LPO01, MDM^{+}04, MED08, MHM09, MMH09, MTHH04, MJ05, MWR^{+}01, MHS08, MCS^{+}08, MEHJ08, MSP^{+}08, MTM09, NFQ07, NTW^{+}04, PMH05, PTN06, PPE03, RL02, RHBS04, Ris00, RBM00, Sch08, SCD^{+}08, SJP02, SJ03, SM07b, SYH^{+}07, SEB^{+}09, SL02b, TFF^{+}07, TJKB06, VBP08, WVT09, WHB03, WSM^{+}08, YASG07, YASG08]. Lake-Dwelling [CBM03]. Lake-Record [WP04b]. Lake-Spawning [CME02]. Lake-Stocked [HV03]. Lakes [BH03a, BCC03, BPB07, BABC08, BPH08, BP01, BHHW02, CJL07, CLB04, CW02, CM05, DHNH05, DHNH07a, DHNH07b, DP02, DV05, Dra07, EWD^{+}06, GWSS02, GSWS04, HBH00, HNE04, HBJE00, HL09, HMM^{+}05, ICB04, IWLB05, JA07, JKHE05, JKS01, KH09, KG05, Mad08, MGH03a, MC00, MC05, MC06, MMR08, Mue03, NBHH01, NBH^{+}03, NB05, OMH05, OBG^{+}00, PP01, PR05b, PWG02, PWK02, PWH02, PW03, PF03, Rad06, RHB03, RHB05, RDME02, SL02a, SH05a, SHN08, SL09, Ste03, TAME03, TAM^{+}03, TP05, VS05, WH09, WPB02, CBK00, DJSG09, DB09b, MOB^{+}08, NC01, PM08, SEB^{+}09, TBB^{+}07, WCE^{+}09, WHQ^{+}08, ZK09]. Lakeshore [RG01]. Lamprey [CFL^{+}03, FGSJ05, MPB07, MMB06, SM09c, SM07b]. Lampreys [DJSG09, JSL05, JLSL06, KMB^{+}09, KG00, KGM01, LZW^{+}08, MB05b, MPR08, SSS08, SLW09, SMW06]. Lampricide [SSS08]. Lance [TKT02]. Land [HTGS08, MFP^{+}08, SHM^{+}08]. Landing [DGAG08]. Landings [PW09]. Landlocked [SRN00]. Lands [KRB^{+}04]. Landscape [BRSA07, CRZ05, DJL09, RPW09b]. Landscape-Level [DJL09]. Landscape-Scale [CRZ05]. Landscapes [KRDC08]. Lapsed [Sha04]. Large [AK05, APS05, BKJ06, BPS00, BMCB09, CPG^{+}08, Col02a, Col02b, FSLS09, GP03, GB03, HP04, HLMB01, HBS^{+}05, KP07, KSB02, LCM06a, MFT08, MP08, Mar01, MMVL05, MD04a, MBHM00, OBG^{+}00, PRP08, QGBS02, Ron03, RH03, SH05b, SH04a, UF06, VPR06, WTH08]. Large-Diameter [APS05]. Large-Lake [GP03]. Large-Mesh [MD04a]. Large-River [FSLS09]. Large-Scale [VPR06]. Largemouth [ATM03, ARMB04, AWM08, BABC08, BSG05, BM04b, BSB03, BH03b, BBS05, CHD03, CRL08, CPSM00, DMB09, DW09, DSM07, GAR^{+}09, GM00, HJ04, HJ06, HRL06, HB05c, HA02, HM08, KB04b, LB08, LF01, Mac04, MMS08, MC00, MCS^{+}08, MNS01, MTAB08, NB05, NLC01, NN02, NN08, OCW04, ON05, RAP09, RP09, RHHFB00, SB06a, SH05a, SWG^{+}06, SD06, SM09b, SWPO06, SWD08, SKM^{+}03, SN02, TFF^{+}07, TAME03, TAM^{+}03, TCD^{+}08, WN04]. Largescale [Ric07]. Larvae [CGP00, MB05b, MPB07]. Larval [AIHRdSD09, BFH^{+}08, BHVS06, CSW05, DJSG09, DS08, HHM05, IHW02, NH07, NH09, PMBM07, PMBM08, PSP^{+}06, QGBS02, SSS08, SLW09, SMW06, WCSJ04]. Last [WCR^{+}05]. Later [Cla02]. Lateral [RRS07]. Lates [DGAG08, GG09a]. Latitudinal [QGSS03, SWD08]. Laurentian [MOB^{+}08, SEB^{+}09, WHQ^{+}08, ZK09]. Lavage [BDLR02, FM08b, HJ04, Wan06, WKAP04]. Lawrence [BF06, SGD09, SFUS07, MMLM06]. Leaf [ACPHT09, RG01, Rad06]. Lean [WHB03]. Learned [MSW^{+}08]. Learning [ARPP06]. Lehigh [HHAK02]. Length [AP00b, BM00, BM01b, BWB^{+}00, BWKS02, BHL03, BT03, Bon02, BE01, Box02b, CWK09, DMB09, DH09, DTMS06, FHG^{+}01, GHAS04, GASH05, HF00, HN05, HMM09, Hor01, HJ02, HV03, IK05a, IV05, Ise07a, Ise07b, IWBL07, Ker08, LSK^{+}05, LZW^{+}08, LM02, MMQ01, MP08, MA00a, MSR08, MA00b, MHW00, NB05, NH00, NMJ01, OW09, PGR^{+}04, PMBM07, PWG02, RAP09, RHGW05, SL02a, SCHK03, SL02b, Sul02, Und00, VJK^{+}07]. Length- [NMJ01]. Length-at-Age [BE01]. Length-Class [BM00, MHW00]. Length-Related [GHAS04, GASH05]. Length-Stratified [NB05]. Lengths [BSB09, FJ02, MNB09]. Lepisosteidae [FI01]. Less [SSS07]. Lessons [MSW^{+}08]. Level [BKZA08, CRC^{+}08, DJL09, GAM^{+}08, GGG08, OGK07, TAM^{+}03, BCM^{+}05]. Levels [BK09b, DeV07, DM07, FWEP08, MP04a, MB05a, McH07, TWPP06]. Lewis [SM09b]. Libby [PKW01]. Liberalized [LF05]. Lice [MR06]. Licensees [TMS05]. Life [BBZW04, CBWB02, GHASLC09, HBFL08, KMCM01, MCPM02, MEL09, NL00, Sha04]. Lifts [WBH02]. Light [FHS06, HHWL00, PSP^{+}06, TBL04]. Lightly [MHM09]. Lights [HBC08, JBG05, RCB07]. Likelihood [JRNS09]. Limestone [HDB00]. Limit [Ach01, AP00b, BCC03, BWKS02, Box02b, FS06, HB05a, HJ02, IWBL07, Jac05, KBF^{+}00, LM02, MA00a, NH00, ROS04, SL02b, VC02]. Limitations [PSB^{+}07]. Limited [KFG^{+}01, MBHK02]. Limitfor [FHG^{+}01]. Limits [CKC^{+}09, CA06, HMM09, Ise07b, IWBL07, MP08, MA00b, PSWD07, PWG02, PPSC03, RAP09, Sul02]. Limnological [NBHH01]. Limulus [GMH09]. Line [HB06, RRS07]. Linear [Vok03]. Lingcod [DO02, PRHS03]. Linking [ARPP06, BTB^{+}02, KRDC08]. Lip [GG09a]. Lip-Gripping [GG09a]. Lipid [CRD05]. Listed [PF09]. Little [CPW^{+}06, RC01]. Littoral [BG05, Bar04, CSW05, CJL07, DJC^{+}06, HBJE00, PSP^{+}01, PCG^{+}01, RUKR07, Ste03]. Live [BTI05, KSM^{+}03, OSC06, SWG^{+}06, SKM^{+}03]. Live-Release [BTI05, KSM^{+}03, SKM^{+}03]. Live-Well [SWG^{+}06]. Lived [Stu06]. Loading [FWEP08, Mar01]. Lobster [Ach01, MMLM06, SNO02, SLG02, TMS05]. Lobsters [HSB^{+}00, O’M08]. Locate [SPR05c]. Locating [KP07, Pie04]. Location [AB02, HKT00, JSMV07, LSK^{+}04, PR03]. Locations [BBK09, BF06, Fri01, GCM^{+}04]. Lochloosa [TAM^{+}03]. Lock [EYI08, HNTM00, JBG05]. Locks [MDH00]. Locomotory [CPSM00]. Log [HS08]. Logarithmic [MR02]. Logbook [GCM^{+}03a, GCM^{+}03b, PW09]. Logbook-Reported [PW09]. Loggerhead [AHC^{+}09, AHYC09]. Loggerheads [AHC^{+}09]. Logging [Ste03]. Logistic [FDM^{+}06]. Long [Dar03, HR06, JRWK08, MSES03, Stu06, ZAF05]. Long-Lived [Stu06]. Long-Term [Dar03, HR06, JRWK08, MSES03, ZAF05]. Longevity [SLW09]. Longfin [MJ04, MTL09]. Longline [ALA07, BJ09, Dia08, Ker08, KG08, MFRS09, PH03b, RGP^{+}07, TKW00]. Longspine [HS03]. Look [MHO02]. Looking [MDKM06, SRGC05]. Loss [Cur06, GCN^{+}08, IYJS02, IK05b, LD04, RHJ^{+}06, SMP04]. Losses [GSV^{+}09]. Lotic [DS05c, LB09]. Louisiana [GBPT08, SW04]. Louse [MRW05, MRK08, Orr07]. Louver [KH01]. Low [BMCB09, CPW^{+}05, FH02, LKS00, MDH00, MH05, PJC00, SGP05, SPL02, SVPP08, vdWF03]. Low-Elevation [MDH00]. Low-Flow [SPL02]. Low-Frequency [FH02, PJC00]. Low-Head [SGP05, SVPP08]. Lower [KJR^{+}09, MSW^{+}01, SSS07, SAC05, BERH02, BLA^{+}07, Bar05, BG00, DS05a, DS05b, DBG00, FSF06, Har03, HBS^{+}08, HZG^{+}09, HLH05, JZMS05, JAAW05, MSW^{+}08, MW04, MSW^{+}01, NBN04, PPAR06, Rei07, SMSM01, SM07a, SMWS02, Spr05b, SPR05c, VRK00]. Lowland [EM06, JJ07]. LR [Kol08]. LR-24 [Kol08]. Lures [DCB09, DCSM01].

Mackerel [HKL08]. Macquaria [COS^{+}09]. Macroinvertebrate [HMW^{+}09]. Macroinvertebrates [SGP05]. Macrophthalmia [MPB07]. Macrophyte [MSW^{+}08]. Macroscale [BHO02]. Made [MBD02]. Madison [SSWC05]. Magdalena [FPRRHQ09, AIHRdSD09]. magellanicus [GDH04, HR06]. Magnitude [DRD01]. Mail [TMS05]. Main [DHGP03, RDN00, SAC05, Spr05b]. Main-Stem [DHGP03, RDN00]. Maine [ML06, OJL08, Ach01, BBJ^{+}00, GTZM09, HZG^{+}09, LI09, McC01, RD09, SNO02, SABW08]. Maintaining [Hil02]. Major [BBK09]. Make [YS05, Dew09]. Malaclemys [CH01]. Male [CPSM00, JSL05, KGM01, MJ04, PJBT09, YMR03, ZSH03]. Manage [Rad03, SL02b]. Managed [GSV^{+}09, KRB^{+}04, NBGK09]. Management [Ach01, Aus01, BLA^{+}07, BWJ^{+}08, BRCC03, BI02, BSS06, BTHE04, Cas04, CKC^{+}09, CP04, CW02, CWP03, DO01, FF04, FHSM08, Ger07, GGG08, GD02, Goe02, GSC02a, HGF06, HKE^{+}01, HB07, ISBC02, JPB^{+}07, JM00, JML02, Ker08, LMA^{+}03, MP04b, MKP05, MJ04, MS03b, Mun02, NFQ07, NTH08, PLP01, PCIB08, PH03a, PH05, PM01, PPSC03, QSGS04, QBH^{+}09, SIJH06, SIB02, SHB^{+}04, SCD^{+}08, Sul03a, VBP08, WT09, WB03, WHCL02, YS05]. Managing [CH04, SWMH06, Sul03a]. Mandible [IV05]. Manipulation [Bin04, HA02]. Manistee [CHAO08]. Manitowish [WH09]. Maps [FJ02]. Marcos [PF09]. Margaret [RBM00]. Margin [DRC03, WF04]. Marginally [BKJ06]. marginatus [O’M08]. Marine [FSK04, HBS^{+}08, HCS09, Li00, MP04a, Orr07, PR03, SH02a, SS09, SFS04, WG08, WP03]. Mark [BPL02, CPWM06, Cur06, DS08, DHNH05, GM04, HLO08, MC05, MZ02, New06, RD05b, RD06, TP06, Zho02, Zho04]. Mark-Recapture [RD06]. Mark-Selective [New06, Zho02, Zho04]. Marked [AB03, IS01, MPB07, MRS00, MMF02, RB09]. Markers [WPV^{+}04]. Market [CBK00]. Markets [WHC09]. Marking [BM04a, BPL02, BSR07, CCBW01, COG^{+}07, COS^{+}09, GM04, HOFM06, IBSC02, LPB04, Luc02, Mau08, MB08, MMLL03, OV01, PF09, SCPC06, SMB07, THE05, vdWF03]. Marks [COG^{+}07, HKFM08, JDB^{+}02, Moh03, NT04, RA04, THE05]. Marlin [Dia08, GH08]. Martin [HM08]. Maryland [SHM^{+}08]. Marys [Fie00]. Mass [Luc02]. Massive [GSV^{+}09]. Masu [AMK^{+}05, MHF^{+}03, MSF^{+}09, SWA^{+}08]. Mating [Dew08, Dew09, JLSL06]. Mattaponi [BHO02]. Matter [GH08]. Mattole [WHHR01]. Maturation [MHO02, NN08, SJ05, TPC^{+}09, WCE^{+}09]. Maturational [EFS04]. Mature [LKYM09, MHO01, PJBT09, YMR03]. Maturing [HF00]. Maturity [HKL08, SHNE05]. Maximum [CJPB09, DSR03, GHDM09, Pow05]. May [She04]. Mayfield [Tip01]. Mazatlán [GCQVMAVP09]. McConaughy [PPE03]. Mean [BM01b, GSWS04, NB05]. Means [BMK^{+}00, Pie04, SB06a]. Measure [GCM^{+}03a, MR06]. Measured [BBZW04, FJ02]. Measurement [DeV07, HBH05, KS02, MB05a, McH07, PGR^{+}04]. Measurement-Based [DeV07, MB05a, McH07]. Measurements [BW09, CH00b, IV05, TKT02]. Measures [CP04, WLG03]. Measuring [CJMP08, HWH05, Ker08, KLW00, LD04, MRS00, PALdlN05]. Mechanical [UHRP01]. Mediated [JMP07]. Medium [MZ02]. Medium-Sized [MZ02]. Melania [MHB07, MB09]. Melanoides [MHB07]. Menhaden [HVN06, HHM05, HLV^{+}07, dSCT01]. Menominee [BKD^{+}09]. Meredith [Mun02]. Meristic [QBH^{+}09]. Merrimack [Spr05b]. Mesh [IHW02, JB00, MD04a, PSW06, RUKR07, SCHK03]. Mesocosm [MJ07]. Mesohabitat [KVMM07]. Mesohaline [SBR^{+}05]. Meta [Dor02]. Meta-Analysis [Dor02]. Meter [PR01]. Method [BDLR02, Cas04, CJPB09, FM09, GCM^{+}03b, GASH05, Ger09, HJ04, HN05, MSP^{+}08, PWPC07, PHHC01, ROS04, SW09, Wan06]. Methodological [BSRS04]. Methodologies [KPL09]. Methodology [MMB06]. Methods [ACBS05, ARK09, Bry00, Chi09, CHJ07, COS^{+}09, DJSG09, DJC^{+}06, EH01, IMW02, KP07, KC00, LL06, LCM06a, LR05, MEHJ08, NC01, PBT05, PALdlN05, Sha04, SBSL06, SB04, Zho02]. Methyltestosterone [LKS00]. Metrics [BABC08, DHNH07a, MRSR01, SPM06]. Mexican [MSES08]. Mexico [AKARF08, AIHRdSD09, Dia08, FPRRHQ09, GCM^{+}03a, GCM^{+}03b, GCQVMAVP09, GGCF08, GLA00, GHASLC09, HJC09, KG08, RPD09, WCR^{+}05, dSCT01]. Michigan [DB09b, PM08, AAF^{+}06, BSRS04, BSEM05, BKD^{+}09, BHM^{+}07, CHAO08, CKC^{+}09, CRZ05, DS05a, DS05b, EH01, GRM07, HL08, HRB^{+}04, HP09, Hor01, LL06, LSK^{+}05, LZS^{+}09, MR04, MTHH04, MSP^{+}08, NB05, PNSB01, RHBS04, RM01, SHH^{+}02, SHH^{+}04, SL02a, SM00b, SB05b, WME05, WHB03, WBE^{+}05, Wil06]. Microchemistry [CSI^{+}04]. Microhabitat [ACB07, WSH09, WTH08]. Microsatellite [BWJ^{+}08, BCW^{+}09, FGSJ05, MGGL04]. Microsatellite-Based [BWJ^{+}08]. Microsatellites [BWW^{+}08, NPE^{+}06]. Microwire [MAC00]. Mid [RPJB05, WGSW07]. Mid-Atlantic [RPJB05, WGSW07]. Midas [BM04a]. Middle [GMH09, PHB^{+}09, TPC^{+}09, UGM^{+}07]. Midlake [MTHH04]. Midorbital [HF00]. Midwater [AAF^{+}06]. Midwest [GM05, TPG02]. Midwestern [CPG^{+}08, SVPP08]. Migrant [HMS^{+}03, MSW^{+}01, PHC^{+}09, SMWS02, WSM01]. Migrating [LRD^{+}04, OGE01, SM00a]. Migration [BBO03, CL04, DBG07, EV09, HMS^{+}03, KPRB04, KPJ^{+}04, KMB^{+}09, KCR^{+}05, MS03a, MDHK08, MTM09, PK01, PPAR06, SAC05, SJP02, Und00, ZSH03]. Migrational [CSB03a]. Migratory [CSB00, GTZM09, Neg03, SSR07, SWMH06, TBL04, TRW00, TGR06, VRK00]. Mille [PGR^{+}04, RB07]. Milltown [Sch03]. Minimal [LD04]. Minimization [Pow05]. Minimum [AP00b, BWKS02, Box02b, FHG^{+}01, HMM09, HJ02, IWBL07, KBF^{+}00, LM02, MP08, MA00a, MSRP01, NH00, SL02b]. Minnesota [BNV02, BP01, BMW^{+}02, CM05, DP02, DV05, Dra07, FPK07a, FPK07b, GP03, GSWS04, Ise07b, Jac05, JA07, KG05, MC00, MC05, MC06, MMR08, Neg03, PP01, PR05b, PC00, PTN06, Rad03, RHJ^{+}06, Rad06, RB07, SL09, TP05]. Minnow [ARK09, BBZW04, Bry00, CCG09, HRK03a]. Miramichi [SEJCC04]. Mississippi [KQP^{+}09, KQH09, PHB^{+}09, Pit02, SMPM02, TPC^{+}09, ZDK^{+}03, DMM02, DR07, OMH05, RPW^{+}08, RRB^{+}07, SGG03, SJ05]. Missouri [BFH^{+}08, DBG00, DR07, SJP02, AWWG08, BRSA07, Col02a, Col02b, DBG00, ESPW03, Mea04, MWN08, QGP^{+}02, SBSL06, SZW02, TC01, VR05]. Misspecified [Wil02]. Mitigation [HDB00]. Mitochondrial [BW06, GMHF03, TWN^{+}07]. Mixed [BRCC03, BKD^{+}09, GMHF03]. Mixed-Stock [BKD^{+}09, GMHF03]. Mixture [MTN09]. mm [BG06, Box02b, CRL08, CRK^{+}05, HJ02, JB00, Tat09]. Mode [PAAK07]. Model [CPWM06, CWP03, DB09b, DA05, FHN05, FDM^{+}06, GMHX08, HDPS07, HTA^{+}09, HFT08, HRK03b, JCW02, KF09, Kha06, LRBJ08, LPL05, New06, PLP01, PK01, PVMR01, PMPS03, PM01, SLG02, SBT^{+}09, SRN00, YGG04, Zho04]. Model-5.0 [PVMR01]. Model-7.5 [PVMR01]. Model-Based [CWP03, DB09b]. Modeled [SZL00]. Modeling [AHH00, AWWG08, BPK^{+}00, CRC^{+}08, CW02, JMHB08, KLW00, MBS^{+}08, PFWC08, RDN00, SHN08, TGR02]. Models [BB08, BBJ^{+}00, GHM^{+}07, GAR^{+}09, GH08, HPSW05, HPSW08, HWB08, JPB^{+}07, Mac02, Mac07, MH08a, MTL09, NL06, PRP00, RPJB05, WMSG09, WTH08]. Modern [WHB03]. Modification [FSF06]. Modifications [ZK00]. Modified [BKNB04, BG08, BCP03, CME02, FS06, LPSL08a, LPSL08b, RC01, RO06]. Modulation [SBCH07]. Mokelumne [MSW^{+}08]. Molecular [WPV^{+}04]. Mollusk [FPRRHQ09]. Monitor [SCB00]. Monitoring [ACBS05, ALA07, Dau05, DV00, DS04, FM09, LR08, MSW^{+}08, PW09, QGBH06, RJY08, RHGW05, SMM^{+}07, UGM^{+}07, YASG08]. Monofilament [MHOH06]. Montana [BFH^{+}08, BMCB09, MBM01a, MBM01b, Muh02, MGH^{+}03b, MM05, MBS^{+}08, SM00a, Sch00, Sch01, SPL02, Sch03, She04, SSWC05]. Montane [MSES08, PRSA06, SA04]. More-Precise [PCIB08]. Morphological [AKARF08, GAR^{+}09, SGPK09]. Morphology [HF00, LB09, MB05b]. Morphometric [QBH^{+}09]. Mortality [ARMB04, AWM08, AKKC08, BWB05, BKZA08, BTI05, BVH00, BAC05, Box02b, BFF^{+}02, BTB^{+}02, CDPA09, CH00a, CSB^{+}03b, CA06, Dar03, DO01, DO02, DP04, DBG00, DMH02, DM04, DD04, DK04, DP07, DCSM01, ENJO04a, FSSR08, GGCF08, GHASLC09, GRM07, GSWS05, GBFJA07, HB05c, HRK03a, HRK03b, HW09, IYJS02, IBSC02, IC08, JSMV07, JBHP07, JPB^{+}07, KWA00, KMCM01, KWP06, LPWH01, LSK^{+}04, Mac07, MHO02, MP08, Mic05, MMKC05, Mir02, MTL09, MMB06, MP02, MC01, NLC01, New06, OSC06, OZR08, PRHS03, QSGS04, RW03, RB07, RB09, Rul07, SIJH06, SP07, SB06a, SWG^{+}06, SR02a, SFS04, SM06, TWH01, TWRH07, TKW00, UBK04, VW07, WWB04, WMB^{+}00, WSP02, WPS03, WT09, Wil02, YI04, Zho02, Zho04]. Motility [BRS02]. Motivation [Arl06]. Motorboat [Rei07]. Mountain [BH03a, BLK^{+}08, MEL09, MNS01, MH05, SWA^{+}08, SN02, MLS06, PN03]. Mountains [BWEM08, DHM01, KM05, PPS03]. Mouth [HP09]. Movement [BKJ06, BCAK07, BRB09, BB04, BLM^{+}09, CSI^{+}04, CPE09, CJMP08, CSB03a, DS05b, GCM^{+}04, GL08, HGP04, HS09, HV03, JEB07, JDB^{+}04, KG00, KHJ^{+}08, MP01, MBM01a, MGH^{+}03b, MM05, PR05b, PWB04, PBT05, PN05, RHHFB00, RD05a, SSR07, SM09d, Tip01, VBR04, VS05, VR05, WH09, ZDK^{+}03]. Movements [ARC^{+}07, BWB05, CKH05, DHG02, FM09, HZG^{+}09, JH01, JZMS05, KHS05, LH04b, NFQ07, PMH05, RRB^{+}07, SHH^{+}02, Sch01, Sch03, SA04, Spr05b, SSR09]. Moxostoma [MRC09]. MS [SSS07]. MS-222 [SSS07]. Mt. [Sch02]. Much [HK00]. Mud [VB08]. Mudflats [SABW08]. Mudsnail [ROS04]. Mudsnails [SVW08a]. Mullet [BWB05, LL08]. Multifilament [MHOH06]. Multilevel [GM05, PLP01]. Multilocus [FGSJ05]. Multiobjective [PLP01]. Multiple [AWWG08, CCG09, CBK00, CWDH03, HHPH03, KM00, MSR08, NBGK09, OS05, PHHC02, RPW09b, SGP05, SBT^{+}09, SVPP08, VWDK00, YASG07]. Multiple-Component [HHPH03, PHHC02]. Multiple-Pass [OS05]. Multiple-Release [SBT^{+}09]. Multispecies [ALA07, JML02]. Multistage [TFW00]. Mummichogs [SCPC06]. Muscle [KHJC06]. Muskegon [HP09]. Muskellunge [Fay03, FPK07a, FPK07b, MA00a, MP04b, OSC06, RDME02, SH05a, Tip01, WJKW07, WESW09, WH09]. Mussel [PM08]. Mussels [LH04a]. Mysis [CLB04]. Myxobolus [BLA^{+}07, GGZH08, NTSJ03, SLS^{+}04, WAB^{+}02, ZB00]. Myxospores [GGZH08].

Nass [BCM^{+}05]. National [FPK07a, FPK07b, HMW^{+}09, KM05, LPO01, SMP04]. Native [CHR^{+}08, GAG^{+}01, HKT00, HDB00, HTGS08, MBS^{+}08, NPE^{+}06, PPS03, PFWC08, PTN06, QBH^{+}09, RC01, SIJH06, SMBM03, She04, SEB^{+}09, WH04, WCR^{+}05, ZK09]. Natural [BWEM08, BKNB04, BFH^{+}08, BRSA07, CA06, GHASLC09, HCFK08, JBHP07, JPB^{+}07, LPWH01, MC06, PWG02, PWH02, PW03, PJBT09, TWRH07, VH00, WT09, Wil02]. Natural-Origin [HCFK08]. Naturalized [MH05]. Naturally [BBTC07, CRD05, RD05a]. Navajo [GLA00]. Navigable [NH07, NH09]. Navigation [Fie00, JBG05, MDH00]. nDNA [WCR^{+}05]. Near [HZG^{+}09, RGP^{+}07]. Near-Surface [RGP^{+}07]. Nearshore [Dra07, HBJE00, MC01, PSKP09]. Nebraska [Bau02, HJ02, PWG02, PWK02, PWH02, PW03, PPE03]. Necessary [PBT05]. Nechako [HI01]. Needed [BG05, HKH^{+}02, MPR08, PHRG00]. Needs [PH09]. Nematode [MRSR01]. Nest [PLW^{+}09, RP09, WN04]. Nesting [CPSM00, SWPO06]. Net [BG08, BFF^{+}02, CK09, DO01, DO02, DAJB09, FSK04, FS07, FCM^{+}09, FRW09, MC06, MTM09, PR04b, Rul07, SSG05, Sto05]. Net-Pens [FSK04, MTM09]. Nets [AK05, BM08, CSW05, CJL07, DGAG08, DBG00, GSWS04, GOG09, JB00, KP07, LKYM09, MHOH06, MS02, PMBM08, PBT05, SH05b, SCHK03, WVT09]. Netting [BR06, Col02a, MC05, RH04, RHB03, RUKR07]. Network [MTF^{+}08]. Networks [MTF^{+}08, Zho03]. Neural [Zho03]. Neuse [SEC01]. Nevada [HMW^{+}09]. Newburgh [SMM^{+}07]. Newfoundland [SPR^{+}05a]. Newly [LPB04, MMQ01]. Niagara [HLH05]. Niche [HHB05]. Nicola [CTB07]. Night [GL02a, TPG06]. Nile [HBJ05]. Nine [BK09b]. Nitrophenol [SSS08]. Nomogram [CJPB09]. Noncompliance [PGR^{+}04]. Nongame [BWB^{+}00]. Nonindigenous [SN07]. Nonlethal [BPL02, KHJC06, NT04]. Nonlinear [Mac07]. Nonmarked [MMF02]. Nonnative [HKT00, HK04, JMHB08, LWHP07, MBS^{+}08, PPS03, PFWC08, SB05a, SMBM03, She04, WP08]. Nonwinterkill [ICB04]. North [BDS^{+}02, GAG^{+}01, HJC09, MGAC00, MRC09, BWB05, BPBS08, CLB04, Goe02, HB00, HW07, How06, NDH04, OZR08, QGSS03, RT09]. North-Central [NDH04]. Northeast [SFS04, LS08]. Northeastern [LGA02, SLS^{+}04]. Northern [AKKC08, BF06, FSC03, GWSS02, HP04, LEJ06, LI09, MP04a, MGH03a, MS03b, MBS^{+}08, NH01, PW03, Pea04, Pea08, PB03, PC00, PTM03, PTD03, Pie04, PSP^{+}06, SH05a, SFUS07, WFS07, AKARF08, BRCC03, DHNH05, DHNH07a, DHNH07b, FHG^{+}01, Fay03, FHN05, GG05, HBH00, HNE04, HBH05, HBJ05, HL09, JKHE05, Mad08, MGH03a, MW07, NBHH01, NBH^{+}03, PBM03, RHHFB00, RHB03, RHB05, RDME02, SHB^{+}04, SHN08, TPG02, VS05]. Northwest [CJM^{+}02, MFRS09, RBB^{+}02, Ron03]. Northwestern [BPK^{+}00, SHH^{+}02, SHH^{+}04]. Norwegian [AHH00]. Nosed [IBSC02]. Note [BM01b, WTW^{+}04]. Nova [GM02, FF04, MFT08]. Novel [BWM08]. Novice [DCSM01]. Nuclear [WCR^{+}05]. Nucleotide [STSS05]. Nuisance [Ris00]. Number [MPM09a, PH03b]. Numbers [Dia08, NC01]. Nursery [BHO02, BPB07]. Nutrient [PRSA06].

O [MFW^{+}09]. Objectives [QSGS04]. Observation [NL06]. Observations [SOM09, SCB00, YH04]. Observer [GGCF08, PH03b]. Observers [PHHC02]. Obtained [SH05b]. Occupied [WP08]. Occurrence [DSR03, FDM^{+}06, MH05, NBH^{+}03, RPJB05]. Ocean [GDH04, VTK^{+}07, WF04, CJM^{+}02, HJC09, Ian02]. Ocean-Type [WF04]. Oconee [WCSJ04]. Ocular [FHS06]. Off [NDH04, BWJ^{+}08, FF04, GCQVMAVP09, GBFJA07, HPB05, HKL08, MRK08, MC01, OZR08, SW04, VTK^{+}07, WB03]. Off-Season [NDH04]. offs [PH03b]. Offset [ASD04]. Offshore [DB03, LCM06a]. Ohio [BCAK07, BHVS06, FH02, HNTM00, WVT09]. Oil [PP00, SSS07].

Oklahoma [Box02b, BSG05, Cun04, LF05, SZL00]. Oligochaete [ZB00]. Oligotrophication [MB01]. On-Site [AFM^{+}06, CH00b]. One [AP00a, HHPH03, PR01, SD05, Wil06]. One-Square-Meter [PR01]. Oneida [VFJ^{+}05, VBJR07, VJK^{+}07, VFJ08]. Ontario [HL09, AW04, KF09, KJR^{+}09, LMA^{+}03, MJ05, PBM03, PF03, SCMF02, SJ03, SKM^{+}03]. Ontogenetic [NTW^{+}04]. Open [GDH04]. Open-Ocean [GDH04]. Opening [CGM^{+}06, ZS03]. Operated [OMH05]. Operating [SMH02]. Operation [CBM00]. Operational [PHC^{+}09]. Operations [HTA^{+}09, MHS08]. Opinion [EWD^{+}06]. Optimal [Bro02, EO03, FHN05, JA07, JR03, LR08]. Optimization [CJKS05]. Optimize [HKE^{+}01]. Optimizing [CWDH03, LCL05]. Optimum [EC00, Li00]. Option [PFWC08]. Orange [TAM^{+}03]. Order [BLM^{+}09, TPG06]. Oregon [ACB07, Bai03, BHO^{+}08, BKZA08, HCFK08, HKH^{+}02, KSB02, ML04, NJP^{+}09, Rei07, RHK^{+}03, RVM^{+}08, SMB03, SDPB00, SLS^{+}04, TWP03, VTK^{+}07, ZSH03]. Oreille [DBG07, HHL^{+}08, KB04b]. Organisms [HBS^{+}08, You09]. Organochlorine [SVS^{+}09]. Orientation [Arl06]. Origin [BHO^{+}08, HCSCT06, HCFK08, NRJ05, PJBT09]. Osmotic [COG^{+}07, COS^{+}09, Moh03]. Other [BG08, JMP07, KG05, KWP06, MHB07, TFF^{+}07]. Otolith [BH05, CSI^{+}04, Chi09, CLB^{+}07, DB06, MEHJ08, SH08, vdWF03]. Otoliths [BIBP02, BH03b, DB06, DW08, HWKN00, HAW01, LF01, MNB09, RDBD08, RWBG04, Ryp08, SI08, Stu06, SB06b, WMO08]. Otter [BPH08, Pin00]. Outfalls [JED^{+}03]. Outlet [JKS01]. Outline [WG03]. Outplanted [BHO^{+}08]. Output [MK09, MS05]. Outputs [PVMR01]. Overfished [BMP07a]. Overlap [HK04, SN02]. Overwinter [MHF^{+}03, PF02]. Overwintering [CQ05, KBPC08]. Oviduct [Pie04]. Ovulating [JSL05, JLSL06]. Owikeno [MWR^{+}01]. Oxygen [MFW^{+}09]. Oxytetracycline [BM04a, BPL02, DS08, IBSC02, JDB^{+}02, LPB04, Luc02, Mau08]. Oxytetracycline-Induced [JDB^{+}02]. Oyster [BPAA03, PAAK07, SBR^{+}05]. Oysters [GBPT08]. Ozark [DP03b, HA02, QK00, WW04, WWB04].

Pacific [BCW^{+}09, CH02, CH04, CFL^{+}03, DO01, HWB^{+}03, HPM08, HKE^{+}01, Ian02, JED^{+}03, KMB^{+}09, LZW^{+}08, MPPB02, MPB07, MRW05, MRK08, MPR08, MMB06, PSKP09, RBB^{+}02, Ron03, SLW09, SMW06, TKT02, TKW00, WB03]. Paddlefish [BCAK07, DBG00, FSF06, Fri01, PS02, SJP02, ZDK^{+}03]. Pallette [NH00]. Pallid [BFH^{+}08, GGG08, HSH04, KH01, TKK^{+}01, Wan06, WLHT08]. Palmetto [Mau08]. Pamunkey [BHO02]. Panasoffkee [CA06]. Pandalus [LI09]. Panel [MD04a]. Panfish [CM05]. Panjet [DC06]. Panulirus [O’M08]. Parallel [RPW^{+}08]. Parameter [Fie04, NL06]. Parameters [MCPM02, NB05, YS05]. Parasite [ZB00]. Parasitism [MRSR01]. Parastichopus [CPE09]. Parentage [BHO^{+}08, KB04a, RDBD08, SDSQ09]. Park [FPK07a, FPK07b, HMW^{+}09, KM05, LPO01]. Parr [ERO^{+}09, MTM09, PB00, SJ03]. Parry [RAH01]. Participation [FD01, HD02, HBB07, TF06]. Party [VJK^{+}07]. Pass [OS05, RJY08]. Passage [BM01a, BMCB09, BJ06, CSB00, DKM07, HHWL00, Kha06, LPSL08b, MDH00, Pea08, PHC^{+}09, RW06, SPL02, SOM09, SMH02, SBT^{+}09, Spr05b, WBH02, Wer07]. Passed [HLMB01]. Passing [Ded07, MSWS01]. Passive [ARK09, BG06, BGB09, BJ06, Dar03, DB09a, DH09, FSF06, FHSM08, GL02b, IC08, KJS^{+}09, LRD^{+}04, MBS05, MBHM00, MMB06, MSW^{+}01, PB03, RFLN01, Tat09, WJKW07, WB03]. past [SBNN03]. Path [GTZM09]. Pathways [BPK^{+}00]. Pattern [HCSCT06, TWP03]. Patterns [BPBS08, BRB09, CSI^{+}04, CRZ05, DS05b, DBG07, GCM^{+}03b, HN05, HD02, ICBM04, JH01, JR03, JEB07, KHJ^{+}08, MRW05, MTM09, NFQ07, SHB^{+}04, SHPCQ00, SGPK09, SJP02]. Peak [PBI03]. Pearl [RRB^{+}07]. Pectoral [BIBP02, HSH04, Mic05, MNB09, PHK03, SB06b, WTW^{+}04]. Pelagic [Dia08, Ker08, KG08, PH03b, RGP^{+}07]. Pelvic [ZM06]. Penaeid [GGCF08]. Pend [HHL^{+}08, DBG07, KB04b]. Peninsula [RPD09]. Pennsylvania [BCAK07, KC01a, KC01b]. Penobscot [HZG^{+}09, GTZM09]. Pens [FSK04, MTM09]. Percentile [RR08]. Perceptions [RAP09]. Perch [BRB09, BPL02, COG^{+}07, COS^{+}09, HL08, HOFM06, Hor01, Ian02, IHW02, IWLB05, IWBL07, LSK^{+}05, MBS05, MR04, NB05, PMBM07, PMBM08, PWK02, PTN06, PR04b, RJB07, RM01, SM00b, Stu06, VBP08, WDB06, WBE^{+}05]. Percid [PTN06]. Percy [SB04]. Performance [AGB^{+}04, BMP07b, BLB^{+}05, BJ06, CDPA09, CFL^{+}03, CJMP08, CSA^{+}04, DMYC02, FHS06, FHSM08, GP03, LB08, MTHH04, MMB06, Pea08, SH02a, SRGC05, SMB07, WAB^{+}02, WH04]. Period [Ano06, BB06, PCG^{+}01, TP06]. Persistence [BB02, DS08, GAG^{+}01, PFWC08, WHNC09]. Perspective [HMSF08, PH03a, ZK09]. Perspectives [BSS06, JML02, SCD^{+}08]. Pest [SWMH06]. Petroleum [SW04]. pH [MC08]. Phase [McC01]. Phenotypic [SLB^{+}09]. Pheromonal [YMR03]. Pheromone [JLSL06]. Photography [Sha01]. Physical [BPT08, CCG09, FSH06, Gol09, KS02, KSB02, NBH^{+}03, RHB08, RPW09a, RPW09b, TPG02]. Physicochemical [GHA^{+}02]. Physiological [BTI05, CCG09, DMYC02, FRW09, GC08, KSM^{+}03, MVH^{+}09, PHC^{+}09, SKM^{+}03, TRW00, WW08]. Physiology [CFL^{+}03, KRK^{+}02, Rei07, TCD^{+}08]. Piecewise [BB08, Mac07, MH08a]. Piedmont [BKNB04]. Piers [BKNB04]. Pigeon [GAG^{+}01]. Pike [AKKC08, BF06, MGH03a, MS03b, MBS^{+}08, NH01, PW03, Pea04, Pea08, PC00, PTM03, PTD03, Pie04, PSP^{+}06, SH05a, SFUS07, WFS07]. Pikeminnow [BW06, FSC03, MWC02, PB03]. Pinaleno [PN03]. Pinatubo [Sch02]. Pink [CP04, HPSW05]. Pinopolis [DIC04]. Pipeline [New06, Zho04]. Pipes [APS05]. Pirarucu [Cas04]. Piscivores [PSP^{+}01]. Piscivorous [LPL04, RFLN01]. PIT [APS05, CJMP08, CRK^{+}05, FHSM08, HMS^{+}03, MS03a, PR03, PB03, SHD^{+}05]. PIT-Tag [APS05]. PIT-Tagged [CJMP08, HMS^{+}03, MS03a]. Pitfalls [GHM^{+}07]. Placement [BBTC07, DH09, EM06, IYJS02, RB03, Ron03]. Placopecten [GDH04, HR06]. Plain [WSH09]. Plains [OGK07, PS02, PHRG00, PWB04]. Planktivorous [PWPC07]. Planning [NWC^{+}03]. Plant [HLV^{+}07]. Planted [HHPH03]. Planting [DV00]. Plants [BBD02]. Plasma [WBH02]. Plastic [DCB09, PSW06]. Plate [RMBZ08]. Plateau [JGW^{+}09]. Platform [SW04]. Please [Dew09]. Ploidy [JT07, SSB01]. Poaching [RPD09]. Point [BH05, HBKJ08, JJ07, LCM06b, MB01, SMR08]. Points [BSS06, HPM02, OJL08]. Policies [BMP07a, Wil02]. Policy [CH04, Ral02]. Pollution [Mea04]. Polychlorinated [BRS02]. Polymictic [OGK07]. Polymorphism [LZW^{+}08, TWN^{+}07]. Polymorphisms [STSS05]. polyphemus [GMH09]. Pond [BMP07b, OGE01, SB06a]. Pond-Cultured [BMP07b]. Ponds [PPBM04]. Pool [BHDA05, BSG05, KS02, ML04, RH03, SAC05, Pit02]. Pools [Bai03, HRL06, WM05, KSB02]. Poor [VB08]. Pop [PMBM08]. Popular [VW07]. Population [AP00b, Bau02, BMW^{+}02, BND^{+}07, Box02a, BSSS02, BTM07, CJM^{+}02, Car01, CPWM06, CPW^{+}06, CPG^{+}08, Col02b, CBM03, CRD05, DS05a, DAJB09, DHMHJ06, DSM07, DBG07, EWD^{+}06, EYI08, FPK07a, FPK07b, GWBH02, GHDM09, GM00, HWB^{+}03, HHL^{+}08, HMSM08, HMM09, HLH05, Ise07b, IWBL07, JGW^{+}09, JZMS05, KDSF07, KD00, KC01a, LI09, LHSS07, LLCBF07, LM02, MED08, MHM09, MTN09, MGGL04, MJ04, MC05, MS03b, NPE^{+}06, NWC^{+}03, NL06, PSWD07, PW03, PTD03, PPF07, RD09, RJB07, RA01, SB00, SIB02, SPM06, SOM09, SMB03, SCMF02, SMM^{+}07, TP06, TP05, VRS01, WME05, WHQ^{+}08, WCR^{+}05, YGG04, vPP05, GJB09]. Populations [ACBS05, Ano06, BMP07a, BH03a, BWEM08, BKNB04, BGB09, BK09b, Bon02, BB06, Bry00, BW09, BTHE04, CKSG09, CMN08, DMMW08, ENJO04b, GAG^{+}01, HR06, HLV^{+}07, Hil02, KBPC08, KQP^{+}09, KWP06, LPL05, Luc02, MMQ01, MTN09, Mic09, NBH^{+}03, OS05, PMH05, PWK02, PTM03, QGP^{+}02, SIJH06, SHB^{+}04, SL02a, SD06, SHN08, SH04a, SZL00, SRN00, TKK^{+}01, WCE^{+}09, WHNC09, ZHBB09]. Porbeagle [CJM^{+}02]. Port [FRW09]. Portable [CRK^{+}05, LaV02, MBHM00, PBB03]. Portion [MMLM06, RC01]. Portrait [HS03]. Portuguese [OFM^{+}09]. Position [GLA00, SVPP08]. Positional [PW09]. Positive [MFP^{+}08]. Possible [HLO08]. Post [PKW01]. Post-Libby [PKW01]. Postfire [ST00]. Postharvest [BHL03]. Postmortem [CWK09]. Postocular [Clo00]. Postrelease [CSA^{+}04, GC08, GH08, KG08, SMP04]. Postspawning [MP02]. Poststocking [BSD09, IBSC02, MMF^{+}03, SB06a]. Posttournament [YI06]. Posttreatment [DSW01]. Potamopyrgus [ROS04, VB08]. Potential [AHH00, ACPHT09, BJ09, Cur06, DP04, EWD^{+}06, Fie00, GMH09, HMM09, HBFL08, IK05b, KMG09, McC01, QGP^{+}02, SIB02, SGG03, SMPM02, SSWC05, VC02, VEK09, WHNC09, WHCL02]. Potomac [HS09]. Pots [AHYC09, ZK00]. Power [BCW04, Ger07, HLV^{+}07, MK09, MD04b, SPR^{+}05a, WVT09, You09, YDJM09]. Powerhouse [JAAW05]. Practicable [Pow05]. Practical [GBLC02]. Practice [RGP^{+}07]. Prairie [MP08, Mea04, UF06, VR01]. Precautionary [Aus01]. Precise [PCIB08]. Precision [Chi09, HAW01, HSH04, KQH09, LF01, MB05b, MNB09, MSP^{+}08, PMBM08, PEW07, QJBH07, RCK05, SZW02, SMR08, SH08, VBP08, YS04]. Precocious [ZSH03]. Precociously [PJBT09]. Predation [BMPU03, BBS05, CCBW01, FS00, FCM^{+}09, FSC03, GP05, HBB04, LGA02, MNS01, MMF02, MTL09, NBN04, PS02, PWK02, RB09, SB06a, SZL00, SM07b, TBL04, TFF^{+}07, WEC07]. Predator [AGB^{+}04, LI09, MC04b]. Predators [BR06, OGK07]. Predatory [JMHB08]. Predict [KF09]. Predictability [AP00a]. Predicted [HWB08, IWBL07, TGR06]. Predicting [BHVS06, FDM^{+}06, HB00, HBB07, IV05, NBH^{+}03, SB04]. Prediction [Che04]. Predictions [YGG04]. Predictive [HFT08, HRK03b, ISBC02, PK01, PRP00, RPJB05]. Preferences [BCD02, HB07, IWLB05, MJ07, RM01]. Preferred [EC00]. Preliminary [OFM^{+}09, VGM00]. Preparing [KQ07, PGS01]. Prepositioned [SH04b, WFW02]. Preseason [HPSW05, HPSW08]. Presence [MHO02, TCGR06]. Preserved [CGP00, PMBM07]. Prespawn [HZG^{+}09]. Pressure [ČSB05, PJC00, YH04]. Pressure-Sensitive [ČSB05]. Prevalence [SWG^{+}06, SLS^{+}04]. Prevent [GZC08, JKS01, Ris00]. Previously [vPP05]. Prey [AAF^{+}06, Bau02, KMG09, NH01, PSP^{+}01, VFJ08, VH00]. Priest [SB04]. Priorities [BPRG08]. Prioritizing [BPRG08, RBB^{+}02]. Priors [MBHK02]. Probabilistic [JGW^{+}09]. Probabilities [BTS02, TF06, WP04a]. Probability [DF07, GFPJ09, GOG09, KJR^{+}09, PPSC03]. Probability-Based [PPSC03]. Probable [MJ04, YMR03]. Procedure [BMB03, CHCD02, FI01, MR02]. Procedures [PP00, SWPO06]. Process [MMW08]. Processes [BHDA05, LH04b]. Processing [IMSW03, VBP08]. Processors [LL08]. Produce [KB04a, LKS00]. Produced [BSRS04, CWDH03, MB08]. Producing [Moh03]. Product [HKE^{+}01]. Production [COG^{+}07, GRM07, JCW02, Mac02, MJ05, MHRR07, MTL09, NTSJ03, Orr07, PTD03, PPBM04, UWBL05]. Productivity [BPB07, Ian02, JM00, Mac02, Mic09, RRCJ08, YS05]. Profile [BBZW04, dSCT01]. Profile-Bar [BBZW04]. Profiles [DMMW08, DHNH07b, MS07, SJS00]. Progeny [SBCH07]. Program [BS01, DR07, GWBH02, IK05a, Ise07a, MRC09, YS04]. Programming [PLP01]. Programs [BK09b, HGHG04, LCL05, PT07, PH03b]. Project [ATM03, MLS06, SBT^{+}09, Cla04]. Projects [MPPB02]. Promote [MK09]. Propagated [AFM^{+}06]. Propeller [KMCM01]. Propeller-Induced [KMCM01]. Properties [BMB03, HH01c]. Proportion [MHO01]. Proportional [BWM08]. Proportions [STSS05]. Proposed [BWB^{+}00, HH01b, NHS08, OW09]. Prosser [HHWL00]. Protect [MPR08, PBM03]. Protected [Sul02]. Protecting [ARC06]. Protocol [BPRG08, UF06, VR01]. Protocols [MJ07]. Prototype [JAAW05]. Provide [STSS05, VTK^{+}07]. Providing [VH00]. Public [BBD02, HBJ05]. Puerto [NLC01, NN08]. Puget [TCSS07]. Pulse [SCMF02]. Pulsed [KWA00, Car01]. Pulsed-DC [Car01]. Pump [CBM00, HLMB01, WBH02]. Pumpkinseed [CSB^{+}03b]. Purse [Yul00]. Purse-Seine [Yul00]. Push [CSW05, PMBM08]. Put [ZS03]. Put-and-Take [ZS03]. Pymatuning [KC01a, KC01b].

Quadrat [LDMW08, PR01]. Quahogs [MP04a]. Quality [BBD02, HKE^{+}01, HHWL00, JH01, MGAC00, OFM^{+}09, PWK02, SGP05, SGG03]. Quantification [BTM07, SS09]. Quantify [ČSB05]. Quantifying [DeV07, IMW02, MB05a, McH07, TGR02, UWBL05]. Quantile [CTP08]. Quantitative [WTH08]. Quantities [FB07]. Quartiles [Ger09]. Quatrefoil [PSP^{+}06]. Queen [BWJ^{+}08]. Queets [WMSG09]. Questions [WSM^{+}08].

Raceway [Pea08]. Racial [HD02]. Rack [KH01]. Radio [BBO03, BM01a, BSSC04, BCP03, BJ06, CFL^{+}03, CB01, FH02, HMS^{+}03, JEB03, KPRB04, MS03a, PASP07, PRP08, Pie04, PN05, RW03, SH07, SL09, SHW00, SMF^{+}09, SVP^{+}04, VRK00, ZDK^{+}03]. Radio- [MS03a]. Radio-Tagged [BM01a, HMS^{+}03, SHW00, VRK00, ZDK^{+}03]. Radio-Tracking [BSSC04]. Radio-Triangulation [JEB03]. Radiotelemetry [BGH04, BMW^{+}02, GBLC02, PRP08]. Raft [Sta01]. Raftable [HKH^{+}02]. Rainbow [BKJ06, BMPU03, BTS02, BSD09, BFR04, BK09b, BB02, CHCD02, Clo00, CJ02, CW02, Ded07, DST00, FWEP08, GLA00, HB05a, HB05b, HBC08, HWN09, HKT00, HB06, HM09, HWKN00, HBB04, HW07, Jen03, KCR^{+}05, KM00, MJ05, MRSR01, MSRP01, MHS08, Mek04, MZ02, MMLL03, Neg03, NT04, NMP07, PP00, QK00, RD05b, RD06, RPM08, SE00, SH04a, SGPK09, SHW00, SHMR03, SSS07, SWWR04, TP06, TSMD03, VB08, WW04, WWB04, WMO08, WHC09, YWM^{+}00, vPP05, vdWF03]. Rainfall [EO03]. Raised [RPM08, SSC07]. Rake [MP04a]. Ranched [SVS^{+}09]. Random [DS04, QGBH06]. Range [DS05b, DRC03, HTGS08, ML04, MSES03, PWB04, PN05, SM09d, WW01]. Ranges [Vok03]. Rangitata [UWBL05]. Rangitikei [BTS02]. Ranking [JMHB08]. Rapid [BWJ^{+}08, RJY08, STSS05, SLB^{+}09, UF06]. Rare [DBM04, HWB08]. Rate [CCBW01, DHNH07b, GHM^{+}07, GHDM09, GRM07, GBFJA07, HNE04, HHPH03, HPM02, HMS^{+}03, Jon05, MWN08, Mic09, PHHC01, PHHC02, PM01, SMP04, SCHK03]. Rates [ATM03, ARPP06, BWB05, BGB09, BPK^{+}00, BP01, Bro02, BSB03, BTM07, DP04, Dia08, Dor02, FHN05, GHASLC09, HBH00, JB00, JPB^{+}07, KPL09, KPRB04, Mac07, MHOH06, MS03a, Mek04, NH01, PHB^{+}09, PHC^{+}09, PH05, RA04, RHB05, RBB08, RT09, SGR02, SR02a, SR02b, Sch04b, SH05b, SBCH07, Sto05, TWH01, TWPP06, TKW00, VFJ^{+}05, VJK^{+}07]. Rathbun [Mit02]. Ratio [Dew08, Dew09, MPBM06, SMPM02, WT09]. Ratios [VC02]. Ray [PHK03, SH08, ZM06]. Rayburn [DSM07]. Rays [HSH04, KQ07, Moh03, SB06b, WMO08, WTW^{+}04]. Razorback [BG00, MMF^{+}03, SM07a]. Re [ZK09]. Re-establishing [ZK09]. Reach [ERO^{+}09, FM08a, FJ02, HCSCT06, RD05a, TCGR06]. Reach-Scale [FM08a]. Reaches [FSF06, MP08]. Reaching [EYI08]. Reading [MBHM00]. Realistic [QSGS04]. Realized [HHB05]. Reallocation [SL01]. Really} [MR02]. Reanalysis [BS00]. Reared [BTS01, BBTC07, BG00, CLB^{+}07, FSK04, HHAK02, MHF^{+}03, OBG^{+}00, PB00, RFL04, SB06a, SBNN03, WWB04]. Rearing [BSD09, EH01, HCSCT06, PPAR06, RTBF09, RMF^{+}02, SBCH07, SAC05, SJ03, TGR02, TGR06]. Reasons [RPD09]. Rebuilding [BMP07a, YS05]. Recapture [CPWM06, Cur06, DHNH05, GM04, HV03, MC05, MZ02, RHB05, RD05b, RD06, TP06]. Recently [BB02]. Reception [JLSL06, PRP08]. Recolonization [HS09, Mea04, RD05a]. Reconnecting [Sch03]. Record [WP04a, WP04b]. Record-Size [WP04a]. Recording [Win04]. Recordings [DKM07]. Recoveries [ARC^{+}07, HCS09, JBHP07]. Recovering [HLH05, LHSS07, Sul03a]. Recovery [BPK^{+}00, GJB09, HLO08, IDK03, JZMS05, MPM09b, NWC^{+}03, PH08, SEB^{+}09, TP06, VEK09, YGG04]. Recreational [ARC^{+}07, Arl06, AKKC08, BDS^{+}02, CBS^{+}05, CW02, CWP03, CBM03, DAJB09, FD01, GMHF03, Goo07, GBFJA07, LMA^{+}03, MSF^{+}09, NRJ05, PGR^{+}04, PH03a, PC00, PMPS03, Rad03, RB07, SH02a, SL01, SMR08, TMS05, TF06, WB03]. Recruit [Dor02, KJS^{+}09, LRBJ08, Wil02]. Recruitment [AP00b, BPBS08, BPK^{+}00, BJ04, Box02b, BSG05, BHVS06, CDPA09, Che04, CH02, CHSS08, DA05, FHN05, Fie02, FPK07b, GHDM09, IMW02, MR02, Mac03, Mac04, MM02, MC01, NBHH00, NBHH01, PBH00, Pit02, Qui07, RHBS04, SBIC02, SGR02, SM00b, SOR05, TP05]. Red [BPBS08, BSR07, CKSG09, CPE09, COG^{+}07, COS^{+}09, DS08, Dew08, Dew09, EM06, GGCF08, GHASLC09, IS01, JDB^{+}02, JDB^{+}04, KSB^{+}08, KBF^{+}00, LPWH01, LRBJ08, MHB07, MB09, MC01, RDBD08, SWA^{+}08, SMP04, SSP08, VC02, VW07]. Red-Rim [MHB07, MB09]. Red-Spotted [SWA^{+}08]. Redband [DJL09, MBM01a, MBM01b, Muh02, SLEM07, ZAF05, ZC06]. Redbreast [SM09a]. Redd [ACBS05, DRD01, GG05, GHA^{+}02, HW07, JGW^{+}09, Sch00, WTH08]. Redds [FWEP08, GG05, GG09b, GCR08, MRJL05, MSK05, MPM09a, TRS04]. Redear [CA06, SPM06]. Redhorse [GFPJ09, MRC09, RR08, WCSJ04]. Redhorses [MJ07]. Redside [RJY08]. Redspotted [RWBG04]. Reduce [MD04a, MA00b, Mir05, PR04b, SB06a, SBCH07, ZK00]. Reduced [FPK07b, Jac05, KD00, MSRP01, SWPO06]. Reduces [GASH05]. Reducing [AHYC09, BFF^{+}02, JBG05, PSW06, SSS07]. Reduction [CH01, HPB05, PMBM07, WMJBC04]. Reductions [BCC03, IWBL07, PSWD07]. Reef [CRBD06, MDM^{+}04]. Reefs [GBPT08]. Reestablishing [BND^{+}07, BSSS02]. Reevaluation [PAAK07]. Reexamination [HNE04]. Reference [FB07, HPM02, KRB^{+}04, OJL08]. Refining [Mir05]. Refuge [MED08, MTHH04, SMP04]. Refuges [Rei07].

Regime [JML02]. Regimes [SHB^{+}04, SEGG05]. Region [Mar01, WCE^{+}09, ML06]. Regional [DPW05, FB07, Loo06]. Regression [BB08, CTP08, FDM^{+}06, MH08a]. Regulated [DHG02, MSRP01, MS04a, MS04b]. Regulation [ENJO04b, LF05, PTD03]. Regulations [AHH00, ARPP06, BTS02, CK09, DMM02, EWD^{+}06, HW09, KM05, PMPS03, RAP09, SL02a, VC02]. Regulatory [BM02, Fay03]. Regurgitation [KPRB04]. Rehabilitate [RJB07]. Rehabilitated [QK00]. Rehabilitation [HMSF08, HBFL08, MDM^{+}04, MED08, PH09, RHB08, Sch08, SCD^{+}08, SEB^{+}09]. Reintroduction [ACPHT09, DR07, MRC09, MW04, PT07]. Related [DO02, ENJO04a, GHAS04, GASH05, Mac07, MEL09, MC01, RW03]. Relation [ERO^{+}09, JH01, MD04b, MRW05, NBHH01, PRHS03, SOR05, TPG02, TAM^{+}03, WHHR01, WC01, YS04, YGG04]. Relations [CH00b]. Relationship [BNV02, BB06, BP01, HNE04, HRB^{+}04, KVMM07, MTF^{+}08, PP01, RH03, SLG02, SM00b, SJ08, WP04b, Ano06]. Relationships [BB08, DTMS06, Dor02, FM08a, HB00, LSK^{+}05, LB09, MR02, MH08a, ML04, NMJ01, OW09, RP09, RHB05, RPW09a, RPW09b, VFJ08]. Relative [BRS02, Bon02, BAC05, BMB03, BRSA07, BHM^{+}07, BHHW02, CMN08, DP03b, DST00, EYI08, HB05b, HN05, HLH05, HH01c, MS04b, PASP07, PF03, QGBH06, RHGW05, SHMR03, TSMD03, VFJ08, VB03]. Release [ASD04, ARPP06, BTI05, CPSM00, FSSR08, Fay03, GLA00, Goo07, HV03, JSMV07, Jen03, JPB^{+}07, KRK^{+}02, KSM^{+}03, LCL05, Mek04, MMKC05, MRS00, MTAB08, NLC01, OZR08, PB00, RAP09, Rul07, SL02a, SRGC05, SBT^{+}09, SM06, SKM^{+}03, TWH01, VW07, WG08, WW08]. Released [AMK^{+}05, BG00, CSA^{+}04, Dia08, GCM^{+}04, HPM08, KSB^{+}08, LSK^{+}04, MHF^{+}03, MTM09, NRJ05]. Releases [BMPU03, Joh04, PH03b, WF04]. Releasing [AB03, O’M08]. Relevance [RHHFB00]. Relict [WGSW07]. relicta [CLB04]. Remaining [WCR^{+}05]. Remote [HL09, KB04a]. Remote-Site [KB04a]. Removal [Bry00, CBP07, GDH05, GBPT08, IMSW03, MLS06, MB08, PHK03, RMD02, RD05b, RD06, SLB^{+}06, ZM06]. Removals [KM00]. Removing [KP07]. Renibacterium [JMP07]. Repatriation [MKP05]. Repeated [FRW09, GW04, GL08]. Replacement [SBNN03]. Reported [MC04a, PW09]. Reporting [HHPH03, PHHC01, PHHC02, TWPP06]. Reproduction [DB03, OCW04, TPC^{+}09]. Reproductive [AIHRdSD09, BHO^{+}08, BTS01, CRC^{+}08, HA02, MMS08, McC01, NTH08, SWPO06, TKK^{+}01]. Required [LCM06b, MOH07]. Requirements [MD04b, PWH02, RHK^{+}03, VRS01, Vok03]. Resampling [DS04]. Rescinding [Box02b]. Research [CB04, MS03b, TJKB06]. Reserve [WP03]. Reserves [Li00, WP03]. Reservoir [BRS02, BKZA08, BP02, Box02b, BSG05, BBS05, CPW^{+}05, CRD05, DD02, DSM07, GGG08, GLA00, GSC02b, HBJ05, HA02, ISBC02, JH01, MB01, MHS08, NBN04, NLC01, NN02, NTSJ03, OGK07, ON05, PS02, PRSA06, QGBS02, RMD02, RWZ09, SB00, SGR02, SZL00, SB04, TWRH07, TGR06, Tip01, WN04, WP04b, ZHBB09, BFR04, HB05a, Mun02, SMB03]. Reservoirs [BG05, BKNB04, Bar04, Bau02, BHVS06, CRC^{+}08, Col02a, Col02b, Cun04, DD02, DS04, GM00, HAW01, HJ02, IBSC02, JM00, LM02, Mac03, PFS^{+}06, QGSS03, QSGS04, RHGW05, SBIC02, SIB02, Sch04a, SSG05, SGR02, SR02a, SR02b, Sch04b, SBSL06, TTRF05, TSMD03, VB03, YWM^{+}00]. Residence [BM01a]. Residency [CBM00]. Resident [BB02, BMW^{+}02, DK04, Hil02, PN05, RBM00, RDN00]. Residual [BUH07, KS02, MP01]. Residualism [SBCH07]. Residuals [CDPA09, Mac04]. Residue [BPL02]. Residues [BRS02]. Resistance [WAB^{+}02]. Resource [Goo07, SN02]. Respect [BLM^{+}09]. Respond [WHCL02]. Response [Ach01, BW08, BWEM08, Box02a, BW09, BM02, CHCD02, CP04, DHG02, FRW09, GJB09, GFPJ09, Han03, LR08, MH08a, McH07, MHW00, MS04b, PBT05, PSKP09, RCB07, SB00, SBIC02, SPR^{+}05a, SCMF02, Ste03, SEGG05, WBH02, WW08, ZJ02]. Responses [AHH00, AP00b, BL08, BTI05, BPS00, BMK^{+}01, CCG09, DMYC02, HBC08, HS08, IDK03, KSM^{+}03, Mic09, ML03, PSWD07, Ron03, SEC01, SSS07]. Restoration [BW08, BWEM08, BPRG08, BHM^{+}07, HDB00, KRDC08, LR08, MHOH06, MJ05, ML03, OM04, PRSA06, RAH01, RBB^{+}02, SGB07, WHB03]. Restoring [Hil02]. Restricted [ST00, Sta01]. Restrictions [Han03, ISBC02, SPM06]. Restrictive [HW09]. Resulting [AWWG08, GG09a]. Results [PCIB08]. Retention [BGB09, BBTC07, BI00, CCBW01, CPE09, Clo00, CH00a, Dar03, DB09a, DH09, Fri01, GL02b, HJ06, HKFM08, IF00, IS01, IC08, JDB^{+}02, JRWK08, LH04a, MPB07, NT04, O’M08, PR03, PSWD07, RW03, RA04, SJP05, SVP^{+}04, VW01, WJKW07, WW04]. Retrieved [BJ09]. Retrospective [DKM07, HPSW05, HPSW08]. Return [DST00, JPB^{+}07, LKS00, Sch04b, YI04]. Returning [KLH^{+}08]. Returns [AP00a, BSB09, MBD02, SHH^{+}04]. Reveal [NPE^{+}06]. Revetments [GTRC02]. Review [BPRG08, DJL09, KWP06, MH08b, MS03b, Mun02, RBB^{+}02, RHB08, WC01]. Revisited [Cla02]. Reward [PHHC01, PHHC02, TWPP06]. Rewards [SHH^{+}04]. Ribbon [KNL02]. Richmond [GWBH02]. Richness [BG05, Bar04, DP03b, HKH^{+}02, PHRG00, RHK^{+}03]. Rico [NLC01, NN08]. Riffle [BTK04, LDMW08, PR01]. Riffle-Dwelling [LDMW08, PR01]. Riffles [HWN09]. Rim [BCW^{+}09, MHB07, MB09]. Ring [SI08]. Rings [GAKB04, RT09]. Rio [ARK09, CCG09, SSC07]. Riparian [BNV02, EM06, OM04]. Riprap [GTRC02]. Risk [AFM^{+}06, HRK03b, MBHK02, SHN05]. River [AHJ^{+}06, AB02, APS05, BE02, BKJ06, BTS02, Bar05, BPRG08, BM01a, BP02, BB04, BH05, BMW^{+}02, BFH^{+}08, BG00, CAM06, CBP07, CHAO08, CR04, CPG^{+}08, CKH05, CHR^{+}08, CTB07, DHG02, DS05a, DS05b, DBG00, DBG07, DW08, EYI08, FSH06, FSLS09, FSC03, GAG^{+}01, GCM^{+}04, GAM^{+}08, GTZM09, GSC02b, HCSCT06, Har03, HBS^{+}08, HP04, HKT00, HBKJ08, HBS^{+}05, HCFK08, HZG^{+}09, HP09, HLH05, JGW^{+}09, JR03, JJ07, JAAW05, JMHB08, KP07, KB04b, KPJ^{+}04, KMB^{+}09, KVMM07, KGM01, KDSF07, KRB^{+}04, KJR^{+}09, KLH^{+}08, KRDC08, KHJ^{+}08, LSK^{+}04, MHO01, MHO02, MP08, MS07, McC01, MSRP01, MB05b, MSES08, MSW^{+}08, MZ02, MSF^{+}09, MDHK08, MPR08, MRC09, MGH^{+}03b, MM05, MBS^{+}08, MS04b, MPM09a, MPM09b, NCK^{+}06, NJP^{+}09, NBGK09, NH07, NH09, PMH05, PH08, PH09, PSB^{+}07]. River [PR04a, PRP08, PPAR06, PPBM04, PHC^{+}09, PN05, PvPR^{+}06, PM01, QK00, QGBH06, QJBH07, QBH^{+}09, Rei07, RPM08, RFLN01, SSR07, SMSM01, SGP05, Sch00, Sch01, SPL02, Sch03, SM07a, SGPK09, SEC01, SMB03, SSWC05, SMH^{+}03, SB05b, SMF^{+}09, SPR05c, SEJCC04, TBL04, TWP03, Und00, UBK04, UGM^{+}07, UWBL05, WHHR01, WPV^{+}04, WCSJ04, ZB00, ZSH03, AWWG08, BCAK07, BLA^{+}07, BWW^{+}08, BVH00, BS00, BPB07, BND^{+}07, BF06, BKD^{+}09, BUH07, BTB^{+}02, CPW^{+}06, CRL08, CB03, CBYB03, CSB03a, CP04, CSA^{+}04, DHGP03, Fie00, FSF06, FH02, GTRC02, GHA^{+}02, GAM^{+}08, GMHX08, GCR08, HNTM00, HTA^{+}09, HRL06, HHAK02, HHWL00, HM09, HI01, How06, JZMS05, KP07, KBPC08, KQP^{+}09, KQH09, MC04b, MS03a, MVH^{+}09, MJ04, MRSR01, MRJL05, MS04a, MMKC05, MW04, MBM01a, MBM01b, MSWS01, NPE^{+}06, NBN04]. River [PK01, PKW01, PJBT09, PBH00, PHB^{+}09, Pit02, PJC00, QGP^{+}02, RWJD04, RB03, RC01, RRB^{+}07, SP07, SGD09, SH04a, SMPM02, SMWS02, SMH^{+}03, STSS05, SFUS07, SBSL06, SWWR04, Spr05b, SJP02, SMM^{+}07, TRW00, TRS04, TGR06, TCGR06, TC01, TPC^{+}09, VRK00, VPR06, WMSG09, WF04, YDJM09, YS05, ZDK^{+}03]. Riverbed [WLG03]. Riverine [BWB^{+}00, BRSA07, GFPJ09, QJBH07]. Rivers [AK05, BCM^{+}05, BHO02, BAC05, BTHE04, CRZ05, GB03, HKH^{+}02, KWP06, LCM06a, LWHP07, MOH07, OS05, SM09a, SL09, SOR05, UF06, WME05, WP08, Wil06, WTH08, YH04, BERH02, CBWB02, DR07, EWK^{+}08, ESPW03, HMS^{+}03, KPRB04, MSW^{+}01, RDN00, WSM01]. Roads [HL09]. Robust [GFPJ09, HPM02, MJ07]. Robustness [Ian02]. Rock [CPDS01, TMS05]. Rockfish [Dor02, HPB05, HPM08, HPM02, PR03, PSKP09]. Rockfishes [SHJ00]. Rocky [DHM01, DJC^{+}06, PPS03, MLS06]. Rod [SH04b]. Roe [LL08]. Rogue [Rei07]. Role [Fie04, KMG09, PW03, SHB^{+}04]. Roosevelt [BMPU03, MHS08]. Root [Kol08, MK09, MS05, PVMR01]. Rotenone [HMW^{+}09, LEJ06, RBK09, SN07, TAME03]. Round [BMW^{+}02, DTMS06, HEB09, DJC^{+}06]. Route [SBT^{+}09]. Route-Specific [SBT^{+}09]. Roving [MC04a, ZS03]. Royale [LPO01]. Rubber [MMW^{+}05]. Ruffe [HBHB00, OW09]. Run [BTK04, HCFK08, KLH^{+}08, MMQ01, MVH^{+}09, WM05].

S [COG^{+}07, COS^{+}09, SH07]. Sablefish [DP04]. Sacramento [WF04, FSH06, MC04b, SMB07]. Sagebrush [ZC06]. Saginaw [Fie02].

Sagittae [JDB^{+}02]. Sailfish [KG08]. Salamander [PF09]. Salamanders [Ron03]. Salmon [AP00a, ARC^{+}07, AMK^{+}05, AGB^{+}04, AGM^{+}09, BHO^{+}08, BK09a, BCM^{+}05, BWW^{+}08, BWJ^{+}08, BCW^{+}09, BMK^{+}00, BCD02, BM01a, BTK04, BTS01, BT03, BPB07, BND^{+}07, BPS00, BUH07, BW09, BFF^{+}02, BTB^{+}02, BJ06, CKC^{+}09, CR04, CBWB02, CB03, CBYB03, CSB03a, CP04, CSA^{+}04, CHJ07, CTB07, Dar03, DHGP03, DC06, DV00, EM06, EH01, ERO^{+}09, FS00, FSK04, FHS06, FCM^{+}09, FSC03, GM04, GG05, GCM^{+}04, GTRC02, GHA^{+}02, GAM^{+}08, GMHX08, GJB09, GTZM09, GL02a, GL02b, GCR08, HCSCT06, HPSW05, HDPS07, HPSW08, HWB^{+}03, HF00, HTA^{+}09, HLMB01, HGHG04, HI01, HMS^{+}03, HCFK08, HZG^{+}09, HCS09, HOFM06, HMM^{+}05, JR03, JED^{+}03, Joh04, JAAW05, KPRB04, KPJ^{+}04, KJR^{+}09, KRK^{+}02, KLH^{+}08, KJS^{+}09, LT01, LSK^{+}04, MPPB02, MFT08, MFW^{+}09, MC04b, MS03a, MVH^{+}09, MPBM06, MJ05, MWR^{+}01]. Salmon [MRJL05, MMVL05, MHRR07, MS04a, MSW^{+}08, MSK05, MHF^{+}03, MSF^{+}09, MMF02, Moh03, MRW05, MR06, MRK08, MDKM06, MSW^{+}01, MW07, MTM09, MPM09a, MPM09b, NCK^{+}06, NJP^{+}09, Neg03, NT04, NWC^{+}03, NFQ07, New06, Orr07, PBI03, PCIB08, PSB^{+}07, PT07, PJBT09, PB00, PB03, PJC00, PPBM04, PHC^{+}09, PM01, RW03, RB03, RFL04, RVM^{+}08, RRCJ08, SWA^{+}08, SP07, SHPCQ00, SPR^{+}05a, Sha01, SH04c, SHD^{+}05, SMWS02, SMH^{+}03, STSS05, SHN05, SJ03, SSR09, SVS^{+}09, SEJCC04, SBNN03, SLB^{+}06, TBL04, TFF^{+}07, TRW00, TGR02, TRS04, TGR06, TCGR06, TTM^{+}05, UBK04, UGM^{+}07, UWBL05, VTK^{+}07, VBHT05, VGM00, VRK00, WMSG09, WBH02, WF04, WME05, WHHR01, WHC09, WM05, Zho02, Zho03, Zho04, ZSH03, CR04, KJR^{+}09]. Salmonid [BKZA08, BS00, DeV07, DS07, HB00, HGF06, HHWL00, HH01c, JBG05, MB05a, McH07, Rei07, RF00, RO06, SLS^{+}04, Yul00, ZB00]. Salmonid-Bearing [DeV07, MB05a, McH07]. Salmonids [BE02, BERH02, BBGVM05, Bin04, CQ05, CME02, CJMP08, CSB00, Har03, JMP07, LRD^{+}04, ML04, MBS^{+}08, MSWS01, MH05, NBN04, NL00, PPS03, PBT05, RHGW05, RFL04, RFLN01, TFF^{+}07, WW08, WSM01]. salmoninarum [JMP07]. Salmonines [CLB^{+}07]. Salt [MB09]. Sam [DSM07]. Sample [FSF06, FM08b, Ger07, Mue03, NH07, PWH02, PHB^{+}09, VRS01]. Sample-Size [VRS01]. Sampled [SWPO06]. Sampler [HWH05, HHM05, PR01]. Samples [FM08b, NB05, NTSJ03, PGS01]. Sampling [BW08, BPT08, BG08, BSEM05, BABC08, BPH08, BS09, Chi09, CSW05, CJL07, Col02a, DJL09, DKM07, DJC^{+}06, Dra07, DS04, DRD01, FS07, FSLS09, FRW09, GL02a, HBS^{+}05, HB06, HH01c, JGW^{+}09, JJ07, KHJC06, LCM06a, LCM06b, LDMW08, MC00, MS02, MJ07, PMBM08, Pin00, PPF07, QGBH06, RWJD04, RO06, RUKR07, SSG05, SH04a, SSP08, Sta01, SMR08, SLB^{+}06, TAME03, TTRF05, UF06, VR01, WFW02, WCE^{+}09, WSH09, WKAP04, WPS03, Win04]. Sampling-Induced [WCE^{+}09]. San [SMSM01, PF09]. Sanctuary [KC01a, KC01b]. Sand [HDB00, KVMM07, MSK05, RM01, TKT02]. Sand-Bed [KVMM07].

Sandbed [FSF06]. Sander [ZHBB09]. Sandhill [PWK02]. Sandy [MP04a, SLN07]. sapidus [CH01, GAKB04, RT09]. Sardis [DMM02]. Satisfaction [Arl06, BBD02]. Sauger [Box02a, GWBH02, PS02, Pit02]. Saugers [AHJ^{+}06, BVH00, JZMS05, KHJ^{+}08]. Saugeye [GWBH02]. Saugeyes [Box02a]. Savannah [RWJD04]. Scale [CRZ05, DB06, ERO^{+}09, FM08a, HCSCT06, KHJC06, LMA^{+}03, MEHJ08, SDSQ09, SH08, TWP03, VPR06, WG03]. Scales [BE01, BW09, HWKN00, HOFM06, HAW01, LF01, MSR08, Moh03, NBGK09, RRS07, RBM00, RPW09b, SB06b]. Scallop [GDH04, GBFJA07, HR06]. Scalloped [HJC09]. Scallops [GDH05]. Scaphirhynchus [GOG09]. Scenarios [TGR06]. Scented [DCSM01]. Schedules [WCE^{+}09]. Science [NMP07]. Science-Based [NMP07]. Score [FM08a]. Scores [KRDC08]. Scotia [FF04, MFT08, GM02]. Scour [SHPCQ00]. Screen [DMYC02, DS07]. Screens [BBZW04, GZC08, Pea04, RMBZ08, SLB^{+}09, ZJ02]. Scuba [HLH05, Mue03]. Sculpin [HS09, KHJC06]. Sea [AIHRdSD09, AHC^{+}09, AHYC09, CPE09, DJSG09, GDH04, GDH05, HR06, JSL05, JLSL06, KG00, KGM01, MCPM02, MRW05, MR06, MRK08, Orr07, RBB08, SSS08, SSP08, SM09c, SM07b, SVS^{+}09, AMK^{+}05, JML02]. Sea-Ranched [SVS^{+}09]. Seabass [ASD04]. Seal [FCM^{+}09]. Seals [MSK05]. Search [Bro02]. Season [DH09, HCS09, NDH04, SMP04, WN04, WP04b, YWM^{+}00]. Seasonal [ARC^{+}07, AMK^{+}05, BB04, BSG05, BLM^{+}09, DS05b, DMB09, HN05, HBJE00, HBHB00, ICB04, IWLB05, JH01, JZMS05, JEB07, MMW^{+}05, MM05, NL00, PHB^{+}09, Sch01, SSS08, SH05b, SPR^{+}05a, SHN05, SJP02, TWRH07, WW01, vPP05]. Seasonality [JSMV07]. Seasonally [JDB^{+}04, VS05]. Seasons [MP02]. Seatrout [JSMV07, SM06]. Seaward [CSB03a]. Seawater [DO01, DO02]. Second [BLM^{+}09]. Second-Order [BLM^{+}09]. Sectioned [DB06, LF01]. Sections [KJR^{+}09]. Sedation [FRW09]. Sediment [BLK^{+}08, DeV07, FWEP08, HWN09, KCA06, LD04, ML04, MB05a, McH07, NTSJ03, ST00, WC01]. Sedimentation [HWH05]. Sediments [ML06, WC01]. Seeded [LKYM09]. Segment [BRSA07]. Segmentation [CBK00]. Seine [BPBS08, WFW02, Yul00]. Seines [CJL07, PMBM08, Win04]. Seining [BG05, GL02a, MSES08, PCG^{+}01]. Selawik [Und00]. Selected [GL08, KMCM01, SL09, TWP03]. Selecting [MSP^{+}08]. Selection [FS06, FS07, GHA^{+}02, GTZM09, LL08, MW07, RP09, SIB02, SGG03]. Selective [HPB05, IDK03, MBS05, New06, SWMH06, Zho02, Zho04]. Selectively [KD00]. Selectivity [AK05, AKKC08, BS09, CBS^{+}05, HK04, IWLB05, JBHP07, MD00, RBB08]. Selenium [AFM^{+}06]. Semiautomated [BWC08]. Sensitive [ČSB05, NMP07]. Sensitivity [SSS08]. Served [SM07b]. Serves [MHB07]. Set [Ker08, WVT09]. Seton [PHC^{+}09]. Setting [BPRG08, MHOH06]. Settings [DMH02]. Seven [WME05]. Several [GG05]. Severity [GC08].

Sex [CWG04, Dew08, Dew09, FI01, WT09]. Sex-Specific [WT09]. Sexes [MMH09]. Sexually [PN03, YMR03]. Shad [Bau02, BHO02, BSG05, EYI08, HHAK02, HLO08, IDK03, KD00, MHO01, MHO02, MHOH06, MH08b, MDH00, SGR02, Spr05b, ZHBB09]. Shads [MH08b]. Shallow [CRK^{+}05, LCM06a, MP04a, MRJL05, NH07, PWH02, PW03, PTD03]. Shark [dSCT01]. Sharks [BJ09, FRW09]. shasta [MVH^{+}09]. Shedding [DP07]. Sheepscot [BBJ^{+}00]. Sheepshead [AKARF08]. Shelf [LS08, SFS04, LGA02]. Shellfish [WC01]. Shield [Ste03]. Shifts [NTW^{+}04]. Shiner [HRK03b]. Ship [PGS01]. Shitike [BHO^{+}08]. Shoal [SM09d]. Shoepack [FPK07a, FPK07b]. Shoreline [CJL07, RP09, RM01, TCGR06, TCSS07]. Short [ARC06, CH00a, GGG01, HW09, MP04a, MMB06, WWB04]. Short-Duration [GGG01]. Short-Term [ARC06, CH00a, HW09, MMB06, WWB04]. Shortnose [CL04, DIC04, IF00, KH01]. Shovelnose [AWWG08, BFH^{+}08, CWG04, ESPW03, GGG08, KDSF07, KQP^{+}09, PHK03, QGP^{+}02, TKK^{+}01, TPC^{+}09, WTW^{+}04]. Shrimp [BJ04, CGM^{+}06, EO03, GCM^{+}03a, GCM^{+}03b, GCN^{+}08, GGCF08, LI09, RPW^{+}08, WMJBC04]. Shrinkage [CGP00]. Sibling [HDPS07]. Sicklefin [MRC09]. Side [AMK^{+}05, BFH^{+}08, MRJL05, MDKM06, RRCJ08]. Side-Looking [MDKM06]. Signals [WG03]. Significance [ACBS05]. Silver [McC01, MD04a, RBK09]. Silver-Phase [McC01]. Silverside [CGP00]. Silvery [ARK09, CCG09]. Simcoe [AW04]. Similarities [Gol09]. Simple [BRB09, Bon02, CJPB09, DS04, GCM^{+}03a, PPSC03, ROS04, VH00, WCSJ04]. Simulated [ARMB04, CHJ07, FPK07b, GG09b, IYJS02, NL00, PTD03, RMD02, SD06, SRGC05, SZL00]. Simulating [Hil02]. Simulation [BG05, Ian02, McC01, RD09, YS05]. Sinaloa [GCQVMAVP09]. Single [OS05, RJY08, STSS05]. Single- [OS05]. Single-Pass [RJY08]. Site [AFM^{+}06, BHM^{+}07, CH00b, GHA^{+}02, HBB07, JH01, KB04a, MS04a, NN02, PWB04, QGBH06, RP09, SGG03, SJP02, WTH08]. Sites [Buc01, CWDH03, LR08, MW07, NH07, PLW^{+}09, SLS^{+}04]. Situ [PWH02, TKT02, HBC08]. Six [BABC08, CWK09, PF03, MDM^{+}04]. Size [AKKC08, Bon02, BWC08, BWM08, BS09, BHVS06, CH02, CBS^{+}05, CA06, DP04, DTMS06, EWD^{+}06, EYI08, FS07, Ger07, GSWS04, ISBC02, IWLB05, JB00, KMB^{+}09, KBF^{+}00, LSK^{+}05, LRBJ08, LCL05, Mac07, MC05, Mek04, MSES03, MWN08, MD00, Mir02, MHF^{+}03, MDHK08, MPR08, NBHH00, PWH02, PHB^{+}09, PTM03, RA01, Sch04b, SH05b, SCHK03, SJ08, Sta01, Tat09, TP05, VJK^{+}07, VC02, VRS01, WF04, WW01, WP04a, YWM^{+}00, ZAF05]. Size-Based [LRBJ08, Mir02]. Size-Dependent [MHF^{+}03]. Size-Related [Mac07]. Size-Specific [Sch04b]. Sized [BSD09, MZ02]. Sizes [BHHW02, DW09, IHW02, KH09, Pin00, PF03, WDB06, Yul00]. Skagit [CP04]. Skate [GHDM09, LS08]. Skeena [BCM^{+}05]. Skewed [WT09]. Skiatook [LF05]. Skimmer [WMJBC04]. Slimy [KHJC06]. Slope [MSES08, NGWK08]. Slot [Kha06]. Small [BG05, BM00, BWKS02, BM04b, Buc01, CRK^{+}05, DP02, DV05, Dra07, ENJO04a, ENJO04b, FM09, HWN09, HBS^{+}05, HK00, KHJ^{+}08, KM00, MLS06, MS02, MWN08, MD04a, MB09, MBM01a, MHW00, MMLL03, OBG^{+}00, PHRG00, PFWC08, RF00, Ron03, RD05b, RD06, RMF^{+}02, RH03, RO06, RUKR07, SWA^{+}08, SA04, SMBM03, SH04b, SDPB00, Ste03, TP06]. Small-Bodied [HBS^{+}05]. Small-Mesh [MD04a, RUKR07]. Small-River [KHJ^{+}08]. Smallmouth [BRSA07, BM02, CH00a, CBM00, CB01, DF07, DCSM01, FS00, GC08, LF01, LF05, NBN04, NH00, OS05, SH05a, SOR05, TFF^{+}07, VBR04]. Smelt [HBC08, WMO08]. Smith [Kol08, MK09, MNS01, MS05, PVMR01, SM09b, SN02]. Smith-Root [Kol08, MK09, MS05, PVMR01]. Smoky [KM05]. Smolt [BSS06, GTZM09, KJS^{+}09, MHF^{+}03, MDKM06, SLN07, SMH02]. Smolt-to-Adult [KJS^{+}09]. Smoltification [MC04b, Neg03]. Smolts [DS07, FSC03, NC01, OGE01, RO06]. Snail [MHB07, MC08, MB09]. Snails [VB08]. Snake [BERH02, BS00, EWK^{+}08, HKT00, JAAW05, KPRB04, MSW^{+}01, NBN04, PPAR06, SGPK09, SMWS02, VRK00, BTB^{+}02, CBWB02, CB03, CBYB03, CSB03a, CSA^{+}04, GCM^{+}04, GCR08, Har03, HMS^{+}03, MS03a, MSWS01, NBGK09, RDN00, SP07, SMB03, SMH^{+}03, WSM01, YS05]. Snakeheads [LEJ06]. Snapper [GCQVMAVP09, GGCF08, OZR08]. Snook [BLB^{+}05, TWH01, TWPP06]. Snorkel [TPG06]. Snorkeling [BM00, MHW00]. Soak [BR06, SLG02]. Social [MP04b, MTF^{+}08]. Socioeconomic [DMM02]. Sockeye [AP00a, BCM^{+}05, BPB07, BPS00, FS00, FHS06, HDPS07, HPSW08, HF00, HGHG04, HMM^{+}05, MFW^{+}09, MPBM06, MWR^{+}01, MHRR07, MDKM06, MTM09, NCK^{+}06, NWC^{+}03, NFQ07, PSB^{+}07, PHC^{+}09, RW03, TBL04, WHC09]. Soft [BI00, DCB09, Rik00]. Softer [KRK^{+}02]. Some [Fie00, KRB^{+}04, WW08]. Sonar [BWC08, MS07, MDKM06, MMW08, TRS04]. Sound [GM02, PJC00, RAH01, WHCL02, PR04b, RPW^{+}08, RRB^{+}07, TCSS07]. Sound-Pressure [PJC00]. Source [DMMW08, NBHH00]. Sources [BJ09, ČSB05, Cur06, DMMW08, DRD01, Fie02]. South [GWBH02, RBB08, SL02b, WPB02, BWKS02, BTHE04, DIC04, HWB08, HKT00, ICB04, IWLB05, IWBL07, JEB07, JDB^{+}04, LPWH01, LB08, Luc02, PFS^{+}06, RDBD08, SGPK09, SAC05, VC02]. South-Central [SAC05]. Southeast [BCM^{+}05, Bry00, BCW04, BLM^{+}09, HWB^{+}03, HJ02, MTN09, TWPP06]. Southeastern [BRS02, RMD02, TWRH07, BK09b, MCPM02, SKM^{+}03]. Southern [DRC03, Bar04, GAG^{+}01, HL08, HWKN00, Hor01, KG08, KM00, LL06, MOH07, MMR08, PNSB01, SM00b]. Southwest [MFP^{+}08, RVM^{+}08]. Southwestern [KVMM07, WH04, AMK^{+}05, MR04, RM01, She04, WBE^{+}05, ZAF05, ZC06]. sp [MRC09]. Spatial [CBP07, CGM^{+}06, CRZ05, CM05, Dew08, Dew09, FF04, GCM^{+}03b, Han03, HTA^{+}09, Hor01, HBB07, MMQ01, MCPM02, MSR08, NTSJ03, NBGK09, RHGW05, RPW09b, SVPP08, SSWC05, YI04]. Spatially [GDH04, KF09, RD09, TGR02]. Spatiotemporal [BPBS08]. Spawn [HKE^{+}01, HCFK08]. Spawn-on-Kelp [HKE^{+}01]. Spawned [CLB^{+}07]. Spawner [BM08, EO03, GCM^{+}04, HI01]. Spawners [NRJ05, Wil02]. Spawning [BBK09, BBTC07, BHO02, BF06, BSSS02, BSSC04, BHVS06, CHAO08, CME02, CP04, CL04, CRD05, DHMHJ06, DIC04, Fie00, FSF06, GAM^{+}08, GMHX08, GM00, HTA^{+}09, HKT00, HCFK08, HW07, HLH05, KF09, KGM01, KCR^{+}05, KJR^{+}09, MFT08, MMS08, MMVL05, MS04a, MSW^{+}08, Muh02, MW07, MTM09, MPM09b, MP02, NDH04, PK01, PKW01, Pie04, RM01, RDME02, SSR07, SHPCQ00, SH04a, SDSQ09, SMPM02, SB05b, Und00, VC02, WN04]. Spearing [BRCC03, HBH00, HBH05, PC00]. Specialization [HB07]. Species [BG05, Bar04, BWB^{+}00, BC05, BBJ^{+}00, CWK09, CHR^{+}08, DHM01, DHG02, DP03b, FHE05, FGSJ05, GM05, GMH09, HWB08, HKH^{+}02, HH01c, KMCM01, LGA02, MKP05, MMH09, MHRR07, MHB07, PR03, PHRG00, PT07, PRP00, PSKP09, QBH^{+}09, RHK^{+}03, SCHK03, SWMH06, WCSJ04, YASG07]. Specific [CPG^{+}08, Gol09, KPJ^{+}04, NMJ01, PPF07, Sch04b, SBT^{+}09, WT09]. Specifications [Kol08]. Specimens [BW06]. Sperm [WLHT08]. Spermiating [JSL05]. Spillways [MSWS01]. Spine [Mic05]. Spines [BE01, BIBP02, KQ07, Log07, MNB09]. Spinners [DD04]. Spiny [O’M08, Rul07]. Spirit [LPL04, PSP^{+}01]. Splake [RJB07]. Splittail [DMYC02, HLMB01, SMB07]. Sport [AHH00, Ano06, BB06, BKD^{+}09, DGAG08, KRK^{+}02, LSK^{+}04, SB00, SZL00]. Sportfishing [RHHFB00]. Spotfin [HRK03a]. Spotted [GAR^{+}09, GM00, HM08, JSMV07, LF01, LF05, SWA^{+}08, SM09b, SM06]. spp [GOG09, ZHBB09]. Spread [BLA^{+}07, ROS04, Ris00]. Spring [BHO^{+}08, BM01a, CK09, CLB04, CHJ07, Dar03, EV09, Fie00, HCFK08, KPJ^{+}04, KLH^{+}08, KJS^{+}09, LSK^{+}04, MC05, MPM09a, MPM09b, PJBT09, SHW00, VBHT05, VGM00]. Spring- [KLH^{+}08]. Springs [WW01]. Square [PR01]. Squid [MTL09]. St [BF06, DS05a, DS05b, MMLM06, SFUS07]. St. [Fie00, SGD09]. Stability [TF06]. Stable [KHJC06, MFW^{+}09]. Stage [LT01]. Stages [BBZW04, GHASLC09, KMCM01, PT07]. Standard [BWB^{+}00, DBM04, GHAS04, GASH05, HH01b, NHS08, OW09, Ric07, RR08, SWPO06]. Standard-Weight [GASH05, HH01b]. Standardization [MK09]. Standardized [BABC08, FI01, VR01]. Standardizing [HI01]. Starch [HSB^{+}00]. State [GM00, BBD02, FB07, OMH05]. State-Operated [OMH05]. States [BLK^{+}08, HTGS08, MCPM02, SB05a, SMU05, SFS04]. Statewide [BS01, SZW02]. Station [GM02, PHK03]. Station-Holding [PHK03]. Stationary [CJMP08]. Stations [You09, YDJM09]. Statistical [BMB03, BCW04, HH01c, TTRF05]. Statistically [GCN^{+}08]. Statistics [BP02, SZW02]. Statoliths [MB05b]. Status [BTM07, CPW^{+}06, DJL09, EFS04, GC08, HTGS08, JGW^{+}09, JRNS09, KCR^{+}05, KJR^{+}09, KHR00, LLCBF07, MDM^{+}04, Neg03, SMU05, TTM^{+}05]. Steamboat [Bai03]. Steelhead [Bai03, BSRS04, BG06, BM01a, BBTC07, BSB09, BJ06, EFS04, EWK^{+}08, GG05, GRM07, HRB^{+}04, JAAW05, KPRB04, KBPC08, KLS01, LKS00, MMQ01, MMW^{+}05, MTN09, MP01, MSW^{+}01, NPE^{+}06, NRJ05, NC01, NBGK09, PPAR06, RTBF09, RVM^{+}08, SDSQ09, SBCH07, SMWS02, ST00, SAC05, Tat09, TWP03, Wer07, WPV^{+}04, YS05]. Stem [DHGP03, RDN00, SAC05, Spr05b]. Step [MHOH06]. Sterile [KG00, KGM01]. Sterility [LKS00]. Steroid [EFS04]. Still [Sha01]. Stillaguamish [PPBM04]. Stochasticity [NL06]. Stock [AKARF08, BCM^{+}05, BWW^{+}08, BWJ^{+}08, BCW^{+}09, BJ04, BH05, BKD^{+}09, BHVS06, CH02, DS05a, DA05, Dor02, FHN05, GMHF03, GHDM09, HJC09, HLO08, Ian02, JCW02, KPJ^{+}04, LCL05, MR02, NPE^{+}06, PW09, PAAK07, VTK^{+}07, WG03, WPV^{+}04, YS04]. Stock-Recruit [Dor02]. Stock-Recruitment [MR02]. Stock-Specific [KPJ^{+}04]. Stocked [BH03a, BCAK07, BB02, BHHW02, BBS05, CRL08, DW09, DST00, HBB04, HB05c, HV03, IBSC02, JDB^{+}04, JKS01, KH09, Luc02, MTHH04, MNS01, MWN08, MSF^{+}09, MAC00, NN02, OBG^{+}00, PTN06, PF03, RL02, SMP04, SJ03, VBJR07, VB03, WW04, WWB04, WBJ02, Wil06, WCR^{+}05, ZHBB09]. Stocking [AW04, BKJ06, BK09b, Box02a, BSB03, BBS05, CR04, CRL08, CWDH03, EH01, FHN05, FHE05, GTZM09, HRL06, HBKJ08, JA07, JKHE05, Joh04, LT01, MCS^{+}08, MWN08, Mic09, Mit02, PP01, SL02a, SM07a, SM09b, SRN00, YWM^{+}00]. Stocks [BW06, LB08, Mac02, MHOH06, TWN^{+}07, WGSW07]. Stomach [FM08b, HH01a, PWPC07, WKAP04]. Storage [ML04]. Stored [CWK09]. Straight [Dia08]. Strain [BSRS04, YWM^{+}00]. Strains [MTHH04, Wil06]. Strait [BMK^{+}00, SBNN03]. Strand [SVP^{+}04]. Stranding [BKZA08, SHN05]. Strategies [AW04, AIHRdSD09, BHM^{+}07, CK09, DKM07, Fie04, GTZM09, Hil02, ISBC02, PFWC08, PTD03, SRN00]. Strategy [BLA^{+}07, MKP05, RBB^{+}02]. Stratification [AAF^{+}06]. Stratified [NB05]. Straying [BSRS04]. Stream [BW08, BPT08, BL08, BRSA07, BCW04, BLM^{+}09, CQ05, CHR^{+}08, DHM01, DM07, DD04, DK04, EM06, FJ02, GL08, GLA00, GWSS02, Gol09, HWN09, HK00, Hil02, HW07, HDB00, HS09, JEB07, KRB^{+}04, KM00, LH04b, LLCBF07, LPSL08a, LPSL08b, MFT08, Mea04, MLS06, MBM01a, Muh02, MW07, NGWK08, NL00, PPS03, PBT05, PBB03, RPJB05, RTBF09, RD05a, RBB^{+}02, RHB08, SWA^{+}08, SP07, SA04, SHPCQ00, SW09, SVPP08, SSWC05, SOR05, SAC05, Vok03, WFW02, WW04, WWB04]. Stream-Channel [GLA00]. Stream-Dwelling [BPT08, PBT05]. Stream-Resident [DK04]. Stream-Type [SP07]. Streams [ACB07, BM00, BWEM08, BCD02, BPK^{+}00, Bin04, Bry00, BLK^{+}08, CJPB09, CH00b, CRK^{+}05, DP03b, DF07, DeV07, DST00, DSR03, DRC03, EV09, FM08a, FM09, GG05, GM05, GL02a, HMW^{+}09, HB00, HWH05, HK00, HWKN00, HS08, KSB02, LWHP07, LZS^{+}09, MFP^{+}08, Mar01, ML04, MB05a, McH07, ML03, MHW00, NMP07, OFM^{+}09, OM04, PHRG00, PB00, PFWC08, PBM03, QGBH06, RHK^{+}03, RFL04, RF00, Ron03, RVM^{+}08, RD05b, RD06, RMF^{+}02, RH03, RO06, RPW09a, RPW09b, SLN07, SB05a, SLEM07, SH04b, Sha01, She04, SJ05, SSR09, TPG02, TP06, TPG06, TFW00, VR01, WSH09, WP08, Win04, YGG04]. Streamside [CH00b]. Strength [BPBS08, CDPA09, DD02, ON05, PP01, Qui07, SB04, TKT02]. Stress [CCG09, CHCD02, DMYC02, DO01, OCW04, RCB07, SSS07, VGM00, WBH02]. Stressors [CCG09]. Striking [Arl06]. Striped [BWB05, BBK09, BRGO03, BTI05, BTHE04, CK09, IYJS02, JH01, JPB^{+}07, LL08, MNS01, MMKC05, PR04b, RMD02, RWJD04, SJS00, SM09b, SRN00, SN02, TWRH07, VW01, WMB^{+}00, WCR^{+}05, YI04, YI06, YDJM09]. Strobe [FHS06, HBC08, JBG05, RCB07]. Strongylocentrotus [SSP08]. Structure [AKARF08, Bon02, CBP07, EWD^{+}06, FHS06, GCQVMAVP09, HBJE00, JB00, MGGL04, MSES03, ML03, MC01, NWC^{+}03, PHB^{+}09, PTM03, QSGS04, RD09, RHGW05, RRCJ08, SOM09, SGB07, SVPP08, TP05, VWDK00, VRS01, WHQ^{+}08, ZAF05]. Structured [CPWM06, YS04]. Structures [BHDA05, CQ05, DTMS06, IMSW03, KQH09, LPSL08a, LPSL08b, MFT08, MSP^{+}08, NH09, PEW07, QJBH07, Ris00, RW06, VBP08]. Structuring [FSH06]. Studies [BK09a, BS00, CKSG09, Cur06, DHNH05, GP05, GM04, GB03, RR05]. Study [BG05, DO02, FD01, GJB09, HI01, LR08, MHO01, RD09, RPD09, Sch04a, SVS^{+}09, UWBL05, VGM00, WVT09]. Studying [BTSB07]. Stunted [SL02a]. Sturgeon [AWWG08, BPK08, BSEM05, BMW^{+}02, BKD^{+}09, BFH^{+}08, CHAO08, CWG04, CL04, DR07, DIC04, ESPW03, GP05, GGG08, HFT08, HR02, HLH05, HSH04, IF00, ICBM04, KP07, KDSF07, KQP^{+}09, KCA06, KH01, PK01, PKW01, PH08, PHK03, PBH00, PHB^{+}09, QGP^{+}02, RRB^{+}07, SB05b, SFS04, SVP^{+}04, SMM^{+}07, TKK^{+}01, TPC^{+}09, VEK09, Wan06, WLHT08, WHQ^{+}08, WTW^{+}04, WGSW07]. Sturgeons [BDLR02]. Style [AB02]. Subadult [BJ09, HH01a, LPWH01, MGH^{+}03b, MM05, PASP07, VW07]. Subcutaneous [SMB07]. Subjected [SD06]. Sublegal [MFRS09]. Sublethal [KWA00]. Submergence [HPM08]. Subsampled [BM01b]. Subsamples [NB05]. Subsequent [How06]. Subsidies [SS09]. Substitutability [SD05]. Substitution [HBB07]. Substrate [PSW06, PWH02, RM01]. Substructure [NPE^{+}06]. Subtidal [BSSC04]. Subwatersheds [HTGS08]. Subyearling [CB03, CBYB03, CSB03a, CSA^{+}04, GTRC02, SMH^{+}03, TRW00, TGR02, TGR06, TCGR06]. Success [BHO^{+}08, BTS01, BBGVM05, CRC^{+}08, Dew08, Dew09, FWEP08, HPM08, HFY00, HBKJ08, HA02, MMS08, NBH^{+}03, PHC^{+}09, RDME02, SWPO06, SJ03, SRN00]. Successful [HDB00]. Sucker [MMF^{+}03, MW04, SB06b]. Suckers [BG00, NHS08, PR04a, Pea08, QBH^{+}09, Ric07, SM07a]. Suitability [BBJ^{+}00, GBLC02, HFT08, KRDC08]. Suitable [BKJ06, CB04]. Sulfide [DHM01]. Summary [PH03a]. Summer [ACB07, Bai03, BFR04, BMCB09, CHJ07, HWN09, KPJ^{+}04, KBPC08, KHJ^{+}08, ML04, MBM01b, MH05, PBM03, RBM00, SA04, SWG^{+}06, SVW^{+}08b]. Sunfish [CA06, KHJC06, RWBG04, SPM06, SM09a]. Sunlight [HKFM08]. Superior [SCD^{+}08, BSSS02, CHSS08, DB09b, HMSF08, HBHB00, KHS05, KG00, KGM01, LHSS07, LPO01, NSH^{+}08, RHBS04, Sch08, SYH^{+}07, SEB^{+}09, WHB03, WSM^{+}08, YASG07, YASG08]. Supersaturated [AGM^{+}09, RDN00]. Supplement [MCS^{+}08]. Supplemental [JKHE05]. Supplementation [BSD09, Hil02, PT07]. Supplemented [CKSG09, SMR08]. Supplier [LL08]. Supply [Box02b, HK04]. Support [PLP01]. Suppression [PFWC08]. Surface [BMK^{+}00, GCR08, HPM08, HRB^{+}04, IHW02, JAAW05, Kha06, LRD^{+}04, PSB^{+}07, RGP^{+}07, Wer07, WP08]. Surface-Released [HPM08]. Surficial [ML06]. Surgical [BG06, GL02b, LaV02]. Surgically [SJP05, SH07]. Surplus [JCW02, MTL09]. Surrogates [RWJD04]. Survey [BPBS08, BM08, BJ09, BCW04, DMB09, DHNH07a, Dew08, Dew09, GMH09, LPO01, Loo06, PBT05, PAAK07, Sha04, SBSL06, TMS05, TJKB06, VPR06]. Surveys [AAF^{+}06, GW04, MC04a, SVW^{+}08b, SPR05c, SZW02, SMR08, SYH^{+}07, TTRF05, VWDK00, WVT09, ZS03]. Survival [ASD04, AW04, BTS02, BG06, BBZW04, BS00, BT03, BHM^{+}07, BHHW02, CSB00, CBYB03, Ded07, DS08, DW09, GBPT08, GH08, GL02b, HWN09, HLMB01, HM09, HMS^{+}03, HCS09, HMM^{+}05, IS01, Joh04, JMP07, KH09, KBPC08, KLS01, KG08, KRK^{+}02, KJS^{+}09, KCA06, LT01, ML04, MRJL05, MPB07, MSK05, MFRS09, MC08, MHF^{+}03, MSW^{+}01, MSWS01, MAC00, NN02, NRJ05, OBG^{+}00, PF02, PF03, Rik00, RA04, RPM08, SM09a, SD06, SHD^{+}05, SMH02, SBT^{+}09, SWD08, SMWS02, SMH^{+}03, SM09c, TSMD03, TCD^{+}08, VTK^{+}07, VBHT05, WAB^{+}02, WJKW07, WBJ02, WSM01, Wil06, YI06, You09, YDJM09, ZM06]. Susceptibility [WEC07]. Suspended [WC01]. Susquehanna [KRDC08]. Sustainability [SHN08]. Sustainable [GHM^{+}07, Pow05]. Sustaining [PMPS03]. Suture [SVP^{+}04]. Suwannee [BAC05]. Swab [HSB^{+}00]. Swamp [IS01, RSF06, SN07]. Swellsharks [FRW09]. Swimming [AGB^{+}04, BMP07b, BTSB07, CFL^{+}03, DMYC02, LB09, MMB06, SRGC05, SHMR03, SMB07, WH04]. Sympatric [PFWC08, SWA^{+}08, TKK^{+}01]. Symposium [PH03a]. Synthesis [SEB^{+}09]. System [APS05, BCAK07, BRCC03, CLB^{+}07, CHR^{+}08, DHGP03, FHS06, GAG^{+}01, GM02, HBJ05, JEB03, KHJ^{+}08, MS03a, MBHM00, MM05, MBS^{+}08, NBN04, PJC00, QGP^{+}02, Ryp08, UWBL05, Wer07, WSM01]. Systems [CJMP08, DS05c, FDM^{+}06, FHSM08, HBS^{+}05, MSWS01, NGWK08, PW09, PRP08, PBB03, RPW^{+}08, TFW00, YDJM09].

T [PR05a]. T-Bar [PR05a]. Table [LaV02]. Tackle [CPDS01, Jen03, RHJ^{+}06]. Tag [ARC^{+}07, BGB09, BLB^{+}05, BI00, Buc01, BSB09, CPE09, CRK^{+}05, Cur06, FHSM08, GL02b, HJ06, HHPH03, HCS09, IS01, IK05b, JPB^{+}07, LH04a, MPB07, MBD02, O’M08, PCIB08, PR03, PHHC01, PHHC02, RW03, RB09, Rul07, SHH^{+}04, SMF^{+}09, TWPP06, WJKW07, WW04, APS05].

Tag-Related [RW03]. Tagged [BM01a, HMS^{+}03, LRD^{+}04, MSW^{+}01, MAC00, PB03, PN05, SHW00, VRK00, ZDK^{+}03, CJMP08, HMS^{+}03, MS03a]. Tagging [BPK08, BBO03, BMW^{+}02, CPE09, JBHP07, MMB06, PH03b, PHHC01, PHHC02, SHD^{+}05, SHJ00, VBHT05, Zho02]. Tags [BJ06, CCBW01, Dar03, DB09a, DC06, FSK04, Fri01, GL02b, HHPH03, IF00, JRWK08, KJS^{+}09, LRD^{+}04, MPB07, MBHM00, MMLL03, MAC00, NH01, OGS04, PR03, PR05a, PB03, RB09, Rik00, RFLN01, SL09, SLW09, SMW06, VW01, VSB02]. Tagson [BT03]. Tailrace [DIC04, MMVL05]. Tailwater [BB02, CLB^{+}07, DHM01, HBB04, QK00, SHW00, SCB00]. Tailwaters [HH01a, HB07]. Take [ZS03]. Tale [CMN08]. Talquin [MCS^{+}08]. Tandem [MS02]. Tanner [ZK00]. Target [BM08, TKT02]. Targets [MHOH06, PPSC03]. Tattoos [DC06]. Technique [AHYC09, Bon02, GL02a, HSB^{+}00, Jen03, KQ07, LDMW08, SMF^{+}09]. Techniques [BG08, BSEM05, Col02a, DeV07, Ger07, JEB03, MB05a, McH07, Qui07, RPW^{+}08, RF00, RBB^{+}02, RHB08, RD05b, RD06, SCPC06, SHJ00, THE05, VSB02, YS05]. Technological [SPL03]. Technologies [GB03]. Technology [ALA07, Dau05]. Telemetry [BBO03, BBTC07, CJKS05, DMMW08, JEB03, RR05, YI04]. Temperature [Bar05, BNV02, BP01, CBYB03, DO01, DO02, DSR03, DRC03, EC00, GAM^{+}08, GCR08, HRB^{+}04, JMP07, MC04b, NMP07, PSB^{+}07, SSWC05, SHM^{+}08, WMB^{+}00]. Temperature-Mediated [JMP07]. Temperature-Sensitive [NMP07]. Temperatures [CJPB09, EC00, KC01b, MH05, WHHR01]. Tempering [CBM00]. Temporal [AWM08, DHNH07b, Hor01, MCPM02, MRW05, MTAB08, RHGW05, WVT09, YI04]. Temporary [NJP^{+}09]. Ten [Cla02]. Tennessee [BB02, BB04, BVH00, HB05c, HB07, IBSC02, ISBC02, SBIC02, SIB02, SB04, VB03]. Term [ARC06, CH00a, Dar03, HR06, HW09, JRWK08, MSES03, MMB06, WWB04, ZAF05]. Terminal [CPDS01]. Terrapin [CH01, CH01]. Territory [SHB^{+}04]. Test [CQ05, FCM^{+}09, LCM06b, WCSJ04]. Testing [BMB03, Ian02, TF06, AHYC09]. Tests [HBH05]. Texas [BP02, BBS05, CKSG09, CGM^{+}06, DSM07, DS04, DW08, HD02, JSMV07, KSB^{+}08, LLCBF07, Mun02, SM06, WW01]. Texoma [BBK09]. Their [BTB^{+}02, DMMW08, HBFL08, HH01c, LZS^{+}09, MFT08, MSES03, PTD03, QBH^{+}09, She04, YGG04]. Theory [RGP^{+}07]. There [HPM02, SHD^{+}05]. Thermal [CJPB09, CBM00, HHB05, PBM03, Rei07]. Thermally [RC01]. Thiamine [KCR^{+}05, KJR^{+}09]. Third [TPG06]. Third-Order [TPG06]. Thirteen [PTN06]. Thomas [CKH05]. Thornyhead [HS03]. Threadfin [BSG05]. Threatened [BC05, HRK03a]. Threats [JMHB08]. Three [BG05, Bar04, BGH04, BHHW02, BW09, BSB03, CH01, CHR^{+}08, CMN08, Cun04, DD02, GCM^{+}04, GH08, HH01a, IMSW03, KSB02, KG00, Kha06, MMH09, NC01, NH07, OGK07, PP01, PT07, PBT05, PHB^{+}09, PHC^{+}09, PSKP09, RF00, RBB08, SPL02, SSG05, SM07a, SH04a, She04, SCHK03, SOR05, SBNN03, VBP08, WTH08]. Three-Dimensional [Kha06]. Thresholds [Aus01, SRGC05]. Throat [SCHK03]. Tidal [GM02, GSV^{+}09].

Tiger [Tip01]. Tilapia [HBJ05]. Time [AKKC08, BMP07a, BR06, CPDS01, GM00, HGP04, MC00, PRHS03, PSB^{+}07, PHC^{+}09, SLG02, SSG05, Sha04, SMWS02, SMH^{+}03, You09]. Time-Lapsed [Sha04]. Times [BM01a, IMSW03, VBP08]. Timing [BBGVM05, Dew08, Dew09, EV09, HKT00, HCFK08, HBB07, KPJ^{+}04, MMQ01, MPM09b, ZHBB09]. Tissue [BK09a, BRS02, BPL02, KHJC06]. Tohopekaliga [Ano06, BB06]. Tolerances [TBB^{+}07]. Tool [BS01, Ger07, LD04]. Tools [BWM08, HPB05, JCW02, QBH^{+}09]. Top [Box02a, OGK07]. Top-Down [Box02a]. Top-Level [OGK07]. Total [DTMS06, IV05, MP02, PGR^{+}04, TMS05]. Tournament [ARMB04, GSWS05, GC08, HM08, SD06, TC01, WPS03]. Tournament-Associated [ARMB04]. Tournament-Caught [GSWS05, GC08]. Tournament-Displaced [HM08]. Tournaments [DSM07, ENJO04a, ENJO04b, JSMV07, KSM^{+}03, NLC01, RHHFB00, SWG^{+}06, SKM^{+}03, WSP02]. Tow [CSW05, PRHS03]. Towed [DO02]. Towing [DO01]. Toxicant [RSF06]. Toxicity [LEJ06, RBK09, SN07]. Track [CRK^{+}05]. Tracking [BSSC04, DMMW08, SHJ00]. Trade [PH03b]. Trade-offs [PH03b]. Traditional [PR04b]. Trail [BKZA08]. Training [WHCL02]. Traits [GM05]. Trammel [GOG09]. Trampling [GG09b]. Transformation [MR02]. Transit [BPL02]. Translocations [HFY00]. Transmission [PRP08]. Transmitter [BBTC07, BTSB07, CJKS05, PASP07, WDB06, YI04]. Transmitter-Implanted [CJKS05]. Transmitters [AGB^{+}04, BCP03, CFL^{+}03, CB01, FH02, IYJS02, KPRB04, Pie04, PN05, SJP05, SH07, SVP^{+}04]. Transponder [BGB09, BJ06, Dar03, DB09a, DH09, FHSM08, GL02b, KJS^{+}09, LRD^{+}04, MBHM00, PB03, RFLN01]. Transponders [ARK09, BG06, IC08, MMB06, MSW^{+}01, Tat09, WJKW07]. Transport [CCG09, Sch03]. Transportation [BSS06, EWK^{+}08]. Trap [Ach01, BM08, CME02, DS07, Har03, JB00, MC05, MC06, MRS00, RH04, RO06, RBB08, RT09, SNO02, SLG02, SLN07, SH05b, SCHK03, SWMH06]. Trap-Net [MC06]. Trap-Netting [MC05, RH04]. Trapping [CHCD02]. Traps [Bry00, GAKB04, HBS^{+}08, JSL05, PSP^{+}06]. Trauma [BE02, BERH02, DD04, DK04]. Travel [SMWS02, SMH^{+}03]. Trawl [AAF^{+}06, FSLS09, GMH09, Han03, HPB05, HBS^{+}05, LRD^{+}04, PRHS03, Rul07, SPR05c, SYH^{+}07, YASG08]. Trawl-Caught [PRHS03]. Trawling [BPH08, GCM^{+}03b]. Trawls [BMK^{+}00, IHW02, MD04a, Pin00, WMJBC04]. Treated [LKS00]. Treatment [IDK03, IYJS02]. Treatments [MHB07, MC08, MB09, RWZ09, SL02a]. Trematodes [MHB07]. Trends [AWM08, Bar05, CPW^{+}06, DMB09, GSWS04, HMSM08, MTAB08, SEJCC04, WVT09]. Trestle [DHMHJ06]. Triangulation [JEB03]. Tributaries [BBK09, BSRS04, HBHB00, How06, KG00, MJ05, NC01, Sch00, SPL02, SJ03, WHHR01, ZDK^{+}03]. Tributary [BM08, BLA^{+}07, BSEM05, CLB^{+}07, GRM07, HHAK02, KF09, NJP^{+}09, SVPP08]. Tributary-Spawned [CLB^{+}07]. Trifluoromethyl [SSS08]. Triggered [Dau05]. Trinity [KLH^{+}08]. Trip [HBB07, VJK^{+}07]. Triploid [BBD02, DST00, HM09, NN08, TSMD03]. Trips [KPL09]. Troll [BWJ^{+}08, SPL03]. Trophic [GM00, JM00, KD00, PTD03, SN02]. Trophy [CHD03, DMB09, KMG09]. Tropical [NN02, ON05, WN04]. Trotline [AB02]. Trout [ACBS05, ACB07, ACPHT09, ARC06, BH03a, BKJ06, BM00, BMPU03, BPT08, BTS02, BSD09, BFR04, BGB09, BK09b, BB02, BB04, BC05, BSSS02, BHM^{+}07, BLM^{+}09, BTM07, BMCB09, Car01, CHCD02, Clo00, CJ02, CKH05, CHSS08, CH00b, CW02, CBM03, DMMW08, DJL09, DCB09, DHM01, DHG02, Ded07, DB03, DH09, DST00, DHMHJ06, DD04, DK04, DP07, DBG07, DRD01, DSR03, DRC03, DSW01, EC00, FTWS06, FM08a, FWEP08, GGG01, GAG^{+}01, GZC08, GLA00, GG09b, HB05a, HB05b, HHL^{+}08, HFY00, HWN09, HH01a, HKT00, HB06, HMSF08, HMSM08, HM09, HK00, Hil02, HK04, HWKN00, HBB04, HW07, HOFM06, HKFM08, HBFL08, HDB00, HTGS08, HB07, HH01b, HV03, JGW^{+}09, JEB07, Jen03, JM00, JMP07, JKS01, JRWK08, KHS05, KWA00, KPL09, KCR^{+}05, KHR00, KM00, LKYM09]. Trout [LH04b, LHSS07, MDM^{+}04, MED08, MSR08, MTHH04, MJ05, MRSR01, MSRP01, MHS08, MP01, Mek04, MSES03, MLS06, MZ02, MMF02, MDHK08, MBHM00, MBM01a, MBM01b, Muh02, MGH^{+}03b, MM05, MHW00, MMLL03, MS04b, Neg03, NT04, NMP07, NTW^{+}04, OV01, OGE01, PSWD07, PT07, PFWC08, PBM03, PN05, PN03, PMPS03, PP00, QK00, RAH01, RJB07, RHBS04, RA01, RD05a, RBM00, RD05b, RD06, RMF^{+}02, RPM08, SSR07, SH07, SE00, SLEM07, SHH^{+}02, SHH^{+}04, Sch00, Sch01, Sch03, SRGC05, Sch08, SCD^{+}08, SH04a, SGPK09, SSC07, She04, SMU05, SHW00, SHMR03, SSS07, SSWC05, SWWR04, SM07b, SH08, SHM^{+}08, SCB00, TP06, TSMD03, TPG06, TWN^{+}07, VS05, VB08, WAB^{+}02, WW04, WWB04, WBJ02, WHB03, WHC09, WHNC09, Wil06, WSM^{+}08, WTH08, YMR03, YH04, YGG04]. Trout [YWM^{+}00, ZAF05, ZC06, ZJ02, ZM06, vPP05, vdWF03]. tuberculata [MHB07]. tuberculosa [FPRRHQ09]. Tubes [BTI05, WKAP04]. Tubifex [ZB00]. Tuckasegee [MRC09]. Tuna [BDS^{+}02, MGAC00]. Turbidity [SH02b, TBB^{+}07]. Turbine [Ded07, SMH02]. Turbines [MSWS01, PJC00]. Turbulence [ERO^{+}09]. Turf [BSSS02]. Turtle [AHYC09, GCN^{+}08]. Turtles [AHC^{+}09]. Two [ACBS05, AW04, ARK09, BBK09, BHDA05, BPH08, CJMP08, CRK^{+}05, DHM01, DHG02, DGAG08, EH01, FSH06, FSK04, GW04, HMW^{+}09, HBHB00, HJ02, HMM^{+}05, IHW02, KH09, KPL09, LB08, MP04a, MTHH04, ML03, OSC06, PFS^{+}06, Pin00, PF03, RFL04, RA04, SPL02, TAME03, TSMD03, TGR02, TGR06, VSB02, VS05, VR05, WDB06, Wil06, YH04, YDJM09]. Two-Dimensional [TGR02]. Type [ASD04, AKKC08, AB02, BNV02, DD04, DK04, HK04, Mek04, PH05, SP07, SD05, WF04]. Types [BHDA05, CSB^{+}03b, OSC06, PHB^{+}09, RBB08, TCSS07, VW07, WDB06, ZJ02].

U.S [LGA02, LS08, Bar04, LWHP07]. U.S. [Dia08, GD02, Han03, HPB05, HR06, KHS05, MGAC00, RBB08, SS09, STSS05, WB03, dSCT01]. Ultrasonic [SJP05]. Ultrasound [CWG04, EFS04]. Umatilla [ZSH03]. Unaccounted [Wil02]. Unaltered [GTRC02]. Unavailable [MHO01]. Uncertainties [Zho02]. Uncertainty [BPB07, Fie04, MPBM06, MBD02, NL06, PBI03]. Uncured [SLW09, SMW06]. Undergoing [RPM08]. Undersized [OZR08]. Understanding [ACBS05, Kol08, MS05]. Underwater [BGH04, CQ05, CB04, HB05b, PPF07, Sha01]. Unexploited [KC00, KC01a, vPP05]. Unimpounded [KHJ^{+}08]. Unique [DBG07, FPK07a]. Unit [AWWG08, PBB03, VWDK00]. United [BLK^{+}08, HTGS08, MCPM02, SB05a, SMU05, SFS04]. Units [VEK09]. Unmanaged [FB07]. Unmarked [Zho02]. Unregulated [GG05]. Unscreened [UWBL05]. Unsectioned [Log07]. Unsuccessful [MLS06]. Updated [MH08b]. Upland [SOM09]. Upper [KJR^{+}09, BTS02, BF06, BFH^{+}08, CP04, GM05, KDSF07, KQP^{+}09, KQH09, MVH^{+}09, MGH^{+}03b, MM05, MBS^{+}08, PMH05, Pit02, QGBH06, QJBH07, QBH^{+}09, SMPM02, SB05b, SFUS07, WF04, ZB00, ZDK^{+}03]. Upstream [CJMP08, FDM^{+}06, KPRB04, KMB^{+}09, KG00, MP01, Sch03]. Urban [FM09, NFQ07]. Urbanization [RPM08, SVPP08]. Urbanized [JEB07]. Urchins [SSP08]. USA [GBFJA07, LZS^{+}09, MHB07, PH09, PPBM04, WP08, WM05, WGSW07]. Use [ACB07, BM08, BPT08, BCAK07, BFR04, BHO^{+}08, BK09a, BGH04, BTK04, BE01, BG00, BS01, BB08, BSSC04, ČSB05, CWG04, CB04, CRZ05, CRBD06, CRK^{+}05, DHG02, DS05b, DS05c, DMB09, DS08, DS04, EFS04, GG05, GTRC02, GGG08, HJ04, HEB09, HHPH03, HGF06, HGP04, HTGS08, JZMS05, JEB07, JT07, JKS01, KB04a, KB04b, KG00, KHJ^{+}08, LKYM09, LRBJ08, LSK^{+}04, LH04b, Log07, Loo06, Mac02, MFW^{+}09, Mac07, MH08a, MFP^{+}08, MA00b, MB09, MBM01a, MBM01b, MGH^{+}03b, MM05, MPM09b, NJP^{+}09, NH09, OGS04, PRP08, PPF07, PHHC02, PR04b, QK00, QSGS04, Rei07, RWJD04, RGP^{+}07, RDBD08, RRS07, RVM^{+}08, RBM00, SL02a, SSG05, SH04b, SDSQ09, SMPM02, SHN05, SL02b, SM09d, SMM^{+}07, WMSG09, WW01, WCSJ04, YWM^{+}00, ZDK^{+}03]. Used [Bai03, CHJ07, FHSM08, HB06, IMSW03, KQH09, MD04a, Qui07, QJBH07, SSS07, SCB00, TTRF05, TMS05, WMJBC04]. Uses [WHCL02]. Using [AFM^{+}06, BBO03, BCW04, CJPB09, Dau05, DKM07, DB09b, FHN05, FGSJ05, GHM^{+}07, Ger09, GMH09, HHB05, IC08, JGW^{+}09, JEB03, JBHP07, JRNS09, KH01, MNB09, MBS^{+}08, NTSJ03, PW09, PR03, PHB^{+}09, PBM03, RPW^{+}08, SOM09, Sha01, Sha04, SLB^{+}09, TWN^{+}07, TGR02, TFW00, YS05, YASG07, Zho02, ZS03, KPL09, SJS00]. Utah [HB05a]. Utility [JGW^{+}09, MP08, TPG06, TC01]. Utilization [CSW05, KFG^{+}01, WTH08].

Vaki [SH04c]. Valid [MR06]. Validate [PW09, SDSQ09]. Validation [BPBS08, BH03b, CHJ07, DB06, DV05, DW08, HWKN00, RD05b, RD06, SI08, Stu06, TPG06]. Validity [BIBP02]. Valley [CH00b, JR03, WM05].

Value [CPW^{+}05, HFT08, KNL02, Loo06, SVW^{+}08b]. Vancouver [HS03, MRK08]. Vantage [BSS06]. Varden [BLM^{+}09]. Variability [BRGO03, FPK07b, IMW02, Ise07b, IWBL07, KBPC08, MMH09, MSRP01, MA00b, SOR05]. Variable [AP00b, Box02b, CAM06, CDPA09, DA05, Mac07, Mic09]. Variables [BPK08, DRC03, FM08a, MC00, NBHH01, RPW09b, SWWR04, TCGR06]. Variance [WPS03]. Variation [BCM^{+}05, BPH08, BW06, BTHE04, CCKF^{+}09, CRD05, DPW05, GMHF03, GSC02b, HPB05, HBJE00, Hor01, Mac04, MGGL04, MSR08, NBHH00, NBHH01, Qui07, RHGW05, SBIC02, SIB02, SM09a, SSS08, VR05, WME05, WW01]. Variations [ZHBB09]. Vector [MHB07]. Vegetated [TAME03]. Vegetation [BTK04, PWH02, RG01, Rad06]. Velocity [BCD02, BBZW04, LB09, Pea04]. Verification [BM08, HNTM00, Mac03]. Verifying [Mac04]. versus [FSH06, HMM^{+}05, MMVL05, MMF02, OS05, QGBH06, RGP^{+}07]. Vertebrates [HHB05, LWHP07]. Vertical [Kha06, KH01, SYH^{+}07]. Vertical-Slot [Kha06]. Vessel [PW09]. Vessels [TJKB06]. via [BH05, HCS09, JMHB08, Moh03, PRSA06]. Viability [CRD05]. VIalpha [Rik00]. Video [ALA07, CQ05, CB04, Dau05, DKM07, HHWL00, Sha04]. Vilas [WH09]. Violating [MTN09]. Virginia [BHO02, BM04b, DD02, HBS^{+}08, HS09, MHO01, MHO02, MHOH06, MNS01, OS05, PMH05, SOR05, SN02]. virginica [GBPT08, SBR^{+}05]. Virtual [CPWM06]. Virtual-Population-Analysis-Based [CPWM06]. Virus [SWG^{+}06, SD06]. Visibility [Cur06, FSK04, HB05b, JRWK08, RB09]. Visible [BSR07, Clo00, CJ02, Cur06, HHWL00, IS01, JRWK08, OV01, OGS04, PF09, RB09, SLW09, SCPC06, SMW06, WW04]. Visitation [CPW^{+}05]. Visual [Bon02, BI00, DeV07, FSK04, FHS06, HJ06, MB05a, McH07, PPF07]. Vital [BTM07]. Volume [PASP07]. Volumes [FB07, KS02]. Voluntary [DCB09, Fay03, MTAB08]. Volunteer [BS01, DMB09]. Voyageurs [FPK07a, FPK07b]. Vulnerability [CCBW01, GP05, MVH^{+}09].

W [HN05]. Wabash [KDSF07]. Wadeable [RPW09a, RPW09b]. Waders [GGZH08]. Walker [IDK03, KD00]. Walleye [BMPU03, BRCC03, BCC03, BH05, Box02a, CRD05, DHNH07a, DHNH07b, FHG^{+}01, FHN05, FHE05, Fie02, GWBH02, GP03, GMHF03, Goe02, HNE04, HN05, HBKJ08, IMSW03, IV05, Ise07b, JA07, JKHE05, KH09, KF09, KC01a, LPB04, Log07, Luc02, Mad08, Mit02, Mun02, NBHH00, NBHH01, NBH^{+}03, OBG^{+}00, PGR^{+}04, PS02, PP01, Pit02, PPE03, QGSS03, RHJ^{+}06, RHB05, SHB^{+}04, SL02a, SHNE05, SHN08, SCMF02, Sul03a, Sul03b, THE05, WCE^{+}09, WG03, WH09]. Walleyes [BHL03, BE01, BP01, BHHW02, DHNH05, Fay03, FHE05, FS06, GSWS05, HBH00, HBH05, IK05b, Jon05, KMG09, KSM^{+}03, KC00, KC01b, MS04b, NHNE00, PMH05, Pea08, PTN06, PF02, PF03, QGSS03, QSGS04, RB07, RHB03, SH05a, Sch02, SL02b, Sul02, VFJ^{+}05, VBJR07, VFJ08, VB03, WVT09, YWM^{+}00]. Wallula [GTRC02].

Wanding [VSB02]. Wands [VSB02]. Warming [EV09]. Warmwater [BWB^{+}00, CHR^{+}08, DM07, DM04, GL08, HS08, SVPP08]. Washington [FS00, NTW^{+}04, TFF^{+}07, BMPU03, BCD02, BPK^{+}00, BBD02, BND^{+}07, EM06, FS00, FB07, FSC03, Har03, HHWL00, JAAW05, MMQ01, MHS08, Mue03, MTM09, MPM09b, NFQ07, NTW^{+}04, PPAR06, PPBM04, RFL04, RF00, SHPCQ00, TBL04, Tip01, VTK^{+}07, Wer07]. Wastewater [MB08]. Water [Bar05, BKZA08, BNV02, BBD02, BP01, BCP03, CAM06, CJPB09, CRC^{+}08, DSR03, GGG08, GSV^{+}09, GCR08, HRB^{+}04, KRK^{+}02, MC04b, Mau08, MC08, MB09, MH05, NH07, PSB^{+}07, Pea04, PHC^{+}09, RW06, SMSM01, SGP05, SGB07, SEGG05, TAM^{+}03, WHHR01, ZJ02]. Water-Level [CRC^{+}08, GGG08]. Waterbody [PH05]. Watercourse [CSI^{+}04]. Watermilfoil [UHRP01]. Waters [BPAA03, BKD^{+}09, EH01, GW04, KHS05, KHR00, LCM06a, LPWH01, LSK^{+}05, LHSS07, LPO01, RHBS04, Sta01, WVT09, WHB03, Yul00]. Watershed [AHJ^{+}06, LR08, MVH^{+}09, MH05, SSB01, SGPK09]. Watersheds [KRB^{+}04, RBB^{+}02, VR05]. Wave [PRP08]. Way [SMR08]. Weakfish [NTH08]. Web [MHRR07]. Weight [BWB^{+}00, BMB03, CMN08, DBM04, GHAS04, GASH05, HN05, HH01b, HH01c, NHS08, NMJ01, OW09, PMBM07, Ric07, RR08, SHMR03, VFJ08]. Weight- [NMJ01]. Weights [WP04b]. Weinberg [MTN09]. Weirs [McC01, RVM^{+}08]. Well [BBGVM05, SWG^{+}06]. Wenatchee [MTM09, MPM09a]. West [Ano06, Han03, HPB05, WB03, BB06, Bro02, Dor02, HS03, HPM02, Ian02, Mac02, MB01, MC01, Ral02, Sch04a]. West-Central [MC01, Sch04a]. Western [TMS05, BLK^{+}08, CHSS08, EM06, HB00, HJC09, Ise07b, KSB02, LPSL08a, LPSL08b, LWHP07, Mue03, NSH^{+}08, RHK^{+}03, RF00, SSR07, SB05a, SHPCQ00, SMU05, WP08, WBJ02]. Westslope [BK09b, DSW01, GZC08, MSR08, Sch00, Sch01, Sch03, She04, SMU05, SSWC05]. Wetland [TBB^{+}07]. Wetlands [HGF06]. Wheel [BUH07, Dau05, UBK04]. whether [MFW^{+}09]. While [HBB07]. Whirling [ZB00]. White [ASD04, Bau02, BJ04, BP02, BWKS02, Box02a, Box02b, BSG05, Col02a, Col02b, DD02, Dia08, GP05, GH08, GSC02a, GSC02b, HR02, JB00, KP07, KCA06, LM02, MP02, PK01, PKW01, PH08, Pea08, PPE03, PR04b, QGBS02, QBH^{+}09, RL02, SIB02, SGR02, SR02a, SR02b, Sch04b, SH02b, SB06b, WPB02]. Whitefish [AW04, DB09b, HEB09, MEL09, MEHJ08, MSP^{+}08, PNSB01, PM08, RV08]. Whole [LF01, MPBM06]. Whole-Lake [MPBM06]. Wild [BPK08, BMP07b, Bin04, Car01, CCKF^{+}09, CHCD02, CBWB02, CB03, CBYB03, CSB03a, CMN08, DD04, DK04, FSC03, HKT00, HMM^{+}05, KB04a, MMQ01, MMF02, MBHM00, MRW05, MRK08, NRJ05, PR03, RFL04, SJS00, SBCH07, SSC07, SBNN03, TWP03, WHB03, Wil06, WTH08]. Wild-Caught [BMP07b]. Wildfire [How06]. Wildlife [SMP04]. Will [AHYC09, CRBD06]. Willow [SEGG05]. Wilson [HMM09, MHM09, MMH09]. Wind [AHJ^{+}06]. Windy [NTSJ03]. Winter [BHDA05, CBM00, DHG02, HH01a, IWLB05, KB04b, KLS01, MBM01a, MGH^{+}03b, NRJ05, PALdlN05, RF00]. Winterkill [ICB04]. Wire [ARC^{+}07, BT03, BLB^{+}05, Buc01, BSB09, Fri01, HCS09, IF00, LH04a, MPB07, MMLL03, NH01, PCIB08, VSB02, VBHT05]. Wisconsin [BRCC03, BCC03, CBP07, DHNH05, DHNH07a, DHNH07b, EH01, FHG^{+}01, Fay03, FHN05, FS06, FM08a, HBH00, HNE04, HBH05, HN05, HBJE00, JKHE05, KH09, LHSS07, LZS^{+}09, Mad08, MA00a, MGH03a, MP04b, MFP^{+}08, NBHH01, NBH^{+}03, NHNE00, NH00, RHB03, RHB05, RDME02, SHB^{+}04, SH05a, SHNE05, SHN08, UHRP01, WH09]. within [BMP07a, DSR03, DRC03, HTGS08, KRB^{+}04, SA04, SMU05, SSWC05, VR05]. without [CLB04, MSWS01]. Wood [FB07]. Woodruff [EYI08]. Woody [BKNB04, BPK^{+}00, KSB02, MFT08, Mar01, Ron03, RH03]. Woss [MHRR07]. Woundfin [CCKF^{+}09]. Wyoming [AHJ^{+}06, BG05, Bin04, CKH05, CHR^{+}08, CH00b, GW04, HB05a, HH01a, KHR00, LH04b, PHRG00, SGPK09, SHW00, Yul00].

Yakima [HHWL00, BND^{+}07, PJBT09]. Yampa [JMHB08]. Yankee [MDM^{+}04]. Year [Ano06, BPBS08, BB06, BMW^{+}02, CDPA09, DD02, IBSC02, JDB^{+}04, MCS^{+}08, ON05, PP01, Qui07, RL02, RHBS04, SWD08, SFUS07, SB04]. Year-Class [BPBS08, CDPA09, DD02, IBSC02, JDB^{+}04, ON05, PP01, Qui07, RL02, SFUS07, SB04]. Year-Classes [MCS^{+}08, RHBS04]. Year-Round [BMW^{+}02]. Yearling [AMK^{+}05, CJ02, CSA^{+}04, HMS^{+}03, KH01, MSW^{+}01, SMWS02, VBJR07, WSM01]. Yearlings [GCM^{+}04]. Years [AP00a, BK09b, BS00, BH03b, Cla02, FSK04, LR08, WME05]. Yeast [BSD09]. Yellow [BRB09, BPL02, HL08, HOFM06, Hor01, IHW02, IWLB05, IWBL07, LSK^{+}05, MBS05, MR04, NB05, PMBM07, PMBM08, PWK02, PTN06, RJB07, RM01, SM00b, VBP08, WDB06, WBE^{+}05]. Yellowstone [ESPW03, FSF06, JZMS05, KNL02, KHR00, MSES03, SGPK09]. Yield [DA05, MMLM06, MHF^{+}03, Pow05]. York [HBS^{+}08, BWEM08, CR04, CBK00, HLH05, KJR^{+}09, MHO01, MHO02, MJ05, MM02, VFJ^{+}05, VBJR07, VJK^{+}07, VFJ08]. Young [GP05, HP09, JBHP07]. Yukon [UGM^{+}07, BWW^{+}08, BUH07, STSS05, UBK04].

Zealand [MJ04, BTS02, MJ04, ROS04, SVW08a, UWBL05, VB08, YH04]. Zero [AWWG08]. Zero-Inflated [AWWG08]. Zone [AGM^{+}09, BG05, GHA^{+}02, HBJE00].

References

Adams:2006:IDA

[AAF^{+}06] Jean V. Adams, Ray L. Argyle, Guy W. Fleischer, Gary L. Curtis, and Richard G. Stickel. Improving the design of acoustic and midwater trawl surveys through stratification, with an application to Lake Michigan prey fishes. North American Journal of Fisheries Management, 26(3):612–621, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Arterburn:2002:EHS

[AB02] John E. Arterburn and Charles R. Berry, Jr. Effect of hook style, bait type, and river location on trotline catches of flathead and channel catfish. North American Journal of Fisheries Management, 22(2):573–578, May 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Absolon:2003:CRM

[AB03] Randall F. Absolon and Dean A. Brege. Canister for releasing marked fish at depth in hydroelectric dam gatewells and forebays. North American Journal of Fisheries Management, 23 (2):605–610, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Al-Chokhachy:2007:SMU

[ACB07] Robert Al-Chokhachy and Phaedra Budy. Summer microhabitat use of fluvial bull trout in Eastern Oregon streams. North American Journal of Fisheries Management, 27(4): 1068–1081, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Al-Chokhachy:2005:USR

[ACBS05] Robert Al-Chokhachy, Phaedra Budy, and Howard Schaller. Understanding the significance of redd counts: a comparison between two methods for estimating the abundance of and monitoring bull trout populations. North American Journal of Fisheries Management, 25(4):1505–1512, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Acheson:2001:CGM

[Ach01] James M. Acheson. Confounding the goals of management: Response of the Maine lobster industry to a trap limit. North American Journal of Fisheries Management, 21(2):404–416, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Al-Chokhachy:2009:ERP

[ACPHT09] Robert Al-Chokhachy, Mary Peacock, Lisa G. Heki, and Gary Thiede. Evaluating the reintroduction potential of Lahontan cutthroat trout in Fallen Leaf Lake, California. North American Journal of Fisheries Management, 29(5):1296–1313, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allert:2006:USB

[AFM^{+}06] Ann L. Allert, James F. Fairchild, Tom W. May, Linda C. Sappington, Nathan Darnall, and Maureen Wilson. Using onsite bioassays to determine selenium risk to propagated endangered fishes. North American Journal of Fisheries Management, 26(2):308–316, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anglea:2004:EAT

[AGB^{+}04] Steven M. Anglea, David R. Geist, Richard S. Brown, Katherine A. Deters, and Robert D. McDonald. Effects of acoustic transmitters on swimming performance and predator avoidance of juvenile Chinook salmon. North American Journal of Fisheries Management, 24(1):162–170, February 2004. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Arntzen:2009:IHZ

[AGM^{+}09] Evan V. Arntzen, David R. Geist, Katherine J. Murray, John Vavrinec, Earl M. Dawley, and Dennis E. Schwartz. Influence of the hyporheic zone on supersaturated gas exposure to incubating chum salmon. North American Journal of Fisheries Management, 29(6):1714–1727, December 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Avissar:2009:LGD

[AHC^{+}09] Naomi Avissar, Kristen M. Hart, Larry B. Crowder, Janet Gannon, and Jesse C. Marsh. At loggerheads: Gear Damage in the blue crab fishery attributed to loggerhead sea turtles. North American Journal of Fisheries Management, 29 (1):163–169, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Aas:2000:ARP

[AHH00] Øystein Aas, Wolfgang Haider, and Len Hunt. Angler responses to potential harvest regulations in a Norwegian sport fishery: a conjoint-based choice modeling approach. North American Journal of Fisheries Management, 20(4):940–950, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Amadio:2006:AAS

[AHJ^{+}06] Craig J. Amadio, Wayne A. Hubert, Kevin Johnson, Dennis Oberlie, and David Dufek. Abundance of adult saugers across the Wind River Watershed, Wyoming. North American Journal of Fisheries Management, 26(1):156–162, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Avissar:2009:WIF

[AHYC09] Naomi Avissar, Elliott Hazen, Nancy Young, and Larry Crowder. Will it float? Testing a new technique for reducing loggerhead sea turtle damage to crab pots. North American Journal of Fisheries Management, 29(1):170–175, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Avendano-Ibarra:2009:RSS

[AIHRdSD09] Raymundo Avendaño-Ibarra, Martín E. Hernández-Rivas, and Roxana de Silva-Dávila. Reproductive strategies of sea basses based on larval abundance in Magdalena Bay, Mexico, 1982–1986. North American Journal of Fisheries Management, 29(1):205–215, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Argent:2005:ESG

[AK05] David G. Argent and William G. Kimmel. Efficiency and selectivity of gill nets for assessing fish community composition of large rivers. North American Journal of Fisheries Management, 25(4):1315–1320, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2008:GAS

[AKARF08] Joel D. Anderson, William J. Karel, Kathryn A. Anderson, and Pilar A. Roper-Foo. Genetic assessment of sheepshead stock structure in the Northern Gulf of Mexico: Morphological divergence in the face of gene flow. North American Journal of Fisheries Management, 28(2):592–606, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Arlinghaus:2008:SSI

[AKKC08] Robert Arlinghaus, Thomas Klefoth, Alexander Kobler, and Steven J. Cooke. Size selectivity, injury, handling time, and determinants of initial hooking mortality in recreational angling for northern pike: The influence of type and size of bait. North American Journal of Fisheries Management, 28 (1):123–134, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ames:2007:EVT

[ALA07] Robert T. Ames, Bruce M. Leaman, and Kelly L. Ames. Evaluation of video technology for monitoring of multispecies longline catches. North American Journal of Fisheries Management, 27(3):955–964, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ando:2005:SDY

[AMK^{+}05] Daisei Ando, Mahito Miyamoto, Kiyoshi Kasugai, Yasuyuki Miyakoshi, and Mitsuhiro Nagata. Seasonal distribution of yearling masu salmon released from the Sea of Japan side of Southwestern Hokkaido, Japan. North American Journal of Fisheries Management, 25(1):22–37, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Amezcua:2009:ICJ

[AMVA09] Felipe Amezcua, Juan Madrid-Vera, and Hugo Aguirre. Incidental capture of juvenile fish from an artisanal fishery in a coastal lagoon in the Gulf of California. North American Journal of Fisheries Management, 29(1):245–255, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2002:E

[Ano02] Anonymous. Errata. North American Journal of Fisheries Management, 22(2):702, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2003:Ea

[Ano03a] Anonymous. Errata. North American Journal of Fisheries Management, 23(2):640, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2003:Eb

[Ano03b] Anonymous. Errata. North American Journal of Fisheries Management, 23(4):1369, November 2003. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2004:E

[Ano04] Anonymous. Errata. North American Journal of Fisheries Management, 24(1):326, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2006:ERB

[Ano06] Anonymous. Erratum: Relationship between Habitat and Sport Fish Populations over a 20-Year Period at West Lake Tohopekaliga, Florida. North American Journal of Fisheries Management, 26(1):e1, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic). See [BB06].

Anonymous:2008:Ea

[Ano08a] Anonymous. Erratum. North American Journal of Fisheries Management, 28(5):1648, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2008:Eb

[Ano08b] Anonymous. Erratum. North American Journal of Fisheries Management, 28(6):e1, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Adkison:2000:PBB

[AP00a] Milo D. Adkison and Randall M. Peterman. Predictability of Bristol Bay, Alaska, sockeye salmon returns one to four years in the future. North American Journal of Fisheries Management, 20(1):69–80, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:2000:DFP

[AP00b] Micheal S. Allen and William E. Pine III. Detecting fish population responses to a minimum length limit: Effects of variable recruitment and duration of evaluation. North American Journal of Fisheries Management, 20(3):672–682, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Axel:2005:PTD

[APS05] G. A. Axel, E. F. Prentice, and B. P. Sandford. PIT-Tag detection system for large-diameter juvenile fish bypass pipes at Columbia River Basin hydroelectric dams. North American Journal of Fisheries Management, 25(2):646–651, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Avenetti:2006:STE

[ARC06] Lorraine D. Avenetti, Anthony T. Robinson, and Christopher J. Cantrell. Short-term effectiveness of constructed barriers at protecting Apache trout. North American Journal of Fisheries Management, 26(1):213–216, February 2006. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Adlerstein:2007:ESM

[ARC^{+}07] Sara A. Adlerstein, Edward S. Rutherford, David Clapp, John A. Clevenger, and James E. Johnson. Estimating seasonal movements of Chinook salmon in Lake Huron from efficiency analysis of coded wire tag recoveries in recreational fisheries. North American Journal of Fisheries Management, 27(3):792–803, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Archdeacon:2009:CTM

[ARK09] Thomas P. Archdeacon, W. Jason Remshardt, and Tammy L. Knecht. Comparison of two methods for implanting passive integrated transponders in Rio Grande silvery minnow. North American Journal of Fisheries Management, 29(2):346–351, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Arlinghaus:2006:ASD

[Arl06] Robert Arlinghaus. On the apparently striking disconnect between motivation and satisfaction in recreational fishing: The case of catch orientation of German anglers. North American Journal of Fisheries Management, 26(3):592–605, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:2004:SIT

[ARMB04] M. S. Allen, M. W. Rogers, R. A. Myers, and W. M. Bivin. Simulated impacts of tournament-associated mortality on largemouth bass fisheries. North American Journal of Fisheries Management, 24(4):1252–1261, November 2004. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Askey:2006:LAC

[ARPP06] Paul J. Askey, Shane A. Richards, John R. Post, and Eric A. Parkinson. Linking angling catch rates and fish learning under catch-and-release regulations. North American Journal of Fisheries Management, 26(4):1020–1029, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Aalbers:2004:ECR

[ASD04] Scott A. Aalbers, Greg M. Stutzer, and Mark A. Drawbridge. The effects of catch-and-release angling on the growth and survival of juvenile white seabass captured on offset circle and J-type hooks. North American Journal of Fisheries Management, 24(3):793–800, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:2003:LBA

[ATM03] Mike S. Allen, Kimberly I. Tugend, and Marty J. Mann. Largemouth bass abundance and angler catch rates following a habitat enhancement project at Lake Kissimmee, Florida. North American Journal of Fisheries Management, 23(3): 845–855, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Auster:2001:DTP

[Aus01] Peter J. Auster. Defining thresholds for precautionary habitat management actions in a fisheries context. North American Journal of Fisheries Management, 21(1):1–9, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Amtstaetter:2004:SGL

[AW04] Frank Amtstaetter and Campbell C. Willox. Survival and growth of lake whitefish from two stocking strategies in Lake Simcoe, Ontario. North American Journal of Fisheries Management, 24(4):1214–1220, November 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:2008:TTL

[AWM08] Micheal S. Allen, Carl J. Walters, and Randall Myers. Temporal trends in largemouth bass mortality, with fishery implications. North American Journal of Fisheries Management, 28(2):418–427, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Arab:2008:ZIM

[AWWG08] Ali Arab, Mark L. Wildhaber, Christopher K. Wikle, and Casey N. Gentry. Zero-inflated modeling of fish catch per unit area resulting from multiple gears: Application to channel catfish and shovelnose sturgeon in the Missouri River. North American Journal of Fisheries Management, 28(4): 1044–1058, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2008:CLB

[BABC08] Kimberly I. Bonvechio, Micheal S. Allen, Timothy F. Bonvechio, and Tim P. Coughlin. Comparison of largemouth bass assessment metrics between standardized and historical sampling designs at six Florida lakes. North American Journal of Fisheries Management, 28(4):1132–1137, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2005:RAG

[BAC05] Timothy F. Bonvechio, Micheal S. Allen, and Richard L. Cailteux. Relative abundance, growth, and mortality of suwannee bass in four Florida rivers. North American Journal of Fisheries Management, 25(1):275–283, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baigun:2003:CDP

[Bai03] Claudio R. M. Baigún. Characteristics of deep pools used by adult summer steelhead in Steamboat Creek, Oregon.

North American Journal of Fisheries Management, 23(4): 1167–1174, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barwick:2004:SRC

[Bar04] D. Hugh Barwick. Species richness and centrarchid abundance in littoral habitats of three Southern U.S. reservoirs. North American Journal of Fisheries Management, 24(1): 76–81, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bartholow:2005:RWT

[Bar05] John M. Bartholow. Recent water temperature trends in the Lower Klamath River, California. North American Journal of Fisheries Management, 25(1):152–162, February 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bauer:2002:WBP

[Bau02] Daryl L. Bauer. White bass population differences in Nebraska reservoirs with gizzard shad or alewife prey bases. North American Journal of Fisheries Management, 22(2): 665–670, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bettinger:2002:FDP

[BB02] Jason M. Bettinger and Phillip W. Bettoli. Fate, dispersal, and persistence of recently stocked and resident rainbow trout in a Tennessee tailwater. North American Journal of Fisheries Management, 22(2):425–432, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bettinger:2004:SMB

[BB04] Jason M. Bettinger and Phillip W. Bettoli. Seasonal movement of brown trout in the Clinch River, Tennessee. North American Journal of Fisheries Management, 24(4):1480– 1485, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2006:RBH

[BB06] Kimberly I. Bonvechio and Timothy F. Bonvechio. Relationship between habitat and sport fish populations over a 20-year period at West Lake Tohopekaliga, Florida. North American Journal of Fisheries Management, 26(1):124–133, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic). See erratum [Ano06].

Brenden:2008:CUP

[BB08] Travis O. Brenden and James R. Bence. Comment: Use of piecewise regression models to estimate changing relationships in fisheries. North American Journal of Fisheries Management, 28(3):844–846, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonar:2002:ETG

[BBD02] Scott A. Bonar, Bruce Bolding, and Marc Divens. Effects of triploid grass carp on aquatic plants, water quality, and public satisfaction in Washington State. North American Journal of Fisheries Management, 22(1):96–105, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bernier-Bourgault:2005:NIA

[BBGVM05] Isabel Bernier-Bourgault, François Guillemette, Charles Vallée, and Pierre Magnan. A new incubator for the assessment of hatching and emergence success as well as the timing of emergence in salmonids. North American Journal of Fisheries Management, 25(1):16–21, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:2000:HSI

[BBJ^{+}00] Stephen K. Brown, Kenneth R. Buja, Steven H. Jury, Mark E. Monaco, and Arnold Banner. Habitat suitability index models for eight fish and invertebrate species in Casco and Sheepscot Bays, Maine. North American Journal of Fisheries Management, 20(2):408–435, May 2000. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baker:2009:DSB

[BBK09] William P. Baker, Jeff Boxrucker, and Kurt E. Kuklinski. Determination of striped bass spawning locations in the two major tributaries of Lake Texoma. North American Journal of Fisheries Management, 29(4):1006–1014, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baxter:2003:ACF

[BBO03] James S. Baxter, Gary J. Birch, and W. R. Olmsted. Assessment of a constructed fish migration barrier using radio telemetry and floy tagging. North American Journal of Fisheries Management, 23(4):1030–1035, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buckmeier:2005:IPS

[BBS05] David L. Buckmeier, Robert K. Betsill, and J. Warren Schlechte. Initial predation of stocked fingerling largemouth bass in a Texas reservoir and implications for improving stocking efficiency. North American Journal of Fisheries Management, 25(2):652–659, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Berejikian:2007:ETT

[BBTC07] Barry A. Berejikian, Richard S. Brown, Christopher P. Tatara, and Steven J. Cooke. Effects of telemetry transmitter placement on egg retention in naturally spawning, captively reared steelhead. North American Journal of Fisheries Management, 27(2):659–664, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bestgen:2004:EHV

[BBZW04] Kevin R. Bestgen, Jay M. Bundy, Koreen A. Zelasko, and Tony L. Wahl. Effectiveness of high-velocity inclined profilebar fish screens measured by exclusion and survival of early life stages of fathead minnow. North American Journal of Fisheries Management, 24(4):1228–1239, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brenkman:2005:EAB

[BC05] Samuel J. Brenkman and Stephen C. Corbett. Extent of anadromy in bull trout and implications for conservation of a threatened species. North American Journal of Fisheries Management, 25(3):1073–1081, August 2005. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barry:2007:MHU

[BCAK07] Patrick M. Barry, Robert F. Carline, David G. Argent, and William G. Kimmel. Movement and habitat use of stocked juvenile paddlefish in the Ohio River system, Pennsylvania. North American Journal of Fisheries Management, 27(4): 1316–1325, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beard:2003:IDB

[BCC03] T. Douglas Beard, Jr., Sean P. Cox, and Stephen R. Carpenter. Impacts of daily bag limit reductions on angler effort in Wisconsin walleye lakes. North American Journal of Fisheries Management, 23(4):1283–1293, November 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beecher:2002:EDV

[BCD02] Hal A. Beecher, Brad A. Caldwell, and S. Brett DeMond. Evaluation of depth and velocity preferences of juvenile Coho salmon in Washington streams. North American Journal of Fisheries Management, 22(3):785–795, August 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beacham:2005:DLV

[BCM^{+}05] Terry D. Beacham, John R. Candy, Brenda McIntosh, Cathy MacConnachie, Amy Tabata, Kristina M. Miller, and Ruth E. Withler. DNA-Level variation of sockeye salmon in Southeast Alaska and the Nass and Skeena Rivers, British Columbia, with applications to stock identification. North American Journal of Fisheries Management, 25(3):763–776, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bunt:2003:MDD

[BCP03] Christopher M. Bunt, Steven J. Cooke, and David P. Philipp. A modified downrigger for detecting radio transmitters in deep water. North American Journal of Fisheries Management, 23(2):258–263, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bryant:2004:ESH

[BCW04] Mason D. Bryant, John P. Caouette, and Brenda E. Wright. Evaluating stream habitat survey data and statistical power using an example from Southeast Alaska. North American Journal of Fisheries Management, 24(4):1353–1362, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beacham:2009:MSI

[BCW^{+}09] Terry D. Beacham, John R. Candy, C. Wallace, Shigehiko Urawa, Shunpei Sato, Natalia V. Varnavskaya, Khai D. Le, and Michael Wetklo. Microsatellite stock identification of chum salmon on a Pacific rim basis. North American Journal of Fisheries Management, 29(6):1757–1776, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brosse:2002:EGL

[BDLR02] Laurent Brosse, Pierre Dumont, Mario Lepage, and Eric Rochard. Evaluation of a gastric lavage method for sturgeons. North American Journal of Fisheries Management, 22(3):955–960, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bohnsack:2002:EIR

[BDS^{+}02] B. L. Bohnsack, R. B. Ditton, J. R. Stoll, R. J. Chen, R. Novak, and L. S. Smutko. The economic impacts of the recreational bluefin tuna fishery in Hatteras, North Carolina. North American Journal of Fisheries Management, 22(1): 165–176, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Borkholder:2001:CUD

[BE01] Brian D. Borkholder and Andrew J. Edwards. Comparing the use of dorsal fin spines with scales to back-calculate length-at-age estimates in walleyes. North American Journal of Fisheries Management, 21(4):935–942, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Backman:2002:GBTa

[BE02] Thomas W. H. Backman and Allen F. Evans. Gas bubble trauma incidence in adult salmonids in the Columbia River Basin. North American Journal of Fisheries Management, 22 (2):579–584, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Backman:2002:GBTb

[BERH02] Thomas W. H. Backman, Allen F. Evans, Michael S. Robertson, and Michael A. Hawbecker. Gas bubble trauma incidence in juvenile salmonids in the Lower Columbia and Snake Rivers. North American Journal of Fisheries Management, 22(3):965–972, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bosworth:2006:GDA

[BF06] Aaron Bosworth and John M. Farrell. Genetic divergence among northern pike from spawning locations in the Upper St. Lawrence River. North American Journal of Fisheries Management, 26(3):676–684, August 2006. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buchanan:2002:RGN

[BFF^{+}02] Seana Buchanan, Anthony P. Farrell, Jake Fraser, Patricia Gallaugher, Ruth Joy, and Rick Routledge. Reducing gill-net mortality of incidentally caught Coho salmon. North American Journal of Fisheries Management, 22(4): 1270–1275, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Braaten:2008:DDL

[BFH^{+}08] Patrick J. Braaten, David B. Fuller, Landon D. Holte, Ryan D. Lott, William Viste, Tyrel F. Brandt, and Robert G. Legare. Drift dynamics of larval pallid sturgeon and shovelnose sturgeon in a natural side channel of the Upper Missouri River, Montana. North American Journal of Fisheries Management, 28(3):808–826, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barwick:2004:SHU

[BFR04] D. Hugh Barwick, Jeffrey W. Foltz, and Daniel M. Rankin. Summer habitat use by rainbow trout and brown trout in Jocassee Reservoir. North American Journal of Fisheries Management, 24(2):735–740, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bradford:2000:HUH

[BG00] Richard H. Bradford and Scott D. Gurtin. Habitat use by hatchery-reared adult razorback suckers released into the Lower Colorado River, California–Arizona. North American Journal of Fisheries Management, 20(1):154–167, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bailey:2005:SEN

[BG05] Paul E. Bailey and Kenneth G. Gerow. Seining effort needed to estimate species richness of small littoral zone fishes: a simulation study in three Wyoming reservoirs. North American Journal of Fisheries Management, 25(4):1310–1314, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bateman:2006:SGA

[BG06] Douglas S. Bateman and Robert E. Gresswell. Survival and growth of age-0 steelhead after surgical implantation of 23mm passive integrated transponders. North American Journal of Fisheries Management, 26(3):545–550, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beesley:2008:MDN

[BG08] L. S. Beesley and J. P. Gilmour. A modified drop net for sampling fish communities in complex habitats: a description and comparison with other techniques. North American Journal of Fisheries Management, 28(5):1214–1222, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bateman:2009:PIT

[BGB09] Douglas S. Bateman, Robert E. Gresswell, and Aaron M. Berger. Passive integrated transponder tag retention rates in headwater populations of coastal cutthroat trout. North American Journal of Fisheries Management, 29(3):653–657, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beeman:2004:CTU

[BGH04] John W. Beeman, Cam Grant, and Philip V. Haner. Comparison of three underwater antennas for use in radiotelemetry. North American Journal of Fisheries Management, 24 (1):275–281, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bailey:2003:FAS

[BH03a] Paul E. Bailey and Wayne A. Hubert. Factors associated with stocked cutthroat trout populations in high-mountain lakes. North American Journal of Fisheries Management, 23 (2):611–618, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buckmeier:2003:VOE

[BH03b] David L. Buckmeier and Robert G. Howells. Validation of otoliths for estimating ages of largemouth bass to 16 years. North American Journal of Fisheries Management, 23(2): 590–593, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bickford:2005:SIW

[BH05] Nate Bickford and Robyn Hannigan. Stock identification of walleye via otolith chemistry in the Eleven Point River, Arkansas. North American Journal of Fisheries Management, 25(4):1542–1549, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barrineau:2005:WIP

[BHDA05] Christina E. Barrineau, Wayne A. Hubert, Paul D. Dey, and Thomas C. Annear. Winter ice processes and pool habitat associated with two types of constructed instream structures. North American Journal of Fisheries Management, 25(3): 1022–1033, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brooks:2002:RST

[BHHW02] Ronald C. Brooks, Roy C. Heidinger, R. John H. Hoxmeier, and David H. Wahl. Relative survival of three sizes of walleyes stocked into Illinois lakes. North American Journal of Fisheries Management, 22(3):995–1006, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:2003:PLC

[BHL03] Brian G. Blackwell, Matthew J. Hubers, and Robert G. Losco. Postharvest length changes of angler-caught walleyes. North American Journal of Fisheries Management, 23(3): 770–778, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bronte:2007:RAS

[BHM^{+}07] Charles R. Bronte, Mark E. Holey, Charles P. Madenjian, Jory L. Jonas, Randall M. Claramunt, Patrick C. McKee, Retired Michael L. Toneys, Mark P. Ebener, Brian Breidert, Guy W. Fleischer, Retired Richard Hess, Archie W. Martell, Jr., and Erik J. Olsen. Relative abundance, site fidelity, and survival of adult lake trout in Lake Michigan from 1999 to 2001: Implications for future restoration strategies. North American Journal of Fisheries Management, 27(1):137–155, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bilkovic:2002:MAA

[BHO02] Donna Marie Bilkovic, Carl H. Hershner, and John E. Olney. Macroscale assessment of American shad spawning and nursery habitat in the Mattaponi and Pamunkey Rivers, Virginia. North American Journal of Fisheries Management, 22 (4):1176–1192, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baumsteiger:2008:UPA

[BHO^{+}08] Jason Baumsteiger, David M. Hand, Douglas E. Olson, Robert Spateholts, Geoff FitzGerald, and William R. Ardren. Use of parentage analysis to determine reproductive success of hatchery-origin spring Chinook salmon outplanted into Shitike Creek, Oregon. North American Journal of Fisheries Management, 28(5):1472–1485, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bunnell:2006:PCR

[BHVS06] David B. Bunnell, R. Scott Hale, Michael J. Vanni, and Roy A. Stein. Predicting crappie recruitment in Ohio reservoirs with spawning stock size, larval density, and chlorophyll concentrations. North American Journal of Fisheries Management, 26(1):1–12, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buckmeier:2000:ESV

[BI00] David L. Buckmeier and Elise R. Irwin. An evaluation of soft visual implant tag retention compared with anchor tag retention in channel catfish. North American Journal of Fisheries Management, 20(1):296–298, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boxrucker:2002:CCM

[BI02] Jeff Boxrucker and Elise Irwin. Challenges of crappie management continuing into the 21st century. North American Journal of Fisheries Management, 22(4):1334–1339, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buckmeier:2002:VOP

[BIBP02] David L. Buckmeier, Elise R. Irwin, Robert K. Betsill, and John A. Prentice. Validity of otoliths and pectoral spines for estimating ages of channel catfish. North American Journal of Fisheries Management, 22(3):934–942, August 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Binns:2004:EHM

[Bin04] N. Allen Binns. Effectiveness of habitat manipulation for wild salmonids in Wyoming streams. North American Journal of Fisheries Management, 24(3):911–921, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Belcher:2004:ESR

[BJ04] Carolyn N. Belcher and Cecil A. Jennings. Evaluation of stock–recruitment curves for white shrimp in Georgia. North American Journal of Fisheries Management, 24(2):654–661, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burke:2006:PPI

[BJ06] Brian J. Burke and Michael A. Jepson. Performance of passive integrated transponder tags and radio tags in determining dam passage behavior of adult Chinook salmon and steelhead. North American Journal of Fisheries Management, 26 (3):742–752, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Belcher:2009:PSS

[BJ09] Carolyn N. Belcher and Cecil A. Jennings. Potential sources of survey bias associated with hand-retrieved longline catches of subadult sharks in Georgia estuaries. North American Journal of Fisheries Management, 29(6):1676–1685, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baumsteiger:2009:ESC

[BK09a] Jason Baumsteiger and Jacob L. Kerby. Effectiveness of salmon carcass tissue for use in DNA extraction and amplification in conservation genetic studies. North American Journal of Fisheries Management, 29(1):40–49, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bennett:2009:LIW

[BK09b] Stephen N. Bennett and Jeffrey L. Kershner. Levels of introgression in westslope cutthroat trout populations nine years after changes to rainbow trout stocking programs in Southeastern British Columbia. North American Journal of Fisheries Management, 29(5):1271–1282, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bott:2009:MSA

[BKD^{+}09] Kristin Bott, Gregory W. Kornely, Michael C. Donofrio, Robert F. Elliott, and Kim T. Scribner. Mixed-stock analysis of lake sturgeon in the Menominee River sport harvest and adjoining waters of Lake Michigan. North American Journal of Fisheries Management, 29(6):1636–1643, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baird:2006:GMC

[BKJ06] Owen E. Baird, Charles C. Krueger, and Daniel C. Josephson. Growth, movement, and catch of brook, rainbow, and brown trout after stocking into a large, marginally suitable Adirondack river. North American Journal of Fisheries Management, 26(1):180–189, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barwick:2004:FPA

[BKNB04] Robert D. Barwick, Thomas J. Kwak, Richard L. Noble, and D. Hugh Barwick. Fish populations associated with habitatmodified piers and natural woody debris in Piedmont Carolina reservoirs. North American Journal of Fisheries Management, 24(4):1120–1133, November 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bell:2008:SFS

[BKZA08] Ethan Bell, Sharon Kramer, David Zajanc, and Jennifer Aspittle. Salmonid fry stranding mortality associated with daily water level fluctuations in Trail Bridge Reservoir, Oregon. North American Journal of Fisheries Management, 28(5): 1515–1528, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bayley:2008:SFR

[BL08] Peter B. Bayley and Hiram W. Li. Stream fish responses to grazing exclosures. North American Journal of Fisheries Management, 28(1):135–147, February 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bartholomew:2007:EMS

[BLA^{+}07] Jerri L. Bartholomew, Harriet V. Lorz, Stephen D. Atkinson, Sascha L. Hallett, Donald G. Stevens, Richard A. Holt, Kenneth Lujan, and Antonio Amandi. Evaluation of a management strategy to control the spread of Myxobolus cerebralis in a Lower Columbia River tributary. North American Journal of Fisheries Management, 27(2):542–550, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brennan:2005:ECW

[BLB^{+}05] Nathan P. Brennan, Kenneth M. Leber, H. Lee Blankenship, John M. Ransier, and Roger DeBruler, Jr. An evaluation of coded wire and elastomer tag performance in juvenile common snook under field and laboratory conditions. North American Journal of Fisheries Management, 25(2):437–445, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bryce:2008:DBB

[BLK^{+}08] Sandra A. Bryce, Gregg A. Lomnicky, Philip R. Kaufmann, Lynne S. McAllister, and Ted L. Ernst. Development of biologically based sediment criteria in mountain streams of the Western United States. North American Journal of Fisheries Management, 28(6):1714–1724, December 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bryant:2009:SMD

[BLM^{+}09] M. D. Bryant, M. D. Lukey, J. P. McDonell, R. A. Gubernick, and R. S. Aho. Seasonal movement of Dolly Varden and cutthroat trout with respect to stream discharge in a secondorder stream in Southeast Alaska. North American Journal of Fisheries Management, 29(6):1728–1742, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bakke:2000:CSA

[BM00] Paul D. Bakke and Kevin A. Meyer. Comment: Snorkeling as an alternative to depletion electrofishing for estimating abundance and length-class frequencies of trout in small streams. North American Journal of Fisheries Management, 20(3):838–839, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beeman:2001:RTD

[BM01a] John W. Beeman and Alec G. Maule. Residence times and diel passage distributions of radio-tagged juvenile spring Chinook salmon and steelhead in a gatewell and fish collection channel of a Columbia River Dam. North American Journal of Fisheries Management, 21(3):455–463, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bettoli:2001:CNA

[BM01b] Phillip W. Bettoli and Leandro E. Miranda. Cautionary note about estimating mean length at age with subsampled data. North American Journal of Fisheries Management, 21(2): 425–428, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:2002:RSB

[BM02] Gerard L. Buynak and Bill Mitchell. Response of smallmouth bass to regulatory and environmental changes in Elkhorn Creek, Kentucky. North American Journal of Fisheries Management, 22(2):500–508, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barker:2004:IMJ

[BM04a] Jennifer M. Barker and Kenneth R. McKaye. Immersion marking of juvenile Midas cichlids with oxytetracycline. North American Journal of Fisheries Management, 24(1): 262–269, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brenden:2004:EAA

[BM04b] Travis O. Brenden and Brian R. Murphy. Experimental assessment of age-0 largemouth bass and juvenile bluegill competition in a small impoundment in Virginia. North American Journal of Fisheries Management, 24(3):1058–1070, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baldwin:2008:UGN

[BM08] Casey M. Baldwin and Jason G. McLellan. Use of gill nets for target verification of a hydroacoustic fisheries survey and comparison with Kokanee spawner escapement estimates from a tributary trap. North American Journal of Fisheries Management, 28(6):1744–1757, December 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brenden:2003:SPR

[BMB03] Travis O. Brenden, Brian R. Murphy, and Jeffrey B. Birch. Statistical properties of the relative weight (W_{r}) index and an alternative procedure for testing W_{r} differences between groups. North American Journal of Fisheries Management, 23(4):1136–1151, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burford:2009:ATP

[BMCB09] D. Drake Burford, Thomas E. McMahon, Joel E. Cahoon, and Matthew Blank. Assessment of trout passage through culverts in a large Montana drainage during summer low flow. North American Journal of Fisheries Management, 29 (3):739–752, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beamish:2000:EAJ

[BMK^{+}00] R. J. Beamish, D. McCaughran, J. R. King, R. M. Sweeting, and G. A. McFarlane. Estimating the abundance of juvenile Coho salmon in the Strait of Georgia by means of surface trawls. North American Journal of Fisheries Management, 20(2):369–375, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buynak:2001:RAF

[BMK^{+}01] Gerard L. Buynak, Bill Mitchell, Lewis Kornman, Albert F. Surmont, Brian Reeder, and Steve Malvestuto. Responses to artificial fertilization at Grayson Lake, Kentucky. North American Journal of Fisheries Management, 21(2):393–403, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Babcock:2007:CHC

[BMP07a] Elizabeth A. Babcock, Murdoch K. McAllister, and Ellen K. Pikitch. Comparison of harvest control policies for rebuilding overfished populations within a fixed rebuilding time frame. North American Journal of Fisheries Management, 27(4): 1326–1342, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beecham:2007:CSP

[BMP07b] Rachel Venn Beecham, C. Douglas Minchew, and Glenn R. Parsons. Comparative swimming performance of juvenile pond-cultured and wild-caught channel catfish. North American Journal of Fisheries Management, 27(3):729–734, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baldwin:2003:WPH

[BMPU03] Casey M. Baldwin, Jason G. McLellan, Matt C. Polacek, and Keith Underwood. Walleye predation on hatchery releases of Kokanees and rainbow trout in Lake Roosevelt, Washington. North American Journal of Fisheries Management, 23 (3):660–676, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Borkholder:2002:EYR

[BMW^{+}02] Brian D. Borkholder, Stephen D. Morse, Heath T. Weaver, Roger A. Hugill, Al T. Linder, Larry M. Schwarzkopf, Terry E. Perrault, Mark J. Zacher, and John A. Frank. Evidence of a year-round resident population of lake sturgeon in the Kettle River, Minnesota, based on radiotelemetry and tagging. North American Journal of Fisheries Management, 22(3):888–894, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bosch:2007:EFR

[BND^{+}07] William J. Bosch, Todd H. Newsome, James L. Dunnigan, Joel D. Hubble, Douglas Neeley, David T. Lind, David E. Fast, Linda L. Lamebull, and Joseph W. Blodgett. Evaluating the feasibility of reestablishing a Coho salmon population in the Yakima River, Washington. North American Journal of Fisheries Management, 27(1):198–214, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blann:2002:RRB

[BNV02] Kristen Blann, Julia Frost Nerbonne, and Bruce Vondracek. Relationship of riparian buffer type to water temperature in the Driftless Area ecoregion of Minnesota. North American Journal of Fisheries Management, 22(2):441–451, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonar:2002:RLF

[Bon02] Scott A. Bonar. Relative length frequency: a simple, visual technique to evaluate size structure in fish populations. North American Journal of Fisheries Management, 22(4): 1086–1094, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boxrucker:2002:IGW

[Box02a] Jeff Boxrucker. Improved growth of a white crappie population following stocking of saugeyes (sauger × walleye): a topdown, density-dependent growth response. North American Journal of Fisheries Management, 22(4):1425–1437, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boxrucker:2002:RMM

[Box02b] Jeff Boxrucker. Rescinding a 254-mm minimum length limit on white crappies at Ft. Supply Reservoir, Oklahoma: The influence of variable recruitment, compensatory mortality, and angler dissatisfaction. North American Journal of Fisheries Management, 22(4):1340–1348, November 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Borkholder:2001:RBE

[BP01] Brian D. Borkholder and Bradford G. Parsons. Relationship between electrofishing catch rates of age-0 walleyes and water temperature in Minnesota lakes. North American Journal of Fisheries Management, 21(2):318–325, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Betsill:2002:CCS

[BP02] Robert K. Betsill and Veronica M. Pitman. Comparison of creel statistics for river and reservoir components of a Texas white bass fishery. North American Journal of Fisheries Management, 22(2):659–664, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Banta:2003:EDE

[BPAA03] Sarah E. Banta, Eric N. Powell, and Kathryn A. Ashton-Alcox. Evaluation of dredging effort by the Delaware Bay oyster fishery in New Jersey waters. North American Journal of Fisheries Management, 23(3):732–741, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bodtker:2007:AUE

[BPB07] Karin M. Bodtker, Randall M. Peterman, and Michael J. Bradford. Accounting for uncertainty in estimates of escapement goals for Fraser River sockeye salmon based on productivity of nursery lakes in British Columbia, Canada. North American Journal of Fisheries Management, 27(1):286–302, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bacheler:2008:RJR

[BPBS08] Nathan M. Bacheler, Lee M. Paramore, Jeffrey A. Buckel, and Frederick S. Scharf. Recruitment of juvenile red drum in North Carolina: Spatiotemporal patterns of year-class strength and validation of a seine survey. North American Journal of Fisheries Management, 28(4):1086–1098, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2008:VBE

[BPH08] Timothy F. Bonvechio, William F. Pouder, and Marty M. Hale. Variation between electrofishing and otter trawling for sampling black crappies in two Florida lakes. North American Journal of Fisheries Management, 28(1):188–192, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beechie:2000:MRR

[BPK^{+}00] Timothy J. Beechie, George Pess, Paul Kennard, Robert E. Bilby, and Susan Bolton. Modeling recovery rates and pathways for woody debris recruitment in Northwestern Washington streams. North American Journal of Fisheries Management, 20(2):436–452, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baker:2008:ECH

[BPK08] Daniel W. Baker, Stephan J. Peake, and James D. Kieffer. The effect of capture, handling, and tagging on hematological variables in wild adult lake sturgeon. North American Journal of Fisheries Management, 28(1):296–300, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:2002:TOM

[BPL02] Michael L. Brown, Jennifer L. Powell, and David O. Lucchesi. In-transit oxytetracycline marking, nonlethal mark detection, and tissue residue depletion in yellow perch. North American Journal of Fisheries Management, 22(2):236–242, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beechie:2008:SRR

[BPRG08] T. Beechie, G. Pess, P. Roni, and G. Giannico. Setting river restoration priorities: a review of approaches and a general protocol for identifying and prioritizing actions. North American Journal of Fisheries Management, 28(3):891–905, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bradford:2000:BRS

[BPS00] Michael J. Bradford, Brian J. Pyper, and Ken S. Shortreed. Biological responses of sockeye salmon to the fertilization of Chilko Lake, a large lake in the interior of British Columbia. North American Journal of Fisheries Management, 20(3): 661–671, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Banish:2008:PBS

[BPT08] Nolan P. Banish, James T. Peterson, and Russell F. Thurow. Physical, biotic, and sampling influences on diel habitat use by stream-dwelling bull trout. North American Journal of Fisheries Management, 28(1):176–187, February 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Breen:2006:GBF

[BR06] Matthew J. Breen and Carl R. Ruetz III. Gear bias in fyke netting: Evaluating soak time, fish density, and predators. North American Journal of Fisheries Management, 26(1): 32–41, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bauer:2009:DMP

[BRB09] William F. Bauer, Nicholas B. Radabaugh, and Michael L. Brown. Diel movement patterns of yellow perch in a simple and a complex lake basin. North American Journal of Fisheries Management, 29(1):64–71, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beard:2003:EMS

[BRCC03] T. Douglas Beard, Jr., Paul W. Rasmussen, Sean Cox, and Stephen R. Carpenter. Evaluation of a management system for a mixed walleye spearing and angling fishery in Northern Wisconsin. North American Journal of Fisheries Management, 23(2):481–491, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bergey:2003:VAC

[BRGO03] Lauren L. Bergey, Roger A. Rulifson, Margie L. Gallagher, and Anthony S. Overton. Variability of Atlantic Coast striped bass egg characteristics. North American Journal of Fisheries Management, 23(2):558–572, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brodziak:2002:SOH

[Bro02] Jon Brodziak. In search of optimal harvest rates for West Coast groundfish. North American Journal of Fisheries Management, 22(2):258–271, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bayne:2002:RMF

[BRS02] David R. Bayne, Eric Reutebuch, and Wendy C. Seesock. Relative motility of fishes in a southeastern reservoir based on tissue polychlorinated biphenyl residues. North American Journal of Fisheries Management, 22(1):122–131, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brewer:2007:NLS

[BRSA07] Shannon K. Brewer, Charles F. Rabeni, Scott P. Sowa, and Gust Annis. Natural landscape and stream segment attributes influencing the distribution and relative abundance of riverine smallmouth bass in Missouri. North American Journal of Fisheries Management, 27(1):326–341, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bryant:2000:EFP

[Bry00] Mason D. Bryant. Estimating fish populations by removal methods with minnow traps in Southeast Alaska streams. North American Journal of Fisheries Management, 20(4): 923–930, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bickford:2000:RIY

[BS00] Shane A. Bickford and John R. Skalski. Reanalysis and interpretation of 25 years of snake–Columbia River juvenile salmonid survival studies. North American Journal of Fisheries Management, 20(1):53–68, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bray:2001:ESV

[BS01] Gregory S. Bray and Harold L. Schramm, Jr. Evaluation of a statewide volunteer angler diary program for use as a fishery assessment tool. North American Journal of Fisheries Management, 21(3):606–615, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buckmeier:2009:CES

[BS09] David L. Buckmeier and J. Warren Schlechte. Capture efficiency and size selectivity of channel catfish and blue catfish sampling gears. North American Journal of Fisheries Management, 29(2):404–416, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buckmeier:2003:SFL

[BSB03] David L. Buckmeier, J. Warren Schlechte, and Robert K. Betsill. Stocking fingerling largemouth bass to alter genetic composition: Efficacy and efficiency of three stocking rates. North American Journal of Fisheries Management, 23(2):523–529, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bumgarner:2009:RHS

[BSB09] Joseph D. Bumgarner, Mark L. Schuck, and H. Lee Blankenship. Returns of hatchery steelhead with different fin clips and coded wire tag lengths. North American Journal of Fisheries Management, 29(4):903–913, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barnes:2009:PHC

[BSD09] Michael E. Barnes, Greg Simpson, and Dan J. Durben. Poststocking harvest of catchable-sized rainbow trout enhanced by dietary supplementation with a fully fermented commercial yeast culture during hatchery rearing. North American Journal of Fisheries Management, 29(5):1287–1295, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Benson:2005:EST

[BSEM05] Angela C. Benson, Trent M. Sutton, Robert F. Elliott, and Thomas G. Meronek. Evaluation of sampling techniques for age-0 lake sturgeon in a Lake Michigan tributary. North American Journal of Fisheries Management, 25(4):1378– 1385, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boxrucker:2005:EED

[BSG05] Jeff C. Boxrucker, Greg L. Summers, and Eugene R. Gilliland. Effects of the extent and duration of seasonal flood pool inundation on recruitment of threadfin shad, white crappies, and largemouth bass in Hugo Reservoir, Oklahoma. North American Journal of Fisheries Management, 25(2): 709–716, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bushon:2007:EVI

[BSR07] Amanda M. Bushon, Gregory W. Stunz, and Megan M. Reese. Evaluation of visible implant elastomer for marking juvenile red drum. North American Journal of Fisheries Management, 27(2):460–464, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bartron:2004:MBE

[BSRS04] Meredith L. Bartron, Dave R. Swank, Edward S. Rutherford, and Kim T. Scribner. Methodological bias in estimates of strain composition and straying of hatchery-produced steelhead in Lake Michigan tributaries. North American Journal of Fisheries Management, 24(4):1288–1299, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buchanan:2006:EES

[BSS06] Rebecca A. Buchanan, John R. Skalski, and Steven G. Smith. Estimating the effects of smolt transportation from different vantage points and management perspectives. North American Journal of Fisheries Management, 26(2):460–472, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brousseau:2004:ATR

[BSSC04] Lorne J. Brousseau, Matthew Sclafani, David R. Smith, and David B. Carter. Acoustic-tracking and radio-tracking of horseshoe crabs to assess spawning behavior and subtidal habitat use in Delaware Bay. North American Journal of Fisheries Management, 24(4):1376–1384, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bronte:2002:RSP

[BSSS02] Charles R. Bronte, Stephen T. Schram, James H. Selgeby, and Bruce L. Swanson. Reestablishing a spawning population of lake trout in Lake Superior with fertilized eggs in artificial turf incubators. North American Journal of Fisheries Management, 22(3):796–805, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blankenship:2003:ELD

[BT03] H. Lee Blankenship and Daniel A. Thompson. The effect of 1.5-length and double-length coded wire tagson Coho salmon survival, growth, homing, and electronic detection. North American Journal of Fisheries Management, 23(1):60–65, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Budy:2002:ELD

[BTB^{+}02] Phaedra Budy, Gary P. Thiede, Nick Bouwes, C. E. Petrosky, and Howard Schaller. Evidence linking delayed mortality of Snake River salmon to their earlier hydrosystem experience. North American Journal of Fisheries Management, 22 (1):35–51, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bulak:2004:GVM

[BTHE04] James S. Bulak, Christopher S. Thomason, Kaiping Han, and Bert Ely. Genetic variation and management of striped bass populations in the coastal rivers of South Carolina. North American Journal of Fisheries Management, 24(4): 1322–1329, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bettinger:2005:HMP

[BTI05] Jason M. Bettinger, Joseph R. Tomasso, Jr., and J. Jeffery Isely. Hooking mortality and physiological responses of striped bass angled in freshwater and held in live-release tubes. North American Journal of Fisheries Management, 25(4):1273–1280, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beland:2004:URR

[BTK04] K. F. Beland, J. G. Trial, and J. F. Kocik. Use of riffle and run habitats with aquatic vegetation by juvenile Atlantic salmon. North American Journal of Fisheries Management, 24(2):525–533, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Budy:2007:QVR

[BTM07] Phaedra Budy, Gary P. Thiede, and Peter McHugh. Quantification of the vital rates, abundance, and status of a critical, endemic population of Bonneville cutthroat trout. North American Journal of Fisheries Management, 27(2):593–604, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Berejikian:2001:RBB

[BTS01] B. A. Berejikian, E. P. Tezak, and S. L. Schroder. Reproductive behavior and breeding success of captively reared Chinook salmon. North American Journal of Fisheries Management, 21(1):255–260, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barker:2002:ECF

[BTS02] Richard J. Barker, Peter H. Taylor, and Stephen Smith. Effect of a change in fishing regulations on the survival and capture probabilities of rainbow trout in the Upper Rangitikei River, New Zealand. North American Journal of Fisheries Management, 22(2):465–473, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:2007:ENC

[BTSB07] Richard S. Brown, Christopher P. Tatara, John R. Stephenson, and Barry A. Berejikian. Evaluation of a new coded electromyogram transmitter for studying swimming behavior and energetics in fish. North American Journal of Fisheries Management, 27(3):765–772, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buckmeier:2001:CWT

[Buc01] David L. Buckmeier. Coded wire tag insertion sites for small fingerling black bass. North American Journal of Fisheries Management, 21(3):696–698, August 2001. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bromaghin:2007:REF

[BUH07] Jeffrey F. Bromaghin, Tevis J. Underwood, and Raymond F. Hander. Residual effects from fish wheel capture and handling of Yukon River fall chum salmon. North American Journal of Fisheries Management, 27(3):860–872, August 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bettoli:2000:HMS

[BVH00] Phillip W. Bettoli, Chris S. Vandergoot, and Paul T. Horner. Hooking mortality of saugers in the Tennessee River. North American Journal of Fisheries Management, 20(3):833–837, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Borley:2006:MDV

[BW06] Kimberly Borley and Matthew M. White. Mitochondrial DNA variation in the endangered Colorado pikeminnow: a comparison among hatchery stocks and historic specimens. North American Journal of Fisheries Management, 26(4): 916–920, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baldigo:2008:DRF

[BW08] B. P. Baldigo and D. R. Warren. Detecting the response of fish assemblages to stream restoration: Effects of different sampling designs. North American Journal of Fisheries Management, 28(3):919–934, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bryant:2009:RSP

[BW09] M. D. Bryant and R. D. Woodsmith. The response of salmon populations to geomorphic measurements at three scales.

North American Journal of Fisheries Management, 29(3): 549–559, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bister:2000:PSW

[BWB^{+}00] Timothy J. Bister, David W. Willis, Michael L. Brown, Stephen M. Jordan, Robert M. Neumann, Michael C. Quist, and Christopher S. Guy. Proposed standard weight (W_{s}) equations and standard length categories for 18 warmwater nongame and riverine fish species. North American Journal of Fisheries Management, 20(2):570–574, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bacheler:2005:MMR

[BWB05] Nathan M. Bacheler, Richard A. Wong, and Jeffrey A. Buckel. Movements and mortality rates of striped mullet in North Carolina. North American Journal of Fisheries Management, 25(1):361–373, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boswell:2008:SAE

[BWC08] Kevin M. Boswell, Matthew P. Wilson, and James H. Cowan, Jr. A semiautomated approach to estimating fish size, abundance, and behavior from dual-frequency identification sonar (DIDSON) data. North American Journal of Fisheries Management, 28(3):799–807, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baldigo:2008:RFP

[BWEM08] Barry P. Baldigo, Dana R. Warren, Anne G. Ernst, and Christiane I. Mulvihill. Response of fish populations to natural channel design restoration in streams of the Catskill Mountains, New York. North American Journal of Fisheries Management, 28(3):954–969, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beacham:2008:ARM

[BWJ^{+}08] Terry D. Beacham, Ivan Winther, Kimberly L. Jonsen, Michael Wetklo, Langtuo Deng, and John R. Candy. The application of rapid microsatellite-based stock identification to management of a Chinook salmon troll fishery off the Queen Charlotte Islands, British Columbia. North American Journal of Fisheries Management, 28(3):849–855, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bister:2002:ECM

[BWKS02] Timothy J. Bister, David W. Willis, Allen D. Knapp, and Todd R. St. Sauver. Evaluation of a 23-cm minimum length limit for black and white crappies in a small South Dakota impoundment. North American Journal of Fisheries Management, 22(4):1364–1368, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brenden:2008:NTA

[BWM08] Travis O. Brenden, Tyler Wagner, and Brian R. Murphy. Novel tools for analyzing proportional size distribution index data. North American Journal of Fisheries Management, 28 (5):1233–1242, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beacham:2008:AMS

[BWW^{+}08] T. D. Beacham, M. Wetklo, C. Wallace, J. B. Olsen, B. G. Flannery, J. K. Wenburg, W. D. Templin, A. Antonovich, and L. W. Seeb. The application of microsatellites for stock identification of Yukon River Chinook salmon. North American Journal of Fisheries Management, 28(1):283–295, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crawford:2006:FNM

[CA06] S. Crawford and M. S. Allen. Fishing and natural mortality of bluegills and redear sunfish at Lake Panasoffkee, Florida: Implications for size limits. North American Journal of Fisheries Management, 26(1):42–51, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Catalano:2006:EVF

[CAM06] Matthew J. Catalano, Micheal S. Allen, and Debra J. Murie. Effects of variable flows on water chemistry gradients and fish communities at the Hillsborough River, Florida. North American Journal of Fisheries Management, 26(1):108–118, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carline:2001:EHF

[Car01] Robert F. Carline. Effects of high-frequency Pulsed-DC electrofishing on a wild brown trout population. North American Journal of Fisheries Management, 21(3):571–579, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Castello:2004:MCP

[Cas04] Leandro Castello. A method to count pirarucu Arapaima gigas: Fishers, assessment, and management. North American Journal of Fisheries Management, 24(2):379–389, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooke:2001:AIE

[CB01] Steven J. Cooke and Christopher M. Bunt. Assessment of internal and external antenna configurations of radio transmitters implanted in smallmouth bass. North American Journal of Fisheries Management, 21(1):236–241, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connor:2003:GWS

[CB03] William P. Connor and Howard L. Burge. Growth of wild subyearling fall Chinook salmon in the Snake River. North American Journal of Fisheries Management, 23(2):594–599, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooke:2004:CJB

[CB04] Steven J. Cooke and Christopher M. Bunt. Construction of a junction box for use with an inexpensive, commercially available underwater video camera suitable for aquatic research. North American Journal of Fisheries Management, 24(1):253–257, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connelly:2000:MMS

[CBK00] Nancy A. Connelly, Tommy L. Brown, and Barbara A. Knuth. A multiple market segmentation of Great Lakes anglers in New York. North American Journal of Fisheries Management, 20(2):399–407, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooke:2000:WRS

[CBM00] Steven J. Cooke, Christopher M. Bunt, and R. Scott McKinley. Winter residency of smallmouth bass in a thermal discharge canal: Implications for tempering pump operation. North American Journal of Fisheries Management, 20(1): 288–295, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Curry:2003:ERF

[CBM03] R. Allen Curry, Charles Brady, and George E. Morgan. Effects of recreational fishing on the population dynamics of lake-dwelling brook trout. North American Journal of Fisheries Management, 23(1):35–47, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Catalano:2007:EDR

[CBP07] Matthew J. Catalano, Michael A. Bozek, and Thomas D. Pellett. Effects of dam removal on fish assemblage structure and spatial distributions in the Baraboo River, Wisconsin. North American Journal of Fisheries Management, 27(2):519–530, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooke:2005:ICH

[CBS^{+}05] Steven J. Cooke, Brandon L. Barthel, Cory D. Suski, Michael J. Siepker, and David P. Philipp. Influence of circle hook size on hooking efficiency, injury, and size selectivity of bluegill with comments on circle hook conservation benefits in recreational fisheries. North American Journal of Fisheries Management, 25(1):211–219, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connor:2002:JLH

[CBWB02] William P. Connor, Howard L. Burge, Russell Waitt, and Theodore C. Bjornn. Juvenile life history of wild fall Chinook salmon in the Snake and Clearwater Rivers. North American Journal of Fisheries Management, 22(3):703–712, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connor:2003:IFT

[CBYB03] William P. Connor, Howard L. Burge, John R. Yearsley, and Theodore C. Bjornn. Influence of flow and temperature on survival of wild subyearling fall Chinook salmon in the Snake River. North American Journal of Fisheries Management, 23 (2):362–375, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Catalano:2001:EIF

[CCBW01] Matthew J. Catalano, Steven R. Chipps, Michelle A. Bouchard, and David H. Wahl. Evaluation of injectable fluorescent tags for marking centrarchid fishes: Retention rate and effects on vulnerability to predation. North American Journal of Fisheries Management, 21(4):911–917, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cho:2009:PSR

[CCG09] Sung Jin Cho, Colleen A. Caldwell, and William R. Gould. Physiological stress responses of Rio Grande silvery minnow: Effects of individual and multiple physical stressors of handling, confinement, and transport. North American Journal of Fisheries Management, 29(6):1698–1706, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chen:2009:GCV

[CCKF^{+}09] Yongjiu Chen, Carole Conway, Connie Keeler-Foster, Roger Hamman, and Steve Meismer. Genetic characterization of variation in captive and wild woundfin. North American Journal of Fisheries Management, 29(4):843–849, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Catalano:2009:EVM

[CDPA09] Matthew J. Catalano, Andrew C. Dutterer, William E. Pine III, and Micheal S. Allen. Effects of variable mortality and recruitment on performance of catch-curve residuals as indicators of fish year-class strength. North American Journal of Fisheries Management, 29(2):295–305, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Close:2003:EII

[CFL^{+}03] David A. Close, Martin S. Fitzpatrick, Christopher M. Lorion, Hiram W. Li, and Carl B. Schreck. Effects of intraperitoneally implanted radio transmitters on the swimming performance and physiology of Pacific lamprey. North American Journal of Fisheries Management, 23(4):1184–1192, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cole:2006:SAS

[CGM^{+}06] John G. Cole, Benny J. Gallaway, Larry R. Martin, James M. Nance, and Michael Longnecker. Spatial allocation of shrimp catch based on fishing effort: Adjusting for the effects of the Texas opening. North American Journal of Fisheries Management, 26(4):789–792, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cunningham:2000:SIS

[CGP00] Mandy K. Cunningham, William F. Granberry, Jr., and Kevin L. Pope. Shrinkage of inland silverside larvae preserved in ethanol and formalin. North American Journal of Fisheries Management, 20(3):816–818, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooke:2000:ERG

[CH00a] Steven J. Cooke and W. James Hogle. Effects of retention gear on the injury and short-term mortality of adult smallmouth bass. North American Journal of Fisheries Management, 20(4):1033–1039, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Covington:2000:RBS

[CH00b] J. Scott Covington and Wayne A. Hubert. Relations between on-site and aerial measurements of streamside features and cover for trout in alluvial valley streams in Wyoming. North American Journal of Fisheries Management, 20(3):627–633, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cole:2001:ETB

[CH01] Randall V. Cole and Thomas E. Helser. Effect of three bycatch reduction devices on diamondback terrapin Malaclemys terrapin capture and blue crab Callinectes sapidus harvest in Delaware Bay. North American Journal of Fisheries Management, 21(4):825–833, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2002:ECS

[CH02] William G. Clark and Steven R. Hare. Effects of climate and stock size on recruitment and growth of Pacific halibut. North American Journal of Fisheries Management, 22(3):852–862, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2004:CCC

[CH04] William G. Clark and Steven R. Hare. A conditional constant catch policy for managing the Pacific halibut fishery. North American Journal of Fisheries Management, 24(1):106–113, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chiotti:2008:LSS

[CHAO08] Justin A. Chiotti, J. Marty Holtgren, Nancy A. Auer, and Stephanie A. Ogren. Lake sturgeon spawning habitat in the Big Manistee River, Michigan. North American Journal of Fisheries Management, 28(4):1009–1019, August 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clements:2002:ETP

[CHCD02] Shaun P. Clements, Brendan J. Hicks, John F. Carragher, and Michel Dedual. The effect of a trapping procedure on the stress response of wild rainbow trout. North American Journal of Fisheries Management, 22(3):907–916, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chen:2003:EEI

[CHD03] R. J. Chen, K. M. Hunt, and R. B. Ditton. Estimating the economic impacts of a trophy largemouth bass fishery: Issues and applications. North American Journal of Fisheries Management, 23(3):835–844, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chen:2004:BBC

[Che04] D. G. Chen. Bias and bias correction in fish recruitment prediction. North American Journal of Fisheries Management, 24(2):724–730, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chih:2009:EOS

[Chi09] Ching-Ping Chih. The effects of otolith sampling methods on the precision of growth curves. North American Journal of Fisheries Management, 29(6):1519–1528, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Copeland:2007:CMU

[CHJ07] Timothy Copeland, Matthew W. Hyatt, and June Johnson. Comparison of methods used to age spring–summer Chinook salmon in Idaho: Validation and simulated effects on estimated age composition. North American Journal of Fisheries Management, 27(4):1393–1401, November 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Compton:2008:IFT

[CHR^{+}08] Robert I. Compton, Wayne A. Hubert, Frank J. Rahel, Michael C. Quist, and Michael R. Bower. Influences of fragmentation on three species of native warmwater fishes in a Colorado River Basin headwater stream system, Wyoming. North American Journal of Fisheries Management, 28(6): 1733–1743, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Corradin:2008:RDL

[CHSS08] Lisa M. Corradin, Michael J. Hansen, Donald R. Schreiner, and Michael J. Seider. Recruitment dynamics of lake trout in Western Lake Superior during 1988–1995. North American Journal of Fisheries Management, 28(3):663–677, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Close:2002:DVI

[CJ02] Tracy L. Close and Thomas S. Jones. Detection of visible implant elastomer in fingerling and yearling rainbow trout.

North American Journal of Fisheries Management, 22(3): 961–964, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clements:2005:OAT

[CJKS05] Shaun Clements, David Jepsen, Mark Karnowski, and Carl B. Schreck. Optimization of an acoustic telemetry array for detecting transmitter-implanted fish. North American Journal of Fisheries Management, 25(2):429–436, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2007:CSS

[CJL07] Sandra J. Clark, John R. Jackson, and Steve E. Lochmann. A comparison of shoreline seines with fyke nets for sampling littoral fish communities in floodplain lakes. North American Journal of Fisheries Management, 27(2):676–680, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Campana:2002:PDP

[CJM^{+}02] Steven E. Campana, Warren Joyce, Linda Marks, Lisa J. Natanson, Nancy E. Kohler, Christopher F. Jensen, Joseph J. Mello, Harold L. Pratt, Jr., and Sigmund Myklevoll. Population dynamics of the porbeagle in the Northwest Atlantic Ocean. North American Journal of Fisheries Management, 22(1):106–121, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connolly:2008:MPT

[CJMP08] Patrick J. Connolly, Ian G. Jezorek, Kyle D. Martens, and Earl F. Prentice. Measuring the performance of two stationary interrogation systems for detecting downstream and upstream movement of PIT-Tagged salmonids. North American Journal of Fisheries Management, 28(2):402–417, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chu:2009:ESM

[CJPB09] Cindy Chu, Nicholas E. Jones, Andrew R. Piggott, and James M. Buttle. Evaluation of a simple method to classify the thermal characteristics of streams using a nomogram of daily maximum air and water temperatures. North American Journal of Fisheries Management, 29(6):1605–1619, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2009:ADS

[CK09] John H. Clark and Desmond M. Kahn. Amount and disposition of striped bass discarded in Delaware’s spring striped bass gill-net fishery during 2002 and 2003: Effects of regulations and fishing strategies. North American Journal of Fisheries Management, 29(3):576–585, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Claramunt:2009:EIC

[CKC^{+}09] Randall M. Claramunt, Tracy L. Kolb, David F. Clapp, Daniel B. Hayes, James L. Dexter, Jr., and David M. Warner. Effects of increasing Chinook salmon bag limits on alewife abundance: Implications for Lake Michigan management goals. North American Journal of Fisheries Management, 29 (3):829–842, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colyer:2005:MFB

[CKH05] Warren T. Colyer, Jeffrey L. Kershner, and Robert H. Hilderbrand. Movements of fluvial Bonneville cutthroat trout in the Thomas Fork of the Bear River, Idaho–Wyoming. North American Journal of Fisheries Management, 25(3):954–963, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carson:2009:GSH

[CKSG09] Evan W. Carson, S. Karlsson, E. Saillant, and John R. Gold. Genetic studies of hatchery-supplemented populations of red drum in four Texas bays. North American Journal of Fisheries Management, 29(5):1502–1510, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooke:2004:IMI

[CL04] Douglas W. Cooke and Steven D. Leach. Implications of a migration impediment on shortnose sturgeon spawning. North American Journal of Fisheries Management, 24(4): 1460–1468, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2002:RTY

[Cla02] William G. Clark. F_{35%} revisited ten years later. North American Journal of Fisheries Management, 22(2):251–257, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clarkson:2004:EEF

[Cla04] Robert W. Clarkson. Effectiveness of electrical fish barriers associated with the Central Arizona Project. North American Journal of Fisheries Management, 24(1):94–105, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clarke:2004:ASE

[CLB04] Lance R. Clarke, Paul S. Letizia, and David H. Bennett. Autumn-to-spring energetic and diet changes among Kokanee from North Idaho lakes with and without Mysis relicta. North American Journal of Fisheries Management, 24(2):597–608, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coghlan:2007:OCD

[CLB^{+}07] Stephen M. Coghlan, Jr., Michael S. Lyerly, Thomas R. Bly, Jeffrey S. Williams, Darrell Bowman, and Robyn Hannigan. Otolith chemistry discriminates among hatchery-reared and tributary-spawned salmonines in a tailwater system. North American Journal of Fisheries Management, 27(2):531–541, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Close:2000:DRP

[Clo00] Tracy L. Close. Detection and retention of postocular visible implant elastomer in fingerling rainbow trout. North American Journal of Fisheries Management, 20(2):542–545, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cross:2005:SHD

[CM05] Timothy K. Cross and Michael C. McInerny. Spatial habitat dynamics affecting bluegill abundance in Minnesota bass– panfish lakes. North American Journal of Fisheries Management, 25(3):1051–1066, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chotkowski:2002:IME

[CME02] Michael A. Chotkowski, J. Ellen Marsden, and Brian J. Ellrott. An inexpensive modified emergent-fry trap for lakespawning salmonids. North American Journal of Fisheries Management, 22(4):1321–1324, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Copeland:2008:IRW

[CMN08] Timothy Copeland, Brian R. Murphy, and John J. Ney. Interpretation of relative weight in three populations of wild bluegills: a cautionary tale. North American Journal of Fisheries Management, 28(2):368–377, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crook:2007:PEF

[COG^{+}07] David A. Crook, Damien O’Mahony, Bronwyn M. Gillanders, Andrew R. Munro, and Andrew C. Sanger. Production of external fluorescent marks on golden perch fingerlings through osmotic induction marking with alizarin red S. North American Journal of Fisheries Management, 27(2):670–675, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colvin:2002:CGN

[Col02a] Michael A. Colvin. A comparison of gill netting and electrofishing as sampling techniques for white bass in Missouri’s large reservoirs. North American Journal of Fisheries Management, 22(2):690–702, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colvin:2002:PFC

[Col02b] Michael A. Colvin. Population and fishery characteristics of white bass in four large Missouri reservoirs. North American Journal of Fisheries Management, 22(2):677–689, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crook:2009:DEM

[COS^{+}09] David A. Crook, Damien J. O’Mahony, Andrew C. Sanger, Andrew R. Munro, Bronwyn M. Gillanders, and Stephen Thurstan. Development and evaluation of methods for osmotic induction marking of golden perch Macquaria ambigua with calcein and alizarin red S. North American Journal of Fisheries Management, 29(2):279–287, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connor:2004:CDD

[CP04] Edward J. Connor and David E. Pflug. Changes in the distribution and density of pink, chum, and Chinook salmon spawning in the Upper Skagit River in response to flow management measures. North American Journal of Fisheries Management, 24(3):835–852, August 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooke:2001:ITT

[CPDS01] Steven J. Cooke, David P. Philipp, Karen M. Dunmall, and Jason F. Schreer. The influence of terminal tackle on injury, handling time, and cardiac disturbance of rock bass. North American Journal of Fisheries Management, 21(2):333–342, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cieciel:2009:TRE

[CPE09] Kristin Cieciel, Brian J. Pyper, and Ginny L. Eckert. Tag retention and effects of tagging on movement of the giant red sea cucumber Parastichopus californicus. North American Journal of Fisheries Management, 29(2):288–294, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colombo:2008:GSP

[CPG^{+}08] Robert E. Colombo, Quinton E. Phelps, James E. Garvey, Roy C. Heidinger, and Tom Stefanavage. Gear-specific population demographics of channel catfish in a large Midwestern river. North American Journal of Fisheries Management, 28 (1):241–246, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooke:2000:LIN

[CPSM00] Steven J. Cooke, David P. Philipp, Jason F. Schreer, and R. Scott McKinley. Locomotory impairment of nesting male largemouth bass following catch-and-release angling. North American Journal of Fisheries Management, 20(4):968–977, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chizinski:2005:EVA

[CPW^{+}05] Christopher J. Chizinski, Kevin L. Pope, David B. Willis, Gene R. Wilde, and Edwin J. Rossman. Economic value of angling at a reservoir with low visitation. North American Journal of Fisheries Management, 25(1):98–104, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coggins:2006:ATS

[CPW^{+}06] Lewis G. Coggins, Jr., William E. Pine III, Carl J. Walters, David R. Van Haverbeke, David Ward, and Helene C. Johnstone. Abundance trends and status of the Little Colorado River population of humpback chub. North American Journal of Fisheries Management, 26(1):233–245, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coggins:2006:ASM

[CPWM06] Lewis G. Coggins, Jr., William E. Pine III, Carl J. Walters, and Steven J. D. Martell. Age-structured mark–recapture analysis: a virtual-population-analysis-based model for analyzing age-structured capture–recapture data. North American Journal of Fisheries Management, 26(1):201–205, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carlson:2005:EEI

[CQ05] Leah D. Carlson and Michael S. Quinn. Evaluating the effectiveness of instream habitat structures for overwintering stream salmonids: a test of underwater video. North American Journal of Fisheries Management, 25(1):130–137, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coghlan:2004:CAS

[CR04] Stephen M. Coghlan, Jr. and Neil H. Ringler. A comparison of Atlantic salmon embryo and fry stocking in the Salmon River, New York. North American Journal of Fisheries Management, 24(4):1385–1397, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Creque:2006:IYB

[CRBD06] Sara M. Creque, Matthew J. Raffenberg, Wayne A. Brofka, and John M. Dettmers. If you build it, will they come? Fish and angler use at a freshwater artificial reef. North American Journal of Fisheries Management, 26(3):702–713, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2008:WLF

[CRC^{+}08] Mark E. Clark, Kenneth A. Rose, James A. Chandler, Tracy J. Richter, Donald J. Orth, and Webster Van Winkle. Water-level fluctuation effects on centrarchid reproductive success in reservoirs: a modeling analysis. North American Journal of Fisheries Management, 28(4):1138–1156, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Czesny:2005:IVE

[CRD05] Sergiusz Czesny, Jacques Rinchard, and Konrad Dabrowski. Intrapopulation variation in egg lipid and fatty acid composition and embryo viability in a naturally spawning walleye population from an inland reservoir. North American Journal of Fisheries Management, 25(1):122–129, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cucherousset:2005:UTN

[CRK^{+}05] Julien Cucherousset, Jean-Marc Roussel, Rachel Keeler, Richard A. Cunjak, and Roland Stump. The use of two new portable 12-mm PIT tag detectors to track small fish in shallow streams. North American Journal of Fisheries Management, 25(1):270–274, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colvin:2008:SCG

[CRL08] N. Elizabeth Colvin, Christopher L. Racey, and Steve E. Lochmann. Stocking contribution and growth of largemouth bass stocked at 50 and 100 mm into backwaters of the Arkansas River. North American Journal of Fisheries Management, 28(2):434–441, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Creque:2005:UGD

[CRZ05] Sara M. Creque, Edward S. Rutherford, and Troy G. Zorn. Use of GIS-Derived landscape-scale habitat features to explain spatial patterns of fish density in Michigan rivers. North American Journal of Fisheries Management, 25(4): 1411–1425, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connor:2004:PPH

[CSA^{+}04] William P. Connor, Steven G. Smith, Todd Andersen, Steven M. Bradbury, Douglas C. Burum, Eric E. Hockersmith, Mark L. Schuck, Glen W. Mendel, and Robert M. Bugert. Postrelease performance of hatchery yearling and subyearling fall Chinook salmon released into the Snake River. North American Journal of Fisheries Management, 24(2):545–560, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connor:2000:FSP

[CSB00] William P. Connor, R. Kirk Steinhorst, and Howard L. Burge. Forecasting survival and passage of migratory juvenile salmonids. North American Journal of Fisheries Management, 20(3):651–660, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connor:2003:MBS

[CSB03a] William P. Connor, R. Kirk Steinhorst, and Howard L. Burge. Migrational behavior and seaward movement of wild subyearling fall Chinook salmon in the Snake River. North American Journal of Fisheries Management, 23(2):414–430, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooke:2003:IMI

[CSB^{+}03b] Steven J. Cooke, Cory D. Suski, Brandon L. Barthel, Kenneth G. Ostrand, Bruce L. Tufts, and David P. Philipp. Injury and mortality induced by four hook types on bluegill and pumpkinseed. North American Journal of Fisheries Management, 23(4):883–893, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cada:2005:UPS

[ČSB05] Glenn Čada, John Smith, and Jessica Busey. Use of pressuresensitive film to quantify sources of injury to fish. North American Journal of Fisheries Management, 25(1):57–66, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cairns:2004:MPA

[CSI^{+}04] D. K. Cairns, J. C. Shiao, Y. Iizuka, W-N. Tzeng, and C. D. MacPherson. Movement patterns of American eels in an impounded watercourse, as indicated by otolith microchemistry. North American Journal of Fisheries Management, 24(2): 452–458, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Claramunt:2005:CPN

[CSW05] Randall M. Claramunt, Daniel E. Shoup, and David H. Wahl. Comparison of push nets and tow nets for sampling larval fish with implications for assessing littoral habitat utilization. North American Journal of Fisheries Management, 25 (1):86–92, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cowen:2007:EAC

[CTB07] Laura Cowen, Nicole Trouton, and Richard E. Bailey. Effects of angling on Chinook salmon for the Nicola River, British Columbia, 1996–2002. North American Journal of Fisheries Management, 27(1):256–267, February 2007. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cade:2008:EFB

[CTP08] Brian S. Cade, James W. Terrell, and Mark T. Porath. Estimating fish body condition with quantile regression. North American Journal of Fisheries Management, 28(2):349–359, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cunningham:2004:ECB

[Cun04] Kenneth K. Cunningham. Efficacy of a chase boat for electrofishing flathead catfish in three Oklahoma reservoirs. North American Journal of Fisheries Management, 24(4):

1427–1430, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Curtis:2006:VIE

[Cur06] Janelle M. R. Curtis. Visible implant elastomer color determination, tag visibility, and tag loss: Potential sources of error for mark–recapture studies. North American Journal of Fisheries Management, 26(2):327–337, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cox:2002:MER

[CW02] Sean P. Cox and Carl Walters. Modeling exploitation in recreational fisheries and implications for effort management on British Columbia rainbow trout lakes. North American Journal of Fisheries Management, 22(1):21–34, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cowley:2003:OAH

[CWDH03] David E. Cowley, Frank A. Ward, Robert Deitner, and Michael D. Hatch. Optimizing the allocation of hatcheryproduced fish among multiple stocking sites. North American Journal of Fisheries Management, 23(2):216–229, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colombo:2004:UUI

[CWG04] Robert E. Colombo, Paul S. Wills, and James E. Garvey. Use of ultrasound imaging to determine sex of shovelnose sturgeon. North American Journal of Fisheries Management, 24(1):322–326, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chesnes:2009:PLC

[CWK09] Thomas C. Chesnes, Raymond E. Waldner, and Cecilia S. Krahforst. Postmortem length changes in six Florida fish species stored on ice. North American Journal of Fisheries Management, 29(5):1242–1252, October 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cox:2003:MBE

[CWP03] Sean P. Cox, Carl J. Walters, and John R. Post. A modelbased evaluation of active management of recreational fishing effort. North American Journal of Fisheries Management, 23 (4):1294–1302, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Donald:2005:SYM

[DA05] David B. Donald and William Aitken. Stock–yield model for a fish with variable annual recruitment. North American Journal of Fisheries Management, 25(4):1226–1238, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dotson:2009:ICG

[DAJB09] Jason R. Dotson, Micheal S. Allen, William E. Johnson, and John Benton. Impacts of commercial gill-net bycatch and recreational fishing on a Florida black crappie population. North American Journal of Fisheries Management, 29(5): 1454–1465, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dare:2003:MLT

[Dar03] Matthew R. Dare. Mortality and long-term retention of passive integrated transponder tags by spring Chinook salmon. North American Journal of Fisheries Management, 23(4): 1015–1019, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daum:2005:MFW

[Dau05] David W. Daum. Monitoring fish wheel catch using eventtriggered video technology. North American Journal of Fisheries Management, 25(1):322–328, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Desorcie:2003:EOL

[DB03] Timothy J. Desorcie and Charles A. Bowen II. Evidence of offshore lake trout reproduction in Lake Huron. North American Journal of Fisheries Management, 23(4):1253– 1256, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DeCicco:2006:DVA

[DB06] Alfred L. DeCicco and Randy J. Brown. Direct validation of annual growth increments on sectioned otoliths from adult Arctic grayling and a comparison of otolith and scale ages. North American Journal of Fisheries Management, 26(3): 580–586, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2009:RPI

[DB09a] Daniel J. Daugherty and David L. Buckmeier. Retention of passive integrated transponder tags in flathead catfish. North American Journal of Fisheries Management, 29(2):343–345, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Deroba:2009:DMB

[DB09b] Jonathan J. Deroba and James R. Bence. Developing modelbased indices of lake whitefish abundance using commercial fishery catch and effort data in Lakes Huron, Michigan, and Superior. North American Journal of Fisheries Management, 29(1):50–63, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dieterman:2000:MPH

[DBG00] Douglas J. Dieterman, Michael S. Baird, and David L. Galat. Mortality of paddlefish in hoop nets in the Lower Missouri River, Missouri. North American Journal of Fisheries Management, 20(1):226–230, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DuPont:2007:UAM

[DBG07] Joseph M. DuPont, Richard S. Brown, and David R. Geist. Unique allacustrine migration patterns of a bull trout population in the Pend Oreille River Drainage, Idaho. North American Journal of Fisheries Management, 27(4):1268– 1275, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Didenko:2004:SWE

[DBM04] Alexander V. Didenko, Scott A. Bonar, and William J. Matter. Standard weight (W_{s}) equations for four rare desert fishes. North American Journal of Fisheries Management, 24(2):697–703, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dietrich:2006:EIC

[DC06] Jason P. Dietrich and Richard A. Cunjak. Evaluation of the impacts of carlin tags, fin clips, and panjet tattoos on juvenile Atlantic salmon. North American Journal of Fisheries Management, 26(1):163–169, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Danner:2009:VIS

[DCB09] G. Russell Danner, Jim Chacko, and Francis Brautigam. Voluntary ingestion of soft plastic fishing lures affects brook trout growth in the laboratory. North American Journal of Fisheries Management, 29(2):352–360, April 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunmall:2001:ESL

[DCSM01] Karen M. Dunmall, Steven J. Cooke, Jason F. Schreer, and R. Scott McKinley. The effect of scented lures on the hooking injury and mortality of smallmouth bass caught by novice and experienced anglers. North American Journal of Fisheries Management, 21(1):242–248, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DiCenzo:2002:IRI

[DD02] Victor J. DiCenzo and Michael C. Duval. Importance of reservoir inflow in determining white bass year-class strength in three Virginia reservoirs. North American Journal of Fisheries Management, 22(2):620–626, May 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DuBois:2004:EHTa

[DD04] Robert B. DuBois and Richard R. Dubielzig. Effect of hook type on mortality, trauma, and capture efficiency of wild stream trout caught by angling with spinners. North American Journal of Fisheries Management, 24(2):609–616, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dedual:2007:SJR

[Ded07] Michel Dedual. Survival of juvenile rainbow trout passing through a Francis turbine. North American Journal of Fisheries Management, 27(1):181–186, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DeVries:2007:CCV

[DeV07] Paul DeVries. Comment: a comparison of visual and measurement-based techniques for quantifying cobble embeddedness and fine-sediment levels in salmonid-bearing streams. North American Journal of Fisheries Management, 27(3): 1025–1026, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dew:2008:RKC

[Dew08] C. Braxton Dew. Red king crab mating success, sex ratio, spatial distribution, and abundance estimates as artifacts of survey timing in Bristol Bay, Alaska. North American Journal of Fisheries Management, 28(5):1618–1637, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dew:2009:RKC

[Dew09] C. Braxton Dew. Red king crab mating success, sex ratio, spatial distribution, and abundance estimates as artifacts of survey timing in Bristol Bay, Alaska: Please make the following corrections in a recent issue of this journal. North American Journal of Fisheries Management, 29(2):527–528, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dauwalter:2007:ECP

[DF07] Daniel C. Dauwalter and William L. Fisher. Electrofishing capture probability of smallmouth bass in streams. North American Journal of Fisheries Management, 27(1):162–171, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DeLestang:2008:ETD

[DGAG08] Paul De Lestang, Roland Griffin, Quentin Allsop, and Blair S. Grace. Effects of two different landing nets on injuries to the barramundi Lates calcarifer, an iconic Australian sport fish. North American Journal of Fisheries Management, 28 (6):1911–1915, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dieterman:2009:IEP

[DH09] Douglas J. Dieterman and R. John H. Hoxmeier. Instream evaluation of passive integrated transponder retention in brook trout and brown trout: Effects of season, anatomical placement, and fish length. North American Journal of Fisheries Management, 29(1):109–115, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dare:2002:CHA

[DHG02] Matthew R. Dare, Wayne A. Hubert, and Kenneth G. Gerow. Changes in habitat availability and habitat use and movements by two trout species in response to declining discharge in a regulated river during winter. North American Journal of Fisheries Management, 22(3):917–928, August 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dauble:2003:ICR

[DHGP03] D. D. Dauble, T. P. Hanrahan, D. R. Geist, and M. J. Parsley. Impacts of the Columbia River hydroelectric system on main-stem habitats of fall Chinook salmon. North American Journal of Fisheries Management, 23(3):641–659, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dare:2001:ISF

[DHM01] Matthew R. Dare, Wayne A. Hubert, and Joseph S. Meyer. Influence of stream flow on hydrogen sulfide concentrations and distributions of two trout species in a rocky mountains tailwater. North American Journal of Fisheries Management, 21(4):971–975, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Downs:2006:SDJ

[DHMHJ06] Christopher C. Downs, Dona Horan, Erin Morgan-Harris, and Robert Jakubowski. Spawning demographics and juvenile dispersal of an adfluvial bull trout population in Trestle Creek, Idaho. North American Journal of Fisheries Management, 26(1):190–200, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Deroba:2005:EAM

[DHNH05] Jonathan J. Deroba, Michael J. Hansen, Nancy A. Nate, and Joseph M. Hennessy. Evaluating assumptions of mark– recapture studies for estimating angling exploitation of walleyes in Northern Wisconsin lakes. North American Journal of Fisheries Management, 25(3):890–896, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Deroba:2007:ECS

[DHNH07a] Jonathan J. Deroba, Michael J. Hansen, Nancy A. Nate, and Joseph M. Hennessy. Evaluating creel survey efficiency for estimating walleye fishery metrics in Northern Wisconsin lakes. North American Journal of Fisheries Management, 27(2): 707–716, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Deroba:2007:TPW

[DHNH07b] Jonathan J. Deroba, Michael J. Hansen, Nancy A. Nate, and Joseph M. Hennessy. Temporal profiles of walleye angling effort, harvest rate, and harvest in Northern Wisconsin lakes. North American Journal of Fisheries Management, 27(2): 717–727, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Diaz:2008:ECH

[Dia08] Guillermo A. Diaz. The effect of circle hooks and straight (j) hooks on the catch rates and numbers of white marlin and blue marlin released alive by the U.S. pelagic longline fleet in the Gulf of Mexico. North American Journal of Fisheries Management, 28(2):500–506, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Duncan:2004:ESS

[DIC04] Michelle S. Duncan, J. Jeffery Isely, and Douglas W. Cooke. Evaluation of shortnose sturgeon spawning in the Pinopolis Dam tailrace, South Carolina. North American Journal of Fisheries Management, 24(3):932–938, August 2004. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Diana:2006:CMS

[DJC^{+}06] Christine M. Diana, Jory L. Jonas, Randall M. Claramunt, John D. Fitzsimons, and J. Ellen Marsden. A comparison of methods for sampling round goby in rocky littoral areas. North American Journal of Fisheries Management, 26(3): 514–522, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dambacher:2009:LLS

[DJL09] Jeffrey M. Dambacher, Kim K. Jones, and David P. Larsen. Landscape-level sampling for status review of Great Basin redband trout. North American Journal of Fisheries Management, 29(4):1091–1105, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dawson:2009:AAD

[DJSG09] Heather A. Dawson, Michael L. Jones, Kim T. Scribner, and Stacy A. Gilmore. An assessment of age determination methods for Great Lakes larval sea lampreys. North American Journal of Fisheries Management, 29(4):914–927, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DuBois:2004:EHTb

[DK04] Robert B. DuBois and Kurt E. Kuklinski. Effect of hook type on mortality, trauma, and capture efficiency of wild, streamresident trout caught by active baitfishing. North American Journal of Fisheries Management, 24(2):617–623, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Davies:2007:RSS

[DKM07] Trevor D. Davies, Daniel G. Kehler, and Ken R. Meade. Retrospective sampling strategies using video recordings to estimate fish passage at fishways. North American Journal of Fisheries Management, 27(3):992–1003, August 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dolan:2004:IMW

[DM04] C. R. Dolan and L. E. Miranda. Injury and mortality of warmwater fishes immobilized by electrofishing. North American Journal of Fisheries Management, 24(1):118–127, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dinger:2007:EHL

[DM07] Eric C. Dinger and J. C. Marks. Effects of high levels of antimycin a on aquatic invertebrates in a warmwater Arizona stream. North American Journal of Fisheries Management, 27(4):1243–1256, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DeJesus:2009:UVA

[DMB09] Marcos J. De Jesus, Stephan J. Magnelia, and C. Craig Bonds. Use of a volunteer angler survey for assessing length distribution and seasonal catch trends of trophy largemouth bass. North American Journal of Fisheries Management, 29 (5):1225–1231, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dolan:2002:ECE

[DMH02] Chad R. Dolan, L. E. Miranda, and T. B. Henry. Electrofishing for crappies: Electrical settings influence immobilization efficiency, injury, and mortality. North American Journal of Fisheries Management, 22(4):1442–1451, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dorr:2002:SBE

[DMM02] Brian Dorr, Ian A. Munn, and Keith O. Meals. A socioeconomic and biological evaluation of current and hypothetical crappie regulations in Sardis Lake, Mississippi: An integrated approach. North American Journal of Fisheries Management, 22(4):1376–1384, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DAmelio:2008:TCB

[DMMW08] Silvia D’Amelio, Jamie Mucha, Rob Mackereth, and Chris C. Wilson. Tracking coaster brook trout to their sources: Combining telemetry and genetic profiles to determine source populations. North American Journal of Fisheries Management, 28(5):1343–1349, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Danley:2002:SPP

[DMYC02] Melody L. Danley, Shawn D. Mayr, Patiencia S. Young, and Joseph J. Cech, Jr. Swimming performance and physiological stress responses of splittail exposed to a fish screen. North American Journal of Fisheries Management, 22(4): 1241–1249, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Davis:2001:SDM

[DO01] Michael W. Davis and Bori L. Olla. Stress and delayed mortality induced in Pacific halibut by exposure to hooking, net towing, elevated seawater temperature and air: Implications for management of bycatch. North American Journal of Fisheries Management, 21(4):725–732, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Davis:2002:MLT

[DO02] Michael W. Davis and Bori L. Olla. Mortality of lingcod towed in a net as related to FishLength, seawater temperature, and air exposure: ALaboratory bycatch study. North American Journal of Fisheries Management, 22(4): 1095–1104, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dorn:2002:AWC

[Dor02] Martin W. Dorn. Advice on West Coast rockfish harvest rates from Bayesian meta-analysis of stock-recruit relationships. North American Journal of Fisheries Management, 22 (2):280–300, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Drake:2002:DFB

[DP02] Melissa T. Drake and Donald L. Pereira. Development of a fish-based index of biotic integrity for small inland lakes in Central Minnesota. North American Journal of Fisheries Management, 22(4):1105–1123, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dauwalter:2003:EEE

[DP03a] Daniel C. Dauwalter and Edmund J. Pert. Effect of electrofishing effort on an index of biotic integrity. North American Journal of Fisheries Management, 23(4):1247–1252, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dauwalter:2003:EEF

[DP03b] Daniel C. Dauwalter and Edmund J. Pert. Electrofishing effort and fish species richness and relative abundance in Ozark highland streams of Arkansas. North American Journal of Fisheries Management, 23(4):1152–1166, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Davis:2004:FSE

[DP04] Michael W. Davis and Steven J. Parker. Fish size and exposure to air: Potential effects on behavioral impairment and mortality rates in discarded sablefish. North American Journal of Fisheries Management, 24(2):518–524, May 2004. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DuBois:2007:HSM

[DP07] Robert B. DuBois and Julie M. Pleski. Hook shedding and mortality of deeply hooked brook trout caught with bait on barbed and barbless hooks. North American Journal of Fisheries Management, 27(4):1203–1207, November 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Durham:2005:FAR

[DPW05] Bart W. Durham, Kevin L. Pope, and Gene R. Wilde. Factors affecting regional variation in growth of channel catfish. North American Journal of Fisheries Management, 25(1): 93–97, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Drauch:2007:GEL

[DR07] Andrea M. Drauch and Olin E. Rhodes, Jr. Genetic evaluation of the lake sturgeon reintroduction program in the Mississippi and Missouri Rivers. North American Journal of Fisheries Management, 27(2):434–442, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Drake:2007:ESE

[Dra07] Melissa T. Drake. Estimating sampling effort for biomonitoring of nearshore fish communities in small Central Minnesota lakes. North American Journal of Fisheries Management, 27 (4):1094–1111, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunham:2003:ITE

[DRC03] Jason Dunham, Bruce Rieman, and Gwynne Chandler. Influences of temperature and environmental variables on the distribution of bull trout within streams at the southern margin of its range. North American Journal of Fisheries Management, 23(4):894–904, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunham:2001:SMS

[DRD01] Jason Dunham, Bruce Rieman, and Kevin Davis. Sources and magnitude of sampling error in redd counts for bull trout. North American Journal of Fisheries Management, 21(2): 343–352, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dumont:2004:URE

[DS04] Spencer C. Dumont and Warren Schlechte. Use of resampling to evaluate a simple random sampling design for general monitoring of fishes in Texas reservoirs. North American Journal of Fisheries Management, 24(2):408–416, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2005:PAS

[DS05a] Daniel J. Daugherty and Trent M. Sutton. Population abundance and stock characteristics of flathead catfish in the Lower St. Joseph River, Michigan. North American Journal of Fisheries Management, 25(3):1191–1201, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2005:SMP

[DS05b] Daniel J. Daugherty and Trent M. Sutton. Seasonal movement patterns, habitat use, and home range of flathead catfish in the Lower St. Joseph River, Michigan. North American Journal of Fisheries Management, 25(1):256–269, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2005:UCB

[DS05c] Daniel J. Daugherty and Trent M. Sutton. Use of a chase boat for increasing electrofishing efficiency for flathead catfish in lotic systems. North American Journal of Fisheries Management, 25(4):1528–1532, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dexter:2007:DCI

[DS07] Daniel J. Dexter and Donald V. Schliep. Design of a compound inclined screen trap for anadromous salmonid smolts. North American Journal of Fisheries Management, 27(3): 885–890, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Denson:2008:UOM

[DS08] Michael R. Denson and Theodore I. J. Smith. Use of oxytetracycline to mark larval red drum: Survival and mark persistence. North American Journal of Fisheries Management, 28(6):1779–1789, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

deSilva:2001:PSB

[dSCT01] Janaka A. de Silva, Richard E. Condrey, and Bruce A. Thompson. Profile of shark bycatch in the U.S. Gulf of Mexico menhaden fishery. North American Journal of Fisheries Management, 21(1):111–124, February 2001. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Driscoll:2007:ITL

[DSM07] M. T. Driscoll, J. L. Smith, and R. A. Myers. Impact of tournaments on the largemouth bass population at Sam Rayburn Reservoir, Texas. North American Journal of Fisheries Management, 27(2):425–433, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunham:2003:IMW

[DSR03] Jason Dunham, Robert Schroeter, and Bruce Rieman. Influence of maximum water temperature on occurrence of Lahontan cutthroat trout within streams. North American Journal of Fisheries Management, 23(4):1042–1049, August 2003.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dillon:2000:RRC

[DST00] Jeff C. Dillon, Daniel J. Schill, and David M. Teuscher. Relative return to creel of triploid and diploid rainbow trout stocked in eighteen Idaho streams. North American Journal of Fisheries Management, 20(1):1–9, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dwyer:2001:EBE

[DSW01] William P. Dwyer, Bradley B. Shepard, and Robert G. White. Effect of backpack electroshock on westslope cutthroat trout injury and growth 110 and 250 days posttreatment. North American Journal of Fisheries Management, 21(3):646–650, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dietrich:2006:ARB

[DTMS06] J. P. Dietrich, A. C. Taraborelli, B. J. Morrison, and T. Schaner. Allometric relationships between size of calcified structures and round goby total length. North American Journal of Fisheries Management, 26(4):926–931, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Donaghy:2000:NDI

[DV00] Michael J. Donaghy and Eric Verspoor. A new design of instream incubator for planting out and monitoring Atlantic salmon eggs. North American Journal of Fisheries Management, 20(2):521–527, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Drake:2005:VAF

[DV05] Melissa T. Drake and Ray D. Valley. Validation and application of a fish-based index of biotic integrity for small Central Minnesota lakes. North American Journal of Fisheries Management, 25(3):1095–1111, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Durham:2008:VDG

[DW08] Bart W. Durham and Gene R. Wilde. Validation of daily growth increment formation in the otoliths of juvenile cyprinid fishes from the Brazos River, Texas. North American Journal of Fisheries Management, 28(2):442–446, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Diana:2009:GSF

[DW09] Matthew J. Diana and David H. Wahl. Growth and survival of four sizes of stocked largemouth bass. North American Journal of Fisheries Management, 29(6):1653–1663, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edsall:2000:OTG

[EC00] Thomas A. Edsall and Joshua Cleland. Optimum temperature for growth and preferred temperatures of age-0 lake trout. North American Journal of Fisheries Management, 20 (3):804–809, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Evans:2004:UUI

[EFS04] Allen F. Evans, Martin S. Fitzpatrick, and Lisbeth K. Siddens. Use of ultrasound imaging and steroid concentrations to identify maturational status in adult steelhead. North American Journal of Fisheries Management, 24(3):967–978, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eggold:2001:CTM

[EH01] Bradley T. Eggold and William H. Horns. A comparison of two methods of rearing and stocking Coho salmon in Wisconsin’s waters of Lake Michigan. North American Journal of Fisheries Management, 21(1):147–155, February 2001. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edmonds:2006:ISC

[EM06] Robert L. Edmonds and Kerri Mikkelsen. Influence of salmon carcass placement in red Alder riparian areas on stream chemistry in lowland Western Washington. North American Journal of Fisheries Management, 26(3):551–558, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edwards:2004:FRM

[ENJO04a] Gordon P. Edwards, Jr., Robert M. Neumann, Robert P. Jacobs, and Eileen B. O’Donnell. Factors related to mortality of black bass caught during small club tournaments in Connecticut. North American Journal of Fisheries Management, 24(3):801–810, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edwards:2004:ISC

[ENJO04b] Gordon P. Edwards, Jr., Robert M. Neumann, Robert P. Jacobs, and Eileen B. O’Donnell. Impacts of small club tournaments on black bass populations in Connecticut and the effects of regulation exemptions. North American Journal of Fisheries Management, 24(3):811–821, August 2004. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Esmaeili:2003:IRO

[EO03] Abdoulkarim Esmaeili and Ishak Omar. Influence of rainfall on optimal spawner catch for the shrimp fishery in Iran. North American Journal of Fisheries Management, 23(2):385–391, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Enders:2009:HCA

[ERO^{+}09] Eva C. Enders, Mathieu L. Roy, Michaël Ovidio, Éric J. Hallot, Claudine Boyer, François Petit, and André G. Roy. Habitat choice by Atlantic salmon parr in relation to turbulence at a reach scale. North American Journal of Fisheries Management, 29(6):1819–1830, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Everett:2003:CAG

[ESPW03] Scott R. Everett, Dennis L. Scarnecchia, Greg J. Power, and Christopher J. Williams. Comparison of age and growth of shovelnose sturgeon in the Missouri and Yellowstone Rivers. North American Journal of Fisheries Management, 23(2): 230–240, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ellis:2009:ESW

[EV09] David Ellis and Jason C. Vokoun. Earlier spring warming of coastal streams and implications for alewife migration timing. North American Journal of Fisheries Management, 29 (6):1584–1589, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Edison:2006:AOP

[EWD^{+}06] Timothy W. Edison, David H. Wahl, Matthew J. Diana, David P. Philipp, and Douglas J. Austen. Angler opinion of potential bluegill regulations on Illinois lakes: Effects of angler demographics and bluegill population size structure. North American Journal of Fisheries Management, 26(4): 800–811, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Evans:2008:TSK

[EWK^{+}08] Allen F. Evans, Robert H. Wertheimer, Matthew L. Keefer, Charles T. Boggs, Christopher A. Peery, and Ken Collis. Transportation of steelhead kelts to increase iteroparity in the Columbia and Snake Rivers. North American Journal of Fisheries Management, 28(6):1818–1827, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ely:2008:PSR

[EYI08] Patrick C. Ely, Shawn P. Young, and J. Jeffery Isely. Population size and relative abundance of adult Alabama shad reaching Jim Woodruff Lock and Dam, Apalachicola River, Florida. North American Journal of Fisheries Management, 28(3):827–831, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fayram:2003:CRV

[Fay03] Andrew H. Fayram. A comparison of regulatory and voluntary release of muskellunge and walleyes in Northern Wisconsin. North American Journal of Fisheries Management, 23 (2):619–624, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fox:2007:RGR

[FB07] Martin Fox and Susan Bolton. A regional and geomorphic reference for quantities and volumes of instream wood in unmanaged forested basins of Washington State. North American Journal of Fisheries Management, 27(1):342–359, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Forrest:2009:EEG

[FCM^{+}09] Keith W. Forrest, Jim D. Cave, Catherine G. J. Michielsens, Martin Haulena, and David V. Smith. Evaluation of an electric gradient to deter seal predation on salmon caught in gillnet test fisheries. North American Journal of Fisheries Management, 29(4):885–894, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fedler:2001:DDS

[FD01] Anthony J. Fedler and Robert B. Ditton. Dropping out and dropping in: a study of factors for changing recreational fishing participation. North American Journal of Fisheries Management, 21(2):283–292, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fransen:2006:LRM

[FDM^{+}06] Brian R. Fransen, Steven D. Duke, L. Guy McWethy, Jason K. Walter, and Robert E. Bilby. A logistic regression model for predicting the upstream extent of fish occurrence based on geographical information systems data. North American Journal of Fisheries Management, 26(4):960–975, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fu:2004:SCM

[FF04] Caihong Fu and L. Paul Fanning. Spatial considerations in the management of Atlantic cod off Nova Scotia, Canada. North American Journal of Fisheries Management, 24(3): 775–784, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Filcek:2005:DLS

[FGSJ05] Kristine B. Filcek, Stacy A. Gilmore, Kim T. Scribner, and Michael L. Jones. Discriminating lamprey species using multilocus microsatellite genotypes. North American Journal of Fisheries Management, 25(2):502–509, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Freund:2002:IDD

[FH02] Jason G. Freund and Kyle J. Hartman. Influence of depth on detection distance of low-frequency radio transmitters in the Ohio River. North American Journal of Fisheries Management, 22(4):1301–1305, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fayram:2005:IBW

[FHE05] Andrew H. Fayram, Michael J. Hansen, and Timothy J. Ehlinger. Interactions between walleyes and four fish species with implications for walleye stocking. North American Journal of Fisheries Management, 25(4):1321–1330, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fayram:2001:EIM

[FHG^{+}01] Andrew H. Fayram, Steven W. Hewett, Stephen J. Gilbert, Scott D. Plaster, and T. Douglas Beard, Jr. Evaluation of a 15-inch minimum length limitfor walleye angling in Northern Wisconsin. North American Journal of Fisheries Management, 21(4):816–824, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fayram:2005:DOS

[FHN05] Andrew H. Fayram, Michael J. Hansen, and Nancy A. Nate. Determining optimal stocking rates using a stock– recruitment model: An example using walleye in Northern Wisconsin. North American Journal of Fisheries Management, 25(4):1215–1225, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flamarique:2006:VPO

[FHS06] Iñigo Novales Flamarique, Steve Hiebert, and Juddson Sechrist. Visual performance and ocular system structure of Kokanee and sockeye salmon following strobe light exposure. North American Journal of Fisheries Management, 26 (2):453–459, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fuller:2008:PCA

[FHSM08] S. Adam Fuller, James P. Henne, John Seals, and Vincent A. Mudrak. Performance of commercially available passive integrated transponder (PIT) tag systems used for fish identification and interjurisdictional fisheries management. North American Journal of Fisheries Management, 28(2):386–393, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ferrara:2001:SPI

[FI01] Allyse M. Ferrara and Elise R. Irwin. A standardized procedure for internal sex identification in lepisosteidae. North American Journal of Fisheries Management, 21(4):956–961, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fielder:2000:LHS

[Fie00] David G. Fielder. Lake herring spawning grounds of the St. Marys River with some potential implications for early spring navigation. North American Journal of Fisheries Management, 20(2):552–561, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fielder:2002:SWR

[Fie02] David G. Fielder. Sources of walleye recruitment in Saginaw Bay, Lake Huron. North American Journal of Fisheries Management, 22(3):1032–1040, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fieberg:2004:RPU

[Fie04] John Fieberg. Role of parameter uncertainty in assessing harvest strategies. North American Journal of Fisheries Management, 24(2):459–474, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Firman:2002:CSR

[FJ02] Julie C. Firman and Steven E. Jacobs. Comparison of stream reach lengths measured in the field and from maps. North American Journal of Fisheries Management, 22(4): 1325–1328, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fayram:2008:RBR

[FM08a] Andrew H. Fayram and Matthew G. Mitro. Relationships between reach-scale habitat variables and biotic integrity score, brook trout density, and brown trout density in Wisconsin streams. North American Journal of Fisheries Management, 28(5):1601–1608, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fowler:2008:NSS

[FM08b] Andy L. Fowler and Joseph E. Morris. A new stomach sample collection device for fish gastric lavage samples. North American Journal of Fisheries Management, 28(6):1775–1778, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ficke:2009:MMM

[FM09] Ashley D. Ficke and Christopher A. Myrick. A method for monitoring movements of small fishes in urban streams. North American Journal of Fisheries Management, 29(5): 1444–1453, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frohnauer:2007:PDA

[FPK07a] Nick K. Frohnauer, Clay L. Pierce, and Larry W. Kallemeyn. Population dynamics and angler exploitation of the unique muskellunge population in Shoepack Lake, Voyageurs National Park, Minnesota. North American Journal of Fisheries Management, 27(1):63–76, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frohnauer:2007:SER

[FPK07b] Nick K. Frohnauer, Clay L. Pierce, and Larry W. Kallemeyn. Simulated effects of recruitment variability, exploitation, and reduced habitat area on the muskellunge population in Shoepack Lake, Voyageurs National Park, Minnesota. North American Journal of Fisheries Management, 27(1): 77–88, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Felix-Pico:2009:GFB

[FPRRHQ09] Esteban F. Félix-Pico, Mauricio Ramírez-Rodríguez, and Oscar Holguín-Quiñones. Growth and fisheries of the Black Ark Anadara tuberculosa, a bivalve mollusk, in Bahía Magdalena, Baja California Sur, Mexico. North American Journal of Fisheries Management, 29(1):231–236, February 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fries:2001:RCW

[Fri01] Joe N. Fries. Retention of coded wire tags in four locations in juvenile paddlefish. North American Journal of Fisheries Management, 21(4):962–966, November 2001. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frick:2009:PRP

[FRW09] Lorenz H. Frick, Richard D. Reina, and Terence I. Walker. The physiological response of port Jackson sharks and Australian swellsharks to sedation, gill-net capture, and repeated sampling in captivity. North American Journal of Fisheries Management, 29(1):127–139, February 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fayram:2000:IPS

[FS00] Andrew H. Fayram and Thomas H. Sibley. Impact of predation by smallmouth bass on sockeye salmon in Lake Washington, Washington. North American Journal of Fisheries Management, 20(1):81–89, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fayram:2006:EMB

[FS06] Andrew H. Fayram and Patrick J. Schmalz. Evaluation of a modified bag limit for walleyes in Wisconsin: Effects of decreased angler effort and lake selection. North American Journal of Fisheries Management, 26(3):606–611, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flammang:2007:EHN

[FS07] Mark K. Flammang and Randall D. Schultz. Evaluation of hoop-net size and bait selection for sampling channel catfish in Iowa impoundments. North American Journal of Fisheries Management, 27(2):512–518, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fresh:2003:PNP

[FSC03] Kurt L. Fresh, Steven L. Schroder, and Mark I. Carr. Predation by northern pikeminnow on hatchery and wild Coho salmon smolts in the Chehalis River, Washington. North American Journal of Fisheries Management, 23(4):1257– 1264, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Firehammer:2006:MPG

[FSF06] Jon A. Firehammer, Dennis L. Scarnecchia, and Steven R. Fain. Modification of a passive gear to sample paddlefish eggs in sandbed spawning reaches of the Lower Yellowstone River. North American Journal of Fisheries Management, 26 (1):63–72, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Feyrer:2006:IFD

[FSH06] Frederick Feyrer, Ted Sommer, and William Harrell. Importance of flood dynamics versus intrinsic physical habitat in structuring fish communities: Evidence from two adjacent engineered floodplains on the Sacramento River, California. North American Journal of Fisheries Management, 26(2):408–417, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

FitzGerald:2004:VVI

[FSK04] Jennifer L. FitzGerald, Timothy F. Sheehan, and John F. Kocik. Visibility of visual implant elastomer tags in Atlantic salmon reared for two years in marine net-pens. North American Journal of Fisheries Management, 24(1):222–227, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Freedman:2009:DEE

[FSLS09] Jonathan A. Freedman, Timothy D. Stecko, Benjamin D. Lorson, and Jay R. Stauffer, Jr. Development and efficacy of an electrified benthic trawl for sampling large-river fish assemblages. North American Journal of Fisheries Management, 29(4):1001–1005, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fabrizio:2008:FAC

[FSSR08] Mary C. Fabrizio, Frederick S. Scharf, Gary R. Shepherd, and John E. Rosendale. Factors affecting catch-and-release mortality of bluefish. North American Journal of Fisheries Management, 28(2):533–546, February 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Faulkner:2006:EEI

[FTWS06] Sean G. Faulkner, William M. Tonn, Marek Welz, and Douglas R. Schmitt. Effects of explosives on incubating lake trout eggs in the Canadian Arctic. North American Journal of Fisheries Management, 26(4):833–842, November 2006. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fudge:2008:EDL

[FWEP08] Thomas S. Fudge, Kerry G. Wautier, Robert E. Evans, and Vince P. Palace. Effect of different levels of fine-sediment loading on the escapement success of rainbow trout fry from artificial redds. North American Journal of Fisheries Management, 28(3):758–765, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Galbreath:2001:PNS

[GAG^{+}01] Peter F. Galbreath, Nathan D. Adams, Stanley Z. Guffey, C. Joseph Moore, and Jerry L. West. Persistence of native Southern Appalachian brook trout populations in the Pigeon River System, North Carolina. North American Journal of Fisheries Management, 21(4):927–934, November 2001. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guillory:2004:ECC

[GAKB04] Vincent Guillory, Ralph Allemand, Kenny King, and Lisa Bare. An evaluation of 5.87-cm and 6.03-cm escape rings for blue crab Callinectes sapidus traps. North American Journal of Fisheries Management, 24(4):1431–1434, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Geist:2008:IRL

[GAM^{+}08] David R. Geist, Evan V. Arntzen, Christopher J. Murray, Kathleen E. McGrath, Yi-Ju Bott, and Timothy P. Hanrahan. Influence of river level on temperature and hydraulic gradients in chum and fall Chinook salmon spawning areas downstream of Bonneville Dam, Columbia River. North American Journal of Fisheries Management, 28(1):30–41, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Godbout:2009:MMI

[GAR^{+}09] Jason D. Godbout, D. Derek Aday, James A. Rice, Max R. Bangs, and Joseph M. Quattro. Morphological models for identifying largemouth bass, spotted bass, and largemouth bass × spotted bass hybrids. North American Journal of Fisheries Management, 29(5):1425–1437, October 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerow:2005:NMC

[GASH05] Kenneth G. Gerow, Richard C. Anderson-Sprecher, and Wayne A. Hubert. A new method to compute standardweight equations that reduces length-related bias. North American Journal of Fisheries Management, 25(4):1288– 1300, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gard:2003:ANT

[GB03] Mark Gard and Ed Ballard. Applications of new technologies to instream flow studies in large rivers. North American Journal of Fisheries Management, 23(4):1114–1125, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Greenawalt-Boswell:2007:MER

[GBFJA07] Jaime Greenawalt-Boswell, Thomas K. Frazer, Charles A. Jacoby, and William S. Arnold. Mortality and exploitation rate estimates for the recreational bay scallop fishery off the Gulf Coast of Florida, USA. North American Journal of Fisheries Management, 27(4):1230–1242, November 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Geist:2002:PAE

[GBLC02] David R. Geist, Richard S. Brown, Ken Lepla, and James Chandler. Practical application of electromyogram radiotelemetry: The suitability of applying laboratory-acquired calibration data to field data. North American Journal of Fisheries Management, 22(2):474–479, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

George:2008:RBD

[GBPT08] Gerald J. George, Kenneth M. Brown, Gary W. Peterson, and Bruce A. Thompson. Removal of black drum on Louisiana reefs to increase survival of Eastern oysters Crassostrea virginica. North American Journal of Fisheries Management, 28(6):1802–1811, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gravel:2008:SBI

[GC08] Marie-Ange Gravel and Steven J. Cooke. Severity of barotrauma influences the physiological status, postrelease behavior, and fate of tournament-caught smallmouth bass. North American Journal of Fisheries Management, 28(2):607–617, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallaway:2003:DSE

[GCM^{+}03a] Benny J. Gallaway, John G. Cole, Larry R. Martin, James M. Nance, and Michael Longnecker. Description of a simple electronic logbook designed to measure effort in the Gulf of Mexico shrimp fishery. North American Journal of Fisheries Management, 23(2):581–589, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallaway:2003:EEL

[GCM^{+}03b] Benny J. Gallaway, John G. Cole, Larry R. Martin, James M. Nance, and Michael Longnecker. An evaluation of an electronic logbook as a more accurate method of estimating spatial patterns of trawling effort and bycatch in the Gulf of Mexico shrimp fishery. North American Journal of Fisheries Management, 23(3):787–809, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Garcia:2004:MSD

[GCM^{+}04] Aaron P. Garcia, William P. Connor, Deborah J. Milks, Stephen J. Rocklage, and R. Kirk Steinhorst. Movement and spawner distribution of hatchery fall Chinook salmon adults acclimated and released as yearlings at three locations in the Snake River Basin. North American Journal of Fisheries Management, 24(4):1134–1144, November 2004. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallaway:2008:SLA

[GCN^{+}08] Benny J. Gallaway, John G. Cole, James M. Nance, Rick A. Hart, and Gary L. Graham. Shrimp loss associated with turtle excluder devices: Are the historical estimates statistically biased. North American Journal of Fisheries Management, 28(1):203–211, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Garcia-Contreras:2009:AGA

[GCQVMAVP09] Oscar Enrique García-Contreras, Casimiro Quiñónez-Velázquez, Ramón Enrique Morán-Angulo, and Maria Candelaria Valdez-Pineda. Age, growth, and age structure of Amarillo snapper off the Coast of Mazatlán, Sinaloa, Mexico. North American Journal of Fisheries Management, 29(1):223–230, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Groves:2008:CTD

[GCR08] Phillip A. Groves, James A. Chandler, and Tracy J. Richter. Comparison of temperature data collected from artificial Chinook salmon redds and surface water in the Snake River. North American Journal of Fisheries Management, 28(3): 766–780, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gillis:2002:CAU

[GD02] Kirk S. Gillis and Robert B. Ditton. A conjoint analysis of U.S. Atlantic billfish fishery management alternatives. North American Journal of Fisheries Management, 22(4): 1218–1228, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gedamke:2004:SEO

[GDH04] Todd Gedamke, William D. DuPaul, and John M. Hoenig. A spatially explicit open-ocean DeLury analysis to estimate gear efficiency in the dredge fishery for sea scallop Placopecten magellanicus. North American Journal of Fisheries Management, 24(2):335–351, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gedamke:2005:IRE

[GDH05] Todd Gedamke, William D. DuPaul, and John M. Hoenig. Index-removal estimates of dredge efficiency for sea scallops on Georges Bank. North American Journal of Fisheries Management, 25(3):1122–1129, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerow:2007:PSS

[Ger07] Kenneth G. Gerow. Power and sample size estimation techniques for fisheries management: Assessment and a new computational tool. North American Journal of Fisheries Management, 27(2):397–404, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerow:2009:CHE

[Ger09] Kenneth G. Gerow. Comment: How to estimate quartiles using the blom method. North American Journal of Fisheries Management, 29(1):176, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grabowski:2009:CPB

[GFPJ09] Timothy B. Grabowski, Tyler D. Ferguson, James T. Peterson, and Cecil A. Jennings. Capture probability and behavioral response of the robust redhorse, a cryptic riverine fish, to electrofishing. North American Journal of Fisheries Management, 29(3):721–729, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallagher:2005:DCS

[GG05] Sean P. Gallagher and Colin M. Gallagher. Discrimination of Chinook salmon, Coho salmon, and steelhead redds and evaluation of the use of redd data for estimating escapement in several unregulated streams in Northern California. North American Journal of Fisheries Management, 25(1):284–300, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gould:2009:IBL

[GG09a] A. Gould and B. S. Grace. Injuries to barramundi Lates calcarifer resulting from lip-gripping devices in the laboratory. North American Journal of Fisheries Management, 29 (5):1418–1424, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gregory:2009:CTS

[GG09b] Jim S. Gregory and Bart L. Gamett. Cattle trampling of simulated bull trout redds. North American Journal of Fisheries Management, 29(2):361–366, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gazey:2008:ACG

[GGCF08] William J. Gazey, Benny J. Gallaway, John G. Cole, and David A. Fournier. Age composition, growth, and densitydependent mortality in juvenile red snapper estimated from observer data from the Gulf of Mexico penaeid shrimp fishery. North American Journal of Fisheries Management, 28 (6):1828–1842, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gaikowski:2001:SDE

[GGG01] Mark P. Gaikowski, William H. Gingerich, and Steve Gutreuter. Short-duration electrical immobilization of lake trout. North American Journal of Fisheries Management, 21 (2):381–392, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerrity:2008:HUJ

[GGG08] Paul C. Gerrity, Christopher S. Guy, and William M. Gardner. Habitat use of juvenile pallid sturgeon and shovelnose sturgeon with implications for water-level management in a downstream reservoir. North American Journal of Fisheries Management, 28(3):832–843, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gates:2008:AMC

[GGZH08] Kiza K. Gates, Christopher S. Guy, Alexander V. Zale, and Travis B. Horton. Adherence of Myxobolus cerebralis Myxospores to waders: Implications for disease dissemination. North American Journal of Fisheries Management, 28(5): 1453–1458, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Graves:2008:DHC

[GH08] John E. Graves and Andrij Z. Horodysky. Does hook choice matter? Effects of three circle hook models on postrelease survival of white marlin. North American Journal of Fisheries Management, 28(2):471–480, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Geist:2002:PCH

[GHA^{+}02] David R. Geist, Timothy P. Hanrahan, Evan V. Arntzen, Geoffrey A. McMichael, Christopher J. Murray, and Yi-Ju Chien. Physicochemical characteristics of the hyporheic zone affect redd site selection by chum salmon and fall Chinook salmon in the Columbia River. North American Journal of Fisheries Management, 22(4):1077–1085, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerow:2004:AAD

[GHAS04] Kenneth G. Gerow, Wayne A. Hubert, and Richard C. Anderson-Sprecher. An alternative approach to detection of length-related biases in standard weight equations. North American Journal of Fisheries Management, 24(3):903–910, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gimenez-Hurtado:2009:NMR

[GHASLC09] Enrique Giménez-Hurtado, Francisco Arreguín-Sánchez, and Salvador E. Lluch-Cota. Natural mortality rates during life history stages of the red grouper on Campeche Bank, Mexico. North American Journal of Fisheries Management, 29 (1):216–222, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gedamke:2009:SRD

[GHDM09] Todd Gedamke, John M. Hoenig, William D. DuPaul, and John A. Musick. Stock–recruitment dynamics and the maximum population growth rate of the barndoor skate on Georges Bank. North American Journal of Fisheries Management, 29(2):512–526, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gedamke:2007:UDM

[GHM^{+}07] Todd Gedamke, John M. Hoenig, John A. Musick, William D. DuPaul, and Samuel H. Gruber. Using demographic models to determine intrinsic rate of increase and sustainable fishing for elasmobranchs: Pitfalls, advances, and applications. North American Journal of Fisheries Management, 27 (2):605–618, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gibson:2009:EAP

[GJB09] A. Jamie F. Gibson, Ross A. Jones, and Heather D. Bowlby. Equilibrium analyses of a Population’s response to recovery activities: a case study with Atlantic salmon. North American Journal of Fisheries Management, 29(4):958–974, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gries:2002:NST

[GL02a] G. Gries and B. H. Letcher. A night seining technique for sampling juvenile Atlantic salmon in streams. North American Journal of Fisheries Management, 22(2):595–601, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gries:2002:TRS

[GL02b] G. Gries and B. H. Letcher. Tag retention and survival of age-0 Atlantic salmon following surgical implantation with passive integrated transponder tags. North American Journal of Fisheries Management, 22(1):219–222, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gatz:2008:ERE

[GL08] A. John Gatz and R. Scott Linder. Effects of repeated electroshocking on condition, growth, and movement of selected warmwater stream fishes. North American Journal of Fisheries Management, 28(3):792–798, June 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gido:2000:SCP

[GLA00] Keith B. Gido, Robert D. Larson, and Lief A. Ahlm. Stream-channel position of adult rainbow trout downstream of Navajo Reservoir, New Mexico, following changes in reservoir release. North American Journal of Fisheries Management, 20(1):250–258, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Greene:2000:ITS

[GM00] James C. Greene and Michael J. Maceina. Influence of trophic state on spotted bass and largemouth bass spawning time and age-0 population characteristics in Alabama reservoirs. North American Journal of Fisheries Management, 20(1):100–108, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gibson:2002:EHF

[GM02] A. Jamie F. Gibson and Ransom A. Myers. Effectiveness of a high-frequency-sound fish diversion system at the Annapolis Tidal Hydroelectric Generating Station, Nova Scotia. North American Journal of Fisheries Management, 22(3):770–784, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gaines:2004:FDM

[GM04] Phillip C. Gaines and Craig D. Martin. Feasibility of dualmarking age-0 Chinook salmon for mark–recapture studies. North American Journal of Fisheries Management, 24(4): 1456–1459, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goldstein:2005:MAF

[GM05] Robert M. Goldstein and Michael R. Meador. Multilevel assessment of fish species traits to evaluate habitat degradation in streams of the Upper Midwest. North American Journal of Fisheries Management, 25(1):180–194, February 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Graham:2009:USC

[GMH09] Larissa J. Graham, Brian R. Murphy, and David Hata. Using species composition data from a trawl survey to determine potential bycatch of the commercial trawl fishery for horseshoe crab Limulus polyphemus in the Middle Atlantic Bight. North American Journal of Fisheries Management, 29(2):478–487, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gatt:2003:MDV

[GMHF03] Michael H. Gatt, Tara L. McParland, Larry C. Halyk, and Moira M. Ferguson. Mitochondrial DNA variation and mixedstock analysis of recreational and commercial walleye fisheries in Eastern Lake Erie. North American Journal of Fisheries Management, 23(2):431–440, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Geist:2008:MEF

[GMHX08] David R. Geist, Christopher J. Murray, Timothy P. Hanrahan, and Yulong Xie. A model of the effects of flow fluctuations on fall Chinook salmon spawning habitat availability in the Columbia River. North American Journal of Fisheries Management, 28(6):1894–1910, December 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goeman:2002:WMN

[Goe02] Timothy J. Goeman. Walleye management in North America. North American Journal of Fisheries Management, 22 (3):973–974, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guy:2009:CCP

[GOG09] Christopher S. Guy, Eric W. Oldenburg, and Paul C. Gerrity. Conditional capture probability of Scaphirhynchus spp. in drifting trammel nets. North American Journal of Fisheries Management, 29(3):817–822, June 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goldstein:2009:ISP

[Gol09] Robert M. Goldstein. Interpreting stream physical characteristics in index of biotic integrity classifications: General similarities and specific differences. North American Journal of Fisheries Management, 29(1):151–162, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goodyear:2007:RCR

[Goo07] C. Phillip Goodyear. Recreational catch and release: Resource allocation between commercial and recreational fishermen. North American Journal of Fisheries Management, 27(4):1189–1194, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gangl:2003:BPI

[GP03] R. Scott Gangl and Donald L. Pereira. Biological performance indicators for evaluating exploitation of Minnesota’s large-lake walleye fisheries. North American Journal of Fisheries Management, 23(4):1303–1311, November 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gadomski:2005:LSV

[GP05] Dena M. Gadomski and Michael J. Parsley. Laboratory studies on the vulnerability of young white sturgeon to predation. North American Journal of Fisheries Management, 25 (2):667–674, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Godby:2007:DFR

[GRM07] Neal A. Godby, Jr., Edward S. Rutherford, and Doran M. Mason. Diet, feeding rate, growth, mortality, and production of juvenile steelhead in a Lake Michigan tributary. North American Journal of Fisheries Management, 27(2):578–592, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guy:2002:EMW

[GSC02a] Christopher S. Guy, Randall D. Schultz, and Michael A. Colvin. Ecology and management of white bass. North American Journal of Fisheries Management, 22(2):606–608, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Guy:2002:VGD

[GSC02b] Christopher S. Guy, Randall D. Schultz, and Carson A. Cox. Variation in gonad development, growth, and condition of white bass in Fall River Reservoir, Kansas. North American Journal of Fisheries Management, 22(2):643–651, May 2002.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grimaldo:2009:FAF

[GSV^{+}09] Lenny F. Grimaldo, Ted Sommer, Nick Van Ark, Gardner Jones, Erika Holland, Peter B. Moyle, Bruce Herbold, and Pete Smith. Factors affecting fish entrainment into massive water diversions in a tidal freshwater estuary: Can fish losses be managed? North American Journal of Fisheries Management, 29(5):1253–1270, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grant:2004:TAM

[GSWS04] Gerold C. Grant, Yaniv Schwartz, Sanford Weisberg, and Dennis H. Schupp. Trends in abundance and mean size of fish captured in gill nets from Minnesota lakes, 1983–1997. North American Journal of Fisheries Management, 24(2):417–428, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Graeb:2005:DMT

[GSWS05] Brian D. S. Graeb, Jamie L. Shepherd, David W. Willis, and Jason S. Sorensen. Delayed mortality of tournament-caught walleyes. North American Journal of Fisheries Management, 25(1):251–255, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Garland:2002:CSF

[GTRC02] Rodney D. Garland, Kenneth F. Tiffan, Dennis W. Rondorf, and Loreley O. Clark. Comparison of subyearling fall Chinook salmon’s use of riprap revetments and unaltered habitats in Lake Wallula of the Columbia River. North American Journal of Fisheries Management, 22(4):1283–1289, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gorsky:2009:ESS

[GTZM09] Dimitry Gorsky, Joan Trial, Joseph Zydlewski, and James McCleave. The effects of smolt stocking strategies on migratory path selection of adult Atlantic salmon in the Penobscot River, Maine. North American Journal of Fisheries Management, 29(4):949–957, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gangl:2004:CFD

[GW04] R. Scott Gangl and Roy A. Whaley. Comparison of fish density estimates from repeated hydroacoustic surveys on two Wyoming waters. North American Journal of Fisheries Management, 24(4):1279–1287, November 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Galinat:2002:ISS

[GWBH02] Gene F. Galinat, David W. Willis, Brian G. Blackwell, and Matthew J. Hubers. Influence of a saugeye (sauger × walleye) introduction program on the black crappie population in Richmond Lake, South Dakota. North American Journal of Fisheries Management, 22(4):1416–1424, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goldstein:2002:DSH

[GWSS02] Robert M. Goldstein, Lizhu Wang, Thomas P. Simon, and Paul M. Stewart. Development of a stream habitat index for the northern lakes and forests ecoregion. North American Journal of Fisheries Management, 22(2):452–464, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gale:2008:EFS

[GZC08] Steven B. Gale, Alexander V. Zale, and Christopher G. Clancy. Effectiveness of fish screens to prevent entrainment of westslope cutthroat trout into irrigation canals. North American Journal of Fisheries Management, 28(5):1541–1553, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunt:2002:EHM

[HA02] Jodee Hunt and Cynthia A. Annett. Effects of habitat manipulation on reproductive success of individual largemouth bass in an Ozark reservoir. North American Journal of Fisheries Management, 22(4):1201–1208, November 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hannah:2003:SCT

[Han03] Robert W. Hannah. Spatial changes in trawl fishing effort in response to footrope diameter restrictions in the U.S. West Coast bottom trawl fishery. North American Journal of Fisheries Management, 23(3):693–702, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harmon:2003:THA

[Har03] Jerrel R. Harmon. A trap for handling adult anadromous salmonids at Lower Granite Dam on the Snake River, Washington. North American Journal of Fisheries Management, 23(4):989–992, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoxmeier:2001:FIP

[HAW01] R. John H. Hoxmeier, D. Derek Aday, and David H. Wahl. Factors influencing precision of age estimation from scales and otoliths of bluegills in Illinois reservoirs. North American Journal of Fisheries Management, 21(2):374–380, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hatfield:2000:PSH

[HB00] Todd Hatfield and James Bruce. Predicting salmonid habitat–flow relationships for streams from Western North America. North American Journal of Fisheries Management, 20(4):1005–1015, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haddix:2005:FLG

[HB05a] Tyler Haddix and Phaedra Budy. Factors that limit growth and abundance of rainbow trout across ecologically distinct areas of Flaming Gorge Reservoir, Utah–Wyoming. North American Journal of Fisheries Management, 25(3):1082– 1094, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hagen:2005:ADC

[HB05b] John Hagen and James S. Baxter. Accuracy of diver counts of fluvial rainbow trout relative to horizontal underwater visibility. North American Journal of Fisheries Management, 25(4):1367–1377, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoffman:2005:GDM

[HB05c] Kevin J. Hoffman and Phillip W. Bettoli. Growth, dispersal, mortality, and contribution of largemouth bass stocked into Chickamauga Lake, Tennessee. North American Journal of Fisheries Management, 25(4):1518–1527, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hetrick:2006:SBH

[HB06] Nicholas J. Hetrick and Jeffrey F. Bromaghin. Sampling bias of hook-and-line gear used to capture rainbow trout in Gertrude Creek, Alaska. North American Journal of Fisheries Management, 26(1):13–23, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hutt:2007:PSM

[HB07] Clifford P. Hutt and Phillip W. Bettoli. Preferences, specialization, and management attitudes of trout anglers fishing in Tennessee tailwaters. North American Journal of Fisheries Management, 27(4):1257–1267, November 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hodgens:2004:GBH

[HBB04] Lynn S. Hodgens, Steven C. Blumenshine, and James C. Bednarz. Great blue Heron predation on stocked rainbow trout in an Arkansas tailwater fishery. North American Journal of Fisheries Management, 24(1):63–75, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunt:2007:PFP

[HBB07] Len M. Hunt, Barry N. Boots, and Peter C. Boxall. Predicting fishing participation and site choice while accounting for spatial substitution, trip timing, and trip context. North American Journal of Fisheries Management, 27(3):832–847, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hamel:2008:BRR

[HBC08] Martin J. Hamel, Michael L. Brown, and Steven R. Chipps. Behavioral responses of rainbow smelt to in situ strobe lights. North American Journal of Fisheries Management, 28(2): 394–401, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Huckins:2008:ELH

[HBFL08] Casey J. Huckins, Edward A. Baker, Kurt D. Fausch, and Jill B. K. Leonard. Ecology and life history of coaster brook trout and potential bottlenecks in their rehabilitation. North American Journal of Fisheries Management, 28(5):1321–1342, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2000:CRC

[HBH00] Michael J. Hansen, T. Douglas Beard, Jr., and Steven W. Hewett. Catch rates and catchability of walleyes in angling and spearing fisheries in Northern Wisconsin lakes. North American Journal of Fisheries Management, 20(1):109–118, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2005:EME

[HBH05] Michael J. Hansen, T. Douglas Beard, Jr., and Steven W. Hewett. Effect of measurement error on tests of density dependence of catchability for walleyes in Northern Wisconsin angling and spearing fisheries. North American Journal of Fisheries Management, 25(3):1010–1015, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Horns:2000:SCR

[HBHB00] William H. Horns, William P. Brown, Scott R. Hulse, and Charles R. Bronte. Seasonal changes in ruffe abundance in two Lake Superior tributaries: Implications for control. North American Journal of Fisheries Management, 20(3):822–826, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Henry:2005:ENT

[HBJ05] Sam D. Henry, Sam W. Barkley, and Ronald L. Johnson. Exploitation of Nile tilapia in a closed-system public fishing reservoir in Northern Arkansas. North American Journal of Fisheries Management, 25(3):853–860, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hatzenbeler:2000:SVF

[HBJE00] Gene R. Hatzenbeler, Michael A. Bozek, Martin J. Jennings, and Edward E. Emmons. Seasonal variation in fish assemblage structure and habitat structure in the nearshore littoral zone of Wisconsin lakes. North American Journal of Fisheries Management, 20(2):360–368, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Henry:2008:ASS

[HBKJ08] Sam D. Henry, Sam W. Barkley, Jeffrey B. Koppelman, and Ronald L. Johnson. Assessment of stocking success of walleye in the Eleven Point River, Arkansas. North American Journal of Fisheries Management, 28(5):1498–1505, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herzog:2005:EBT

[HBS^{+}05] David P. Herzog, Valerie A. Barko, John S. Scheibe, Robert A. Hrabik, and David E. Ostendorf. Efficacy of a benthic trawl for sampling small-bodied fishes in large river systems. North American Journal of Fisheries Management, 25(2):594–603, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Havens:2008:EDB

[HBS^{+}08] Kirk J. Havens, Donna Marie Bilkovic, Dave Stanhope, Kory Angstadt, and Carl Hershner. The effects of derelict blue crab traps on marine organisms in the Lower York River, Virginia. North American Journal of Fisheries Management, 28(5):1194–1200, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoffnagle:2008:RTS

[HCFK08] Timothy L. Hoffnagle, Richard W. Carmichael, Kathryn A. Frenyea, and Patrick J. Keniry. Run timing, spawn timing, and spawning distribution of hatchery- and natural-origin spring Chinook salmon in the Imnaha River, Oregon. North American Journal of Fisheries Management, 28(1):148–164, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holt:2009:SFC

[HCS09] Kendra R. Holt, Sean P. Cox, and Joel Sawada. In-season forecasting of Coho salmon marine survival via coded wire tag recoveries. North American Journal of Fisheries Management, 29(4):1165–1182, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haas-Castro:2006:SPA

[HCSCT06] Ruth E. Haas-Castro, Timothy F. Sheehan, Steven X. Cadrin, and Joan G. Trial. Scale pattern analysis discriminates Atlantic salmon by river reach rearing origin. North American Journal of Fisheries Management, 26(2):289–300, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunt:2002:FFP

[HD02] Kevin M. Hunt and Robert B. Ditton. Freshwater fishing participation patterns of racial and ethnic groups in Texas. North American Journal of Fisheries Management, 22(1): 52–65, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hudy:2000:SRA

[HDB00] Mark Hudy, Daniel M. Downey, and Darrell W. Bowman. Successful restoration of an acidified native brook trout stream through mitigation with limestone sand. North American Journal of Fisheries Management, 20(2):453–466, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haeseker:2007:ISM

[HDPS07] Steven L. Haeseker, Brigitte Dorner, Randall M. Peterman, and Zhenming Su. An improved sibling model for forecasting chum salmon and sockeye salmon abundance. North American Journal of Fisheries Management, 27(2):634–642, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hasselman:2009:DEA

[HEB09] D. J. Hasselman, T. A. Edge, and R. G. Bradford. Discrimination of the endangered Atlantic whitefish from lake whitefish and round whitefish by use of external characters. North American Journal of Fisheries Management, 29(4): 1046–1057, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hamon:2000:CMH

[HF00] Troy R. Hamon and Chris J. Foote. Changes in midorbital to hypural length and morphology in maturing sockeye salmon. North American Journal of Fisheries Management, 20(1): 245–249, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haxton:2008:PVL

[HFT08] Tim J. Haxton, C. Scott Findlay, and R. W. Threader. Predictive value of a lake sturgeon habitat suitability model. North American Journal of Fisheries Management, 28(5): 1373–1383, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harig:2000:FIS

[HFY00] Amy L. Harig, Kurt D. Fausch, and Michael K. Young. Factors influencing success of greenback cutthroat trout translocations. North American Journal of Fisheries Management, 20(4):994–1004, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Henning:2006:JSU

[HGF06] Julie A. Henning, Robert E. Gresswell, and Ian A. Fleming. Juvenile salmonid use of freshwater emergent wetlands in the floodplain and its implications for conservation management. North American Journal of Fisheries Management, 26 (2):367–376, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herrmann:2004:EFC

[HGHG04] Mark Herrmann, Joshua Greenberg, Charles Hamel, and Hans Geier. Extending Federal crop insurance programs to commercial fisheries: The case of Bristol Bay, Alaska, sockeye salmon. North American Journal of Fisheries Management, 24(2):352–366, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Horton:2004:ITI

[HGP04] Travis B. Horton, Christopher S. Guy, and Jeffrey S. Pontius. Influence of time interval on estimations of movement and habitat use. North American Journal of Fisheries Management, 24(2):690–696, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hebdon:2001:BCS

[HH01a] Jason L. Hebdon and Wayne A. Hubert. Body conditions and stomach contents of subadult trout during fall and winter in three Wyoming tailwaters. North American Journal of Fisheries Management, 21(4):897–903, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hyatt:2001:PSW

[HH01b] Matthew W. Hyatt and Wayne A. Hubert. Proposed standard-weight equations for brook trout. North American Journal of Fisheries Management, 21(1):253–254, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hyatt:2001:SPR

[HH01c] Matthew W. Hyatt and Wayne A. Hubert. Statistical properties of relative weight distributions of four salmonid species and their sampling implications. North American Journal of Fisheries Management, 21(3):666–670, August 2001. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hendricks:2002:HHR

[HHAK02] Michael L. Hendricks, Rickalon L. Hoopes, David A. Arnold, and Michael L. Kaufmann. Homing of hatchery-reared American shad to the Lehigh River, a tributary to the Delaware River. North American Journal of Fisheries Management, 22 (2):243–248, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Huff:2005:UFD

[HHB05] David D. Huff, Shannon L. Hubler, and Aaron N. Borisenko. Using field data to estimate the realized thermal niche of aquatic vertebrates. North American Journal of Fisheries Management, 25(1):346–360, February 2005. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2008:DIL

[HHL^{+}08] Michael J. Hansen, Ned J. Horner, Mark Liter, Mike P. Peterson, and Melo A. Maiolie. Dynamics of an increasing lake trout population in Lake Pend Oreille, Idaho. North American Journal of Fisheries Management, 28(4):1160–1171, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hedrick:2005:SCL

[HHM05] James D. Hedrick, L. Ras Hedrick, and F. Joseph Margraf. A sampler for capturing larval and juvenile Atlantic menhaden. North American Journal of Fisheries Management, 25(1):245–250, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hearn:2003:TRR

[HHPH03] William S. Hearn, John M. Hoenig, Kenneth H. Pollock, and Daniel A. Hepworth. Tag reporting rate estimation: 3. use of planted tags in one component of a multiple-component fishery. North American Journal of Fisheries Management, 23(1):66–77, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hiebert:2000:ASP

[HHWL00] Steve Hiebert, Louis A. Helfrich, Diana L. Weigmann, and Charles Liston. Anadromous salmonid passage and video image quality under infrared and visible light at Prosser Dam, Yakima River, Washington. North American Journal of Fisheries Management, 20(3):827–832, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hill:2001:SSE

[HI01] Ryan A. Hill and James R. Irvine. Standardizing spawner escapement data: a case study of the Nechako River Chinook salmon. North American Journal of Fisheries Management, 21(3):651–655, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilderbrand:2002:SSS

[Hil02] Robert H. Hilderbrand. Simulating supplementation strategies for restoring and maintaining stream resident cutthroat trout populations. North American Journal of Fisheries Management, 22(3):879–887, August 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hurley:2002:EMM

[HJ02] Keith L. Hurley and Jeffrey J. Jackson. Evaluation of a 254mm minimum length limit for crappies in two Southeast Nebraska reservoirs. North American Journal of Fisheries Management, 22(4):1369–1375, November 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hakala:2004:EGL

[HJ04] James P. Hakala and Frank D. Johnson. Evaluation of a gastric lavage method for use on largemouth bass. North American Journal of Fisheries Management, 24(4):1398– 1403, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hartman:2006:VIE

[HJ06] Kyle J. Hartman and Eric C. Janney. Visual implant elastomer and anchor tag retention in largemouth bass. North American Journal of Fisheries Management, 26(3):665–669, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hayes:2009:SAS

[HJC09] Christopher G. Hayes, Yan Jiao, and Enric Cortés. Stock assessment of scalloped hammerheads in the Western North Atlantic Ocean and Gulf of Mexico. North American Journal of Fisheries Management, 29(5):1406–1417, October 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilderbrand:2000:CIC

[HK00] Robert H. Hilderbrand and Jeffrey L. Kershner. Conserving inland cutthroat trout in small streams: How much stream is enough? North American Journal of Fisheries Management, 20(2):513–520, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilderbrand:2004:IHT

[HK04] Robert H. Hilderbrand and Jeffrey L. Kershner. Influence of habitat type on food supply, selectivity, and diet overlap of Bonneville cutthroat trout and nonnative brook trout in Beaver Creek, Idaho. North American Journal of Fisheries Management, 24(1):33–40, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hershberger:2001:MPH

[HKE^{+}01] P. K. Hershberger, R. M. Kocan, N. E. Elder, G. D. Marty, and J. Johnson. Management of Pacific herring spawn-on-kelp fisheries to optimize fish health and product quality. North American Journal of Fisheries Management, 21(4): 976–981, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Honeyfield:2008:EAS

[HKFM08] Dale C. Honeyfield, Tom Kehler, John W. Fletcher, and Jerre W. Mohler. Effect of artificial sunlight on the retention of external calcein marks on lake trout. North American Journal of Fisheries Management, 28(5):1243–1248, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hughes:2002:EDN

[HKH^{+}02] Robert M. Hughes, Philip R. Kaufmann, Alan T. Herlihy, Stephen S. Intelmann, Stephen C. Corbett, Meghan C. Arbogast, and Randy C. Hjort. Electrofishing distance needed to estimate fish species richness in raftable Oregon rivers. North American Journal of Fisheries Management, 22(4): 1229–1240, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hwang:2008:AGM

[HKL08] Sun-Do Hwang, Jin-Yeong Kim, and Tae-Won Lee. Age, growth, and maturity of chub mackerel off Korea. North American Journal of Fisheries Management, 28(5):1414– 1425, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Henderson:2000:TLS

[HKT00] Richard Henderson, Jeffrey L. Kershner, and C. Anna Toline. Timing and location of spawning by nonnative wild rainbow trout and native cutthroat trout in the South Fork Snake River, Idaho, with implications for hybridization. North American Journal of Fisheries Management, 20(3):584–596, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Headley:2008:DDG

[HL08] Heath C. Headley and Thomas E. Lauer. Density-dependent growth of yellow perch in Southern Lake Michigan, 1984– 2004. North American Journal of Fisheries Management, 28 (1):57–69, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunt:2009:EFR

[HL09] Len M. Hunt and Nigel Lester. The effect of forestry roads on access to remote fishing lakes in Northern Ontario, Canada. North American Journal of Fisheries Management, 29(3): 586–597, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hughes:2005:AGR

[HLH05] Thomas C. Hughes, Christopher E. Lowie, and James M. Haynes. Age, growth, relative abundance, and scuba capture of a new or recovering spawning population of lake sturgeon in the Lower Niagara River, New York. North American Journal of Fisheries Management, 25(4):1263–1272, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Helfrich:2001:SIS

[HLMB01] Louis A. Helfrich, Charles R. Liston, Brent Mefford, and Ray Bark. Survival and injury of splittail and Chinook salmon passed through a large hidrostal pump. North American Journal of Fisheries Management, 21(3):616–623, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoenig:2008:ESC

[HLO08] John M. Hoenig, Robert J. Latour, and John E. Olney. Estimating stock composition of anadromous fishes from mark– recovery data: Possible application to American shad. North American Journal of Fisheries Management, 28(2):507–515, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heimbuch:2007:ACE

[HLV^{+}07] Douglas G. Heimbuch, Ernesto Lorda, Douglas Vaughan, Lawrence W. Barnthouse, Jim Uphoff, Webster Van Winkle, Andy Kahnle, Byron Young, John Young, and Lisa Kline. Assessing coastwide effects of power plant entrainment and impingement on fish populations: Atlantic menhaden example. North American Journal of Fisheries Management, 27 (2):569–577, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunter:2008:DTD

[HM08] Ryan W. Hunter and Michael J. Maceina. Dispersal of tournament-displaced largemouth bass and spotted bass in Lake Martin, Alabama. North American Journal of Fisheries Management, 28(3):678–687, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

High:2009:SDH

[HM09] Brett High and Kevin A. Meyer. Survival and dispersal of hatchery triploid rainbow trout in an Idaho River. North American Journal of Fisheries Management, 29(6):1797– 1805, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hyatt:2005:EHV

[HMM^{+}05] Kim D. Hyatt, Karin L. Mathias, Donald J. McQueen, Brian Mercer, Patrick Milligan, and D. Paul Rankin. Evaluation of hatchery versus wild sockeye salmon fry growth and survival in two British Columbia lakes. North American Journal of Fisheries Management, 25(3):745–762, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holley:2009:FPC

[HMM09] Michael P. Holley, Matthew D. Marshall, and Michael J. Maceina. Fishery and population characteristics of blue catfish and channel catfish and potential impacts of minimum length limits on the fishery in Lake Wilson, Alabama. North American Journal of Fisheries Management, 29(4):1183–1194, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hockersmith:2003:CMR

[HMS^{+}03] Eric E. Hockersmith, William D. Muir, Steven G. Smith, Benjamin P. Sandford, Russell W. Perry, Noah S. Adams, and Dennis W. Rondorf. Comparison of migration rate and survival between radio-tagged and PIT-Tagged migrant yearling Chinook salmon in the Snake and Columbia Rivers. North American Journal of Fisheries Management, 23(2):404–413, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hewitt:2008:CBT

[HMSF08] Laura E. Hewitt, Karen G. Mumford, Donald R. Schreiner, and Gregory J. Fischer. Coaster brook trout rehabilitation in Lake Superior: a human dimensions perspective. North American Journal of Fisheries Management, 28(5): 1365–1372, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

High:2008:DAP

[HMSM08] Brett High, Kevin A. Meyer, Daniel J. Schill, and Elizabeth R. J. Mamer. Distribution, abundance, and population trends of bull trout in Idaho. North American Journal of Fisheries Management, 28(6):1687–1701, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hamilton:2009:CER

[HMW^{+}09] Bryan T. Hamilton, Steve E. Moore, Tod B. Williams, Neal Darby, and Mark R. Vinson. Comparative effects of rotenone and antimycin on macroinvertebrate diversity in two streams in Great Basin National Park, Nevada. North American Journal of Fisheries Management, 29(6):1620–1635, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2005:MCR

[HN05] Michael J. Hansen and Nancy A. Nate. A method for correcting the relative weight (W_{r}) index for seasonal patterns in relative condition (K_{n}) with length as applied to walleye in Wisconsin. North American Journal of Fisheries Management, 25(4):1256–1262, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2004:RRB

[HNE04] Michael J. Hansen, Steven P. Newman, and Clayton J. Edwards. A reexamination of the relationship between electrofishing catch rate and age-0 walleye density in Northern Wisconsin lakes. North American Journal of Fisheries Management, 24(2):429–439, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hartman:2000:VHE

[HNTM00] K. J. Hartman, B. Nagy, R. C. Tipton, and Scott Morrison. Verification of hydroacoustic estimates of fish abundance in Ohio River lock chambers. North American Journal of Fisheries Management, 20(4):1049–1056, November 2000. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Honeyfield:2006:DCM

[HOFM06] Dale C. Honeyfield, Christian S. Ostrowski, John W. Fletcher, and Jerre W. Mohler. Dietary calcein marking of brook trout, Atlantic salmon, yellow perch, and Coho salmon scales. North American Journal of Fisheries Management, 26 (2):431–437, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Horns:2001:STV

[Hor01] William H. Horns. Spatial and temporal variation in length at age and condition of yellow perch in Southern Lake Michigan during 1986–1988. North American Journal of Fisheries Management, 21(3):580–591, August 2001. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Howell:2006:EWS

[How06] Philip J. Howell. Effects of wildfire and subsequent hydrologic events on fish distribution and abundance in tributaries of North Fork John Day River. North American Journal of Fisheries Management, 26(4):983–994, November 2006. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haxton:2004:ADC

[HP04] Tim Haxton and Kirby Punt. Abundance and demography of channel catfish in a large northern river. North American Journal of Fisheries Management, 24(4):1440–1446, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hook:2009:ECY

[HP09] Tomas O. Höök and Steven A. Pothoven. Energy content of young alewives in Eastern Lake Michigan and Muskegon Lake, a connected drowned river mouth lake. North American Journal of Fisheries Management, 29(2):378–387, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hannah:2005:ESF

[HPB05] Robert W. Hannah, Steven J. Parker, and Troy V. Buell. Evaluation of a selective flatfish trawl and diel variation in rockfish catchability as bycatch reduction tools in the deepwater complex fishery off the U.S. West Coast. North American Journal of Fisheries Management, 25(2):581–593, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilborn:2002:ERR

[HPM02] Ray Hilborn, Ana Parma, and Mark Maunder. Exploitation rate reference points for West Coast rockfish: Are they robust and are there better alternatives? North American Journal of Fisheries Management, 22(2):365–375, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hannah:2008:ESE

[HPM08] Robert W. Hannah, Steven J. Parker, and Keith M. Matteson. Escaping the surface: The effect of capture depth on submergence success of surface-released Pacific rockfish. North American Journal of Fisheries Management, 28(3):694–700, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haeseker:2005:REP

[HPSW05] Steven L. Haeseker, Randall M. Peterman, Zhenming Su, and Chris C. Wood. Retrospective evaluation of preseason forecasting models for pink salmon. North American Journal of Fisheries Management, 25(3):897–918, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haeseker:2008:REP

[HPSW08] Steven L. Haeseker, Randall M. Peterman, Zhenming Su, and Chris C. Wood. Retrospective evaluation of preseason forecasting models for sockeye and chum salmon. North American Journal of Fisheries Management, 28(1):12–29, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holliman:2002:EII

[HR02] F. Michael Holliman and James B. Reynolds. Electroshockinduced injury in juvenile white sturgeon. North American Journal of Fisheries Management, 22(2):494–499, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hart:2006:LTD

[HR06] Deborah R. Hart and Paul J. Rago. Long-term dynamics of U.S. Atlantic sea scallop Placopecten magellanicus populations. North American Journal of Fisheries Management, 26 (2):490–501, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hook:2004:RBS

[HRB^{+}04] Tomas O. Höök, Edward S. Rutherford, Shannon J. Brines, David J. Schwab, and Michael J. McCormick. Relationship between surface water temperature and steelhead distributions in Lake Michigan. North American Journal of Fisheries Management, 24(1):211–221, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holliman:2003:EII

[HRK03a] F. Michael Holliman, James B. Reynolds, and Thomas J. Kwak. Electroshock-induced injury and mortality in the spotfin chub, a threatened minnow. North American Journal of Fisheries Management, 23(4):962–966, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holliman:2003:PRM

[HRK03b] F. Michael Holliman, James B. Reynolds, and Thomas J. Kwak. A predictive risk model for electroshock-induced mortality of the endangered Cape Fear shiner. North American Journal of Fisheries Management, 23(4):905–912, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heitman:2006:SCG

[HRL06] N. Elizabeth Heitman, Christopher L. Racey, and Steve E. Lochmann. Stocking contribution and growth of largemouth bass in pools of the Arkansas River. North American Journal of Fisheries Management, 26(1):175–179, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haigh:2003:LTF

[HS03] Rowan Haigh and Jon T. Schnute. The longspine thornyhead fishery along the West Coast of Vancouver Island, British Columbia, Canada: Portrait of a developing fishery. North American Journal of Fisheries Management, 23(1):120–140, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hrodey:2008:FCR

[HS08] Peter J. Hrodey and Trent M. Sutton. Fish community responses to half-log additions in warmwater streams. North American Journal of Fisheries Management, 28(1):70–80, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hudy:2009:MRP

[HS09] Mark Hudy and Jeremy Shiflet. Movement and recolonization of Potomac sculpin in a Virginia stream. North American Journal of Fisheries Management, 29(1):196–204, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heckman:2000:SIS

[HSB^{+}00] Kacey W. Heckman, Stephen L. Smith, Robert C. Bayer, Edward G. E. Jahngen, and Francis X. Smith. Starch–iodide swab technique for detecting bleach-dipped female lobsters. North American Journal of Fisheries Management, 20(2): 538–541, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hurley:2004:APA

[HSH04] Keith L. Hurley, Robert J. Sheehan, and Roy C. Heidinger. Accuracy and precision of age estimates for pallid sturgeon from pectoral fin rays. North American Journal of Fisheries Management, 24(2):715–718, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hatten:2009:SMA

[HTA^{+}09] James R. Hatten, Kenneth F. Tiffan, Donald R. Anglin, Steven L. Haeseker, Joseph J. Skalicky, and Howard Schaller. A spatial model to assess the effects of hydropower operations on Columbia River fall Chinook salmon spawning habitat. North American Journal of Fisheries Management, 29 (5):1379–1405, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hudy:2008:DSL

[HTGS08] Mark Hudy, Teresa M. Thieling, Nathaniel Gillespie, and Eric P. Smith. Distribution, status, and land use characteristics of subwatersheds within the native range of brook trout in the Eastern United States. North American Journal of Fisheries Management, 28(4):1069–1085, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hyvarinen:2003:LRA

[HV03] Pekka Hyvärinen and Teppo Vehanen. Length at release affects movement and recapture of lake-stocked brown trout. North American Journal of Fisheries Management, 23(4): 1126–1135, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanson:2006:FAH

[HVN06] Peter J. Hanson, Douglas S. Vaughan, and Sridhar Narayan. Forecasting annual harvests of Atlantic and Gulf menhaden. North American Journal of Fisheries Management, 26(3): 753–764, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holecek:2007:SCA

[HW07] Dean E. Holecek and Jody P. Walters. Spawning characteristics of adfluvial rainbow trout in a North Idaho stream: Implications for error in redd counts. North American Journal of Fisheries Management, 27(3):1010–1017, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoxmeier:2009:FIS

[HW09] R. John H. Hoxmeier and David H. Wahl. Factors influencing short-term hooking mortality of bluegills and the implications for restrictive harvest regulations. North American Journal of Fisheries Management, 29(5):1372–1378, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Halupka:2003:CPD

[HWB^{+}03] Karl C. Halupka, Mary F. Willson, Mason D. Bryant, Fred H. Everest, and Anthony J. Gharrett. Conservation of population diversity of Pacific salmon in Southeast Alaska. North American Journal of Fisheries Management, 23(4):1057– 1086, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hayer:2008:EPF

[HWB08] Cari-Ann Hayer, Steven S. Wall, and Charles R. Berry, Jr. Evaluation of predicted fish distribution models for rare fish species in South Dakota. North American Journal of Fisheries Management, 28(5):1259–1269, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hedrick:2005:NSD

[HWH05] Lara B. Hedrick, Stuart A. Welsh, and James D. Hedrick. A new sampler design for measuring sedimentation in streams. North American Journal of Fisheries Management, 25(1): 238–244, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hining:2000:VSO

[HWKN00] K. J. Hining, J. L. West, M. A. Kulp, and A. D. Neubauer. Validation of scales and otoliths for estimating age of rainbow trout from Southern Appalachian streams. North American Journal of Fisheries Management, 20(4):978–985, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harvey:2009:EDF

[HWN09] Bret C. Harvey, Jason L. White, and Rodney J. Nakamoto. The effect of deposited fine sediment on summer survival and growth of rainbow trout in riffles of a small stream. North American Journal of Fisheries Management, 29(2):434–440, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holbrook:2009:MPA

[HZG^{+}09] Christopher M. Holbrook, Joseph Zydlewski, Dimitry Gorsky, Steven L. Shepard, and Michael T. Kinnison. Movements of prespawn adult Atlantic salmon near hydroelectric dams in the Lower Penobscot River, Maine. North American Journal of Fisheries Management, 29(2):495–505, April 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ianelli:2002:SAT

[Ian02] James N. Ianelli. Simulation analyses testing the robustness of productivity determinations from West Coast Pacific Ocean perch stock assessment data. North American Journal of Fisheries Management, 22(2):301–310, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2002:IPM

[IBSC02] Daniel A. Isermann, Phillip W. Bettoli, Steve M. Sammons, and Timothy N. Churchill. Initial poststocking mortality, oxytetracycline marking, and year-class contribution of black-nosed crappies stocked into Tennessee reservoirs. North American Journal of Fisheries Management, 22(4): 1399–1408, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2008:IMR

[IC08] Daniel A. Isermann and Andrew J. Carlson. Initial mortality and retention associated with using passive integrated transponders in black crappies. North American Journal of Fisheries Management, 28(4):1157–1159, August 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2004:SDB

[ICB04] Daniel A. Isermann, Steven R. Chipps, and Michael L. Brown. Seasonal Daphnia biomass in winterkill and nonwinterkill glacial lakes of South Dakota. North American Journal of Fisheries Management, 24(1):287–292, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Israel:2004:GPG

[ICBM04] Joshua A. Israel, Jan F. Cordes, Marc A. Blumberg, and Bernie May. Geographic patterns of genetic differentiation among collections of green sturgeon. North American Journal of Fisheries Management, 24(3):922–931, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Irwin:2003:RGS

[IDK03] Brian J. Irwin, Dennis R. DeVries, and Gene W. Kim. Responses to gizzard shad recovery following selective treatment in Walker County Lake, Alabama, 1996–1999. North American Journal of Fisheries Management, 23(4):1225– 1237, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isely:2000:RCW

[IF00] J. Jeffery Isely and Quenton C. Fontenot. Retention of coded wire tags in juvenile shortnose sturgeon. North American Journal of Fisheries Management, 20(4):1040–1043, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2002:CTM

[IHW02] Daniel A. Isermann, Patrick A. Hanchin, and David W. Willis. Comparison of two mesh sizes for collecting larval yellow perch in surface trawls. North American Journal of Fisheries Management, 22(2):585–589, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2005:CPA

[IK05a] Daniel A. Isermann and Carey T. Knight. A computer program for age–length keys incorporating age assignment to individual fish. North American Journal of Fisheries Management, 25(3):1153–1160, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2005:PEJ

[IK05b] Daniel A. Isermann and Carey T. Knight. Potential effects of jaw tag loss on exploitation estimates for Lake Erie walleyes. North American Journal of Fisheries Management, 25(2): 557–562, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2003:ETD

[IMSW03] Daniel A. Isermann, Jonathan R. Meerbeek, George D. Scholten, and David W. Willis. Evaluation of three different structures used for walleye age estimation with emphasis on removal and processing times. North American Journal of Fisheries Management, 23(2):625–631, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2002:AMQ

[IMW02] Daniel A. Isermann, William L. McKibbin, and David W. Willis. An analysis of methods for quantifying crappie recruitment variability. North American Journal of Fisheries Management, 22(4):1124–1135, November 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isely:2001:TRG

[IS01] J. Jeffery Isely and Patricia E. Stockett. Tag retention, growth, and survival of red swamp crayfish marked with a visible implant tag. North American Journal of Fisheries Management, 21(2):422–424, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2002:PES

[ISBC02] Daniel A. Isermann, Steve M. Sammons, Phillip W. Bettoli, and Timothy N. Churchill. Predictive evaluation of size restrictions as management strategies for Tennessee reservoir crappie fisheries. North American Journal of Fisheries Management, 22(4):1349–1357, November 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2007:CCP

[Ise07a] Daniel Isermann. Comment: a computer program for age– length keys incorporating age assignment to individual fish. North American Journal of Fisheries Management, 27(1): 360, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2007:EWL

[Ise07b] Daniel A. Isermann. Evaluating walleye length limits in the face of population variability: Case histories from Western Minnesota. North American Journal of Fisheries Management, 27(2):551–568, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2005:PWT

[IV05] Daniel A. Isermann and Christopher S. Vandergoot. Predicting walleye total length from head and mandible measurements. North American Journal of Fisheries Management, 25(1):316–321, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2007:YPS

[IWBL07] Daniel A. Isermann, David W. Willis, Brian G. Blackwell, and David O. Lucchesi. Yellow perch in South Dakota: Population variability and predicted effects of creel limit reductions and minimum length limits. North American Journal of Fisheries Management, 27(3):918–931, August 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2005:SHE

[IWLB05] Daniel A. Isermann, David W. Willis, David O. Lucchesi, and Brian G. Blackwell. Seasonal harvest, exploitation, size selectivity, and catch preferences associated with winter yellow perch anglers on South Dakota lakes. North American Journal of Fisheries Management, 25(3):827–840, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isely:2002:EAP

[IYJS02] J. Jeffery Isely, Shawn P. Young, Tucker A. Jones, and Jason J. Schaffler. Effects of antenna placement and antibiotic treatment on loss of simulated transmitters and mortality in hybrid striped bass. North American Journal of Fisheries Management, 22(1):204–207, February 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobson:2007:OSD

[JA07] Peter C. Jacobson and Charles S. Anderson. Optimal stocking densities of walleye fingerlings in Minnesota lakes. North American Journal of Fisheries Management, 27(2):650–658, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2005:EPS

[JAAW05] Gary E. Johnson, Steven M. Anglea, Noah S. Adams, and Timothy O. Wik. Evaluation of a prototype surface flow bypass for juvenile salmon and steelhead at the powerhouse of Lower Granite Dam, Snake River, Washington, 1996–2000. North American Journal of Fisheries Management, 25(1): 138–151, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobson:2005:EAR

[Jac05] Peter C. Jacobson. Experimental analysis of a reduced daily bluegill limit in Minnesota. North American Journal of Fisheries Management, 25(1):203–210, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jackson:2000:SSC

[JB00] Jeffrey J. Jackson and Daryl L. Bauer. Size structure and catch rates of white crappie, black crappie, and bluegill in trap nets with 13-mm and 16-mm mesh. North American Journal of Fisheries Management, 20(3):646–650, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2005:ESL

[JBG05] Peter N. Johnson, Kyle Bouchard, and Frederick A. Goetz. Effectiveness of strobe lights for reducing juvenile salmonid entrainment into a navigation lock. North American Journal of Fisheries Management, 25(2):491–501, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jiang:2007:EFM

[JBHP07] Honghua Jiang, Cavell Brownie, Joseph E. Hightower, and Kenneth H. Pollock. Estimating fishing mortality, natural mortality, and selectivity using recoveries from tagging young fish. North American Journal of Fisheries Management, 27 (3):773–781, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobson:2002:TES

[JCW02] Larry D. Jacobson, Steven X. Cadrin, and James R. Weinberg. Tools for estimating surplus production and F_{MSY} in any stock assessment model. North American Journal of Fisheries Management, 22(2):326–338, February 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jenkins:2002:ROI

[JDB^{+}02] Wallace E. Jenkins, Michael R. Denson, Charles B. Bridgham, Mark R. Collins, and Theodore I. J. Smith. Retention of oxytetracycline-induced marks on sagittae of red drum. North American Journal of Fisheries Management, 22 (2):590–594, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jenkins:2004:YCC

[JDB^{+}04] Wallace E. Jenkins, Michael R. Denson, Charles B. Bridgham, Mark R. Collins, and Theodore I. J. Smith. Yearclass component, growth, and movement of juvenile red drum stocked seasonally in a South Carolina Estuary. North American Journal of Fisheries Management, 24(2):636–647, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

James:2003:UGI

[JEB03] Daniel A. James, Jack W. Erickson, and Bruce A. Barton. Using a geographic information system to assess the accuracy of radio-triangulation techniques for fish telemetry. North American Journal of Fisheries Management, 23 (4):1271–1275, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

James:2007:BTS

[JEB07] Daniel A. James, Jack W. Erickson, and Bruce A. Barton. Brown trout seasonal movement patterns and habitat use in an urbanized South Dakota stream. North American Journal of Fisheries Management, 27(3):978–985, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2003:EJE

[JED^{+}03] Gary E. Johnson, Blaine D. Ebberts, Dennis D. Dauble, Albert E. Giorgi, Paul G. Heisey, Robert P. Mueller, and Duane A. Neitzel. Effects of jet entry at high-flow outfalls on juvenile Pacific salmon. North American Journal of Fisheries Management, 23(2):441–449, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jenkins:2003:ERI

[Jen03] Thomas M. Jenkins, Jr. Evaluating recent innovations in bait fishing tackle and technique for catch and release of rainbow trout. North American Journal of Fisheries Management, 23 (4):1098–1107, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobs:2009:UPS

[JGW^{+}09] Steven E. Jacobs, William Gaeuman, Matt A. Weeber, Stephanie L. Gunckel, and Steven J. Starcevich. Utility of a probabilistic sampling design to determine bull trout population status using redd counts in basins of the Columbia River Plateau. North American Journal of Fisheries Management, 29(6):1590–1604, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jackson:2001:RSB

[JH01] James R. Jackson and Joseph E. Hightower. Reservoir striped bass movements and site fidelity in relation to seasonal patterns in habitat quality. North American Journal of Fisheries Management, 21(1):34–45, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Janac:2007:CPA

[JJ07] Michal Janáĉ and Pavel Jurajda. A comparison of point abundance and continuous sampling by electrofishing for age-0 fish in a channelized lowland river. North American Journal of Fisheries Management, 27(4):1119–1125, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jennings:2005:ESW

[JKHE05] Martin J. Jennings, Jeffrey M. Kampa, Gene R. Hatzenbeler, and Edward E. Emmons. Evaluation of supplemental walleye stocking in Northern Wisconsin lakes. North American Journal of Fisheries Management, 25(3):1171–1178, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Josephson:2001:UOB

[JKS01] Daniel C. Josephson, Charles C. Krueger, and Patrick J. Sullivan. Use of outlet barriers to prevent fall emigration of brook trout stocked in Adirondack lakes. North American Journal of Fisheries Management, 21(4):843–854, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2006:MPR

[JLSL06] Nicholas S. Johnson, Mark A. Luehring, Michael J. Siefkes, and Weiming Li. Mating pheromone reception and induced behavior in ovulating female sea lampreys. North American Journal of Fisheries Management, 26(1):88–96, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2000:TEL

[JM00] Brett M. Johnson and Patrick J. Martinez. Trophic economics of lake trout management in reservoirs of differing productivity. North American Journal of Fisheries Management, 20 (1):127–143, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2008:RPT

[JMHB08] Brett M. Johnson, Patrick J. Martinez, John A. Hawkins, and Kevin R. Bestgen. Ranking predatory threats by nonnative fishes in the Yampa River, Colorado, via bioenergetics modeling. North American Journal of Fisheries Management, 28 (6):1941–1953, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jurado-Molina:2002:MPB

[JML02] Jesús Jurado-Molina and Patricia Livingston. Multispecies perspectives on the Bering Sea groundfish fisheries management regime. North American Journal of Fisheries Management, 22(4):1164–1175, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:2007:TMD

[JMP07] Darin T. Jones, Christine M. Moffitt, and K. Kenneth Peters. Temperature-mediated differences in bacterial kidney disease expression and survival in Renibacterium salmoninarum -challenged bull trout and other salmonids. North American Journal of Fisheries Management, 27(2):695–706, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2004:CSG

[Joh04] James H. Johnson. Comparative survival and growth of Atlantic salmon from egg stocking and fry releases. North American Journal of Fisheries Management, 24(4):1409– 1412, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:2005:ICH

[Jon05] Thomas S. Jones. The influence of circle hooks on the capture efficiency and injury rate of walleyes. North American Journal of Fisheries Management, 25(2):725–731, May 2005.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jiang:2007:TRM

[JPB^{+}07] Honghua Jiang, Kenneth H. Pollock, Cavell Brownie, John M. Hoenig, Robert J. Latour, Brian K. Wells, and Joseph E. Hightower. Tag return models allowing for harvest and catch and release: Evidence of environmental and management impacts on striped bass fishing and natural mortality rates. North American Journal of Fisheries Management, 27(2):387–396, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jager:2003:DOF

[JR03] Henriette I. Jager and Kenneth A. Rose. Designing optimal flow patterns for fall Chinook salmon in a Central Valley, California, river. North American Journal of Fisheries Management, 23(1):1–21, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jiao:2009:GEF

[JRNS09] Yan Jiao, Kevin Reid, Tom Nudds, and Eric Smith. Graphical evaluation of fishery status using a likelihood inference approach. North American Journal of Fisheries Management, 29(4):1106–1118, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Josephson:2008:LTR

[JRWK08] Daniel C. Josephson, Jason M. Robinson, Brian C. Weidel, and Clifford E. Kraft. Long-term retention and visibility of visible implant elastomer tags in brook trout. North American Journal of Fisheries Management, 28(6):1758–1761, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2005:COF

[JSL05] Nicholas S. Johnson, Michael J. Siefkes, and Weiming Li. Capture of ovulating female sea lampreys in traps baited with spermiating male sea lampreys. North American Journal of Fisheries Management, 25(1):67–72, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

James:2007:CRM

[JSMV07] Jason T. James, Gregory W. Stunz, David A. McKee, and Robert R. Vega. Catch-and-release mortality of spotted seatrout in Texas: Effects of tournaments, seasonality, and anatomical hooking location. North American Journal of Fisheries Management, 27(3):900–907, August 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jenkins:2007:UEE

[JT07] Jill A. Jenkins and R. Glenn Thomas. Use of eyeballs for establishing ploidy of Asian carp. North American Journal of Fisheries Management, 27(4):1195–1202, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jaeger:2005:SMH

[JZMS05] Matthew E. Jaeger, Alexander V. Zale, Thomas E. McMahon, and Brad J. Schmitz. Seasonal movements, habitat use, aggregation, exploitation, and entrainment of saugers in the Lower Yellowstone River: An empirical assessment of factors affecting population recovery. North American Journal of Fisheries Management, 25(4):1550–1568, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kaeding:2004:URS

[KB04a] Lynn R. Kaeding and Glenn D. Boltz. Use of remote-site incubators to produce Arctic grayling fry of wild parentage. North American Journal of Fisheries Management, 24(3): 1031–1037, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Karchesky:2004:WHU

[KB04b] Christopher M. Karchesky and David H. Bennett. Winter habitat use by adult largemouth bass in the Pend Oreille River, Idaho. North American Journal of Fisheries Management, 24(2):577–585, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kruse:2000:AMS

[KBF^{+}00] Gordon H. Kruse, Laurence C. Byrne, Fritz C. Funk, Scott C. Matulich, and Jie Zheng. Analysis of minimum size limit for the red king crab fishery in Bristol Bay, Alaska. North American Journal of Fisheries Management, 20(2):307–319, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keefer:2008:ODB

[KBPC08] Matthew L. Keefer, Charles T. Boggs, Christopher A. Peery, and Christopher C. Caudill. Overwintering distribution, behavior, and survival of adult summer steelhead: Variability among Columbia River populations. North American Journal of Fisheries Management, 28(1):81–96, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kocovsky:2000:CME

[KC00] Patrick M. Kocovsky and Robert F. Carline. A comparison of methods for estimating ages of unexploited walleyes. North American Journal of Fisheries Management, 20(4): 1044–1048, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kocovsky:2001:DUW

[KC01a] Patrick M. Kocovsky and Robert F. Carline. Dynamics of the unexploited walleye population of pymatuning sanctuary, Pennsylvania, 1997–1998. North American Journal of Fisheries Management, 21(1):178–187, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kocovsky:2001:IET

[KC01b] Patrick M. Kocovsky and Robert F. Carline. Influence of extreme temperatures on consumption and condition of walleyes in pymatuning sanctuary, Pennsylvania. North American Journal of Fisheries Management, 21(1):198–207, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kock:2006:ESC

[KCA06] Tobias J. Kock, James L. Congleton, and Paul J. Anders. Effects of sediment cover on survival and development of white sturgeon embryos. North American Journal of Fisheries Management, 26(1):134–141, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ketola:2005:TSC

[KCR^{+}05] H. George Ketola, Thomas L. Chiotti, Robert S. Rathman, John D. Fitzsimons, Dale C. Honeyfield, Peter J. Van Dusen, and Graham E. Lewis. Thiamine status of Cayuga Lake rainbow trout and its influence on spawning migration. North American Journal of Fisheries Management, 25 (4):1281–1287, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kim:2000:ESR

[KD00] Gene W. Kim and Dennis R. DeVries. Effects of a selectively reduced gizzard shad population on trophic interactions and age-0 fishes in Walker County Lake, Alabama. North American Journal of Fisheries Management, 20(4): 860–872, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kennedy:2007:PCS

[KDSF07] Anthony J. Kennedy, Daniel J. Daugherty, Trent M. Sutton, and Brant E. Fisher. Population characteristics of shovelnose sturgeon in the Upper Wabash River, Indiana. North American Journal of Fisheries Management, 27(1):52–62, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kerstetter:2008:MLP

[Ker08] David W. Kerstetter. Measuring the length of a pelagic longline set: Applications for management. North American Journal of Fisheries Management, 28(2):378–385, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kelder:2009:SEM

[KF09] Brian F. Kelder and John M. Farrell. A spatially explicit model to predict walleye spawning in an Eastern Lake Ontario tributary. North American Journal of Fisheries Management, 29(6):1686–1697, December 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kirkley:2001:ACC

[KFG^{+}01] James E. Kirkley, Rolf Färe, Shawna Grosskopf, Kenneth McConnell, Dale E. Squires, and Ivar Strand. Assessing capacity and capacity utilization in fisheries when data are limited. North American Journal of Fisheries Management, 21 (3):482–497, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kelso:2000:EUM

[KG00] J. R. M. Kelso and W. M. Gardner. Emigration, upstream movement, and habitat use by sterile and fertile sea lampreys in three Lake Superior tributaries. North American Journal of Fisheries Management, 20(1):144–153, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knapp:2005:CHF

[KG05] Michael L. Knapp and Timothy J. Goeman. Catch and harvest by fishing guides compared with other anglers in five Central Minnesota lakes. North American Journal of Fisheries Management, 25(3):885–889, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kerstetter:2008:PSS

[KG08] D. W. Kerstetter and J. E. Graves. Postrelease survival of sailfish caught by commercial pelagic longline gear in the Southern Gulf of Mexico. North American Journal of Fisheries Management, 28(5):1578–1586, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kelso:2001:IAF

[KGM01] J. R. M. Kelso, W. M. Gardner, and R. B. McDonald. Interactions among fertile male, female, and sterile male sea lampreys during spawning in the Carp River, Lake Superior. North American Journal of Fisheries Management, 21 (4):904–910, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kynard:2001:GYS

[KH01] Boyd Kynard and Martin Horgan. Guidance of yearling shortnose and pallid sturgeon using vertical bar rack and louver arrays. North American Journal of Fisheries Management, 21(3):561–570, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kampa:2009:SGW

[KH09] Jeffrey M. Kampa and Gene R. Hatzenbeler. Survival and growth of walleye fingerlings stocked at two sizes in 24 Wisconsin lakes. North American Journal of Fisheries Management, 29(4):996–1000, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Khan:2006:TDC

[Kha06] Liaqat A. Khan. A three-dimensional computational fluid dynamics (CFD) model analysis of free surface hydrodynamics and fish passage energetics in a vertical-slot fishway. North American Journal of Fisheries Management, 26(2):255–267, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kuhn:2008:HUM

[KHJ^{+}08] Kristopher M. Kuhn, Wayne A. Hubert, Kevin Johnson, Dennis Oberlie, and David Dufek. Habitat use and movement patterns by adult saugers from fall to summer in an unimpounded small-river system. North American Journal of Fisheries Management, 28(2):360–367, February 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kelly:2006:NSS

[KHJC06] Maureen H. Kelly, William G. Hagar, Timothy D. Jardine, and Richard A. Cunjak. Nonlethal sampling of sunfish and slimy sculpin for stable isotope analysis: How scale and fin tissue compare with muscle tissue. North American Journal of Fisheries Management, 26(4):921–925, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kruse:2000:SYC

[KHR00] Carter G. Kruse, Wayne A. Hubert, and Frank J. Rahel. Status of Yellowstone cutthroat trout in Wyoming waters. North American Journal of Fisheries Management, 20(3):693–705, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kapuscinski:2005:MLT

[KHS05] Kevin L. Kapuscinski, Michael J. Hansen, and Stephen T. Schram. Movements of lake trout in U.S. waters of Lake Superior, 1973–2001. North American Journal of Fisheries Management, 25(2):696–708, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ketola:2009:ETS

[KJR^{+}09] H. George Ketola, James H. Johnson, Jacques Rinchard, Francis J. Verdoliva, Mary E. Penney, Andreas W. Greulich, and Russell C. Lloyd. Effect of thiamine status on probability of Lake Ontario Chinook salmon spawning in the upper or lower sections of Salmon River, New York. North American Journal of Fisheries Management, 29(4):895–902, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knudsen:2009:EPI

[KJS^{+}09] Curtis M. Knudsen, Mark V. Johnston, Steven L. Schroder, William J. Bosch, David E. Fast, and Charles R. Strom. Effects of passive integrated transponder tags on smolt-to-adult recruit survival, growth, and behavior of hatchery spring Chinook salmon. North American Journal of Fisheries Management, 29(3):658–669, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kinziger:2008:HBS

[KLH^{+}08] Andrew P. Kinziger, Eric J. Loudenslager, David G. Hankin, Eric C. Anderson, and John Carlos Garza. Hybridization between spring- and fall-run Chinook salmon returning to the Trinity River, California. North American Journal of Fisheries Management, 28(5):1426–1438, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kenaston:2001:EAH

[KLS01] Kenneth R. Kenaston, Robert B. Lindsay, and R. Kirk Schroeder. Effect of acclimation on the homing and survival of hatchery winter steelhead. North American Journal of Fisheries Management, 21(4):765–773, November 2001. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kondolf:2000:MMH

[KLW00] G. Mathias Kondolf, Eric W. Larsen, and John G. Williams. Measuring and modeling the hydraulic environment for assessing instream flows. North American Journal of Fisheries Management, 20(4):1016–1028, November 2000. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kulp:2000:MER

[KM00] Matt A. Kulp and Stephen E. Moore. Multiple electrofishing removals for eliminating rainbow trout in a small Southern Appalachian stream. North American Journal of Fisheries Management, 20(1):259–266, February 2000. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kulp:2005:CHF

[KM05] Matt A. Kulp and Stephen E. Moore. A case history in fishing regulations in Great Smoky Mountains National Park: 1934– 2004. North American Journal of Fisheries Management, 25 (2):510–524, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keefer:2009:EBS

[KMB^{+}09] Matthew L. Keefer, Mary L. Moser, Charles T. Boggs, William R. Daigle, and Christopher A. Peery. Effects of body size and river environment on the upstream migration of adult Pacific lampreys. North American Journal of Fisheries Management, 29(5):1214–1224, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Killgore:2001:EPI

[KMCM01] K. Jack Killgore, Steve T. Maynord, Matthew D. Chan, and Raymond P. Morgan II. Evaluation of propeller-induced mortality on early life stages of selected fish species. North American Journal of Fisheries Management, 21(4):947–955, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kaufman:2009:RCP

[KMG09] Scott D. Kaufman, George E. Morgan, and John M. Gunn. The role of ciscoes as prey in the trophy growth potential of walleyes. North American Journal of Fisheries Management, 29(2):468–477, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kerkvliet:2002:EVG

[KNL02] Joe Kerkvliet, Clifford Nowell, and Scott Lowe. The economic value of the Greater Yellowstone’s blue-ribbon fishery. North American Journal of Fisheries Management, 22(2):418–424, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kolz:2008:USR

[Kol08] A. Lawrence Kolz. Understanding Smith-root LR-24 equipment specifications. North American Journal of Fisheries Management, 28(5):1563–1567, October 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kappenman:2007:GNC

[KP07] Kevin M. Kappenman and Blaine L. Parker. Ghost nets in the Columbia River: Methods for locating and removing derelict gill nets in a large river and an assessment of impact to white sturgeon. North American Journal of Fisheries Management, 27(3):804–809, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keefer:2004:SSM

[KPJ^{+}04] Matthew L. Keefer, Christopher A. Peery, Michael A. Jepson, Kenneth R. Tolotti, Theodore C. Bjornn, and Lowell C. Stuehrenberg. Stock-specific migration timing of adult spring–summer Chinook salmon in the Columbia River Basin. North American Journal of Fisheries Management, 24(4): 1145–1162, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keefe:2009:CTM

[KPL09] Donald G. Keefe, Robert C. Perry, and J. Glenn Luther. A comparison of two methodologies for estimating brook trout catch and harvest rates using incomplete and complete fishing trips. North American Journal of Fisheries Management, 29 (4):1058–1064, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keefer:2004:RRI

[KPRB04] M. L. Keefer, C. A. Peery, R. R. Ringe, and T. C. Bjornn. Regurgitation rates of intragastric radio transmitters by adult Chinook salmon and steelhead during upstream migration in the Columbia and Snake Rivers. North American Journal of Fisheries Management, 24(1):47–54, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koch:2007:TPF

[KQ07] Jeff D. Koch and Michael C. Quist. A technique for preparing fin rays and spines for age and growth analysis. North American Journal of Fisheries Management, 27(3):782–784, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koch:2009:PHS

[KQH09] Jeff D. Koch, Michael C. Quist, and Kirk A. Hansen. Precision of hard structures used to estimate age of bowfin in the Upper Mississippi River. North American Journal of Fisheries Management, 29(2):506–511, April 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koch:2009:ECH

[KQP^{+}09] Jeff D. Koch, Michael C. Quist, Clay L. Pierce, Kirk A. Hansen, and Michael J. Steuck. Effects of commercial harvest on shovelnose sturgeon populations in the Upper Mississippi River. North American Journal of Fisheries Management, 29 (1):84–100, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kershner:2004:ASH

[KRB^{+}04] Jeffrey L. Kershner, Brett B. Roper, Nicolaas Bouwes, Richard Henderson, and Eric Archer. An analysis of stream habitat conditions in reference and managed watersheds on some Federal lands within the Columbia River Basin. North American Journal of Fisheries Management, 24(4):1363– 1375, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kocovsky:2008:LLH

[KRDC08] Patrick M. Kocovsky, Robert M. Ross, David S. Dropkin, and John M. Campbell. Linking landscapes and habitat suitability scores for diadromous fish restoration in the Susquehanna River Basin. North American Journal of Fisheries Management, 28(3):906–918, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kieffer:2002:PSA

[KRK^{+}02] James D. Kieffer, Andrew M. Rossiter, Christine A. Kieffer, Kevin Davidson, and Bruce L. Tufts. Physiology and survival of Atlantic salmon following exhaustive exercise in hard and softer water: Implications for the catch-and-release sport fishery. North American Journal of Fisheries Management, 22(1):132–144, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keim:2002:PAHa

[KS02] Richard F. Keim and Arne E. Skaugset. Physical aquatic habitat. I. Errors associated with measurement and estimation of residual pool volumes. North American Journal of Fisheries Management, 22(1):145–150, February 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keim:2002:PAHb

[KSB02] Richard F. Keim, Arne E. Skaugset, and Douglas S. Bateman. Physical aquatic habitat. II. Pools and cover affected by large woody debris in three Western Oregon streams. North American Journal of Fisheries Management, 22(1):151–164, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Karlsson:2008:GIH

[KSB^{+}08] S. Karlsson, E. Saillant, B. W. Bumguardner, R. R. Vega, and J. R. Gold. Genetic identification of hatchery-released red drum in Texas bays and estuaries. North American Journal of Fisheries Management, 28(5):1294–1304, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Killen:2003:PRW

[KSM^{+}03] Shaun S. Killen, Cory D. Suski, Michael B. Morrissey, Peter Dyment, Marosh Furimsky, and Bruce L. Tufts. Physiological responses of walleyes to live-release angling tournaments. North American Journal of Fisheries Management, 23(4):1238–1246, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kehmeier:2007:RFM

[KVMM07] Jon W. Kehmeier, Richard A. Valdez, C. Nicolas Medley, and Orrin B. Myers. Relationship of fish mesohabitat to flow in a sand-bed southwestern river. North American Journal of Fisheries Management, 27(3):750–764, August 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keefe:2000:IMS

[KWA00] Mary Louise Keefe, Timothy A. Whitesel, and Peter Angelone. Induced mortality and sublethal injuries in embryonic brook trout from pulsed DC electroshocking. North American Journal of Fisheries Management, 20(2):320–327, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kwak:2006:AGM

[KWP06] Thomas J. Kwak, D. Scott Waters, and William E. Pine III. Age, growth, and mortality of introduced flathead catfish in Atlantic rivers and a review of other populations. North American Journal of Fisheries Management, 26(1): 73–87, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

LaVigne:2002:IPS

[LaV02] Henry R. LaVigne. An improved portable surgical table for the field and laboratory. North American Journal of Fisheries Management, 22(2):571–572, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leitner:2008:PDB

[LB08] Jean Leitner and James Bulak. Performance differences between two endemic stocks of largemouth bass in South Carolina. North American Journal of Fisheries Management, 28 (2):516–522, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leavy:2009:RAS

[LB09] Tracy R. Leavy and Timothy H. Bonner. Relationships among swimming ability, current velocity association, and morphology for freshwater lotic fishes. North American Journal of Fisheries Management, 29(1):72–83, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leber:2005:OCE

[LCL05] Kenneth M. Leber, Robert N. Cantrell, and Ping Sun Leung. Optimizing cost-effectiveness of size at release in stock enhancement programs. North American Journal of Fisheries Management, 25(4):1596–1608, November 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lapointe:2006:CMS

[LCM06a] Nicolas W. R. Lapointe, Lynda D. Corkum, and Nicholas E. Mandrak. A comparison of methods for sampling fish diversity in shallow offshore waters of large rivers. North American Journal of Fisheries Management, 26(3):503–513, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lapointe:2006:PSB

[LCM06b] Nicolas W. R. Lapointe, Lynda D. Corkum, and Nicholas E. Mandrak. Point sampling by boat electrofishing: a test of the effort required to assess fish communities. North American Journal of Fisheries Management, 26(4):793–799, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lachance:2004:NTM

[LD04] Stephanie Lachance and Maryse Dubé. A new tool for measuring sediment accumulation with minimal loss of fines. North American Journal of Fisheries Management, 24(1): 303–310, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Larson:2008:EQS

[LDMW08] Eric R. Larson, Robert J. DiStefano, Daniel D. Magoulick, and Jacob T. Westhoff. Efficiency of a quadrat sampling technique for estimating riffle-dwelling crayfish density. North American Journal of Fisheries Management, 28(4):1036– 1043, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lazur:2006:ATR

[LEJ06] Andrew Lazur, Steve Early, and John M. Jacobs. Acute toxicity of 5% rotenone to northern snakeheads. North American Journal of Fisheries Management, 26(3):628–630, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Long:2001:PBL

[LF01] James M. Long and William L. Fisher. Precision and bias of largemouth, smallmouth, and spotted bass ages estimated from scales, whole otoliths, and sectioned otoliths. North American Journal of Fisheries Management, 21(3):636–645, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Long:2005:DAB

[LF05] James M. Long and William L. Fisher. Distribution and abundance of black bass in Skiatook Lake, Oklahoma, after introduction of smallmouth bass and a liberalized harvest regulation on spotted bass. North American Journal of Fisheries Management, 25(1):49–56, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Link:2002:EIB

[LGA02] Jason S. Link, Lance P. Garrison, and Frank P. Almeida. Ecological interactions between elasmobranchs and groundfish species on the Northeastern U.S. Continental Shelf. I. Evaluating predation. North American Journal of Fisheries Management, 22(2):550–562, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Layzer:2004:CWT

[LH04a] James B. Layzer and Jackie R. Heinricher. Coded wire tag retention in ebonyshell mussels Fusconaia ebena. North American Journal of Fisheries Management, 24(1):228–230, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lindstrom:2004:IPA

[LH04b] Jason W. Lindstrom and Wayne A. Hubert. Ice processes affect habitat use and movements of adult cutthroat trout and brook trout in a Wyoming foothills stream. North American Journal of Fisheries Management, 24(4):1341–1352, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Linton:2007:DRL

[LHSS07] Brian C. Linton, Michael J. Hansen, Stephen T. Schram, and Shawn P. Sitar. Dynamics of a recovering lake trout population in Eastern Wisconsin waters of Lake Superior, 1980– 2001. North American Journal of Fisheries Management, 27 (3):940–954, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Li:2000:OHM

[Li00] Eric A. L. Li. Optimum harvesting with marine reserves. North American Journal of Fisheries Management, 20(4): 882–896, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Link:2009:EPC

[LI09] Jason S. Link and Josef S. Idoine. Estimates of predator consumption of the northern shrimp Pandalus borealis with implications for estimates of population biomass in the Gulf of Maine. North American Journal of Fisheries Management, 29(6):1567–1583, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lindsay:2000:LAR

[LKS00] Robert B. Lindsay, Kenneth R. Kenaston, and R. Kirk Schroeder. Low adult return of juvenile steelhead treated with 17α-methyltestosterone to produce sterility. North American Journal of Fisheries Management, 20(3):575–583, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lamansky:2009:UHN

[LKYM09] James A. Lamansky, Jr., Ernest R. Keeley, Michael K. Young, and Kevin A. Meyer. The use of hoop nets seeded with mature brook trout to capture conspecifics. North American Journal of Fisheries Management, 29(1):10–17, February 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

LaBay:2006:EAA

[LL06] Steven R. LaBay and Thomas E. Lauer. An evaluation of the accuracy of age estimation methods for Southern Lake Michigan alewives. North American Journal of Fisheries Management, 26(3):571–579, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lin:2008:SSC

[LL08] Pei-Chun Lin and Kung-Yu Lin. Supplier selection criteria for dried striped mullet Roe processors. North American Journal of Fisheries Management, 28(1):165–175, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Littrell:2007:SIG

[LLCBF07] Bradley M. Littrell, Dijar J. Lutz-Carrillo, Timothy H. Bonner, and Loraine T. Fries. Status of an introgressed guadalupe bass population in a Central Texas stream. North American Journal of Fisheries Management, 27(3):785–791, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lovell:2002:PAM

[LM02] Richard G. Lovell and Michael J. Maceina. Population assessment and minimum length limit evaluations for white bass in four Alabama reservoirs. North American Journal of Fisheries Management, 22(2):609–619, May 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lester:2003:BSA

[LMA^{+}03] Nigel P. Lester, Terry R. Marshall, Kim Armstrong, Warren I. Dunlop, and Bev Ritchie. A broad-scale approach to management of Ontario’s recreational fisheries. North American Journal of Fisheries Management, 23(4):1312–1328, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Logsdon:2007:UUD

[Log07] Dale E. Logsdon. Use of unsectioned dorsal spines for estimating walleye ages. North American Journal of Fisheries Management, 27(4):1112–1118, November 2007. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Loomis:2006:USD

[Loo06] John Loomis. Use of survey data to estimate economic value and regional economic effects of fishery improvements. North American Journal of Fisheries Management, 26(2):301–307, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Logsdon:2004:OMN

[LPB04] Dale E. Logsdon, Bruce J. Pittman, and Gary C. Barnard. Oxytetracycline marking of newly hatched walleye fry. North American Journal of Fisheries Management, 24(3):1071– 1077, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Liao:2004:CDA

[LPL04] Hongsheng Liao, Clay L. Pierce, and Joe G. Larscheid. Consumption dynamics of the adult piscivorous fish community in Spirit Lake, Iowa. North American Journal of Fisheries Management, 24(3):890–902, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Liao:2005:EME

[LPL05] Hongsheng Liao, Clay L. Pierce, and Joe G. Larscheid. An empirical model for estimating annual consumption by freshwater fish populations. North American Journal of Fisheries Management, 25(2):525–532, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lockwood:2001:SAL

[LPO01] Roger N. Lockwood, James Peck, and Jack Oelfke. Survey of angling in Lake Superior waters at Isle Royale National Park, 1998. North American Journal of Fisheries Management, 21 (3):471–481, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Litvan:2008:FAW

[LPSL08a] Mary E. Litvan, Clay L. Pierce, Timothy W. Stewart, and Chris J. Larson. Fish assemblages in a Western Iowa stream modified by grade control structures. North American Journal of Fisheries Management, 28(5):1398–1413, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Litvan:2008:FPW

[LPSL08b] Mary E. Litvan, Clay L. Pierce, Timothy W. Stewart, and Chris J. Larson. Fish passage in a Western Iowa stream modified by grade control structures. North American Journal of Fisheries Management, 28(5):1384–1397, October 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Latour:2001:EFN

[LPWH01] Robert J. Latour, Kenneth H. Pollock, Charles A. Wenner, and John M. Hoenig. Estimates of fishing and natural mortality for subadult red drum in South Carolina waters. North American Journal of Fisheries Management, 21(4):733–744, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lockwood:2005:CIA

[LR05] Roger N. Lockwood and Gerald P. Rakoczy. Comparison of interval and aerial count methods for estimating Fisher boating effort. North American Journal of Fisheries Management, 25(4):1331–1340, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Liermann:2008:MSM

[LR08] Martin Liermann and Phil Roni. More sites or more years? Optimal study design for monitoring fish response to watershed restoration. North American Journal of Fisheries Management, 28(3):935–943, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leaf:2008:EUS

[LRBJ08] Robert T. Leaf, Laura Rogers-Bennett, and Yan Jiao. Exploring the use of a size-based egg-per-recruit model for the red abalone fishery in California. North American Journal of Fisheries Management, 28(5):1638–1647, October 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ledgerwood:2004:STD

[LRD^{+}04] Richard D. Ledgerwood, Brad A. Ryan, Earl M. Dawley, Edward P. Nunnallee, and John W. Ferguson. A surface trawl to detect migrating juvenile salmonids tagged with passive integrated transponder tags. North American Journal of Fisheries Management, 24(2):440–451, May 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Link:2008:EIS

[LS08] Jason S. Link and Katherine Sosebee. Estimates and implications of skate consumption in the Northeast U.S. Continental Shelf Ecosystem. North American Journal of Fisheries Management, 28(3):649–662, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lindsay:2004:HMA

[LSK^{+}04] Robert B. Lindsay, R. Kirk Schroeder, Kenneth R. Kenaston, Robert N. Toman, and Mary A. Buckman. Hooking mortality by anatomical location and its use in estimating mortality of spring Chinook salmon caught and released in a river sport fishery. North American Journal of Fisheries Management, 24(2):367–378, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lauer:2005:YPL

[LSK^{+}05] Thomas E. Lauer, Steven M. Shroyer, John M. Kilpatrick, Thomas S. McComish, and Paul J. Allen. Yellow perch length–fecundity and length–egg size relationships in Indiana waters of Lake Michigan. North American Journal of Fisheries Management, 25(3):791–796, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Letcher:2001:EDS

[LT01] Benjamin H. Letcher and Timothy D. Terrick. Effects of developmental stage at stocking on growth and survival of Atlantic salmon fry. North American Journal of Fisheries Management, 21(1):102–110, February 2001. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lucchesi:2002:ECS

[Luc02] David O. Lucchesi. Evaluating the contribution of stocked walleye fry and fingerlings to South Dakota walleye populations through mass marking with oxytetracycline. North American Journal of Fisheries Management, 22(3):985–994, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lomnicky:2007:DNA

[LWHP07] Gregg A. Lomnicky, Thomas R. Whittier, Robert M. Hughes, and David V. Peck. Distribution of nonnative aquatic vertebrates in Western U.S. streams and rivers. North American Journal of Fisheries Management, 27(4):1082–1093, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lyons:2009:DCC

[LZS^{+}09] John Lyons, Troy Zorn, Jana Stewart, Paul Seelbach, Kevin Wehrly, and Lizhu Wang. Defining and characterizing coolwater streams and their fish assemblages in Michigan and Wisconsin, USA. North American Journal of Fisheries Management, 29(4):1130–1151, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lin:2008:AFL

[LZW^{+}08] Binbin Lin, Ziping Zhang, Yilei Wang, Kenneth P. Currens, Adrian Spidle, Yuji Yamazaki, and David A. Close. Amplified fragment length polymorphism assessment of genetic diversity in Pacific lampreys. North American Journal of Fisheries Management, 28(4):1182–1193, August 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Margenau:2000:EIM

[MA00a] Terry L. Margenau and Steven P. AveLallemant. Effects of a 40-inch minimum length limit on muskellunge in Wisconsin. North American Journal of Fisheries Management, 20 (4):986–993, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:2000:ULL

[MA00b] L. E. Miranda and M. S. Allen. Use of length limits to reduce variability in crappie fisheries. North American Journal of Fisheries Management, 20(3):752–758, August 2000. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Myers:2000:SFB

[MAC00] Randall A. Myers, Micheal S. Allen, and Douglas E. Colle. Survival of fingerling black crappies tagged with microwire tags stocked in a Florida lake. North American Journal of Fisheries Management, 20(2):534–537, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

MacCall:2002:UKB

[Mac02] Alec D. MacCall. Use of known-biomass production models to determine productivity of West Coast groundfish stocks. North American Journal of Fisheries Management, 22(2): 272–279, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:2003:VIH

[Mac03] Michael J. Maceina. Verification of the influence of hydrologic factors on crappie recruitment in Alabama reservoirs. North American Journal of Fisheries Management, 23(2):470–480, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:2004:VRC

[Mac04] Michael J. Maceina. Verifying residuals from catch curves to detect recruitment variation in largemouth bass and crappies. North American Journal of Fisheries Management, 24 (1):231–236, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:2007:UPN

[Mac07] Michael J. Maceina. Use of piecewise nonlinear models to estimate variable size-related mortality rates. North American Journal of Fisheries Management, 27(3):971–977, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madsen:2008:FWA

[Mad08] Eric R. Madsen. Forecasting walleye abundance in Northern Wisconsin lakes. North American Journal of Fisheries Management, 28(2):481–499, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martin:2001:IGF

[Mar01] Douglas J. Martin. The influence of geomorphic factors and geographic region on large woody debris loading and fish habitat in Alaska coastal streams. North American Journal of Fisheries Management, 21(3):429–440, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mauk:2008:EOM

[Mau08] Robert Mauk. Efficacy of oxytetracycline marking of fingerling palmetto bass in hard water. North American Journal of Fisheries Management, 28(1):258–262, February 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:2001:CBB

[MB01] Michael J. Maceina and David R. Bayne. Changes in the black bass community and fishery with oligotrophication in West Point Reservoir, Georgia. North American Journal of Fisheries Management, 21(4):745–755, November 2001. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McHugh:2005:CVM

[MB05a] Peter McHugh and Phaedra Budy. A comparison of visual and measurement-based techniques for quantifying cobble embeddedness and fine-sediment levels in salmonid-bearing streams. North American Journal of Fisheries Management, 25(4):1208–1214, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meeuwig:2005:MAP

[MB05b] Michael H. Meeuwig and Jennifer M. Bayer. Morphology and aging precision of statoliths from larvae of Columbia River Basin lampreys. North American Journal of Fisheries Management, 25(1):38–48, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mohler:2008:RCW

[MB08] Jerre W. Mohler and Kelly M. Bradley. Removal of calcein in wastewater produced from the batch marking of fish. North American Journal of Fisheries Management, 28(4):

1177–1181, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mitchell:2009:UIW

[MB09] Andrew J. Mitchell and Thomas M. Brandt. Use of ice-water and salt treatments to eliminate an exotic snail, the red-rim melania, from small immersible fisheries equipment. North American Journal of Fisheries Management, 29(3):823–828, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:2002:UEE

[MBD02] L. E. Miranda, R. E. Brock, and B. S. Dorr. Uncertainty of exploitation estimates made from tag returns. North American Journal of Fisheries Management, 22(4):1358– 1363, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Myers:2002:IBP

[MBHK02] Ransom A. Myers, N. J. Barrowman, Ray Hilborn, and Daniel G. Kehler. Inferring Bayesian priors with limited direct data: Applications to risk analysis. North American Journal of Fisheries Management, 22(2):351–364, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morhardt:2000:PSR

[MBHM00] J. Emil Morhardt, David Bishir, Cristin I. Handlin, and Samuel D. Mulder. A portable system for reading large passive integrated transponder tags from wild trout. North American Journal of Fisheries Management, 20(1):276–283, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muhlfeld:2001:FWH

[MBM01a] Clint C. Muhlfeld, David H. Bennett, and Brian Marotz. Fall and winter habitat use and movement by Columbia River redband trout in a small stream in Montana. North American Journal of Fisheries Management, 21(1):170–177, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muhlfeld:2001:SHU

[MBM01b] Clint C. Muhlfeld, David H. Bennett, and Brian Marotz. Summer habitat use by Columbia River redband trout in the Kootenai River drainage, Montana. North American Journal of Fisheries Management, 21(1):223–235, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mangan:2005:CPG

[MBS05] Matthew T. Mangan, Michael L. Brown, and Todd R. St. Sauver. A comparison of passive gears for selective yellow perch harvest. North American Journal of Fisheries Management, 25(3):1067–1072, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muhlfeld:2008:UBM

[MBS^{+}08] Clint C. Muhlfeld, David H. Bennett, R. Kirk Steinhorst, Brian Marotz, and Matthew Boyer. Using bioenergetics modeling to estimate consumption of native juvenile salmonids by nonnative northern pike in the Upper Flathead River System, Montana. North American Journal of Fisheries Management, 28(3):636–648, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McInerny:2000:EST

[MC00] Michael C. McInerny and Timothy K. Cross. Effects of sampling time, intraspecific density, and environmental variables on electrofishing catch per effort of largemouth bass in Minnesota lakes. North American Journal of Fisheries Management, 20(2):328–336, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murphy:2001:CAS

[MC01] Michael D. Murphy and Roy E. Crabtree. Changes in the age structure of nearshore adult red drum off West-Central Florida related to recruitment and fishing mortality. North American Journal of Fisheries Management, 21(3):671–678, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mallison:2004:AAR

[MC04a] Craig T. Mallison and Charles E. Cichra. Accuracy of anglerreported harvest in roving creel surveys. North American Journal of Fisheries Management, 24(3):880–889, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marine:2004:EHW

[MC04b] Keith R. Marine and Joseph J. Cech, Jr. Effects of high water temperature on growth, smoltification, and predator avoidance in juvenile Sacramento River Chinook salmon. North American Journal of Fisheries Management, 24(1):198–210, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McInerny:2005:EMR

[MC05] Michael C. McInerny and Timothy K. Cross. An evaluation of mark–recapture estimates of black crappie population size determined by trap-netting during fall and spring in Minnesota lakes. North American Journal of Fisheries Management, 25 (2):475–490, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McInerny:2006:FAT

[MC06] Michael C. McInerny and Timothy K. Cross. Factors affecting trap-net catchability of black crappies in natural Minnesota lakes. North American Journal of Fisheries Management, 26 (3):652–664, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mitchell:2008:SFS

[MC08] Andrew J. Mitchell and Rebecca A. Cole. Survival of the faucet snail after chemical disinfection, pH extremes, and heated water bath treatments. North American Journal of Fisheries Management, 28(5):1597–1600, October 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCleave:2001:SID

[McC01] James D. McCleave. Simulation of the impact of dams and fishing weirs on reproductive potential of silver-phase American eels in the Kennebec River Basin, Maine. North American Journal of Fisheries Management, 21(3):592–605, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McHugh:2007:CVM

[McH07] Peter McHugh. A comparison of visual and measurementbased techniques for quantifying cobble embeddedness and fine-sediment levels in salmonid-bearing streams: Response to comment. North American Journal of Fisheries Management, 27(3):1027–1028, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McGovern:2002:TSD

[MCPM02] John C. McGovern, Mark R. Collins, Oleg Pashuk, and H. Scott Meister. Temporal and spatial differences in life history parameters of black sea bass in the Southeastern United States. North American Journal of Fisheries Management, 22(4):1151–1163, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mesing:2008:SAF

[MCS^{+}08] Charles L. Mesing, Richard L. Cailteux, P. Andrew Strickland, Eric A. Long, and Mark W. Rogers. Stocking of advanced-fingerling largemouth bass to supplement yearclasses in Lake Talquin, Florida. North American Journal of Fisheries Management, 28(6):1762–1774, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:2000:SSC

[MD00] L. E. Miranda and B. S. Dorr. Size selectivity of crappie angling. North American Journal of Fisheries Management, 20(3):706–710, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Milliken:2004:ELM

[MD04a] Henry O. Milliken and Joseph T. DeAlteris. Evaluation of a large-mesh panel to reduce the flatfish bycatch in the smallmesh bottom trawls used in the New England silver hake fishery. North American Journal of Fisheries Management, 24(1):20–32, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:2004:EPR

[MD04b] L. E. Miranda and C. R. Dolan. Electrofishing power requirements in relation to duty cycle. North American Journal of Fisheries Management, 24(1):55–62, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moser:2000:IPE

[MDH00] Mary L. Moser, Ann Marie Darazsdi, and J. Robin Hall. Improving passage efficiency of adult American shad at lowelevation dams with navigation locks. North American Journal of Fisheries Management, 20(2):376–385, May 2000. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Monnot:2008:IBS

[MDHK08] Lauri Monnot, Jason B. Dunham, Tammy Hoem, and Peter Koetsier. Influences of body size and environmental factors on autumn downstream migration of bull trout in the Boise River, Idaho. North American Journal of Fisheries Management, 28(1):231–240, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:2006:ESS

[MDKM06] Anna-Maria Mueller, Don J. Degan, Robert Kieser, and Tim Mulligan. Estimating sockeye salmon smolt flux and abundance with side-looking sonar. North American Journal of Fisheries Management, 26(3):523–534, August 2006. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madenjian:2004:SLT

[MDM^{+}04] Charles P. Madenjian, Timothy J. Desorcie, Jerry R. Mc-Clain, Aaron P. Woldt, Jeffrey D. Holuszko, and Charles A. Bowen II. Status of lake trout rehabilitation on Six Fathom Bank and Yankee Reef in Lake Huron. North American Journal of Fisheries Management, 24(3):1003–1016, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meade:2004:FIR

[Mea04] Richard Meade. Fish and invertebrate recolonization in a Missouri prairie stream after an acute pollution event.

North American Journal of Fisheries Management, 24(1): 7–19, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madenjian:2008:LTP

[MED08] Charles P. Madenjian, Mark P. Ebener, and Timothy J. Desorcie. Lake trout population dynamics at drummond Island refuge in Lake Huron: Implications for future rehabilitation. North American Journal of Fisheries Management, 28(4): 979–992, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muir:2008:CSO

[MEHJ08] Andrew M. Muir, Mark P. Ebener, Ji X. He, and James E. Johnson. A comparison of the scale and otolith methods of age estimation for lake whitefish in Lake Huron. North American Journal of Fisheries Management, 28(3):625–635, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meka:2004:IHT

[Mek04] Julie M. Meka. The influence of hook type, angler experience, and fish size on injury rates and the duration of capture in an Alaskan catch-and-release rainbow trout fishery. North American Journal of Fisheries Management, 24(4): 1309–1321, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2009:EFR

[MEL09] Kevin A. Meyer, F. Steven Elle, and James A. Lamansky, Jr. Environmental factors related to the distribution, abundance, and life history characteristics of mountain whitefish in Idaho. North American Journal of Fisheries Management, 29(3):753–767, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marshall:2008:PEA

[MFP^{+}08] David W. Marshall, Andrew H. Fayram, John C. Panuska, James Baumann, and Joseph Hennessy. Positive effects of agricultural land use changes on coldwater fish communities in Southwest Wisconsin streams. North American Journal of Fisheries Management, 28(3):944–953, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Milliken:2009:SDS

[MFRS09] Henry O. Milliken, Marianne Farrington, Tom Rudolph, and Melissa Sanderson. Survival of discarded sublegal Atlantic cod in the Northwest Atlantic demersal longline fishery. North American Journal of Fisheries Management, 29(4):985–995, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

MacInnis:2008:LWD

[MFT08] Charles MacInnis, Trevor A. Floyd, and Barry R. Taylor. Large woody debris structures and their influence on Atlantic salmon spawning in a stream in Nova Scotia, Canada. North American Journal of Fisheries Management, 28(3):781–791, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Macdonald:2009:USO

[MFW^{+}09] R. W. Macdonald, V. Forsland, R. E. Withler, D. A. Patterson, and A. Demsky. The use of stable oxygen isotope (δ 18O) composition in sockeye salmon body fluid to determine whether a fish has been caught in freshwater. North American Journal of Fisheries Management, 29(3):560–569, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martinez-Garmendia:2000:EHA

[MGAC00] Josué Martínez-Garmendia, James L. Anderson, and Michael T. Carroll. Effect of harvesting alternatives on the quality of U.S. North Atlantic bluefin tuna. North American Journal of Fisheries Management, 20(4):908–922, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matala:2004:MVI

[MGGL04] Andrew P. Matala, A. K. Gray, A. J. Gharrett, and M. S. Love. Microsatellite variation indicates population genetic structure of bocaccio. North American Journal of Fisheries Management, 24(4):1189–1202, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Margenau:2003:ACH

[MGH03a] Terry L. Margenau, Stephen J. Gilbert, and Gene R. Hatzenbeler. Angler catch and harvest of northern pike in Northern Wisconsin lakes. North American Journal of Fisheries Management, 23(2):307–312, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muhlfeld:2003:WDH

[MGH^{+}03b] Clint C. Muhlfeld, Steve Glutting, Rick Hunt, Durae Daniels, and Brian Marotz. Winter diel habitat use and movement by subadult bull trout in the Upper Flathead River, Montana. North American Journal of Fisheries Management, 23 (1):163–171, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mullner:2005:LSW

[MH05] Scott A. Mullner and Wayne A. Hubert. Low summer water temperatures influence occurrence of naturalized salmonids across a mountain watershed. North American Journal of Fisheries Management, 25(3):1034–1040, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:2008:UPR

[MH08a] Michael J. Maceina and Ryan H. Hunter. Use of piecewise regression models to estimate changing relationships in fisheries: Response to comment. North American Journal of Fisheries Management, 28(3):847–848, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McBride:2008:RUA

[MH08b] Richard S. McBride and Jay C. Holder. A review and updated assessment of Florida’s anadromous shads: American shad and hickory shad. North American Journal of Fisheries Management, 28(6):1668–1686, December 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mitchell:2007:ECT

[MHB07] Andrew J. Mitchell, Melissa S. Hobbs, and Thomas M. Brandt. The effect of chemical treatments on red-rim melania melanoides tuberculata, an exotic aquatic snail that serves as a vector of trematodes to fish and other species in the USA. North American Journal of Fisheries Management, 27 (4):1287–1293, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miyakoshi:2003:SDS

[MHF^{+}03] Yasuyuki Miyakoshi, Hirofumi Hayano, Makoto Fujiwara, Mitsuhiro Nagata, and James R. Irvine. Size-dependent smolt yield and overwinter survival of hatchery-reared masu salmon released in fall. North American Journal of Fisheries Management, 23(2):264–269, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marshall:2009:AFC

[MHM09] Matthew D. Marshall, Michael P. Holley, and Michael J. Maceina. Assessment of the flathead catfish population in a lightly exploited fishery in Lake Wilson, Alabama. North American Journal of Fisheries Management, 29(4):869–875, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maki:2001:EPM

[MHO01] Kristin L. Maki, John M. Hoenig, and John E. Olney. Estimating proportion mature at age when immature fish are unavailable for study, with application to American shad in the York River, Virginia. North American Journal of Fisheries Management, 21(4):703–716, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maki:2002:IMD

[MHO02] Kristin L. Maki, John M. Hoenig, and John E. Olney. Interpreting maturation data for American shad in the presence of fishing mortality: a look at historical data from the York River, Virginia. North American Journal of Fisheries Management, 22(4):1209–1217, November 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maki:2006:CHC

[MHOH06] Kristin L. Maki, John M. Hoenig, John E. Olney, and Dennis M. Heisey. Comparing historical catch rates of American shad in multifilament and monofilament nets: a step toward setting restoration targets for Virginia stocks. North American Journal of Fisheries Management, 26(2):282–288, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McQueen:2007:CAS

[MHRR07] Donald J. McQueen, Kim D. Hyatt, D. Paul Rankin, and Charles J. Ramcharan. Changes in algal species composition affect juvenile sockeye salmon production at Woss Lake, British Columbia: a lake fertilization and food Web analysis. North American Journal of Fisheries Management, 27 (2):369–386, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McLellan:2008:ERO

[MHS08] Holly J. McLellan, Stephen G. Hayes, and Allan T. Scholz. Effects of reservoir operations on hatchery coastal rainbow trout in Lake Roosevelt, Washington. North American Journal of Fisheries Management, 28(5):1201–1213, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mullner:2000:SAD

[MHW00] Scott A. Mullner, Wayne A. Hubert, and Thomas A. Wesche. Snorkeling as an alternative to depletion electrofishing for estimating abundance and length-class frequencies of trout in small streams: Response to comment. North American Journal of Fisheries Management, 20(3):839–840, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:2005:DPS

[Mic05] Paul H. Michaletz. Does pectoral spine extraction cause mortality to channel catfish? North American Journal of Fisheries Management, 25(2):533–535, May 2005. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:2009:VRC

[Mic09] Paul H. Michaletz. Variable responses of channel catfish populations to stocking rate: Density-dependent and lake productivity effects. North American Journal of Fisheries Management, 29(1):177–188, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:2002:ESB

[Mir02] L. E. Miranda. Establishing size-based mortality caps. North American Journal of Fisheries Management, 22(2):433–440, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:2005:RBE

[Mir05] L. E. Miranda. Refining boat electrofishing equipment to improve consistency and reduce harm to fish. North American Journal of Fisheries Management, 25(2):609–618, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mitzner:2002:EWF

[Mit02] Larry Mitzner. Effectiveness of walleye fry and fingerling stocking in Rathbun Lake, Iowa, 1990–1999. North American Journal of Fisheries Management, 22(3):1007–1013, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCleave:2004:MDN

[MJ04] James D. McCleave and Donald J. Jellyman. Male dominance in the New Zealand longfin eel population of a New Zealand River: Probable causes and implications for management. North American Journal of Fisheries Management, 24 (2):490–505, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKenna:2005:JRT

[MJ05] James E. McKenna, Jr. and James H. Johnson. Juvenile rainbow trout production in New York tributaries of Lake Ontario: Implications for Atlantic salmon restoration. North American Journal of Fisheries Management, 25(1):391–403, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mosley:2007:FPJ

[MJ07] Diarra L. Mosley and Cecil A. Jennings. Flow preferences for juvenile robust redhorses in an experimental mesocosm: Implications for developing sampling protocols. North American Journal of Fisheries Management, 27(4):1383–1392, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martinez:2009:EPO

[MK09] Patrick J. Martinez and A. Lawrence Kolz. Evaluating the power output of the Smith-root GPP 5.0 electrofisher to promote electrofishing fleet standardization. North American Journal of Fisheries Management, 29(3):570–575, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marsh:2005:RMS

[MKP05] Paul C. Marsh, Brian R. Kesner, and Carol A. Pacey. Repatriation as a management strategy to conserve a critically imperiled fish species. North American Journal of Fisheries Management, 25(2):547–556, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moerke:2003:RFC

[ML03] Ashley H. Moerke and Gary A. Lamberti. Responses in fish community structure to restoration of two Indiana streams. North American Journal of Fisheries Management, 23(3): 748–759, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

May:2004:RAC

[ML04] Christine L. May and Danny C. Lee. The relationships among in-channel sediment storage, pool depth, and summer survival of juvenile salmonids in Oregon Coast range streams. North American Journal of Fisheries Management, 24(3): 761–774, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Methratta:2006:ABS

[ML06] Elizabeth T. Methratta and Jason S. Link. Associations between surficial sediments and groundfish distributions in the Gulf of Maine–Georges Bank Region. North American Journal of Fisheries Management, 26(2):473–489, May 2006. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2006:EUB

[MLS06] Kevin A. Meyer, James A. Lamansky, Jr., and Daniel J. Schill. Evaluation of an unsuccessful brook trout electrofishing removal project in a small Rocky Mountain stream. North American Journal of Fisheries Management, 26(4):849–860, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKeown:2002:FIR

[MM02] Paul E. McKeown and Stephen R. Mooradian. Factors influencing recruitment of crappies in Chautauqua Lake, New York. North American Journal of Fisheries Management, 22 (4):1385–1392, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muhlfeld:2005:SMH

[MM05] Clint C. Muhlfeld and Brian Marotz. Seasonal movement and habitat use by subadult bull trout in the Upper Flathead River System, Montana. North American Journal of Fisheries Management, 25(3):797–810, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:2006:TJP

[MMB06] Robert P. Mueller, Russell A. Moursund, and Matthew D. Bleich. Tagging juvenile Pacific lamprey with passive integrated transponders: Methodology, short-term mortality, and influence on swimming performance. North American Journal of Fisheries Management, 26(2):361–366, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mohler:2002:PCW

[MMF02] Jerre W. Mohler, Michael J. Millard, and John W. Fletcher. Predation by captive wild brook trout on calcein-marked versus nonmarked Atlantic salmon fry. North American Journal of Fisheries Management, 22(2):223–228, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:2003:FIP

[MMF^{+}03] Gordon A. Mueller, Paul C. Marsh, Dean Foster, Manuel Ulibarri, and Tom Burke. Factors influencing poststocking dispersal of razorback sucker. North American Journal of Fisheries Management, 23(2):270–275, February 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marshall:2009:AGV

[MMH09] Matthew D. Marshall, Michael J. Maceina, and Michael P. Holley. Age and growth variability between sexes of three catfish species in Lake Wilson, Alabama. North American Journal of Fisheries Management, 29(5):1283–1286, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Millard:2005:MAC

[MMKC05] Michael J. Millard, Jerre W. Mohler, Andrew Kahnle, and Amanda Cosman. Mortality associated with catch-and-release angling of striped bass in the Hudson River. North American Journal of Fisheries Management, 25(4):1533– 1541, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Munro:2003:EBM

[MMLL03] Andrew R. Munro, Thomas E. McMahon, Stephen A. Leathe, and George Liknes. Evaluation of batch marking small rainbow trout with coded wire tags. North American Journal of Fisheries Management, 23(2):600–604, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:2006:BFY

[MMLM06] Robert J. Miller, Manon Mallet, Marc Lanteigne, and Robert MacMillan. Bottlenecks to fishery yield for American lobster in a portion of the Gulf of St. Lawrence. North American Journal of Fisheries Management, 26(3):765–776, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mackey:2001:CRT

[MMQ01] Gregory Mackey, Jennifer E. McLean, and Thomas P. Quinn. Comparisons of run timing, spatial distribution, and length of wild and newly established hatchery populations of steelhead in Forks Creek, Washington. North American Journal of Fisheries Management, 21(4):717–724, November 2001. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:2008:CHS

[MMR08] Loren M. Miller, Michael C. McInerny, and Jason Roloff. Crappie hybridization in Southern Minnesota lakes and its effects on growth estimates. North American Journal of Fisheries Management, 28(4):1120–1131, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:2008:IEA

[MMS08] Michael J. Maceina, Matthew D. Marshall, and Steven M. Sammons. Impacts of endothall applications on largemouth bass spawning behavior and reproductive success. North American Journal of Fisheries Management, 28(6):1812– 1817, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMichael:2005:FCS

[MMVL05] Geoffrey A. McMichael, C. A. McKinstry, J. A. Vucelick, and J. A. Lukas. Fall Chinook salmon spawning activity versus daylight and flow in the tailrace of a large hydroelectric dam. North American Journal of Fisheries Management, 25 (2):573–580, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Manning:2005:SES

[MMW^{+}05] David J. Manning, Jonathon A. Mann, Sean K. White, Shawn D. Chase, and Ron C. Benkert. Steelhead emigration in a seasonal impoundment created by an inflatable rubber dam. North American Journal of Fisheries Management, 25 (4):1239–1255, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:2008:CSI

[MMW08] Anna-Maria Mueller, Tim Mulligan, and Peter K. Withler. Classifying sonar images: Can a computer-driven process identify eels? North American Journal of Fisheries Management, 28(6):1876–1886, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:2009:APE

[MNB09] Paul H. Michaletz, Dean M. Nicks, and Earl W. Buckner, Jr. Accuracy and precision of estimates of back-calculated channel catfish lengths and growth increments using pectoral spines and otoliths. North American Journal of Fisheries Management, 29(6):1664–1675, December 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaelson:2001:LBP

[MNS01] Daniel P. Michaelson, John J. Ney, and Trent M. Sutton. Largemouth bass predation on stocked striped bass in Smith Mountain Lake, Virginia. North American Journal of Fisheries Management, 21(2):326–332, May 2001. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madenjian:2008:AEA

[MOB^{+}08] Charles P. Madenjian, Robert O’Gorman, David B. Bunnell, Ray L. Argyle, Edward F. Roseman, David M. Warner, Jason D. Stockwell, and Martin A. Stapanian. Adverse effects of alewives on Laurentian Great Lakes fish communities. North American Journal of Fisheries Management, 28(1):263–282, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mohler:2003:PFM

[Moh03] Jerre W. Mohler. Producing fluorescent marks on Atlantic salmon fin rays and scales with calcein via osmotic induction. North American Journal of Fisheries Management, 23 (4):1108–1113, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maret:2007:EER

[MOH07] Terry R. Maret, Douglas S. Ott, and Alan T. Herlihy. Electrofishing effort required to estimate biotic condition in Southern Idaho rivers. North American Journal of Fisheries Management, 27(3):1041–1052, August 2007. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMichael:2001:UMR

[MP01] Geoffrey A. McMichael and Todd N. Pearsons. Upstream movement of residual hatchery steelhead into areas containing bull trout and cutthroat trout. North American Journal of Fisheries Management, 21(4):943–946, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murie:2002:CTM

[MP02] Debra J. Murie and Daryl C. Parkyn. Comparison of total mortality of white grunt from the headboat fishery on the Gulf Coast of Florida during spawning and postspawning seasons. North American Journal of Fisheries Management, 22 (3):806–814, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

MacKenzie:2004:GEM

[MP04a] Clyde L. MacKenzie, Jr. and Robert Pikanowski. Gear effects on marine habitats: Harvesting northern quahogs in a shallow sandy bed at two levels of intensity with a short rake. North American Journal of Fisheries Management, 24 (4):1221–1227, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Margenau:2004:SAM

[MP04b] Terry L. Margenau and Jordan B. Petchenik. Social aspects of muskellunge management in Wisconsin. North American Journal of Fisheries Management, 24(1):82–93, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Makinster:2008:EUM

[MP08] Andrew S. Makinster and Craig P. Paukert. Effects and utility of minimum length limits and mortality caps for flathead catfish in discrete reaches of a large prairie river. North American Journal of Fisheries Management, 28(1):97–108, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meeuwig:2007:STR

[MPB07] Michael H. Meeuwig, Amy L. Puls, and Jennifer M. Bayer. Survival and tag retention of Pacific lamprey larvae and macrophthalmia marked with coded wire tags. North American Journal of Fisheries Management, 27(1):96–102, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maxwell:2006:BAB

[MPBM06] Marla Rae Maxwell, Randall M. Peterman, Michael J. Bradford, and Erland A. MacIsaac. A Bayesian analysis of biological uncertainty for a whole-lake fertilization experiment for sockeye salmon in Chilko Lake, British Columbia, and implications for the benefit–cost ratio. North American Journal of Fisheries Management, 26(2):418–430, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murdoch:2009:NRC

[MPM09a] Andrew R. Murdoch, Todd N. Pearsons, and Travis W. Maitland. The number of redds constructed per female spring Chinook salmon in the Wenatchee River Basin. North American Journal of Fisheries Management, 29(2):441–446, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murdoch:2009:UCR

[MPM09b] Andrew R. Murdoch, Todd N. Pearsons, and Travis W. Maitland. Use of carcass recovery data in evaluating the spawning distribution and timing of spring Chinook salmon in the Chiwawa River, Washington. North American Journal of Fisheries Management, 29(5):1206–1213, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

MacGregor:2002:DAF

[MPPB02] Brice W. MacGregor, Randall M. Peterman, Brian J. Pyper, and Michael J. Bradford. A decision analysis framework for comparing experimental designs of projects to enhance Pacific salmon. North American Journal of Fisheries Management, 22(2):509–527, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moser:2008:GSN

[MPR08] Mary L. Moser, Howard T. Pennington, and Jeremy M. Roos. Grating size needed to protect adult Pacific lampreys in the Columbia River Basin. North American Journal of Fisheries Management, 28(2):557–562, February 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

MacCall:2002:LTR

[MR02] Alec D. MacCall and Stephen Ralston. Is logarithmic transformation Really the best procedure for estimating stockrecruitment relationships? North American Journal of Fisheries Management, 22(2):339–350, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marsden:2004:DYP

[MR04] J. Ellen Marsden and Steven R. Robillard. Decline of yellow perch in Southwestern Lake Michigan, 1987–1997. North American Journal of Fisheries Management, 24(3):952–966, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morton:2006:FCF

[MR06] Alexandra Morton and Richard D. Routledge. Fulton’s condition factor: Is it a valid measure of sea lice impact on juvenile salmon? North American Journal of Fisheries Management, 26(1):56–62, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moyer:2009:GEC

[MRC09] Gregory R. Moyer, J. D. Rousey, and M. A. Cantrell. Genetic evaluation of a conservation hatchery program for reintroduction of sicklefin redhorse Moxostoma sp. in the Tuckasegee River, North Carolina. North American Journal of Fisheries Management, 29(5):1438–1443, October 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMichael:2005:EFC

[MRJL05] Geoffrey A. McMichael, Cindy L. Rakowski, Brenda B. James, and Joe A. Lukas. Estimated fall Chinook salmon survival to emergence in dewatered redds in a shallow side channel of the Columbia River. North American Journal of Fisheries Management, 25(3):876–884, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morton:2008:SLI

[MRK08] Alexandra Morton, Rick Routledge, and Martin Krkosek. Sea louse infestation in wild juvenile salmon and Pacific herring associated with fish farms off the East-Central Coast of Vancouver Island, British Columbia. North American Journal of Fisheries Management, 28(2):523–532, February 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:2000:ADR

[MRS00] Bruce A. Miller, Jeffrey D. Rodgers, and Mario F. Solazzi. An automated device to release marked juvenile fish for measuring trap efficiency. North American Journal of Fisheries Management, 20(1):284–287, February 2000. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKinney:2001:HAA

[MRSR01] T. McKinney, A. T. Robinson, D. W. Speas, and R. S. Rogers. Health assessment, associated metrics, and nematode parasitism of rainbow trout in the Colorado River below Glen Canyon Dam, Arizona. North American Journal of Fisheries Management, 21(1):62–69, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morton:2005:TPS

[MRW05] Alexandra Morton, Richard D. Routledge, and Rob Williams. Temporal patterns of sea louse infestation on wild Pacific salmon in relation to the fallowing of Atlantic salmon farms. North American Journal of Fisheries Management, 25(3): 811–821, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:2002:SCC

[MS02] Paul H. Michaletz and Kevin P. Sullivan. Sampling channel catfish with tandem hoop nets in small impoundments. North American Journal of Fisheries Management, 22(3):870–878, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matter:2003:CMR

[MS03a] Alicia L. Matter and Benjamin P. Sandford. A comparison of migration rates of radio- and PIT-Tagged adult Snake River Chinook salmon through the Columbia River hydropower system. North American Journal of Fisheries Management, 23 (4):967–973, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:2003:RNP

[MS03b] Loren M. Miller and Wansuk Senanan. A review of northern pike population genetics research and its implications for management. North American Journal of Fisheries Management, 23(2):297–306, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Merz:2004:ESH

[MS04a] Joseph E. Merz and Jose D. Setka. Evaluation of a spawning habitat enhancement site for Chinook salmon in a regulated California River. North American Journal of Fisheries Management, 24(2):397–407, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murchie:2004:RAB

[MS04b] Karen J. Murchie and Karen E. Smokorowski. Relative activity of brook trout and walleyes in response to flow in a regulated river. North American Journal of Fisheries Management, 24(3):1050–1057, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:2005:UOS

[MS05] L. E. Miranda and A. B. Spencer. Understanding the output of a Smith-root GPP electrofisher. North American Journal of Fisheries Management, 25(3):848–852, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maxwell:2007:GRB

[MS07] Suzanne L. Maxwell and April V. Smith. Generating river bottom profiles with a dual-frequency identification sonar (DIDSON). North American Journal of Fisheries Management, 27(4):1294–1309, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2003:LTC

[MSES03] Kevin A. Meyer, Daniel J. Schill, F. Steven Elle, and William C. Schrader. A long-term comparison of Yellowstone cutthroat trout abundance and size structure in their historical range in Idaho. North American Journal of Fisheries Management, 23(1):149–162, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mercado-Silva:2008:CSE

[MSES08] Norman Mercado-Silva and David S. Escandón-Sandoval. A comparison of seining and electrofishing for fish community bioassessment in a Mexican Atlantic slope montane river. North American Journal of Fisheries Management, 28(6): 1725–1732, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miyakoshi:2009:IRF

[MSF^{+}09] Yasuyuki Miyakoshi, Yoshitaka Sasaki, Makoto Fujiwara, Keiko Tanaka, Naokazu Matsueda, James R. Irvine, and Shuichi Kitada. Implications of recreational fishing on juvenile masu salmon stocked in a Hokkaido river. North American Journal of Fisheries Management, 29(1):33–39, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2005:SSC

[MSK05] Carolyn B. Meyer, Michael D. Sparkman, and Bernard A. Klatte. Sand seals in Coho salmon redds: Do they improve egg survival? North American Journal of Fisheries Management, 25(1):105–121, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muir:2008:EAE

[MSP^{+}08] Andrew M. Muir, Trent M. Sutton, Paul J. Peeters, Randall M. Claramunt, and Ronald E. Kinnunen. An evaluation of age estimation structures for lake whitefish in Lake Michigan: Selecting an aging method based on precision and a decision analysis. North American Journal of Fisheries Management, 28(6):1928–1940, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McGrath:2008:LVA

[MSR08] Kathleen E. McGrath, J. Michael Scott, and Bruce E. Rieman. Length variation in age-0 westslope cutthroat trout at multiple spatial scales. North American Journal of Fisheries Management, 28(5):1529–1540, October 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKinney:2001:RTR

[MSRP01] Ted McKinney, David W. Speas, Roland S. Rogers, and William R. Persons. Rainbow trout in a regulated river below Glen Canyon Dam, Arizona, following increased minimum flows and reduced discharge variability. North American Journal of Fisheries Management, 21(1):216–222, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muir:2001:SEM

[MSW^{+}01] William D. Muir, Steven G. Smith, John G. Williams, Eric E. Hockersmith, and John R. Skalski. Survival estimates for migrant yearling Chinook salmon and steelhead tagged with passive integrated transponders in the Lower Snake and Lower Columbia Rivers, 1993–1998. North American Journal of Fisheries Management, 21(2):269–282, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Merz:2008:AME

[MSW^{+}08] Joseph E. Merz, James R. Smith, Michelle L. Workman, Jose D. Setka, and Bert Mulchaey. Aquatic macrophyte encroachment in Chinook salmon spawning beds: Lessons learned from gravel enhancement monitoring in the Lower Mokelumne River, California. North American Journal of Fisheries Management, 28(5):1568–1577, October 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muir:2001:SJS

[MSWS01] William D. Muir, Steven G. Smith, John G. Williams, and Benjamin P. Sandford. Survival of juvenile salmonids passing through bypass systems, turbines, and spillways with and without flow deflectors at Snake River dams. North American Journal of Fisheries Management, 21(1):135–146, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Myers:2008:TTV

[MTAB08] Randall Myers, John Taylor, Micheal Allen, and Timothy F. Bonvechio. Temporal trends in voluntary release of largemouth bass. North American Journal of Fisheries Management, 28(2):428–433, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:2008:SNF

[MTF^{+}08] Katrina B. Mueller, William W. Taylor, Kenneth A. Frank, John M. Robertson, and Dennis L. Grinold. Social networks and fisheries: The relationship between a charter fishing network, social capital, and catch dynamics. North American Journal of Fisheries Management, 28(2):447–462, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKee:2004:PTS

[MTHH04] Patrick C. McKee, Michael L. Toneys, Michael J. Hansen, and Mark E. Holey. Performance of two strains of lake trout stocked in the midlake refuge of Lake Michigan. North American Journal of Fisheries Management, 24(4):1101– 1111, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moustahfid:2009:AEP

[MTL09] Hassan Moustahfid, Megan C. Tyrrell, and Jason S. Link. Accounting explicitly for predation mortality in surplus production models: An application to longfin inshore squid. North American Journal of Fisheries Management, 29(6): 1555–1566, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murdoch:2009:MPS

[MTM09] Andrew R. Murdoch, Michael A. Tonseth, and Todd L. Miller. Migration patterns and spawning distribution of adult hatchery sockeye salmon released as parr from net-pens in Lake Wenatchee, Washington. North American Journal of Fisheries Management, 29(2):447–459, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Masuda:2009:EVH

[MTN09] Michele Masuda, Frank Thrower, and Krista M. Nichols. The effects of violating Hardy–Weinberg equilibrium assumptions on a cluster-based population mixture analysis of steelhead populations in Southeast Alaska. North American Journal of Fisheries Management, 29(1):140–150, February 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:2003:CES

[Mue03] Karl W. Mueller. A comparison of electrofishing and scuba diving to sample black bass in Western Washington lakes. North American Journal of Fisheries Management, 23(2): 632–639, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muhlfeld:2002:SCR

[Muh02] Clint C. Muhlfeld. Spawning characteristics of redband trout in a headwater stream in Montana. North American Journal of Fisheries Management, 22(4):1314–1320, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Munger:2002:RWM

[Mun02] Charles Munger. A review of walleye management in Texas with emphasis on Meredith Reservoir. North American Journal of Fisheries Management, 22(3):1064–1075, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maule:2009:PDV

[MVH^{+}09] Alec G. Maule, Scott P. VanderKooi, John B. Hamilton, Richard Stocking, and Jerri Bartholomew. Physiological development and vulnerability to Ceratomyxa shasta of fallrun Chinook salmon in the Upper Klamath River watershed. North American Journal of Fisheries Management, 29(6): 1743–1756, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:2004:RFS

[MW04] Gordon A. Mueller and Richard Wydoski. Reintroduction of the flannelmouth sucker in the Lower Colorado River. North American Journal of Fisheries Management, 24(1): 41–46, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mull:2007:SSS

[MW07] Kristin E. Mull and Margaret A. Wilzbach. Selection of spawning sites by Coho salmon in a Northern California stream. North American Journal of Fisheries Management, 27(4):1343–1354, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morizot:2002:BGC

[MWC02] Donald C. Morizot, J. Holt Williamson, and Gary J. Carmichael. Biochemical genetics of Colorado pikeminnow. North American Journal of Fisheries Management, 22(1): 66–76, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:2008:ESR

[MWN08] Paul H. Michaletz, Michael J. Wallendorf, and Dean M. Nicks. Effects of stocking rate, stocking size, and angler catch inequality on exploitation of stocked channel catfish in small Missouri impoundments. North American Journal of Fisheries Management, 28(5):1486–1497, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKinnell:2001:DOL

[MWR^{+}01] S. M. McKinnell, C. C. Wood, D. T. Rutherford, K. D. Hyatt, and D. W. Welch. The demise of Owikeno Lake sockeye salmon. North American Journal of Fisheries Management, 21(4):774–791, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mitro:2002:EAA

[MZ02] Matthew G. Mitro and Alexander V. Zale. Estimating abundances of age-0 rainbow trout by mark–recapture in a medium-sized river. North American Journal of Fisheries Management, 22(1):188–203, February 2002. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nate:2005:CML

[NB05] Nancy A. Nate and Mary T. Bremigan. Comparison of mean length at age and growth parameters of bluegills, largemouth bass, and yellow perch from length-stratified subsamples and samples in Michigan lakes. North American Journal of Fisheries Management, 25(4):1486–1492, November 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nielsen:2009:SGD

[NBGK09] Jennifer L. Nielsen, Alan Byrne, Sara L. Graziano, and Christine C. Kozfkay. Steelhead genetic diversity at multiple spatial scales in a managed basin: Snake River, Idaho. North American Journal of Fisheries Management, 29(3):680–701, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nate:2003:POS

[NBH^{+}03] Nancy A. Nate, Michael A. Bozek, Michael J. Hansen, Carl W. Ramm, Mary T. Bremigan, and Steven W. Hewett. Predicting the occurrence and success of walleye populations from physical and biological features of Northern Wisconsin lakes. North American Journal of Fisheries Management, 23 (4):1207–1214, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nate:2000:VWA

[NBHH00] Nancy A. Nate, Michael A. Bozek, Michael J. Hansen, and Steven W. Hewett. Variation in walleye abundance with lake size and recruitment source. North American Journal of Fisheries Management, 20(1):119–126, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nate:2001:VAW

[NBHH01] Nancy A. Nate, Michael A. Bozek, Michael J. Hansen, and Steven W. Hewett. Variation of adult walleye abundance in relation to recruitment and limnological variables in Northern Wisconsin lakes. North American Journal of Fisheries Management, 21(3):441–447, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Naughton:2004:PJS

[NBN04] George P. Naughton, David H. Bennett, and Ken B. Newman. Predation on juvenile salmonids by smallmouth bass in the Lower Granite Reservoir System, Snake River. North American Journal of Fisheries Management, 24(2):534–544, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newcomb:2001:ETM

[NC01] Tammy J. Newcomb and Thomas G. Coon. Evaluation of three methods for estimating numbers of steelhead smolts emigrating from Great Lakes tributaries. North American Journal of Fisheries Management, 21(3):548–560, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Naughton:2006:FAS

[NCK^{+}06] George P. Naughton, Christopher C. Caudill, Matthew L. Keefer, Theodore C. Bjornn, Christopher A. Peery, and Lowell C. Stuehrenberg. Fallback by adult sockeye salmon at Columbia River Dams. North American Journal of Fisheries Management, 26(2):380–390, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nagid:2004:SSB

[NDH04] Eric J. Nagid, Kevin J. Dockendorf, and Robert W. Hujik. Off-season spawning of black crappie in a North-Central Florida lake. North American Journal of Fisheries Management, 24(1):299–302, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Negus:2003:DSS

[Neg03] Mary T. Negus. Determination of smoltification status in juvenile migratory rainbow trout and Chinook salmon in Minnesota. North American Journal of Fisheries Management, 23(4):913–927, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newman:2006:CPM

[New06] Ken B. Newman. Comment: a pipeline model for estimating fishing mortality in salmon mark-selective fisheries. North American Journal of Fisheries Management, 26(2):377–379, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newell:2007:APM

[NFQ07] J. C. Newell, K. L. Fresh, and T. P. Quinn. Arrival patterns and movements of adult sockeye salmon in Lake Washington: Implications for management of an urban fishery. North American Journal of Fisheries Management, 27(3):908–917, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neeson:2008:FAA

[NGWK08] Thomas M. Neeson, Ann Marie Gorman, Peter J. Whiting, and Joseph F. Koonce. Factors affecting accuracy of stream channel slope estimates derived from geographical information systems. North American Journal of Fisheries Management, 28(3):722–732, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newman:2000:EIM

[NH00] Steven P. Newman and Michael H. Hoff. Evaluation of a 16inch minimum length limit for smallmouth bass in Pallette Lake, Wisconsin. North American Journal of Fisheries Management, 20(1):90–99, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Niva:2001:ERC

[NH01] Teuvo Niva and Pekka Hyvärinen. Evacuation rates of coded wire tags implanted in prey of northern pike. North American Journal of Fisheries Management, 21(3):692–695, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Niles:2007:CTL

[NH07] Jonathan M. Niles and Kyle J. Hartman. Comparison of three larval fish gears to sample shallow water sites on a navigable river. North American Journal of Fisheries Management, 27 (4):1126–1138, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Niles:2009:LFU

[NH09] Jonathan M. Niles and Kyle J. Hartman. Larval fish use of dike structures on a navigable river. North American Journal of Fisheries Management, 29(4):1035–1045, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newby:2000:CWA

[NHNE00] Justine R. Newby, Michael J. Hansen, Steven P. Newman, and Clayton J. Edwards. Catchability of walleyes to angling in Escanaba Lake, Wisconsin, 1980–1995. North American Journal of Fisheries Management, 20(4):873–881, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neely:2008:PSW

[NHS08] Ben C. Neely, Martin J. Hamel, and Kirk D. Steffensen. A proposed standard weight equation for blue suckers. North American Journal of Fisheries Management, 28(5):1450– 1452, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Naughton:2009:ETT

[NJP^{+}09] George P. Naughton, Michael A. Jepson, Christopher A. Peery, Christopher V. Brun, and Jennifer C. Graham. Effects of temporary tributary use on escapement estimates of adult fall Chinook salmon in the Deschutes River, Oregon. North American Journal of Fisheries Management, 29 (6):1511–1518, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nordwall:2000:SHS

[NL00] Fredrik Nordwall and Per Lundberg. Simulated harvesting of stream salmonids with a seasonal life history. North American Journal of Fisheries Management, 20(2):481–492, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newman:2006:ADE

[NL06] Ken B. Newman and Steven T. Lindley. Accounting for demographic and environmental stochasticity, observation error, and parameter uncertainty in fish population dynamics models. North American Journal of Fisheries Management, 26(3):685–701, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neal:2001:MLB

[NLC01] J. Wesley Neal and Darien Lopez-Clayton. Mortality of largemouth bass during catch-and-release tournaments in a Puerto Rico reservoir. North American Journal of Fisheries Management, 21(4):834–842, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nitschke:2001:CLW

[NMJ01] Paul Nitschke, Martha Mather, and Francis Juanes. A comparison of length-, weight-, and age-specific fecundity relationships for cunner in Cape Cod Bay. North American Journal of Fisheries Management, 21(1):86–95, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nelitz:2007:SBA

[NMP07] Marc A. Nelitz, Erland A. MacIsaac, and Randall M. Peterman. A science-based approach for identifying temperaturesensitive streams for rainbow trout. North American Journal of Fisheries Management, 27(2):405–424, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neal:2002:GSS

[NN02] J. Wesley Neal and Richard L. Noble. Growth, survival, and site fidelity of Florida and intergrade largemouth bass stocked in a tropical reservoir. North American Journal of Fisheries Management, 22(2):528–536, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neal:2008:CDT

[NN08] J. Wesley Neal and Richard L. Noble. Comparison of diploid and triploid largemouth bass growth and maturation through age 1 in Puerto Rico. North American Journal of Fisheries Management, 28(3):688–693, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Narum:2006:MRP

[NPE^{+}06] Shawn R. Narum, Madison S. Powell, Rolf Evenson, Bill Sharp, and AndréJ. Talbot. Microsatellites reveal population substructure of Klickitat River native steelhead and genetic divergence from an introduced stock. North American Journal of Fisheries Management, 26(1):147–155, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nelson:2005:BSW

[NRJ05] T. C. Nelson, M. L. Rosenau, and N. T. Johnston. Behavior and survival of wild and hatchery-origin winter steelhead spawners caught and released in a recreational fishery. North American Journal of Fisheries Management, 25(3):931–943, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Negus:2008:BEF

[NSH^{+}08] Mary T. Negus, Donald R. Schreiner, Theodore N. Halpern, Stephen T. Schram, Michael J. Seider, and Dennis M. Pratt. Bioenergetics evaluation of the fish community in the Western Arm of Lake Superior in 2004. North American Journal of Fisheries Management, 28(6):1649–1667, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Negus:2004:RNE

[NT04] Mary T. Negus and Fred T. Tureson. Retention and nonlethal external detection of calcein marks in rainbow trout and Chinook salmon. North American Journal of Fisheries Management, 24(2):741–747, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nye:2008:RCW

[NTH08] Janet A. Nye, Timothy E. Targett, and Thomas E. Helser. Reproductive characteristics of weakfish in Delaware Bay: Implications for management. North American Journal of Fisheries Management, 28(1):1–11, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nehring:2003:USC

[NTSJ03] R. Barry Nehring, Kevin G. Thompson, David L. Shuler, and Terence M. James. Using sediment core samples to examine the spatial distribution of Myxobolus cerebralis actinospore production in Windy Gap Reservoir, Colorado. North American Journal of Fisheries Management, 23(2):376–384, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nowak:2004:OSH

[NTW^{+}04] Gretchen M. Nowak, Roger A. Tabor, Eric J. Warner, Kurt L. Fresh, and Thomas P. Quinn. Ontogenetic shifts in habitat and diet of cutthroat trout in Lake Washington, Washington. North American Journal of Fisheries Management, 24 (2):624–635, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nelson:2003:PSS

[NWC^{+}03] R. John Nelson, Chris C. Wood, Glenn Cooper, Christian Smith, and Ben Koop. Population structure of sockeye salmon of the Central Coast of British Columbia: Implications for recovery planning. North American Journal of Fisheries Management, 23(3):703–720, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olson:2000:SGI

[OBG^{+}00] Mark H. Olson, Thomas E. Brooking, David M. Green, Anthony J. VanDeValk, and Lars G. Rudstam. Survival and growth of intensively reared large walleye fingerlings and extensively reared small fingerlings stocked concurrently in small lakes. North American Journal of Fisheries Management, 20(2):337–348, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ostrand:2004:ESL

[OCW04] Kenneth G. Ostrand, Steven J. Cooke, and David H. Wahl. Effects of stress on largemouth bass reproduction. North American Journal of Fisheries Management, 24(3):1038– 1045, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Oliveira:2009:PFQ

[OFM^{+}09] João M. Oliveira, Maria T. Ferreira, Paula Morgado, Robert M. Hughes, Amílcar Teixeira, Rui M. Cortes, and Jorge H. Bochechas. A preliminary fishery quality index for Portuguese streams. North American Journal of Fisheries Management, 29(5):1466–1478, October 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olsson:2001:EAP

[OGE01] Ivan C. Olsson, Larry A. Greenberg, and Anders G. Eklöv. Effect of an artificial pond on migrating brown trout smolts. North American Journal of Fisheries Management, 21(3): 498–506, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olson:2007:IAT

[OGK07] Nathan W. Olson, Christopher S. Guy, and Keith D. Koupal. Interactions among three top-level predators in a polymictic Great Plains reservoir. North American Journal of Fisheries Management, 27(1):268–278, February 2007. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olsen:2004:EUV

[OGS04] Esben M. Olsen, Jakob Gjøsæter, and Nils C. Stenseth. Evaluation of the use of visible implant tags in age-0 Atlantic cod. North American Journal of Fisheries Management, 24 (1):282–286, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Overholtz:2008:EAA

[OJL08] W. J. Overholtz, L. D. Jacobson, and J. S. Link. An ecosystem approach for assessment advice and biological reference points for the Gulf of Maine–Georges Bank Atlantic herring complex. North American Journal of Fisheries Management, 28(1):247–257, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Opperman:2004:ERR

[OM04] Jeff J. Opperman and Adina M. Merenlender. The effectiveness of riparian restoration for improving instream fish habitat in four hardwood-dominated California streams. North American Journal of Fisheries Management, 24(3):822–834, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

OMalley:2008:ETR

[O’M08] Joseph M. O’Malley. Evaluations of tag retention and a device for releasing discarded Hawaiian spiny lobsters Panulirus marginatus. North American Journal of Fisheries Management, 28(3):619–624, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olive:2005:CAM

[OMH05] J. A. Olive, L. E. Miranda, and W. D. Hubbard. Centrarchid assemblages in Mississippi State-operated fishing lakes. North American Journal of Fisheries Management, 25(1): 7–15, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ozen:2005:AAY

[ON05] Ozcan Ozen and Richard L. Noble. Assessing age-0 yearclass strength of fast-growing largemouth bass in a tropical reservoir. North American Journal of Fisheries Management, 25(1):163–170, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Orr:2007:ESL

[Orr07] Craig Orr. Estimated sea louse egg production from marine harvest Canada farmed Atlantic salmon in the Broughton Archipelago, British Columbia, 2003–2004. North American Journal of Fisheries Management, 27(1):187–197, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Odenkirk:2005:SVM

[OS05] John Odenkirk and Scott Smith. Single- versus multiple-pass boat electrofishing for assessing smallmouth bass populations in Virginia rivers. North American Journal of Fisheries Management, 25(2):717–724, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ostrand:2006:CET

[OSC06] Kenneth G. Ostrand, Michael J. Siepker, and Steven J. Cooke. Capture efficiencies of two hook types and associated injury and mortality of juvenile muskellunge angled with live baitfish. North American Journal of Fisheries Management, 26(3):622–627, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olsen:2001:EVI

[OV01] Esben Moland Olsen and Leif Asbjørn Vøllestad. An evaluation of visible implant elastomer for marking age-0 brown trout. North American Journal of Fisheries Management, 21(4):967–970, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ogle:2009:RLW

[OW09] Derek H. Ogle and Ian J. Winfield. Ruffe length–weight relationships with a proposed standard weight equation. North American Journal of Fisheries Management, 29(4):850–858, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Overton:2008:RMU

[OZR08] Anthony S. Overton, John Zabawski, and Kenneth L. Riley. Release mortality of undersized fish from the snapper– grouper complex off the North Carolina Coast. North American Journal of Fisheries Management, 28(3):733–739, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Powell:2007:REO

[PAAK07] Eric N. Powell, Kathryn A. Ashton-Alcox, and John N. Kraeuter. Reevaluation of Eastern oyster dredge efficiency in survey mode: Application in stock assessment. North American Journal of Fisheries Management, 27(2):492–511, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Plante:2005:AMM

[PALdlN05] Sébastien Plante, Céline Audet, Yvan Lambert, and Joël de la Noüe. Alternative methods for measuring energy content in winter flounder. North American Journal of Fisheries Management, 25(1):1–6, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Penne:2007:ERV

[PASP07] Christopher R. Penne, Nicholas L. Ahrens, Robert C. Summerfelt, and Clay L. Pierce. Effect of relative volume on radio transmitter expulsion in subadult common carp. North American Journal of Fisheries Management, 27(3):986–991, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peery:2000:DHR

[PB00] C. A. Peery and T. C. Bjornn. Dispersal of hatcheryreared Chinook salmon parr following release into four Idaho streams. North American Journal of Fisheries Management, 20(1):19–27, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Petersen:2003:EPI

[PB03] James H. Petersen and Craig A. Barfoot. Evacuation of passive integrated transponder (PIT) tags from northern pikeminnow consuming tagged juvenile Chinook salmon. North American Journal of Fisheries Management, 23(4): 1265–1270, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Price:2003:PFE

[PBB03] David J. Price, Michael G. Beiting, and Wesley J. Birge. Portable floating electrofishing unit for fish collections in freshwater stream systems. North American Journal of Fisheries Management, 23(4):974–977, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2000:EDR

[PBH00] Douglas L. Peterson, Mark B. Bain, and Nancy Haley. Evidence of declining recruitment of Atlantic sturgeon in the Hudson River. North American Journal of Fisheries Management, 20(1):231–238, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parken:2003:IUA

[PBI03] Charles K. Parken, Richard E. Bailey, and James R. Irvine. Incorporating uncertainty into area-under-the-curve and peak count salmon escapement estimation. North American Journal of Fisheries Management, 23(1):78–90, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Picard:2003:EUS

[PBM03] Chris R. Picard, Michael A. Bozek, and Walter T. Momot. Effectiveness of using summer thermal indices to classify and protect brook trout streams in Northern Ontario. North American Journal of Fisheries Management, 23(2):206–215, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2005:BNN

[PBT05] James T. Peterson, Nolan P. Banish, and Russell F. Thurow. Are block nets necessary?: Movement of stream-dwelling salmonids in response to three common survey methods.

North American Journal of Fisheries Management, 25(2): 732–743, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:2000:RDS

[PC00] Rodney B. Pierce and Mark F. Cook. Recreational darkhouse spearing for northern pike in Minnesota: Historical changes in effort and harvest and comparisons with angling. North American Journal of Fisheries Management, 20(1):239–244, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:2001:IDP

[PCG^{+}01] Clay L. Pierce, Amanda M. Corcoran, Amber N. Gronbach, Sandy Hsia, Benjamin J. Mullarkey, and Aaron J. Schwartzhoff. Influence of diel period on electrofishing and beach seining assessments of littoral fish assemblages. North American Journal of Fisheries Management, 21(4):918–926, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parken:2008:GCW

[PCIB08] Charles K. Parken, John R. Candy, James R. Irvine, and Terry D. Beacham. Genetic and coded wire tag results combine to allow more-precise management of a complex Chinook salmon aggregate. North American Journal of Fisheries Management, 28(1):328–340, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peake:2004:EAV

[Pea04] Stephan Peake. Effect of approach velocity on impingement of juvenile northern pike at water intake screens. North American Journal of Fisheries Management, 24(2):390–396, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peake:2008:BPP

[Pea08] Stephan J. Peake. Behavior and passage performance of northern pike, walleyes, and white suckers in an experimental raceway. North American Journal of Fisheries Management, 28(1):321–327, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phelps:2007:PFS

[PEW07] Quinton E. Phelps, Kris R. Edwards, and David W. Willis. Precision of five structures for estimating age of common carp. North American Journal of Fisheries Management, 27 (1):103–105, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pratt:2002:EFC

[PF02] Thomas C. Pratt and Michael G. Fox. Effect of fin clipping on overwinter growth and survival of age-0 walleyes. North American Journal of Fisheries Management, 22(4): 1290–1294, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pratt:2003:CRS

[PF03] Thomas C. Pratt and Michael G. Fox. Contribution and relative survival of age-0 walleyes stocked at two sizes in six Eastern Ontario lakes. North American Journal of Fisheries Management, 23(4):997–1005, August 2003. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phillips:2009:EVI

[PF09] Catherine T. Phillips and Joe N. Fries. An evaluation of visible implant elastomer for marking the federally listed fountain darter and the San Marcos salamander. North American Journal of Fisheries Management, 29(3):529–532, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paller:2006:EFT

[PFS^{+}06] M. H. Paller, D. E. Fletcher, M. M. Standora, T. B. Grabowski, T. A. Jones, S. A. Dyer, and J. J. Isely. Emigration of fish from two South Carolina cooling reservoirs. North American Journal of Fisheries Management, 26(4):976–982, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2008:WEO

[PFWC08] Douglas P. Peterson, Kurt D. Fausch, James Watmough, and Richard A. Cunjak. When eradication is not an option: Modeling strategies for electrofishing suppression of nonnative brook trout to Foster persistence of sympatric native cutthroat trout in small streams. North American Journal of Fisheries Management, 28(6):1847–1867, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Page:2004:FTL

[PGR^{+}04] Kevin S. Page, Gerold C. Grant, Paul Radomski, Thomas S. Jones, and Richard E. Bruesewitz. Fish total length measurement error from recreational anglers: Causes and contribution to noncompliance for the mille lacs walleye fishery. North American Journal of Fisheries Management, 24(3):939–951, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pepper:2001:APB

[PGS01] Phillip N. Pepper, Thomas L. Girard, and Martin A. Stapanian. An apparatus for preparing benthic samples aboard ship. North American Journal of Fisheries Management, 21 (1):249–252, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pereira:2003:PCR

[PH03a] Donald L. Pereira and Michael J. Hansen. A perspective on challenges to recreational fisheries management: Summary of the symposium on active management of recreational fisheries. North American Journal of Fisheries Management, 23 (4):1276–1282, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Polacheck:2003:DTP

[PH03b] Tom Polacheck and William Hearn. Designing tagging programs for pelagic longline fisheries: Trade-offs between the number of releases and observer coverage. North American Journal of Fisheries Management, 23(3):810–821, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porath:2005:EWT

[PH05] Mark T. Porath and Keith L. Hurley. Effects of waterbody type and management actions on bluegill growth rates. North American Journal of Fisheries Management, 25(3): 1041–1050, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:2008:ERG

[PH08] Vaughn L. Paragamian and Michael J. Hansen. Evaluation of recovery goals for endangered white sturgeon in the Kootenai River, Idaho. North American Journal of Fisheries Management, 28(2):463–470, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:2009:RNB

[PH09] Vaughn L. Paragamian and Michael J. Hansen. Rehabilitation needs for burbot in the Kootenai River, Idaho, USA, and British Columbia, Canada. North American Journal of Fisheries Management, 29(3):768–777, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phelps:2009:SCC

[PHB^{+}09] Quinton E. Phelps, David P. Herzog, Ronald C. Brooks, Valerie A. Barko, David E. Ostendorf, Joseph W. Ridings, Sara J. Tripp, Robert E. Colombo, James E. Garvey, and Robert A. Hrabik. Seasonal comparison of catch rates and size structure using three gear types to sample sturgeon in the Middle Mississippi River. North American Journal of Fisheries Management, 29(5):1487–1495, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pon:2009:CPC

[PHC^{+}09] L. B. Pon, S. G. Hinch, S. J. Cooke, D. A. Patterson, and A. P. Farrell. A comparison of the physiological condition, and fishway passage time and success of migrant adult sockeye salmon at Seton River Dam, British Columbia, under three operational water discharge rates. North American Journal of Fisheries Management, 29(5):1195–1205, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pollock:2001:TRR

[PHHC01] Kenneth H. Pollock, John M. Hoenig, William S. Hearn, and Brian Calingaert. Tag reporting rate estimation: 1. An evaluation of the high-reward tagging method. North American Journal of Fisheries Management, 21(3):521–532, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pollock:2002:TRR

[PHHC02] Kenneth H. Pollock, John M. Hoenig, William S. Hearn, and Brian Calingaert. Tag reporting rate estimation: 2. use of high-reward tagging and observers in multiple-component fisheries. North American Journal of Fisheries Management, 22(3):727–736, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsons:2003:EPF

[PHK03] Glenn R. Parsons, Jan Jeffrey Hoover, and K. Jack Killgore. Effect of pectoral fin ray removal on station-holding ability of shovelnose sturgeon. North American Journal of Fisheries Management, 23(3):742–747, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Patton:2000:ENE

[PHRG00] Tim M. Patton, Wayne A. Hubert, Frank J. Rahel, and Kenneth G. Gerow. Effort needed to estimate species richness in small streams on the Great Plains in Wyoming. North American Journal of Fisheries Management, 20(2):394–398, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:2004:OIR

[Pie04] Rodney B. Pierce. Oviduct insertion of radio transmitters as a means of locating northern pike spawning habitat. North American Journal of Fisheries Management, 24(1):244–248, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pine:2000:CTO

[Pin00] William E. Pine III. Comparison of two otter trawls of different sizes for sampling black crappies. North American Journal of Fisheries Management, 20(3):819–821, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pitlo:2002:EEF

[Pit02] John Pitlo, Jr. Effects of environmental factors on walleye and sauger recruitment in Pool 13, Upper Mississippi River. North American Journal of Fisheries Management, 22(3):

1021–1031, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pearsons:2009:ADP

[PJBT09] Todd N. Pearsons, Christopher L. Johnson, Brenda Ben James, and Gabriel M. Temple. Abundance and distribution of precociously mature male spring Chinook salmon of hatchery and natural origin in the Yakima River. North American Journal of Fisheries Management, 29(3):778–790, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ploskey:2000:ELF

[PJC00] Gene R. Ploskey, Peter N. Johnson, and Thomas J. Carlson. Evaluation of a low-frequency sound-pressure system for guiding juvenile salmon away from turbines at Bonneville Dam, Columbia River. North American Journal of Fisheries Management, 20(4):951–967, November 2000. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:2001:KRW

[PK01] Vaughn L. Paragamian and Gretchen Kruse. Kootenai River white sturgeon spawning migration behavior and a predictive model. North American Journal of Fisheries Management, 21(1):10–21, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paragamian:2001:SHK

[PKW01] Vaughn L. Paragamian, Gretchen Kruse, and Virginia Wakkinen. Spawning habitat of Kootenai River white sturgeon, post-Libby Dam. North American Journal of Fisheries Management, 21(1):22–33, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pan:2001:DSM

[PLP01] Minling Pan, Ping Sun Leung, and Samuel G. Pooley. A decision support model for fisheries management in Hawaii: a multilevel and multiobjective programming approach. North American Journal of Fisheries Management, 21(2):293–309, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phelps:2009:HCB

[PLW^{+}09] Quinton E. Phelps, Adam M. Lohmeyer, Nicholas C. Wahl, John M. Zeigler, and Gregory W. Whitledge. Habitat characteristics of black crappie nest sites in an Illinois impoundment. North American Journal of Fisheries Management, 29 (1):189–195, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prager:2001:HRM

[PM01] Michael H. Prager and Michael S. Mohr. The harvest rate model for Klamath River Fall Chinook salmon, with management applications and comments on model development and documentation. North American Journal of Fisheries Management, 21(3):533–547, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pothoven:2008:CCA

[PM08] Steven A. Pothoven and Charles P. Madenjian. Changes in consumption by alewives and lake whitefish after dreissenid mussel invasions in Lakes Michigan and Huron. North American Journal of Fisheries Management, 28(1):308–320, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paradis:2007:LWR

[PMBM07] Yves Paradis, Pierre Magnan, Philippe Brodeur, and Marc Mingelbier. Length and weight reduction in larval and juvenile yellow perch preserved with dry ice, formalin, and ethanol. North American Journal of Fisheries Management, 27(3):1004–1009, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paradis:2008:CCP

[PMBM08] Yves Paradis, Marc Mingelbier, Philippe Brodeur, and Pierre Magnan. Comparisons of catch and precision of pop nets, push nets, and seines for sampling larval and juvenile yellow perch. North American Journal of Fisheries Management, 28(5):1554–1562, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Palmer:2005:MWC

[PMH05] George C. Palmer, Brian R. Murphy, and Eric M. Hallerman. Movements of walleyes in Claytor Lake and the Upper New River, Virginia, indicate distinct lake and river populations. North American Journal of Fisheries Management, 25(4):1448–1455, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Post:2003:AAH

[PMPS03] John R. Post, Craig Mushens, Andrew Paul, and Michael Sullivan. Assessment of alternative harvest regulations for sustaining recreational fisheries: Model development and application to bull trout. North American Journal of Fisheries Management, 23(1):22–34, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porath:2003:ESD

[PN03] Mark T. Porath and Jennifer L. Nielsen. Evidence of sexually dimorphic introgression in Pinaleno Mountain Apache trout. North American Journal of Fisheries Management, 23 (1):172–180, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Popoff:2005:RMR

[PN05] Nicholas D. Popoff and Robert M. Neumann. Range and movement of resident holdover and hatchery brown trout tagged with radio transmitters in the Farmington River, Connecticut. North American Journal of Fisheries Management, 25(2):413–422, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pothoven:2001:CDB

[PNSB01] Steven A. Pothoven, Thomas F. Nalepa, Philip J. Schneeberger, and Stephen B. Brandt. Changes in diet and body condition of lake whitefish in Southern Lake Michigan associated with changes in benthos. North American Journal of Fisheries Management, 21(4):876–883, November 2001. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Powers:2005:MSY

[Pow05] Joseph E. Powers. Maximum sustainable yield and bycatch minimization “to the extent practicable”. North American Journal of Fisheries Management, 25(3):785–790, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prince:2000:COA

[PP00] Angela Prince and Cynthia Powell. Clove oil as an anesthetic for invasive field procedures on adult rainbow trout. North American Journal of Fisheries Management, 20(4): 1029–1032, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsons:2001:RBW

[PP01] Bradford G. Parsons and Donald L. Pereira. Relationship between walleye stocking and year-class strength in three Minnesota lakes. North American Journal of Fisheries Management, 21(4):801–808, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Plumb:2006:ERI

[PPAR06] John M. Plumb, Russell W. Perry, Noah S. Adams, and Dennis W. Rondorf. The effects of river impoundment and hatchery rearing on the migration behavior of juvenile steelhead in the Lower Snake River, Washington. North American Journal of Fisheries Management, 26(2):438–452, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pollock:2004:IBP

[PPBM04] Michael M. Pollock, George R. Pess, Timothy J. Beechie, and David R. Montgomery. The importance of Beaver ponds to Coho salmon production in the Stillaguamish River Basin, Washington, USA. North American Journal of Fisheries Management, 24(3):749–760, August 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porath:2003:IAI

[PPE03] Mark T. Porath, Edward J. Peters, and Darrol L. Eichner. Impact of alewife introduction on walleye and white bass condition in Lake McConaughy, Nebraska, 1980–1995. North American Journal of Fisheries Management, 23(4):

1050–1055, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pink:2007:UUV

[PPF07] Melissa Pink, Thomas C. Pratt, and Michael G. Fox. Use of underwater visual distance sampling for estimating habitatspecific population density. North American Journal of Fisheries Management, 27(1):246–255, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paul:2003:CAI

[PPS03] Andrew J. Paul, John R. Post, and Jim D. Stelfox. Can anglers influence the abundance of native and nonnative salmonids in a stream from the Canadian rocky mountains? North American Journal of Fisheries Management, 23(1): 109–119, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prager:2003:TLM

[PPSC03] Michael H. Prager, Clay E. Porch, Kyle W. Shertzer, and John F. Caddy. Targets and limits for management of fisheries: a simple probability-based approach. North American Journal of Fisheries Management, 23(2):349–361, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2001:EEO

[PR01] James T. Peterson and Charles F. Rabeni. Evaluating the efficiency of a one-square-meter quadrat sampler for riffledwelling fish. North American Journal of Fisheries Management, 21(1):76–85, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parker:2003:TLR

[PR03] Steven J. Parker and Polly S. Rankin. Tag location and retention in black rockfish: Feasibility of using PIT tags in a wild marine species. North American Journal of Fisheries Management, 23(4):993–996, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paukert:2004:FAC

[PR04a] Craig Paukert and R. Scott Rogers. Factors affecting condition of flannelmouth suckers in the Colorado River, Grand Canyon, Arizona. North American Journal of Fisheries Management, 24(2):648–653, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Price:2004:UTE

[PR04b] A. Blake Price and Roger A. Rulifson. Use of traditional ecological knowledge to reduce striped bass bycatch in the Currituck Sound white perch gill-net fishery. North American Journal of Fisheries Management, 24(3):785–792, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsons:2005:EBA

[PR05a] Bradford G. Parsons and Jeffrey R. Reed. Effect of T-bar anchor tags on growth of black crappies. North American Journal of Fisheries Management, 25(4):1591–1595, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsons:2005:MBC

[PR05b] Bradford G. Parsons and Jeffrey R. Reed. Movement of black crappies and bluegills among interconnected lakes in Minnesota. North American Journal of Fisheries Management, 25(2):689–695, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parker:2003:DMT

[PRHS03] Steven J. Parker, Polly S. Rankin, Robert W. Hannah, and Carl B. Schreck. Discard mortality of trawl-caught lingcod in relation to tow duration and time on deck. North American Journal of Fisheries Management, 23(2):530–542, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porter:2000:PMF

[PRP00] Marc S. Porter, Jordan Rosenfeld, and Eric A. Parkinson. Predictive models of fish species distribution in the blackwater drainage, British Columbia. North American Journal of Fisheries Management, 20(2):349–359, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peters:2008:FIR

[PRP08] Lindsay M. Peters, Ulrich G. Reinhardt, and Mark A. Pegg. Factors influencing radio wave transmission and reception: Use of radiotelemetry in large river systems. North American Journal of Fisheries Management, 28(1):301–307, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perrin:2006:RMR

[PRSA06] C. J. Perrin, Marvin L. Rosenau, T. Brock Stables, and K. I. Ashley. Restoration of a montane reservoir fishery via biomanipulation and nutrient addition. North American Journal of Fisheries Management, 26(2):391–407, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parken:2002:PAP

[PS02] Charles K. Parken and Dennis L. Scarnecchia. Predation on age-0 paddlefish by walleye and sauger in a Great Plains reservoir. North American Journal of Fisheries Management, 22(3):750–759, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Patterson:2007:IWT

[PSB^{+}07] D. A. Patterson, K. M. Skibo, D. P. Barnes, J. A. Hills, and J. S. Macdonald. The influence of water temperature on time to surface for adult sockeye salmon carcasses and the limitations in estimating salmon carcasses in the Fraser River, British Columbia. North American Journal of Fisheries Management, 27(3):878–884, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pribyl:2009:DRD

[PSKP09] Alena L. Pribyl, Carl B. Schreck, Michael L. Kent, and Steven J. Parker. The differential response to decompression in three species of nearshore Pacific rockfish. North American Journal of Fisheries Management, 29(5):1479–1486, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:2001:DLF

[PSP^{+}01] Clay L. Pierce, Mark D. Sexton, Mark E. Pelham, Hongsheng Liao, and Joe G. Larscheid. Dynamics of the littoral fish assemblage in Spirit Lake, Iowa, and implications for prey availability for piscivores. North American Journal of Fisheries Management, 21(4):884–896, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:2006:CLJ

[PSP^{+}06] Rodney B. Pierce, Steve Shroyer, Bruce Pittman, Dale E. Logsdon, and Todd D. Kolander. Catchability of larval and juvenile northern pike in quatrefoil light traps. North American Journal of Fisheries Management, 26(4):908–915, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parkos:2006:EPM

[PSW06] Joseph J. Parkos III, Victor J. Santucci, Jr., and David H. Wahl. Effectiveness of a plastic mesh substrate cover for reducing the effects of common carp on aquatic ecosystems. North American Journal of Fisheries Management, 26(4): 861–866, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parker:2007:BTP

[PSWD07] Brian R. Parker, David W. Schindler, Frank M. Wilhelm, and David B. Donald. Bull trout population responses to reductions in angler effort and retention limits. North American Journal of Fisheries Management, 27(3):848–859, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pearsons:2007:IES

[PT07] Todd N. Pearsons and Gabriel M. Temple. Impacts of early stages of salmon supplementation and reintroduction programs on three trout species. North American Journal of Fisheries Management, 27(1):1–20, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:2003:PDT

[PTD03] Rodney B. Pierce, Cynthia M. Tomcko, and Melissa T. Drake. Population dynamics, trophic interactions, and production of northern pike in a shallow bog lake and their effects on simulated regulation strategies. North American Journal of Fisheries Management, 23(2):323–330, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:2003:DDG

[PTM03] Rodney B. Pierce, Cynthia M. Tomcko, and Terry L. Margenau. Density dependence in growth and size structure of northern pike populations. North American Journal of Fisheries Management, 23(2):331–339, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:2006:IBS

[PTN06] Rodney B. Pierce, Cynthia M. Tomcko, and Mary T. Negus. Interactions between stocked walleyes and native yellow perch in Lake Thirteen, Minnesota: a case history of Percid community dynamics. North American Journal of Fisheries Management, 26(1):97–107, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pope:2001:AOS

[PVMR01] Kevin L. Pope, Brian E. Van Zee, Michael C. Mayo, and Mohammed Rahman. Assessment of outputs from Smithroot model-5.0 GPP and model-7.5 GPP electrofishers. North American Journal of Fisheries Management, 21(2):353–357, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Post:2006:FEI

[PvPR^{+}06] John R. Post, Brett T. van Poorten, Trevor Rhodes, Paul Askey, and Andrew Paul. Fish entrainment into irrigation canals: An analytical approach and application to the Bow River, Alberta, Canada. North American Journal of Fisheries Management, 26(4):875–887, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paukert:2003:PCE

[PW03] Craig P. Paukert and David W. Willis. Population characteristics and ecological role of northern pike in shallow natural lakes in Nebraska. North American Journal of Fisheries Management, 23(2):313–322, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Palmer:2009:UPD

[PW09] Michael C. Palmer and Susan E. Wigley. Using positional data from vessel monitoring systems to validate the logbookreported area fished and the stock allocation of commercial fisheries landings. North American Journal of Fisheries Management, 29(4):928–942, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paukert:2004:MHR

[PWB04] Craig P. Paukert, David W. Willis, and Michelle A. Bouchard. Movement, home range, and site fidelity of bluegills in a Great Plains lake. North American Journal of Fisheries Management, 24(1):154–161, February 2004. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paukert:2002:EAB

[PWG02] Craig P. Paukert, David W. Willis, and Donald W. Gabelhouse, Jr. Effect and acceptance of bluegill length limits in Nebraska natural lakes. North American Journal of Fisheries Management, 22(4):1306–1313, November 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paukert:2002:SSR

[PWH02] Craig P. Paukert, David W. Willis, and Richard S. Holland. Sample size requirements for in situ vegetation and substrate classifications in shallow, natural Nebraska lakes. North American Journal of Fisheries Management, 22(4): 1329–1333, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paukert:2002:EPE

[PWK02] Craig P. Paukert, David W. Willis, and Joel A. Klammer. Effects of predation and environment on quality of yellow perch and bluegill populations in Nebraska sandhill lakes. North American Journal of Fisheries Management, 22(1): 86–95, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phelps:2007:MDS

[PWPC07] Quinton E. Phelps, David W. Willis, Kipp A. Powell, and Steven R. Chipps. A method for determining stomach fullness for planktivorous fishes. North American Journal of Fisheries Management, 27(3):932–935, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quist:2009:MMD

[QBH^{+}09] Michael C. Quist, Michael R. Bower, Wayne A. Hubert, Thomas L. Parchman, and David B. McDonald. Morphometric and meristic differences among bluehead suckers, flannelmouth suckers, white suckers, and their hybrids: Tools for the management of native species in the Upper Colorado River Basin. North American Journal of Fisheries Management, 29 (2):460–467, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quist:2006:RVF

[QGBH06] Michael C. Quist, Kenneth G. Gerow, Michael R. Bower, and Wayne A. Hubert. Random versus fixed-site sampling when monitoring relative abundance of fishes in headwater streams of the Upper Colorado River Basin. North American Journal of Fisheries Management, 26(4):1011–1019, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quist:2002:ELW

[QGBS02] Michael C. Quist, Christopher S. Guy, Randall J. Bernot, and James L. Stephen. Ecology of larval white bass in a large Kansas reservoir. North American Journal of Fisheries Management, 22(2):637–642, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quist:2002:PIH

[QGP^{+}02] Michael C. Quist, Christopher S. Guy, Mark A. Pegg, Patrick J. Braaten, Clay L. Pierce, and Vincent H. Travnichek. Potential influence of harvest on shovelnose sturgeon populations in the Missouri River system. North American Journal of Fisheries Management, 22(2):537–549, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quist:2003:LCW

[QGSS03] Michael C. Quist, Christopher S. Guy, Randall D. Schultz, and James L. Stephen. Latitudinal comparisons of walleye growth in North America and factors influencing growth of walleyes in Kansas reservoirs. North American Journal of Fisheries Management, 23(3):677–692, August 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quist:2007:PHS

[QJBH07] Michael C. Quist, Zachary J. Jackson, Michael R. Bower, and Wayne A. Hubert. Precision of hard structures used to estimate age of riverine catostomids and cyprinids in the Upper Colorado River Basin. North American Journal of Fisheries Management, 27(2):643–649, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:2000:URH

[QK00] Jeffrey W. Quinn and Thomas J. Kwak. Use of rehabilitated habitat by brown trout and rainbow trout in an Ozark tailwater river. North American Journal of Fisheries Management, 20(3):737–751, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quist:2004:ASM

[QSGS04] Michael C. Quist, James L. Stephen, Christopher S. Guy, and Randall D. Schultz. Age structure and mortality of walleyes in Kansas reservoirs: Use of mortality caps to establish realistic management objectives. North American Journal of Fisheries Management, 24(3):990–1002, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quist:2007:ETU

[Qui07] Michael C. Quist. An evaluation of techniques used to index recruitment variation and year-class strength. North American Journal of Fisheries Management, 27(1):30–42, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rieman:2001:EPS

[RA01] B. E. Rieman and F. W. Allendorf. Effective population size and genetic conservation criteria for bull trout. North American Journal of Fisheries Management, 21(4):756–764, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roberts:2004:CIF

[RA04] James H. Roberts and Paul L. Angermeier. A comparison of injectable fluorescent marks in two genera of darters: Effects on survival and retention rates. North American Journal of Fisheries Management, 24(3):1017–1024, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Radomski:2003:IAA

[Rad03] Paul Radomski. Initial attempts to actively manage recreational fishery harvest in Minnesota. North American Journal of Fisheries Management, 23(4):1329–1342, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Radomski:2006:HCA

[Rad06] Paul Radomski. Historical changes in abundance of floatingleaf and emergent vegetation in Minnesota lakes. North American Journal of Fisheries Management, 26(4):932–940, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reid:2001:RLT

[RAH01] David M. Reid, David M. Anderson, and Bryan A. Henderson. Restoration of lake trout in Parry Sound, Lake Huron. North American Journal of Fisheries Management, 21(1): 156–169, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ralston:2002:WCG

[Ral02] Stephen Ralston. West Coast groundfish harvest policy. North American Journal of Fisheries Management, 22(2): 249–250, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Radomski:2009:ELB

[RAP09] Paul Radomski, Charles S. Anderson, and Kevin S. Page. Evaluation of largemouth bass length limits and catch-and-release regulations, with emphasis on the incorporation of biologists’ perceptions of largemouth bass length frequency distributions. North American Journal of Fisheries Management, 29(3):614–625, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reischel:2003:IFP

[RB03] Tami S. Reischel and Theodore C. Bjornn. Influence of fishway placement on fallback of adult salmon at the Bonneville Dam on the Columbia River. North American Journal of Fisheries Management, 23(4):1215–1224, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reeves:2007:FIH

[RB07] Keith A. Reeves and Richard E. Bruesewitz. Factors influencing the hooking mortality of walleyes caught by recreational anglers on mille lacs, Minnesota. North American Journal of Fisheries Management, 27(2):443–452, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reeves:2009:MPT

[RB09] Kerry S. Reeves and David L. Buckmeier. Mortality, predation, and tag visibility of fish marked with visible implant elastomer tags. North American Journal of Fisheries Management, 29(2):323–329, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roni:2002:RSR

[RBB^{+}02] Philip Roni, Timothy J. Beechie, Robert E. Bilby, Frank E. Leonetti, Michael M. Pollock, and George R. Pess. A review of stream restoration techniques and a hierarchical strategy for prioritizing restoration in Pacific Northwest watersheds. North American Journal of Fisheries Management, 22(1): 1–20, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rudershausen:2008:CRS

[RBB08] Paul J. Rudershausen, M. Scott Baker, Jr., and Jeffrey A. Buckel. Catch rates and selectivity among three trap types in the U.S. South Atlantic black sea bass commercial trap fishery. North American Journal of Fisheries Management, 28(4):1099–1107, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rach:2009:TRA

[RBK09] Jeff J. Rach, Michael Boogaard, and Cynthia Kolar. Toxicity of rotenone and antimycin to silver carp and bighead carp. North American Journal of Fisheries Management, 29 (2):388–395, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rooper:2000:USA

[RBM00] Christopher N. Rooper, Mason D. Bryant, and Steven J. Mc-Curdy. Use of scales to assess summer growth of resident cutthroat trout in Margaret Lake, Alaska. North American Journal of Fisheries Management, 20(2):467–480, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robinson:2001:JGN

[RC01] Anthony T. Robinson and Michael R. Childs. Juvenile growth of native fishes in the Little Colorado River and in a thermally modified portion of the Colorado River. North American Journal of Fisheries Management, 21(4):809–815, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richards:2007:SRA

[RCB07] Nathan S. Richards, Steven R. Chipps, and Michael L. Brown. Stress response and avoidance behavior of fishes as influenced by high-frequency strobe lights. North American Journal of Fisheries Management, 27(4):1310–1315, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ross:2005:APA

[RCK05] Jeff R. Ross, Jeffery D. Crosby, and Jarrad T. Kosa. Accuracy and precision of age estimation of crappies. North American Journal of Fisheries Management, 25(2):423–428, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roghair:2005:BTM

[RD05a] Craig N. Roghair and C. Andrew Dolloff. Brook trout movement during and after recolonization of a naturally defaunated stream reach. North American Journal of Fisheries Management, 25(3):777–784, August 2005. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rosenberger:2005:VAE

[RD05b] Amanda E. Rosenberger and Jason B. Dunham. Validation of abundance estimates from mark–recapture and removal techniques for rainbow trout captured by electrofishing in small streams. North American Journal of Fisheries Management, 25(4):1395–1410, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rosenberger:2006:EVA

[RD06] Amanda E. Rosenberger and Jason B. Dunham. Erratum: Validation of abundance estimates from mark-recapture and removal techniques for rainbow trout captured by electrofishing in small streams. North American Journal of Fisheries Management, 26(2):502, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reich:2009:SSE

[RD09] Danielle Ameen Reich and Joseph Thomas DeAlteris. A simulation study of the effects of spatially complex population structure for Gulf of Maine Atlantic cod. North American Journal of Fisheries Management, 29(1):116–126, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robbins:2008:GPA

[RDBD08] Stacey N. Robbins, Tanya L. Darden, Amy O. Ball, and Michael R. Denson. Genetic parentage analysis of red drum in South Carolina and the use of genotypic data from archived otoliths. North American Journal of Fisheries Management, 28(6):1796–1801, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rust:2002:LCI

[RDME02] Ashley J. Rust, James S. Diana, Terry L. Margenau, and Clayton J. Edwards. Lake characteristics influencing spawning success of muskellunge in Northern Wisconsin lakes. North American Journal of Fisheries Management, 22(3): 834–841, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ryan:2000:MES

[RDN00] Brad A. Ryan, Earl M. Dawley, and Richard A. Nelson. Modeling the effects of supersaturated dissolved gas on resident aquatic biota in the main-stem Snake and Columbia Rivers. North American Journal of Fisheries Management, 20(1): 192–204, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reid:2007:IMU

[Rei07] Ian S. Reid. Influence of motorboat use on thermal refuges and implications to salmonid physiology in the Lower Rogue River, Oregon. North American Journal of Fisheries Management, 27(4):1162–1173, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roni:2000:EWS

[RF00] Philip Roni and Andrew Fayram. Estimating winter salmonid abundance in small Western Washington streams: a comparison of three techniques. North American Journal of Fisheries Management, 20(3):683–692, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Riley:2004:EEH

[RFL04] Stephen C. Riley, Howard J. Fuss, and Larry L. LeClair. Ecological effects of hatchery-reared juvenile Chinook and Coho salmon on wild juvenile salmonids in two Washington streams. North American Journal of Fisheries Management, 24(2):506–517, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ryan:2001:DPI

[RFLN01] Brad A. Ryan, John W. Ferguson, Richard D. Ledgerwood, and Edmond P. Nunnallee. Detection of passive integrated transponder tags from juvenile salmonids on piscivorous bird colonies in the Columbia River Basin. North American Journal of Fisheries Management, 21(2):417–421, May 2001. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Radomski:2001:CHL

[RG01] Paul Radomski and Timothy J. Goeman. Consequences of human lakeshore development on emergent and floating-leaf vegetation abundance. North American Journal of Fisheries Management, 21(1):46–61, February 2001. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rice:2007:UCG

[RGP^{+}07] Patrick H. Rice, C. Phillip Goodyear, Eric D. Prince, Derke Snodgrass, and Joseph E. Serafy. Use of catenary geometry to estimate hook depth during near-surface pelagic longline fishing: Theory versus practice. North American Journal of Fisheries Management, 27(4):1148–1161, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rosenfeld:2003:RBL

[RH03] Jordan S. Rosenfeld and Leonardo Huato. Relationship between large woody debris characteristics and pool formation in small coastal British Columbia streams. North American Journal of Fisheries Management, 23(4):928–938, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reynolds:2004:IAE

[RH04] James B. Reynolds and F. Michael Holliman. Injury of American eels captured by electrofishing and trap-netting. North American Journal of Fisheries Management, 24(2):686–689, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rogers:2003:CWF

[RHB03] Mark W. Rogers, Michael J. Hansen, and T. Douglas Beard, Jr. Catchability of walleyes to fyke netting and electrofishing in Northern Wisconsin lakes. North American Journal of Fisheries Management, 23(4):1193–1206, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rogers:2005:RBR

[RHB05] Mark W. Rogers, Michael J. Hansen, and T. Douglas Beard, Jr. Relationships between recapture rates from different gears for estimating walleye abundance in Northern Wisconsin lakes. North American Journal of Fisheries Management, 25(1):195–202, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roni:2008:GRP

[RHB08] Phil Roni, Karrie Hanson, and Tim Beechie. Global review of the physical and biological effectiveness of stream habitat rehabilitation techniques. North American Journal of Fisheries Management, 28(3):856–890, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richards:2004:RDY

[RHBS04] Jessica M. Richards, Michael J. Hansen, Charles R. Bronte, and Shawn P. Sitar. Recruitment dynamics of the 1971–1991 year-classes of lake trout in Michigan waters of Lake Superior. North American Journal of Fisheries Management, 24 (2):475–489, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rhea:2005:TSV

[RHGW05] Darren T. Rhea, Wayne A. Hubert, R. Scott Gangl, and Roy A. Whaley. Temporal and spatial variation in relative abundance and length structure of salmonids in reservoirs: Implications for monitoring. North American Journal of Fisheries Management, 25(4):1301–1309, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richardson-Heft:2000:MLB

[RHHFB00] Carol A. Richardson-Heft, Alan A. Heft, Leon Fewlass, and Stephen B. Brandt. Movement of largemouth bass in Northern Chesapeake Bay: Relevance to sportfishing tournaments. North American Journal of Fisheries Management, 20(2): 493–501, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Radomski:2006:ETL

[RHJ^{+}06] Paul Radomski, Tom Heinrich, Thomas S. Jones, Pat Rivers, and Phil Talmage. Estimates of tackle loss for five Minnesota walleye fisheries. North American Journal of Fisheries Management, 26(1):206–212, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reynolds:2003:EER

[RHK^{+}03] Lou Reynolds, Alan T. Herlihy, Philip R. Kaufmann, Stanley V. Gregory, and Robert M. Hughes. Electrofishing effort requirements for assessing species richness and biotic integrity in Western Oregon streams. North American Journal of Fisheries Management, 23(2):450–461, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richter:2007:DES

[Ric07] Tracy J. Richter. Development and evaluation of standard weight equations for bridgelip suckers and largescale suckers. North American Journal of Fisheries Management, 27 (3):936–939, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rikardsen:2000:EFS

[Rik00] Audun H. Rikardsen. Effects of floy and soft VIalpha tags on growth and survival of juvenile Arctic char. North American Journal of Fisheries Management, 20(3):720–729, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rischbieter:2000:SPS

[Ris00] Douglas B. C. Rischbieter. Structures to prevent the spread of nuisance fish from Lake Davis, California. North American Journal of Fisheries Management, 20(3):784–790, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Retired:2007:ARC

[RJB07] Charles R. Rumsey Retired, Nicholas E. Jones, and Hugh H. Banks. An attempt to rehabilitate a collapsed brook trout population by introducing F_{1} splake to control yellow perch. North American Journal of Fisheries Management, 27(4): 1139–1147, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reid:2008:ESP

[RJY08] S. M. Reid, N. E. Jones, and G. Yunker. Evaluation of singlepass electrofishing and rapid habitat assessment for monitoring redside dace. North American Journal of Fisheries Management, 28(1):50–56, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Racey:2002:YCC

[RL02] Christopher L. Racey and Steve E. Lochmann. Year-class contribution of white crappies stocked into Lake Chicot, Arkansas. North American Journal of Fisheries Management, 22(4):1409–1415, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robillard:2001:SSP

[RM01] Steven R. Robillard and J. Ellen Marsden. Spawning substrate preferences of yellow perch along a sand–cobble shoreline in Southwestern Lake Michigan. North American Journal of Fisheries Management, 21(1):208–215, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rose:2008:FBE

[RMBZ08] Brien P. Rose, Matthew G. Mesa, and Gayle Barbin-Zydlewski. Field-based evaluations of horizontal flat-plate fish screens. North American Journal of Fisheries Management, 28(6):1702–1713, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Raborn:2002:ESR

[RMD02] Scott W. Raborn, Leandro E. Miranda, and M. Todd Driscoll. Effects of simulated removal of striped bass from a southeastern reservoir. North American Journal of Fisheries Management, 22(2):406–417, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rosenfeld:2002:ISS

[RMF^{+}02] Jordan S. Rosenfeld, Scott Macdonald, Donald Foster, Seth Amrhein, Bert Bales, Tim Williams, Fred Race, and Tom Livingstone. Importance of small streams as rearing habitat for coastal cutthroat trout. North American Journal of Fisheries Management, 22(1):177–187, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roth:2006:MBT

[RO06] Robert N. Roth, Jr. and David R. Orvos. A modified box trap design for sampling emigrating salmonid smolts from small streams. North American Journal of Fisheries Management, 26(1):170–174, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roni:2003:RBF

[Ron03] Philip Roni. Responses of benthic fishes and giant salamanders to placement of large woody debris in small Pacific Northwest streams. North American Journal of Fisheries Management, 23(4):1087–1097, November 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richards:2004:SCM

[ROS04] David C. Richards, Pat O’Connell, and Dianne Cazier Shinn. Simple control method to limit the spread of the New Zealand mudsnail Potamopyrgus antipodarum. North American Journal of Fisheries Management, 24(1):114–117, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reed:2009:RBS

[RP09] Jeffrey R. Reed and Donald L. Pereira. Relationships between shoreline development and nest site selection by black crappie and largemouth bass. North American Journal of Fisheries Management, 29(4):943–948, August 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reyes:2009:FRP

[RPD09] Ricardo Bórquez Reyes, Oscar Alberto Pombo, and Germán Ponce Díaz. Fishers’ reasons for poaching abalone (haliotidae): a study in the Baja California Peninsula, Mexico. North American Journal of Fisheries Management, 29(1):237–244, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rashleigh:2005:PHM

[RPJB05] Brenda Rashleigh, Rajbir Parmar, John M. Johnston, and M. Craig Barber. Predictive habitat models for the occurrence of stream fishes in the mid-Atlantic highlands. North American Journal of Fisheries Management, 25(4):1353– 1366, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Runge:2008:SDH

[RPM08] Jonathan P. Runge, James T. Peterson, and Christopher R. Martin. Survival and dispersal of hatchery-raised rainbow trout in a river basin undergoing urbanization. North American Journal of Fisheries Management, 28(3):745–757, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Riedel:2008:UCA

[RPW^{+}08] Ralf Riedel, Harriet Perry, James Warren, Alan Criss, and Thomas van Devender. Using conventional analysis in parallel with geographical information systems techniques to examine the distribution of brown shrimp in the Mississippi Sound. North American Journal of Fisheries Management, 28(5):1439–1449, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rowe:2009:FAR

[RPW09a] David C. Rowe, Clay L. Pierce, and Thomas F. Wilton. Fish assemblage relationships with physical habitat in wadeable Iowa streams. North American Journal of Fisheries Management, 29(5):1314–1332, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rowe:2009:PHF

[RPW09b] David C. Rowe, Clay L. Pierce, and Thomas F. Wilton. Physical habitat and fish assemblage relationships with landscape variables at multiple spatial scales in wadeable Iowa streams. North American Journal of Fisheries Management, 29(5): 1333–1351, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roberts:2005:AAT

[RR05] James J. Roberts and Frank J. Rahel. Accuracy of aerial telemetry in fisheries studies. North American Journal of Fisheries Management, 25(2):660–666, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rypel:2008:EPS

[RR08] Andrew L. Rypel and Tracy J. Richter. Empirical percentile standard weight equation for the blacktail redhorse. North American Journal of Fisheries Management, 28(6):

1843–1846, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rogillio:2007:GSM

[RRB^{+}07] Howard E. Rogillio, Ronald T. Ruth, Elizabeth H. Behrens, Cedric N. Doolittle, Whitney J. Granger, and James P. Kirk. Gulf sturgeon movements in the Pearl River drainage and the Mississippi Sound. North American Journal of Fisheries Management, 27(1):89–95, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rosenfeld:2008:ESC

[RRCJ08] Jordan S. Rosenfeld, Elizabeth Raeburn, Patrick C. Carrier, and Rachel Johnson. Effects of side channel structure on productivity of floodplain habitats for juvenile Coho salmon. North American Journal of Fisheries Management, 28(4): 1108–1119, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roberts:2007:DCG

[RRS07] Nathan M. Roberts, Charles F. Rabeni, and John S. Stanovick. Distinguishing centrarchid genera by use of lateral line scales. North American Journal of Fisheries Management, 27(1):215–219, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reinert:2006:EAT

[RSF06] Thomas R. Reinert, Carrie A. Straight, and Byron J. Freeman. Effectiveness of atimycin-A as a toxicant for control of invasive Asian swamp eels. North American Journal of Fisheries Management, 26(4):949–952, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rudershausen:2009:ECR

[RT09] Paul J. Rudershausen and Marc J. Turano. The effect of cull rings on catch rates of blue crabs Callinectes sapidus in the North Carolina trap fishery. North American Journal of Fisheries Management, 29(4):1152–1164, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Riley:2009:BSF

[RTBF09] Stephen C. Riley, Christopher P. Tatara, Barry A. Berejikian, and Thomas A. Flagg. Behavior of steelhead fry in a laboratory stream is affected by fish density but not rearing environment. North American Journal of Fisheries Management, 29(6):1806–1818, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ruetz:2007:SLF

[RUKR07] Carl R. Ruetz III, Donald G. Uzarski, Damon M. Krueger, and Edward S. Rutherford. Sampling a littoral fish assemblage: Comparison of small-mesh fyke netting and boat electrofishing. North American Journal of Fisheries Management, 27(3):825–831, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rulifson:2007:SDM

[Rul07] Roger A. Rulifson. Spiny dogfish mortality induced by gillnet and trawl capture and tag and release. North American Journal of Fisheries Management, 27(1):279–285, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rennie:2008:DEC

[RV08] Michael D. Rennie and Richard Verdon. Development and evaluation of condition indices for the lake whitefish. North American Journal of Fisheries Management, 28(5):1270– 1293, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roni:2008:ACS

[RVM^{+}08] Phil Roni, Dan Van Slyke, Bruce A. Miller, Joseph L. Ebersole, and George Pess. Adult Coho salmon and steelhead use of boulder weirs in Southwest Oregon streams. North American Journal of Fisheries Management, 28(3):970–978, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ramstad:2003:RTR

[RW03] Kristina M. Ramstad and Carol Ann Woody. Radio tag retention and tag-related mortality among adult sockeye salmon. North American Journal of Fisheries Management, 23(4): 978–982, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rulifson:2006:FBC

[RW06] Roger A. Rulifson and Brian L. Wall. Fish and blue crab passage through water control structures of a coastal bay lake. North American Journal of Fisheries Management, 26 (2):317–326, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roberts:2004:DIF

[RWBG04] Matt E. Roberts, James E. Wetzel II, Ronald C. Brooks, and James E. Garvey. Daily increment formation in otoliths of the redspotted sunfish. North American Journal of Fisheries Management, 24(1):270–274, February 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reinert:2004:UES

[RWJD04] Thomas R. Reinert, Ted A. Will, Cecil A. Jennings, and William T. Davin. Use of egg surrogates to estimate sampling efficiency of striped bass eggs in the Savannah River. North American Journal of Fisheries Management, 24(2):704–710, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ratcliff:2009:EEG

[RWZ09] Donald R. Ratcliff, Wayne A. Wurtsbaugh, and Joseph Zustak. Evaluating the effectiveness of grassbed treatments as habitat for juvenile black bass in a drawdown reservoir. North American Journal of Fisheries Management, 29(4): 1119–1129, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rypel:2008:IIA

[Ryp08] Andrew L. Rypel. An inexpensive image analysis system for fish otoliths. North American Journal of Fisheries Management, 28(1):193–197, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmetterling:2004:SMW

[SA04] David A. Schmetterling and Susan B. Adams. Summer movements within the fish community of a small montane stream. North American Journal of Fisheries Management, 24(4): 1163–1172, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sypitkowski:2008:HEB

[SABW08] Eben Sypitkowski, William G. Ambrose, Jr., Curtis Bohlen, and Joseph Warren. Harvest efficiency of bloodworms on Maine mudflats. North American Journal of Fisheries Management, 28(5):1506–1514, October 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spina:2005:DMR

[SAC05] Anthony P. Spina, Mark A. Allen, and Michael Clarke. Downstream migration, rearing abundance, and pool habitat associations of juvenile steelhead in the lower main stem of a South-Central California stream. North American Journal of Fisheries Management, 25(3):919–930, August 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2000:PDR

[SB00] Steve M. Sammons and Phillip W. Bettoli. Population dynamics of a reservoir sport fish community in response to hydrology. North American Journal of Fisheries Management, 20(3):791–800, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

StJohn:2004:MPA

[SB04] R. Todd St. John and William Patrick Black. Methods for predicting age-0 crappie year-class strength in J. Percy Priest Reservoir, Tennessee. North American Journal of Fisheries Management, 24(4):1300–1308, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schade:2005:DAN

[SB05a] Charles B. Schade and Scott A. Bonar. Distribution and abundance of nonnative fishes in streams of the Western United States. North American Journal of Fisheries Management, 25(4):1386–1394, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2005:CSL

[SB05b] Kregg M. Smith and Edward A. Baker. Characteristics of spawning lake sturgeon in the Upper Black River, Michigan. North American Journal of Fisheries Management, 25 (1):301–307, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schlechte:2006:PEH

[SB06a] J. Warren Schlechte and David L. Buckmeier. A pond evaluation of habituation as a means to reduce initial mortality associated with poststocking predation of hatchery-reared largemouth bass. North American Journal of Fisheries Management, 26(1):119–123, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sylvester:2006:CWS

[SB06b] Ryan M. Sylvester and Charles R. Berry, Jr. Comparison of white sucker age estimates from scales, pectoral fin rays, and otoliths. North American Journal of Fisheries Management, 26(1):24–31, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharpe:2007:GMD

[SBCH07] Cameron S. Sharpe, Brian R. Beckman, Kathleen A. Cooper, and Patrick L. Hulett. Growth modulation during juvenile rearing can reduce rates of residualism in the progeny of wild steelhead broodstock. North American Journal of Fisheries Management, 27(4):1355–1368, November 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2002:RVC

[SBIC02] Steve M. Sammons, Phillip W. Bettoli, Daniel A. Isermann, and Timothy N. Churchill. Recruitment variation of crappies in response to hydrology of Tennessee reservoirs. North American Journal of Fisheries Management, 22(4):1393– 1398, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sweeting:2003:RWC

[SBNN03] R. M. Sweeting, R. J. Beamish, D. J. Noakes, and C. M. Neville. Replacement of wild Coho salmon by hatchery-reared Coho salmon in the Strait of Georgia over the past three decades. North American Journal of Fisheries Management, 23(2):492–502, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2005:ARH

[SBR^{+}05] G. F. Smith, D. G. Bruce, E. B. Roach, A. Hansen, R. I. E. Newell, and A. M. McManus. Assessment of recent habitat conditions of Eastern oyster Crassostrea virginica bars in mesohaline Chesapeake Bay. North American Journal of Fisheries Management, 25(4):1569–1590, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Soupir:2006:CCS

[SBSL06] Craig A. Soupir, Michael L. Brown, Cliff C. Stone, and John P. Lott. Comparison of creel survey methods on Missouri River reservoirs. North American Journal of Fisheries Management, 26(2):338–350, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skalski:2009:MRM

[SBT^{+}09] John R. Skalski, Rebecca A. Buchanan, Richard L. Townsend, Tracey W. Steig, and Steven Hemstrom. A multiple-release model to estimate route-specific and dam passage survival at a hydroelectric project. North American Journal of Fisheries Management, 29(3):670–679, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sutton:2000:OHI

[SCB00] Ronald J. Sutton, Colleen A. Caldwell, and Vicki S. Blazer. Observations of health indices used to monitor a tailwater trout fishery. North American Journal of Fisheries Management, 20(1):267–275, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schreiner:2008:MPC

[SCD^{+}08] Donald R. Schreiner, Ken I. Cullis, Michael C. Donofrio, Gregory J. Fischer, Laura Hewitt, Karen G. Mumford, Dennis M. Pratt, Henry R. Quinlan, and Steven J. Scott. Management perspectives on coaster brook trout rehabilitation in the Lake Superior Basin. North American Journal of Fisheries Management, 28(5):1350–1364, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmetterling:2000:RCF

[Sch00] David A. Schmetterling. Redd characteristics of fluvial westslope cutthroat trout in four tributaries to the Blackfoot River, Montana. North American Journal of Fisheries Management, 20(3):776–783, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmetterling:2001:SMF

[Sch01] David A. Schmetterling. Seasonal movements of fluvial westslope cutthroat trout in the Blackfoot River Drainage, Montana. North American Journal of Fisheries Management, 21 (3):507–520, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schupp:2002:WDM

[Sch02] Dennis H. Schupp. What does Mt. Pinatubo have to do with walleyes? North American Journal of Fisheries Management, 22(3):1014–1020, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmetterling:2003:RFR

[Sch03] David A. Schmetterling. Reconnecting a fragmented river: Movements of westslope cutthroat trout and bull trout after transport upstream of Milltown Dam, Montana. North American Journal of Fisheries Management, 23(3):721–731, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schleiger:2004:FHA

[Sch04a] Steve L. Schleiger. Fish health assessment index study of four reservoirs in West-Central Georgia. North American Journal of Fisheries Management, 24(4):1173–1188, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2004:SSA

[Sch04b] Randall D. Schultz. Size-specific angling return rates of white bass in Kansas reservoirs. North American Journal of Fisheries Management, 24(2):719–723, May 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schreiner:2008:CBT

[Sch08] Donald R. Schreiner. Coaster brook trout rehabilitation in Lake Superior: Introduction. North American Journal of Fisheries Management, 28(5):1305–1306, August 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shoup:2003:CSC

[SCHK03] Daniel E. Shoup, Robert E. Carlson, Robert T. Heath, and Mark W. Kershner. Comparison of the species composition, catch rate, and length distribution of the catch from trap nets with three different mesh and throat size combinations. North American Journal of Fisheries Management, 23(2):462–469, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spencer:2002:RWP

[SCMF02] Stephen C. Spencer, Peter J. Colby, Walter T. Momot, and Mike Fruetel. Response of a walleye population to pulse fishing in Henderson Lake, Ontario. North American Journal of Fisheries Management, 22(3):842–851, August 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skinner:2006:ETM

[SCPC06] Marc A. Skinner, Simon C. Courtenay, W. Roy Parker, and R. Allen Curry. Evaluation of techniques for the marking of mummichogs with emphasis on visible implant elastomer. North American Journal of Fisheries Management, 26(4): 1003–1010, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sutton:2005:SOT

[SD05] Stephen G. Sutton and Robert B. Ditton. The substitutability of one type of fishing for another. North American Journal of Fisheries Management, 25(2):536–546, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schramm:2006:SLB

[SD06] Harold L. Schramm, Jr. and Jonathan G. Davis. Survival of largemouth bass from populations infected with largemouth bass virus and subjected to simulated tournament conditions.

North American Journal of Fisheries Management, 26(4): 826–832, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2000:BSI

[SDPB00] Lance W. Smith, Eric Dittmer, Marc Prevost, and Donald R. Burt. Breaching of a small irrigation dam in Oregon: a case history. North American Journal of Fisheries Management, 20(1):205–219, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seamons:2009:UPA

[SDSQ09] Todd R. Seamons, Michael B. Dauer, John Sneva, and Thomas P. Quinn. Use of parentage assignment and DNA genotyping to validate scale analysis for estimating steelhead age and spawning history. North American Journal of Fisheries Management, 29(2):396–403, April 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schill:2000:HEI

[SE00] Daniel J. Schill and F. Steven Elle. Healing of electroshockinduced hemorrhages in hatchery rainbow trout. North American Journal of Fisheries Management, 20(3):730–736, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stockwell:2009:SCR

[SEB^{+}09] Jason D. Stockwell, Mark P. Ebener, Jeff A. Black, Owen T. Gorman, Thomas R. Hrabik, Ronald E. Kinnunen, William P. Mattes, Jason K. Oyadomari, Stephen T. Schram, Donald R. Schreiner, Michael J. Seider, Shawn P. Sitar, and Daniel L. Yule. A synthesis of cisco recovery in Lake Superior: Implications for native fish rehabilitation in the Laurentian Great Lakes. North American Journal of Fisheries Management, 29(3):626–652, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Selberg:2001:HNR

[SEC01] Carrie D. Selberg, Lisa A. Eby, and Larry B. Crowder. Hypoxia in the Neuse River Estuary: Responses of blue crabs and crabbers. North American Journal of Fisheries Management, 21(2):358–366, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Strakosh:2005:RWW

[SEGG05] Timothy R. Strakosh, Jeffery L. Eitzmann, Keith B. Gido, and Christopher S. Guy. The response of water willow Justicia americana to different water inundation and desiccation regimes. North American Journal of Fisheries Management, 25(4):1476–1485, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swansburg:2004:HTM

[SEJCC04] Erin Swansburg, Nassir El-Jabi, Daniel Caissie, and Gérald Chaput. Hydrometeorological trends in the Miramichi River, Canada: Implications for Atlantic salmon growth. North American Journal of Fisheries Management, 24(2):561–576, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stein:2004:ASM

[SFS04] Andrew B. Stein, Kevin D. Friedland, and Michael Sutherland. Atlantic sturgeon marine bycatch and mortality on the Continental Shelf of the Northeast United States. North American Journal of Fisheries Management, 24(1):171–183, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2007:YCF

[SFUS07] Brian M. Smith, John M. Farrell, H. Brian Underwood, and Stephen J. Smith. Year-class formation of Upper St. Lawrence River northern pike. North American Journal of Fisheries Management, 27(2):481–491, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2007:BIF

[SGB07] Douglas W. Schultz, James E. Garvey, and Ronald C. Brooks. Backwater immigration by fishes through a water control structure: Implications for connectivity and restoration. North American Journal of Fisheries Management, 27(1): 172–180, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schreer:2009:ICB

[SGD09] Jason F. Schreer, Jason Gokey, and Victor J. DeGhett. The incidence and consequences of barotrauma in fish in the St.

Lawrence River. North American Journal of Fisheries Management, 29(6):1707–1713, December 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schramm:2003:IEQ

[SGG03] Harold L. Schramm, Jr., Patrick D. Gerard, and Duane A. Gill. The importance of environmental quality and catch potential to fishing site selection by freshwater anglers in Mississippi. North American Journal of Fisheries Management, 23 (2):512–522, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Santucci:2005:EML

[SGP05] Victor J. Santucci, Jr., Stephen R. Gephard, and Stephen M. Pescitelli. Effects of multiple low-head dams on fish, macroinvertebrates, habitat, and water quality in the Fox River, Illinois. North American Journal of Fisheries Management, 25 (3):975–992, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seiler:2009:MPH

[SGPK09] Steven M. Seiler, Kelly Gunnell, Margaret B. Ptacek, and Ernest R. Keeley. Morphological patterns of hybridization between Yellowstone cutthroat trout and introduced rainbow trout in the South Fork of the Snake River Watershed, Idaho and Wyoming. North American Journal of Fisheries Management, 29(6):1529–1539, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2002:CIG

[SGR02] Randall D. Schultz, Christopher S. Guy, and Douglas A. Robinson, Jr. Comparative influences of gizzard shad catch rates and reservoir hydrology on recruitment of white bass in Kansas reservoirs. North American Journal of Fisheries Management, 22(2):671–676, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schaeffer:2002:PBB

[SH02a] Jeffrey S. Schaeffer and Elizabeth M. Hoffman. Performance of barbed and barbless hooks in a marine recreational fishery. North American Journal of Fisheries Management, 22 (2):229–235, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spier:2002:ETG

[SH02b] Timothy W. Spier and Roy C. Heidinger. Effect of turbidity on growth of black crappie and white crappie. North American Journal of Fisheries Management, 22(4):1438– 1441, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schwanke:2004:ETG

[SH04a] Craig J. Schwanke and Wayne A. Hubert. Evaluation of three gears for sampling spawning populations of rainbow trout in a large Alaskan River. North American Journal of Fisheries Management, 24(3):1078–1082, August 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schwartz:2004:UPA

[SH04b] John S. Schwartz and Edwin E. Herricks. Use of prepositioned areal electrofishing devices with rod electrodes in small streams. North American Journal of Fisheries Management, 24(4):1330–1340, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shardlow:2004:ACA

[SH04c] T. F. Shardlow and K. D. Hyatt. Assessment of the counting accuracy of the vaki infrared counter on chum salmon. North American Journal of Fisheries Management, 24(1):249–252, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schoenebeck:2005:ECW

[SH05a] Casey W. Schoenebeck and Michael J. Hansen. Electrofishing catchability of walleyes, largemouth bass, smallmouth bass, northern pike, and muskellunge in Wisconsin lakes. North American Journal of Fisheries Management, 25(4): 1341–1352, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2005:CSB

[SH05b] Randall D. Schultz and Daniel E. Haines. Comparison of seasonal bluegill catch rates and size distributions obtained with trap nets and electrofishing in a large, heated impoundment. North American Journal of Fisheries Management, 25 (1):220–224, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sanderson:2007:AGC

[SH07] Travis B. Sanderson and Wayne A. Hubert. Assessment of gaseous CO_{2} and AQUI-S as anesthetics when surgically implanting radio transmitters into cutthroat trout. North American Journal of Fisheries Management, 27(4):1053– 1057, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stolarski:2008:EPF

[SH08] Jason T. Stolarski and Kyle J. Hartman. An evaluation of the precision of fin ray, otolith, and scale age determinations for brook trout. North American Journal of Fisheries Management, 28(6):1790–1795, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shardlow:2001:UUS

[Sha01] T. F. Shardlow. Using underwater still photography to estimate salmon escapement in streams. North American Journal of Fisheries Management, 21(3):699–702, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shardlow:2004:UTL

[Sha04] Tom Shardlow. Using time-lapsed video to estimate survey life for area-under-the-curve methods of escapement estimation. North American Journal of Fisheries Management, 24 (4):1413–1420, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sass:2004:RDD

[SHB^{+}04] Greg G. Sass, Steven W. Hewett, T. Douglas Beard, Jr., Andrew H. Fayram, and James F. Kitchell. The role of density dependence in growth patterns of ceded territory walleye populations of Northern Wisconsin: Effects of changing management regimes. North American Journal of Fisheries Management, 24(4):1262–1278, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sigourney:2005:EPT

[SHD^{+}05] Douglas B. Sigourney, Gregg E. Horton, Todd L. Dubreuil, Aimee M. Varaday, and Benjamin H. Letcher. Electroshocking and PIT tagging of juvenile Atlantic salmon: Are there interactive effects on growth and survival? North American Journal of Fisheries Management, 25(3):1016–1021, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shepard:2004:FMI

[She04] Bradley B. Shepard. Factors that may be influencing nonnative brook trout invasion and their displacement of native westslope cutthroat trout in three adjacent Southwestern Montana streams. North American Journal of Fisheries Management, 24(3):1088–1100, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmalz:2002:LTM

[SHH^{+}02] Patrick J. Schmalz, Michael J. Hansen, Mark E. Holey, Patrick C. McKee, and Michael L. Toneys. Lake trout movements in Northwestern Lake Michigan. North American Journal of Fisheries Management, 22(3):737–749, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmalz:2004:ERL

[SHH^{+}04] Patrick J. Schmalz, Michael J. Hansen, Mark E. Holey, Patrick C. McKee, and Michael L. Toneys. Effect of rewards on lake trout tag returns in Northwestern Lake Michigan. North American Journal of Fisheries Management, 24 (1):1–6, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Starr:2000:TTT

[SHJ00] Richard M. Starr, John N. Heine, and Korie A. Johnson. Techniques for tagging and tracking deepwater rockfishes. North American Journal of Fisheries Management, 20(3): 597–609, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stranko:2008:BTD

[SHM^{+}08] Scott A. Stranko, Robert H. Hilderbrand, Raymond P. Morgan II, Mark W. Staley, Andrew J. Becker, Ann Roseberry-Lincoln, Elgin S. Perry, and Paul T. Jacobson. Brook trout declines with land cover and temperature changes in Maryland. North American Journal of Fisheries Management, 28 (5):1223–1232, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Simpkins:2003:ESA

[SHMR03] Darin G. Simpkins, Wayne A. Hubert, Carlos Martinez del Rio, and Daniel C. Rule. Effect of swimming activity on relative weight and body composition of juvenile rainbow trout. North American Journal of Fisheries Management, 23(2): 283–289, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sommer:2005:HUS

[SHN05] Ted R. Sommer, William C. Harrell, and Matthew L. Nobriga. Habitat use and stranding risk of juvenile Chinook salmon on a seasonal floodplain. North American Journal of Fisheries Management, 25(4):1493–1504, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schueller:2008:MSW

[SHN08] Amy M. Schueller, Michael J. Hansen, and Steven P. Newman. Modeling the sustainability of walleye populations in Northern Wisconsin lakes. North American Journal of Fisheries Management, 28(6):1916–1927, December 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schueller:2005:DDW

[SHNE05] Amy M. Schueller, Michael J. Hansen, Steven P. Newman, and Clayton J. Edwards. Density dependence of walleye maturity and fecundity in Big Crooked Lake, Wisconsin, 1997– 2003. North American Journal of Fisheries Management, 25 (3):841–847, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schuett-Hames:2000:PGS

[SHPCQ00] Dave E. Schuett-Hames, N. Phil Peterson, Robert Conrad, and Thomas P. Quinn. Patterns of gravel scour and fill after spawning by chum salmon in a Western Washington stream. North American Journal of Fisheries Management, 20(3): 610–617, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Simpkins:2000:ESF

[SHW00] Darin G. Simpkins, Wayne A. Hubert, and Thomas A. Wesche. Effects of a spring flushing flow on the distribution of radio-tagged juvenile rainbow trout in a Wyoming tailwater. North American Journal of Fisheries Management, 20 (2):546–551, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sakaris:2008:VDR

[SI08] Peter C. Sakaris and Elise R. Irwin. Validation of daily ring deposition in the otoliths of age-0 channel catfish. North American Journal of Fisheries Management, 28(1):212–218, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2002:VPC

[SIB02] Steve M. Sammons, Daniel A. Isermann, and Phillip W. Bettoli. Variation in population characteristics and gear selection between black and white crappies in Tennessee reservoirs: Potential effects on management decisions. North American Journal of Fisheries Management, 22(3):863–869, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sakaris:2006:CNI

[SIJH06] Peter C. Sakaris, Elise R. Irwin, Jeffrey C. Jolley, and Donald Harrison. Comparison of native and introduced flathead catfish populations in Alabama and Georgia: Growth, mortality, and management. North American Journal of Fisheries Management, 26(4):867–874, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stanfield:2003:FIR

[SJ03] Les Stanfield and Michael L. Jones. Factors influencing rearing success of Atlantic salmon stocked as fry and parr in Lake Ontario tributaries. North American Journal of Fisheries Management, 23(4):1175–1183, November 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shephard:2005:CCM

[SJ05] Samuel Shephard and Donald C. Jackson. Channel catfish maturation in Mississippi streams. North American Journal of Fisheries Management, 25(4):1467–1475, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slaughter:2008:GBS

[SJ08] Joe E. Slaughter IV and Brad Jacobson. Gape: Body size relationship of flathead catfish. North American Journal of Fisheries Management, 28(1):198–202, February 2008. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stancill:2002:SMP

[SJP02] Wayne Stancill, George R. Jordan, and Craig P. Paukert. Seasonal migration patterns and site fidelity of adult paddlefish in Lake Francis Case, Missouri River. North American Journal of Fisheries Management, 22(3):815–824, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sakaris:2005:RSI

[SJP05] Peter C. Sakaris, Roman V. Jesien, and Alfred E. Pinkney. Retention of surgically implanted ultrasonic transmitters in the brown bullhead catfish. North American Journal of Fisheries Management, 25(3):822–826, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seaborn:2000:DBC

[SJS00] Gloria T. Seaborn, Michael L. Jahncke, and Theodore I. J. Smith. Differentiation between cultured hybrid striped bass and wild striped bass and hybrid bass using fatty acid profiles. North American Journal of Fisheries Management, 20 (3):618–626, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Suski:2003:PCL

[SKM^{+}03] Cory D. Suski, Shaun S. Killen, Michael B. Morrissey, Susan G. Lund, and Bruce L. Tufts. Physiological changes in largemouth bass caused by live-release angling tournaments in Southeastern Ontario. North American Journal of Fisheries Management, 23(3):760–769, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharma:2001:EIC

[SL01] Khem R. Sharma and Ping Sun Leung. Economic impacts of catch reallocation from the commercial fishery to the recreational fishery in Hawaii. North American Journal of Fisheries Management, 21(1):125–134, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schneider:2002:UWS

[SL02a] James C. Schneider and Roger N. Lockwood. Use of walleye stocking, antimycin treatments, and catch-and-release angling regulations to increase growth and length of stunted bluegill populations in Michigan lakes. North American Journal of Fisheries Management, 22(3):1041–1052, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stone:2002:UML

[SL02b] Clifton Stone and John Lott. Use of a minimum length limit to manage walleyes in Lake Francis Case, South Dakota. North American Journal of Fisheries Management, 22(3): 975–984, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shroyer:2009:DDS

[SL09] Steven M. Shroyer and Dale E. Logsdon. Detection distances of selected radio and acoustic tags in Minnesota lakes and rivers. North American Journal of Fisheries Management, 29(4):876–884, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sweka:2006:ERS

[SLB^{+}06] John A. Sweka, Christopher M. Legault, Kenneth F. Beland, Joan Trial, and Michael J. Millard. Evaluation of removal sampling for basinwide assessment of Atlantic salmon. North American Journal of Fisheries Management, 26(4):995–1002, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stehr:2009:EEF

[SLB^{+}09] Carla M. Stehr, Tiffany L. Linbo, David H. Baldwin, Nathaniel L. Scholz, and John P. Incardona. Evaluating the effects of forestry herbicides on fish development using rapid phenotypic screens. North American Journal of Fisheries Management, 29(4):975–984, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schill:2007:AER

[SLEM07] Daniel J. Schill, George W. LaBar, F. Steven Elle, and Elizabeth R. J. M. Mamer. Angler exploitation of redband trout in eight Idaho desert streams. North American Journal of Fisheries Management, 27(2):665–669, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Saila:2002:MCE

[SLG02] S. B. Saila, D. F. Landers, Jr., and P. Geoghegan. Model comparisons for estimating the relationship between catch and soak time for the American lobster trap fishery. North American Journal of Fisheries Management, 22(3):943–949, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scace:2007:EST

[SLN07] Justin G. Scace, Benjamin H. Letcher, and John Noreika. An efficient smolt trap for sandy and debris-laden streams. North American Journal of Fisheries Management, 27(4): 1276–1286, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sollid:2004:PMC

[SLS^{+}04] Sarah A. Sollid, Harriet V. Lorz, Donald G. Stevens, Paul W. Reno, and Jerri L. Bartholomew. Prevalence of Myxobolus cerebralis at juvenile salmonid acclimation sites in Northeastern Oregon. North American Journal of Fisheries Management, 24(1):146–153, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Silver:2009:DLU

[SLW09] Gregory S. Silver, Christina W. Luzier, and Timothy A. Whitesel. Detection and longevity of uncured and cured visible implant elastomer tags in larval Pacific lampreys. North American Journal of Fisheries Management, 29(5): 1496–1501, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmetterling:2000:ADF

[SM00a] David A. Schmetterling and David H. McEvoy. Abundance and diversity of fishes migrating to a hydroelectric dam in Montana. North American Journal of Fisheries Management, 20(3):711–719, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shroyer:2000:RBA

[SM00b] Steven M. Shroyer and Thomas S. McComish. Relationship between alewife abundance and yellow perch recruitment in Southern Lake Michigan. North American Journal of Fisheries Management, 20(1):220–225, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stunz:2006:CRM

[SM06] Gregory W. Stunz and David A. McKee. Catch-and-release mortality of spotted seatrout in Texas. North American Journal of Fisheries Management, 26(4):843–848, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schooley:2007:SER

[SM07a] Jason D. Schooley and Paul C. Marsh. Stocking of endangered razorback suckers in the Lower Colorado River Basin over three decades: 1974–2004. North American Journal of Fisheries Management, 27(1):43–51, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stapanian:2007:ELT

[SM07b] Martin A. Stapanian and Charles P. Madenjian. Evidence that lake trout served as a buffer against sea lamprey predation on burbot in Lake Erie. North American Journal of Fisheries Management, 27(1):238–245, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2009:VGS

[SM09a] Steven M. Sammons and Michael J. Maceina. Variation in growth and survival of bluegills and redbreast sunfish in Georgia rivers. North American Journal of Fisheries Management, 29(1):101–108, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shepherd:2009:ESB

[SM09b] Michael D. Shepherd and Michael J. Maceina. Effects of striped bass stocking on largemouth bass and spotted bass in Lewis Smith Lake, Alabama. North American Journal of Fisheries Management, 29(5):1232–1241, October 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2009:FAS

[SM09c] Stephen J. Smith and J. Ellen Marsden. Factors affecting sea lamprey egg survival. North American Journal of Fisheries Management, 29(4):859–868, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stormer:2009:HUH

[SM09d] David G. Stormer and Michael J. Maceina. Habitat use, home range, and movement of shoal bass in Alabama. North American Journal of Fisheries Management, 29(3):604–613, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shrader:2003:PDC

[SMB03] Terry M. Shrader, Barron Moody, and Mary Buckman. Population dynamics of channel catfish in Brownlee Reservoir and the Snake River, Oregon. North American Journal of Fisheries Management, 23(3):822–834, August 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sutphin:2007:SPS

[SMB07] Zachary A. Sutphin, Christopher A. Myrick, and Mandi M. Brandt. Swimming performance of Sacramento splittail injected with subcutaneous marking agents. North American Journal of Fisheries Management, 27(4):1378–1382, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2003:EFA

[SMBM03] Andrew A. Schultz, O. Eugene Maughan, Scott A. Bonar, and William J. Matter. Effects of flooding on abundance of native and nonnative fishes downstream from a small impoundment. North American Journal of Fisheries Management, 23(2):503–511, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2009:EGR

[SMF^{+}09] Joseph M. Smith, Martha E. Mather, Holly J. Frank, Robert M. Muth, John T. Finn, and Stephen D. McCormick. Evaluation of a gastric radio tag insertion technique for anadromous river herring. North American Journal of Fisheries Management, 29(2):367–377, April 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skalski:2002:ETO

[SMH02] John R. Skalski, Dilip Mathur, and Paul G. Heisey. Effects of turbine operating efficiency on smolt passage survival. North American Journal of Fisheries Management, 22 (4):1193–1200, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2003:IRC

[SMH^{+}03] Steven G. Smith, William D. Muir, Eric E. Hockersmith, Richard W. Zabel, Ritchie J. Graves, Chris V. Ross, William P. Connor, and Billy D. Arnsberg. Influence of river conditions on survival and travel time of Snake River subyearling fall Chinook salmon. North American Journal of Fisheries Management, 23(4):939–961, August 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sweka:2007:JAS

[SMM^{+}07] John A. Sweka, Jerre Mohler, Michael J. Millard, Thomas Kehler, Andrew Kahnle, Kathryn Hattala, Gregg Kenney, and Amanda Higgs. Juvenile Atlantic sturgeon habitat use in Newburgh and Haverstraw Bays of the Hudson River: Implications for population monitoring. North American Journal of Fisheries Management, 27(4):1058–1067, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sherwood:2004:PRL

[SMP04] Edward T. Sherwood, Debra J. Murie, and Daryl C. Parkyn. Postrelease rate of loss of juvenile red drum stocked out of season in the Chassahowitzka National Wildlife Refuge, Florida. North American Journal of Fisheries Management, 24(4):1469–1479, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slipke:2002:USP

[SMPM02] Jeffrey W. Slipke, Allen D. Martin, John Pitlo, Jr., and Michael J. Maceina. Use of the spawning potential ratio for the Upper Mississippi River channel catfish fishery. North American Journal of Fisheries Management, 22(4): 1295–1300, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steffe:2008:SAP

[SMR08] Aldo S. Steffe, Jeffrey J. Murphy, and Dennis D. Reid. Supplemented access point sampling designs: a cost-effective way of improving the accuracy and precision of fishing effort and harvest estimates derived from recreational fishing surveys. North American Journal of Fisheries Management, 28(4): 1001–1008, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Saiki:2001:EAD

[SMSM01] Michael K. Saiki, Barbara A. Martin, Steven E. Schwarzbach, and Thomas W. May. Effects of an agricultural drainwater bypass on fishes inhabiting the Grassland Water District and the Lower San Joaquin River, California. North American Journal of Fisheries Management, 21(3):624–635, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shepard:2005:SCW

[SMU05] Bradley B. Shepard, Bruce E. May, and Wendi Urie. Status and conservation of westslope cutthroat trout within the Western United States. North American Journal of Fisheries Management, 25(4):1426–1440, November 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stone:2006:DUV

[SMW06] Jen Stone, Michelle McGree, and Timothy A. Whitesel. Detection of uncured visible implant elastomer tags in larval Pacific lampreys. North American Journal of Fisheries Management, 26(1):142–146, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2002:FAT

[SMWS02] Steven G. Smith, William D. Muir, John G. Williams, and John R. Skalski. Factors associated with travel time and survival of migrant yearling Chinook salmon and steelhead in the Lower Snake River. North American Journal of Fisheries Management, 22(2):385–405, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sutton:2002:TRO

[SN02] Trent M. Sutton and John J. Ney. Trophic resource overlap between age-0 striped bass and largemouth bass in Smith Mountain Lake, Virginia. North American Journal of Fisheries Management, 22(4):1250–1259, November 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schofield:2007:TRN

[SN07] Pamela J. Schofield and Leo G. Nico. Toxicity of 5% rotenone to nonindigenous Asian swamp eels. North American Journal of Fisheries Management, 27(2):453–459, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Saila:2002:DLT

[SNO02] Saul B. Saila, Scott W. Nixon, and Candace A. Oviatt. Does lobster trap bait influence the Maine inshore trap fishery? North American Journal of Fisheries Management, 22(2):

602–605, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmidt:2009:OAE

[SOM09] Robert E. Schmidt, Catherine M. O’Reilly, and Daniel Miller. Observations of American eels using an upland passage facility and effects of passage on the population structure. North American Journal of Fisheries Management, 29(3):715–720, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2005:SBR

[SOR05] Scott M. Smith, John S. Odenkirk, and Stephen J. Reeser. Smallmouth bass recruitment variability and its relation to stream discharge in three Virginia rivers. North American Journal of Fisheries Management, 25(3):1112–1121, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schaller:2007:AHI

[SP07] Howard A. Schaller and Charles E. Petrosky. Assessing hydrosystem influence on delayed mortality of Snake River stream-type Chinook salmon. North American Journal of Fisheries Management, 27(3):810–824, August 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmetterling:2002:ETD

[SPL02] David A. Schmetterling, Ronald W. Pierce, and Bradley W. Liermann. Efficacy of three Denil fish ladders for low-flow fish passage in two tributaries to the Blackfoot River, Montana. North American Journal of Fisheries Management, 22 (3):929–933, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharma:2003:TEI

[SPL03] Khem R. Sharma, Naresh C. Pradhan, and Pingsun Leung. Technological and economic interrelationships in Hawaii’s troll and handline fisheries. North American Journal of Fisheries Management, 23(4):869–882, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2006:DPM

[SPM06] Steve M. Sammons, David G. Partridge, and Michael J. Maceina. Differences in population metrics between bluegill and redear sunfish: Implications for the effectiveness of harvest restrictions. North American Journal of Fisheries Management, 26(3):777–787, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scruton:2005:SRJ

[SPR^{+}05a] D. A. Scruton, C. J. Pennell, M. J. Robertson, L. M. N. Ollerhead, K. D. Clarke, K. Alfredsen, A. Harby, and R. S. McKinley. Seasonal response of juvenile Atlantic salmon to experimental hydropeaking power generation in Newfoundland, Canada. North American Journal of Fisheries Management, 25(3):964–974, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sprankle:2005:IMP

[Spr05b] Kenneth Sprankle. Interdam movements and passage attraction of American shad in the Lower Merrimack River main stem. North American Journal of Fisheries Management, 25 (4):1456–1466, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stables:2005:ATS

[SPR05c] T. Brock Stables, Chris J. Perrin, and Marvin L. Rosenau. Acoustic and trawl surveys to locate eulachon aggregations in the Lower Fraser River, British Columbia. North American Journal of Fisheries Management, 25(2):675–688, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2002:EMR

[SR02a] Randall D. Schultz and Douglas A. Robinson, Jr. Exploitation and mortality rates of white bass in Kansas reservoirs. North American Journal of Fisheries Management, 22(2): 652–658, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2002:IAB

[SR02b] Randall D. Schultz and Douglas A. Robinson, Jr. Influence of angler behavior on white bass catch rates and exploitation in Kansas reservoirs. North American Journal of Fisheries Management, 22(4):1146–1150, November 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schreer:2005:SPB

[SRGC05] Jason F. Schreer, Dayna M. Resch, Malachy L. Gately, and Steven J. Cooke. Swimming performance of brook trout after simulated catch-and-release angling: Looking for air exposure thresholds. North American Journal of Fisheries Management, 25(4):1513–1517, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sutton:2000:MAS

[SRN00] Trent M. Sutton, Kenneth A. Rose, and John J. Ney. A model analysis of strategies for enhancing stocking success of landlocked striped bass populations. North American Journal of Fisheries Management, 20(4):841–859, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sharp:2009:QUM

[SS09] Renée Sharp and U. Rashid Sumaila. Quantification of U.S. marine fisheries subsidies. North American Journal of Fisheries Management, 29(1):18–32, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2001:PFG

[SSB01] Summer L. W. Schultz, Ed L. Steinkoenig, and Bonnie L. Brown. Ploidy of feral grass carp in the Chesapeake Bay watershed. North American Journal of Fisheries Management, 21(1):96–101, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shemai:2007:CBH

[SSC07] Barak Shemai, Rossana Sallenave, and David E. Cowley. Competition between hatchery-raised Rio Grande cutthroat trout and wild brown trout. North American Journal of Fisheries Management, 27(1):315–325, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schoenebeck:2005:EBN

[SSG05] Casey W. Schoenebeck, Timothy R. Strakosh, and Christopher S. Guy. Effect of block net use and time of sampling on backpack electrofishing catches in three Kansas reservoirs. North American Journal of Fisheries Management, 25(2): 604–608, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skibo:2008:ESD

[SSP08] Karen M. Skibo, Carl J. Schwarz, and Randall M. Peterman. Evaluation of sampling designs for red sea urchins Strongylocentrotus franciscanus in British Columbia. North American Journal of Fisheries Management, 28(1):219–230, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Saiget:2007:SMB

[SSR07] D. A. Saiget, M. R. Sloat, and G. H. Reeves. Spawning and movement behavior of migratory coastal cutthroat trout on the Western Copper River Delta, Alaska. North American Journal of Fisheries Management, 27(3):1029–1040, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Strobel:2009:SCM

[SSR09] Burke Strobel, Daniel R. Shively, and Brett B. Roper. Salmon carcass movements in forest streams. North American Journal of Fisheries Management, 29(3):702–714, June 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sink:2007:COU

[SSS07] Todd D. Sink, Richard J. Strange, and R. Eric Sawyers. Clove oil used at lower concentrations is less effective than MS-222 at reducing cortisol stress responses in anesthetized rainbow trout. North American Journal of Fisheries Management, 27 (1):156–161, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scholefield:2008:SVS

[SSS08] Ronald J. Scholefield, Karen S. Slaght, and Brian E. Stephens. Seasonal variation in sensitivity of larval sea lampreys to the lampricide 3-trifluoromethyl-4-nitrophenol. North American Journal of Fisheries Management, 28(5): 1609–1617, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sloat:2005:IST

[SSWC05] Matthew R. Sloat, Bradley B. Shepard, Robert G. White, and Steve Carson. Influence of stream temperature on the spatial distribution of westslope cutthroat trout growth potential within the Madison River Basin, Montana. North American Journal of Fisheries Management, 25(1):225–237, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spina:2000:PSD

[ST00] Anthony P. Spina and Daniel R. Tormey. Postfire sediment deposition in geographically restricted steelhead habitat. North American Journal of Fisheries Management, 20 (2):562–569, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stangl:2001:ERS

[Sta01] Michael J. Stangl. An electrofishing raft for sampling intermediate-size waters with restricted boat access. North American Journal of Fisheries Management, 21(3):679–682, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steedman:2003:LFR

[Ste03] Robert J. Steedman. Littoral fish response to experimental logging around small boreal shield lakes. North American Journal of Fisheries Management, 23(2):392–403, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stone:2005:EBH

[Sto05] Dennis M. Stone. Effect of baiting on hoop net catch rates of endangered humpback chub. North American Journal of Fisheries Management, 25(2):640–645, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2005:SNP

[STSS05] Christian T. Smith, William D. Templin, James E. Seeb, and Lisa W. Seeb. Single nucleotide polymorphisms provide rapid and accurate estimates of the proportions of U.S. and Canadian Chinook salmon caught in Yukon River fisheries. North American Journal of Fisheries Management, 25(3):944–953, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stuart:2006:VOD

[Stu06] Ivor G. Stuart. Validation of otoliths for determining age of golden perch, a long-lived freshwater fish of Australia. North American Journal of Fisheries Management, 26(1): 52–55, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2002:IAH

[Sul02] Michael G. Sullivan. Illegal angling harvest of walleyes protected by length limits in Alberta. North American Journal of Fisheries Management, 22(3):1053–1063, August 2002. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2003:AMW

[Sul03a] Michael G. Sullivan. Active management of walleye fisheries in Alberta: Dilemmas of managing recovering fisheries. North American Journal of Fisheries Management, 23(4): 1343–1358, November 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2003:EWC

[Sul03b] Michael G. Sullivan. Exaggeration of walleye catches by Alberta anglers. North American Journal of Fisheries Management, 23(2):573–580, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sutton:2004:EAS

[SVP^{+}04] Trent M. Sutton, Eric T. Volkman, Kevin L. Pangle, Deborah A. Rajchel, and Jeremy P. Duehr. Effects of absorbable suture Strand diameter on retention of external radio transmitters by juvenile lake sturgeon. North American Journal of Fisheries Management, 24(4):1404–1408, November 2004.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slawski:2008:ETS

[SVPP08] Thomas M. Slawski, Francis M. Veraldi, Stephen M. Pescitelli, and Michael J. Pauers. Effects of tributary spatial position, urbanization, and multiple low-head dams on warmwater fish community structure in a Midwestern stream. North American Journal of Fisheries Management, 28(4): 1020–1035, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Svendsen:2009:OFD

[SVS^{+}09] Tore C. Svendsen, Katrin Vorkamp, Jon C. Svendsen, Kim Aarestrup, and Jens-Ole Frier. Organochlorine fingerprinting to determine foraging areas of sea-ranched Atlantic salmon: a case study from Denmark. North American Journal of Fisheries Management, 29(3):598–603, June 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schisler:2008:AHD

[SVW08a] George J. Schisler, Nicole K. M. Vieira, and Peter G. Walker. Application of household disinfectants to control New Zealand mudsnails. North American Journal of Fisheries Management, 28(4):1172–1176, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slacum:2008:VAC

[SVW^{+}08b] H. Ward Slacum, Jr., Jon H. Vølstad, Edward D. Weber, William A. Richkus, Robert J. Diaz, and Christina O. Tallent. The value of applying commercial fishers’ experience to designed surveys for identifying characteristics of essential fish habitat for adult summer flounder. North American Journal of Fisheries Management, 28(3):710–721, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stanley:2004:EHD

[SW04] David R. Stanley and Charles A. Wilson. Effect of hypoxia on the distribution of fishes associated with a petroleum platform off coastal Louisiana. North American Journal of Fisheries Management, 24(2):662–671, May 2004. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sheehan:2009:IMS

[SW09] Kenneth R. Sheehan and Stuart A. Welsh. An interpolation method for stream habitat assessments. North American Journal of Fisheries Management, 29(1):1–9, February 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sato:2008:IHB

[SWA^{+}08] Takuya Sato, Katsutoshi Watanabe, Masahiro Arizono, Seiichi Mori, Makoto Nagoshi, and Yasushi Harada. Intergeneric hybridization between sympatric kirikuchi char and redspotted masu salmon in a small Japanese mountain stream. North American Journal of Fisheries Management, 28(2): 547–556, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slaughter:2008:LIF

[SWD08] Joe E. Slaughter IV, Russell A. Wright, and Dennis R. De-Vries. Latitudinal influence on first-year growth and survival of largemouth bass. North American Journal of Fisheries Management, 28(4):993–1000, August 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schramm:2006:ELW

[SWG^{+}06] Harold L. Schramm, Jr., Aaron R. Walters, John M. Grizzle, Benjamin H. Beck, Larry A. Hanson, and Steven B. Rees. Effects of live-well conditions on mortality and largemouth bass virus prevalence in largemouth bass caught during summer tournaments. North American Journal of Fisheries Management, 26(4):812–825, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stuart:2006:MMP

[SWMH06] Ivor G. Stuart, Alan Williams, John McKenzie, and Terry Holt. Managing a migratory pest species: a selective trap for common carp. North American Journal of Fisheries Management, 26(4):888–893, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Siepker:2006:ERR

[SWPO06] Michael J. Siepker, David H. Wahl, David P. Philipp, and Kenneth G. Ostrand. Evidence of reduced reproductive success of nesting largemouth bass sampled with standard electrofishing procedures. North American Journal of Fisheries Management, 26(3):631–635, August 2006. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Speas:2004:EID

[SWWR04] David W. Speas, Carl J. Walters, David L. Ward, and R. Scott Rogers. Effects of intraspecific density and environmental variables on electrofishing catchability of brown and rainbow trout in the Colorado River. North American Journal of Fisheries Management, 24(2):586–596, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stockwell:2007:VDF

[SYH^{+}07] Jason D. Stockwell, Daniel L. Yule, Thomas R. Hrabik, Jean V. Adams, Owen T. Gorman, and Beth V. Holbrook. Vertical distribution of fish biomass in Lake Superior: Implications for day bottom trawl surveys. North American Journal of Fisheries Management, 27(3):735–749, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Simmonds:2000:MED

[SZL00] Robert L. Simmonds, Jr., Alexander V. Zale, and David M. Leslie, Jr. Modeled effects of double-crested cormorant predation on simulated reservoir sport and forage fish populations in Oklahoma. North American Journal of Fisheries Management, 20(1):180–191, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stanovick:2002:PSD

[SZW02] John S. Stanovick, Richard D. Zweifel, and Bruce Wagner. Precision of statistics derived from Missouri Statewide angler surveys. North American Journal of Fisheries Management, 22(4):1266–1269, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tate:2003:RAL

[TAM^{+}03] William B. Tate, Mike S. Allen, Randall A. Myers, Eric J. Nagid, and James R. Estes. Relation of age-0 largemouth bass abundance to hydrilla coverage and water level at Lochloosa and Orange Lakes, Florida. North American Journal of Fisheries Management, 23(2):251–257, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tate:2003:CER

[TAME03] William B. Tate, Mike S. Allen, Randall A. Myers, and James R. Estes. Comparison of electrofishing and rotenone for sampling largemouth bass in vegetated areas of two Florida lakes. North American Journal of Fisheries Management, 23(1):181–188, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tatara:2009:SIA

[Tat09] Christopher P. Tatara. Size at implantation affects growth of juvenile steelhead implanted with 12-mm passive integrated transponders. North American Journal of Fisheries Management, 29(2):417–422, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trebitz:2007:TTG

[TBB^{+}07] Anett S. Trebitz, John C. Brazner, Valerie J. Brady, Richard Axler, and Danny K. Tanner. Turbidity tolerances of Great Lakes coastal wetland fishes. North American Journal of Fisheries Management, 27(2):619–633, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tabor:2004:ELI

[TBL04] Roger A. Tabor, Gayle S. Brown, and Victoria T. Luiting. The effect of light intensity on sockeye salmon fry migratory behavior and predation by cottids in the Cedar River, Washington. North American Journal of Fisheries Management, 24(1):128–145, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Travnichek:2001:UTD

[TC01] Vincent H. Travnichek and H. Doug Clemons. Utility of tournament data for assessing effects of eliminating commercial catfish harvest in the Missouri River. North American Journal of Fisheries Management, 21(3):688–691, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:2008:PBS

[TCD^{+}08] Lisa A. Thompson, Steven J. Cooke, Michael R. Donaldson, Kyle C. Hanson, Andrew Gingerich, Thomas Klefoth, and Robert Arlinghaus. Physiology, behavior, and survival of angled and air-exposed largemouth bass. North American Journal of Fisheries Management, 28(4):1059–1068, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiffan:2006:VIP

[TCGR06] Kenneth F. Tiffan, Loreley O. Clark, Rodney D. Garland, and Dennis W. Rondorf. Variables influencing the presence of subyearling fall Chinook salmon in shoreline habitats of the Hanford reach, Columbia River. North American Journal of Fisheries Management, 26(2):351–360, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Toft:2007:FDA

[TCSS07] Jason D. Toft, Jeffery R. Cordell, Charles A. Simenstad, and Lia A. Stamatiou. Fish distribution, abundance, and behavior along city shoreline types in Puget Sound. North American Journal of Fisheries Management, 27(2):465–480, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thunberg:2006:TSR

[TF06] Eric M. Thunberg and Charles M. Fulcher. Testing the stability of recreational fishing participation probabilities. North American Journal of Fisheries Management, 26(3):636–644, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tabor:2007:SBL

[TFF^{+}07] Roger A. Tabor, Brian A. Footen, Kurt L. Fresh, Mark T. Celedonia, Francine Mejia, David L. Low, and Linda Park. Smallmouth bass and largemouth bass predation on juvenile Chinook salmon and other salmonids in the Lake Washington Basin. North American Journal of Fisheries Management, 27(4):1174–1188, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Toepfer:2000:MAE

[TFW00] Conrad S. Toepfer, William L. Fisher, and William D. Warde. A multistage approach to estimate fish abundance in streams using geographic information systems. North American Journal of Fisheries Management, 20(3):634–645, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiffan:2002:QFD

[TGR02] Kenneth F. Tiffan, Rodney D. Garland, and Dennis W. Rondorf. Quantifying flow-dependent changes in subyearling fall Chinook salmon rearing habitat using two-dimensional spatially explicit modeling. North American Journal of Fisheries Management, 22(3):713–726, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiffan:2006:PCS

[TGR06] Kenneth F. Tiffan, Rodney D. Garland, and Dennis W. Rondorf. Predicted changes in subyearling fall Chinook salmon rearing and migratory habitat under two drawdown scenarios for John day reservoir, Columbia River. North American Journal of Fisheries Management, 26(4):894–907, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:2005:CEM

[THE05] James M. Thompson, Pradeep S. Hirethota, and Bradley T. Eggold. A comparison of elastomer marks and fin clips as marking techniques for walleye. North American Journal of Fisheries Management, 25(1):308–315, February 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tipping:2001:MTM

[Tip01] Jack M. Tipping. Movement of tiger muskellunge in Mayfield Reservoir, Washington. North American Journal of Fisheries Management, 21(3):683–687, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tyson:2006:IRS

[TJKB06] Jeffrey T. Tyson, Timothy B. Johnson, Carey T. Knight, and Michael T. Bur. Intercalibration of research survey vessels on Lake Erie. North American Journal of Fisheries Management, 26(3):559–570, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tranah:2001:RIS

[TKK^{+}01] Gregory J. Tranah, Harold L. Kincaid, Charles C. Krueger, Donald E. Campton, and Bernie May. Reproductive isolation in sympatric populations of pallid and shovelnose sturgeon. North American Journal of Fisheries Management, 21 (2):367–373, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thomas:2002:EST

[TKT02] G. L. Thomas, Jay Kirsch, and Richard E. Thorne. Ex situ target strength measurements of Pacific herring and Pacific sand lance. North American Journal of Fisheries Management, 22(4):1136–1145, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trumble:2000:EDM

[TKW00] Robert J. Trumble, Stephen M. Kaimmer, and Gregg H. Williams. Estimation of discard mortality rates for Pacific halibut bycatch in groundfish longline fisheries. North American Journal of Fisheries Management, 20(4):931–939, November 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thomson:2005:DIU

[TMS05] Adrian W. Thomson and Roy Melville-Smith. Have different inducements used with the western rock lobster mail survey of recreational licensees had an effect on total catch estimates? North American Journal of Fisheries Management, 25(4):1203–1207, November 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tomcko:2005:BRG

[TP05] Cynthia M. Tomcko and Rodney B. Pierce. Bluegill recruitment, growth, population size structure, and associated factors in Minnesota lakes. North American Journal of Fisheries Management, 25(1):171–179, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Temple:2006:ERP

[TP06] Gabriel M. Temple and Todd N. Pearsons. Evaluation of the recovery period in mark–recapture population estimates of rainbow trout in small streams. North American Journal of Fisheries Management, 26(4):941–948, November 2006. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tripp:2009:MRS

[TPC^{+}09] Sara J. Tripp, Quinton E. Phelps, Robert E. Colombo, James E. Garvey, Brooks M. Burr, David P. Herzog, and Robert A. Hrabik. Maturation and reproduction of shovelnose sturgeon in the Middle Mississippi River. North American Journal of Fisheries Management, 29(3):730–738, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Talmage:2002:RIH

[TPG02] Philip J. Talmage, James A. Perry, and Robert M. Goldstein. Relation of instream habitat and physical conditions to fish communities of agricultural streams in the Northern Midwest. North American Journal of Fisheries Management, 22(3):825–833, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thurow:2006:UVD

[TPG06] Russell F. Thurow, James T. Peterson, and John W. Guzevich. Utility and validation of day and night snorkel counts for estimating bull trout abundance in first- to third-order streams. North American Journal of Fisheries Management, 26(1):217–232, February 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiffan:2004:IFC

[TRS04] Kenneth F. Tiffan, Dennis W. Rondorf, and Joseph J. Skalicky. Imaging fall Chinook salmon redds in the Columbia River with a dual-frequency identification sonar. North American Journal of Fisheries Management, 24(4):1421– 1426, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiffan:2000:PDM

[TRW00] Kenneth F. Tiffan, Dennis W. Rondorf, and Paul G. Wagner. Physiological development and migratory behavior of subyearling fall Chinook salmon in the Columbia River. North American Journal of Fisheries Management, 20(1):28–40, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Teuscher:2003:RSG

[TSMD03] David M. Teuscher, Daniel J. Schill, Douglas J. Megargle, and Jeff C. Dillon. Relative survival and growth of triploid and diploid rainbow trout in two Idaho reservoirs. North American Journal of Fisheries Management, 23(4):983–988, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trudel:2005:IES

[TTM^{+}05] M. Trudel, S. Tucker, J. F. T. Morris, D. A. Higgs, and D. W. Welch. Indicators of energetic status in juvenile Coho salmon and Chinook salmon. North American Journal of Fisheries Management, 25(1):374–390, February 2005. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2005:SSC

[TTRF05] J. Christopher Taylor, Jessica S. Thompson, Peter S. Rand, and Montserrat Fuentes. Sampling and statistical considerations for hydroacoustic surveys used in estimating abundance of forage fishes in reservoirs. North American Journal of Fisheries Management, 25(1):73–85, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2001:CRM

[TWH01] Ronald G. Taylor, James A. Whittington, and Douglas E. Haymans. Catch-and-release mortality rates of common snook in Florida. North American Journal of Fisheries Management, 21(1):70–75, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiano:2007:GIH

[TWN^{+}07] Todd J. Tiano, Catherine M. Willis, Amy A. Noble, Mark R. Luttenton, and Alexey G. Nikitin. Genetic identification of hatchery stocks of brown trout using mitochondrial DNA polymorphism. North American Journal of Fisheries Management, 27(3):965–970, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tattam:2003:SPA

[TWP03] Ian A. Tattam, Timothy A. Whitesel, and Yangdong Pan. Scale pattern analysis of selected scale characteristics and the first annulus for distinguishing wild and hatchery steelhead in the Hood River, Oregon. North American Journal of Fisheries Management, 23(3):856–868, August 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2006:EDR

[TWPP06] Ronald G. Taylor, James A. Whittington, William E. Pine III, and Kenneth H. Pollock. Effect of different reward levels on tag reporting rates and behavior of common snook anglers in Southeast Florida. North American Journal of Fisheries Management, 26(3):645–651, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:2007:SNF

[TWRH07] Jessica S. Thompson, D. Scott Waters, James A. Rice, and Joseph E. Hightower. Seasonal natural and fishing mortality of striped bass in a southeastern reservoir. North American Journal of Fisheries Management, 27(2):681–694, May 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Underwood:2004:EHM

[UBK04] Tevis J. Underwood, Jeffrey F. Bromaghin, and Steve P. Klosiewski. Evidence of handling mortality of adult chum salmon caused by fish wheel capture in the Yukon River, Alaska. North American Journal of Fisheries Management, 24(1):237–243, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Utrup:2006:DRB

[UF06] Nicholas J. Utrup and William L. Fisher. Development of a rapid bioassessment protocol for sampling fish in large prairie rivers. North American Journal of Fisheries Management, 26 (3):714–726, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Underwood:2007:ADE

[UGM^{+}07] Tevis J. Underwood, Judith A. Gordon, Michael J. Millard, Brian R. Lubinski, and Steve P. Klosiewski. Appropriateness of the Darroch estimator for monitoring adult chum salmon on the Middle Yukon River, Alaska. North American Journal of Fisheries Management, 27(1):303–314, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Unmuth:2001:EMH

[UHRP01] Jean M. L. Unmuth, Michael J. Hansen, Paul W. Rasmussen, and Thomas D. Pellett. Effects of mechanical harvesting of Eurasian watermilfoil on angling for bluegills in Fish Lake, Wisconsin. North American Journal of Fisheries Management, 21(3):448–454, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Underwood:2000:ALC

[Und00] Tevis J. Underwood. Abundance, length composition, and migration of spawning inconnu in the Selawik River, Alaska. North American Journal of Fisheries Management, 20(2): 386–393, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Unwin:2005:QPS

[UWBL05] Martin J. Unwin, Mark Webb, Richard J. Barker, and William A. Link. Quantifying production of salmon fry in an unscreened irrigation system: a case study on the Rangitata River, New Zealand. North American Journal of Fisheries Management, 25(2):619–634, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vandergoot:2003:RCS

[VB03] Christopher S. Vandergoot and Phillip W. Bettoli. Relative contribution of stocked walleyes in Tennessee reservoirs. North American Journal of Fisheries Management, 23(4): 1036–1041, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vinson:2008:PGR

[VB08] Mark R. Vinson and Michelle A. Baker. Poor growth of rainbow trout fed New Zealand mud snails Potamopyrgus antipodarum. North American Journal of Fisheries Management, 28(3):701–709, June 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanderHaegen:2005:EAF

[VBHT05] G. E. Vander Haegen, H. L. Blankenship, A. Hoffmann, and D. A. Thompson. The effects of adipose fin clipping and coded wire tagging on the survival and growth of spring Chinook salmon. North American Journal of Fisheries Management, 25(3):1161–1170, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDeValk:2007:CSY

[VBJR07] Anthony J. VanDeValk, Thomas E. Brooking, James R. Jackson, and Lars G. Rudstam. Contribution of stocked yearling walleyes to the fishery in Oneida Lake, New York. North American Journal of Fisheries Management, 27(3): 1018–1024, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vandergoot:2008:LEY

[VBP08] Christopher S. Vandergoot, Michael T. Bur, and Kipp A. Powell. Lake Erie yellow perch age estimation based on three structures: Precision, processing times, and management implications. North American Journal of Fisheries Management, 28(2):563–571, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanArnum:2004:MSB

[VBR04] Christy J. Gunderson VanArnum, Gerard L. Buynak, and Jeffrey R. Ross. Movement of smallmouth bass in Elkhorn Creek, Kentucky. North American Journal of Fisheries Management, 24(1):311–315, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vaughan:2002:EIS

[VC02] Douglas S. Vaughan and John T. Carmichael. Estimating improvement in spawning potential ratios for South Atlantic red drum through bag and size limit regulations. North American Journal of Fisheries Management, 22(3):895–906, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

vanderWalt:2003:OMR

[vdWF03] Bryan van der Walt and Robert A. Faragher. Otolith marking of rainbow trout fry by immersion in low concentrations of alizarin complexone. North American Journal of Fisheries Management, 23(1):141–148, February 2003. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Velez-Espino:2009:RPA

[VEK09] Luis A. Vélez-Espino and Marten A. Koops. Recovery potential assessment for lake sturgeon in Canadian designatable units. North American Journal of Fisheries Management, 29 (4):1065–1090, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDeValk:2005:ACR

[VFJ^{+}05] Anthony J. VanDeValk, John L. Forney, James R. Jackson, Lars G. Rudstam, Thomas E. Brooking, and Scott D. Krueger. Angler catch rates and catchability of walleyes in Oneida Lake, New York. North American Journal of Fisheries Management, 25(4):1441–1447, November 2005. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDeValk:2008:RBR

[VFJ08] Anthony J. VanDeValk, John L. Forney, and James R. Jackson. Relationships between relative weight, prey availability, and growth of walleyes in Oneida Lake, New York. North American Journal of Fisheries Management, 28(6): 1868–1875, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanderKooi:2000:PSG

[VGM00] Scott P. VanderKooi, William L. Gale, and Alec G. Maule. Preliminary study of gill Na^{+}, K^{+}-ATPase activity in juvenile spring Chinook salmon following electroshock or handling stress. North American Journal of Fisheries Management, 20 (2):528–533, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Volpe:2000:SIA

[VH00] John P. Volpe and Gerry Horne. A simple and inexpensive apparatus for providing natural prey in a laboratory environment. North American Journal of Fisheries Management, 20 (3):801–803, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDeValk:2007:IPS

[VJK^{+}07] Anthony J. VanDeValk, James R. Jackson, Scott D. Krueger, Thomas E. Brooking, and Lars G. Rudstam. Influence of party size and trip length on angler catch rates on Oneida Lake, New York. North American Journal of Fisheries Management, 27(1):127–136, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vokoun:2003:KDE

[Vok03] Jason C. Vokoun. Kernel density estimates of linear home ranges for stream fishes: Advantages and data requirements. North American Journal of Fisheries Management, 23(4): 1020–1029, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

vanPoorten:2005:SFD

[vPP05] Brett T. van Poorten and John R. Post. Seasonal fishery dynamics of a previously unexploited rainbow trout population with contrasts to established fisheries. North American Journal of Fisheries Management, 25(1):329–345, February 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Volstad:2006:CCE

[VPR06] Jon H. Vølstad, Kenneth H. Pollock, and William A. Richkus. Comparing and combining effort and catch estimates from aerial–access designs as applied to a large-scale angler survey in the Delaware River. North American Journal of Fisheries Management, 26(3):727–741, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vokoun:2001:SSP

[VR01] Jason C. Vokoun and Charles F. Rabeni. A standardized sampling protocol for channel catfish in prairie streams. North American Journal of Fisheries Management, 21(1):188–197, February 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vokoun:2005:VAM

[VR05] Jason C. Vokoun and Charles F. Rabeni. Variation in an annual movement cycle of flathead catfish within and between two Missouri watersheds. North American Journal of Fisheries Management, 25(2):563–572, May 2005. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Venditti:2000:MBF

[VRK00] David A. Venditti, Dennis W. Rondorf, and John M. Kraut. Migratory behavior and forebay delay of radio-tagged juvenile fall Chinook salmon in a Lower Snake River impoundment. North American Journal of Fisheries Management, 20 (1):41–52, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vokoun:2001:SSR

[VRS01] J. C. Vokoun, C. F. Rabeni, and J. S. Stanovick. Sample-size requirements for evaluating population size structure. North American Journal of Fisheries Management, 21(3):660–665, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Venard:2005:SDM

[VS05] Jacob A. Venard and Dennis L. Scarnecchia. Seasonally dependent movement of lake trout between two Northern Idaho lakes. North American Journal of Fisheries Management, 25 (2):635–639, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanderHaegen:2002:DCW

[VSB02] G. E. Vander Haegen, A. M. Swanson, and H. L. Blankenship. Detecting coded wire tags with handheld wands: Effectiveness of two wanding techniques. North American Journal of Fisheries Management, 22(4):1260–1265, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDoornik:2007:GAP

[VTK^{+}07] Donald M. Van Doornik, David J. Teel, David R. Kuligowski, Cheryl A. Morgan, and Edmundo Casillas. Genetic analyses provide insight into the early ocean stock distribution and survival of juvenile Coho salmon off the coasts of Washington and Oregon. North American Journal of Fisheries Management, 27(1):220–237, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDenAvyle:2001:RIA

[VW01] Michael J. Van Den Avyle and Julie E. Wallin. Retention of internal anchor tags by juvenile striped bass. North American Journal of Fisheries Management, 21(3):656–659, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vecchio:2007:CRM

[VW07] Julie L. Vecchio and Charles A. Wenner. Catch-and-release mortality in subadult and adult red drum captured with popular fishing hook types. North American Journal of Fisheries Management, 27(3):891–899, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vaux:2000:EBE

[VWDK00] Peter D. Vaux, Thomas R. Whittier, Gregory DeCesare, and James P. Kurtenbach. Evaluation of a backpack electrofishing unit for multiple lake surveys of fish assemblage structure. North American Journal of Fisheries Management, 20 (1):168–179, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:2002:SPR

[WAB^{+}02] Eric Wagner, Ronney Arndt, Mark Brough, Chris Wilson, and Gordon Nelson. Survival, performance, and resistance to Myxobolus cerebralis infection of lake trout × brook trout hybrids. North American Journal of Fisheries Management, 22(3):760–769, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wanner:2006:EGL

[Wan06] Greg A. Wanner. Evaluation of a gastric lavage method on juvenile pallid sturgeon. North American Journal of Fisheries Management, 26(3):587–591, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:2003:APM

[WB03] Gregg H. Williams and Calvin L. Blood. Active and passive management of the recreational fishery for Pacific halibut off the U.S. West Coast. North American Journal of Fisheries Management, 23(4):1359–1368, November 2003. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilberg:2005:YPD

[WBE^{+}05] Michael J. Wilberg, James R. Bence, Bradley T. Eggold, Daniel Makauskas, and David F. Clapp. Yellow perch dynamics in Southwestern Lake Michigan during 1986–2002. North American Journal of Fisheries Management, 25(3): 1130–1152, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weber:2002:PCS

[WBH02] Edward D. Weber, Sandra M. Borthwick, and Louis A. Helfrich. Plasma cortisol stress response of juvenile Chinook salmon to passage through Archimedes lifts and a hidrostal pump. North American Journal of Fisheries Management, 22 (2):563–570, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilberg:2002:SJL

[WBJ02] Michael J. Wilberg, James R. Bence, and James E. Johnson. Survival of juvenile lake trout stocked in Western Lake Huron during 1974–1992. North American Journal of Fisheries Management, 22(1):213–218, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilber:2001:BES

[WC01] Dara H. Wilber and Douglas G. Clarke. Biological effects of suspended sediments: a review of suspended sediment impacts on fish and shellfish with relation to dredging activities in estuaries. North American Journal of Fisheries Management, 21(4):855–875, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wang:2009:MSW

[WCE^{+}09] Hui-Yu Wang, H. Andrew Cook, Donald W. Einhouse, David G. Fielder, Kevin A. Kayle, Lars G. Rudstam, and Tomas O. Höök. Maturation schedules of walleye populations in the Great Lakes region: Comparison of maturation indices and evaluation of sampling-induced biases. North American Journal of Fisheries Management, 29(6):1540–1554, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wirgin:2005:IND

[WCR^{+}05] Isaac Wirgin, Diane Currie, Nirmal Roy, Lorraine Maceda, and John R. Waldman. Introgression of nuclear DNA (nDNA) alleles of stocked Atlantic Coast striped bass with the last remaining native Gulf of Mexico population. North American Journal of Fisheries Management, 25(2):464–474, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wirgin:2004:DUS

[WCSJ04] Isaac Wirgin, Diane D. Currie, Joseph Stabile, and Cecil A. Jennings. Development and use of a simple DNA test to distinguish larval redhorse species in the Oconee River, Georgia. North American Journal of Fisheries Management, 24 (1):293–298, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weimer:2006:CTE

[WDB06] Eric J. Weimer, Jeremy P. Duehr, and Michael L. Brown. Comparison of two external transmitter types on two sizes of bluegills and yellow perch. North American Journal of Fisheries Management, 26(3):670–675, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wahl:2007:EEB

[WEC07] David H. Wahl, Lisa M. Einfalt, and Sean P. Callahan. Effects of electroshock on bluegill feeding and susceptibility to predation. North American Journal of Fisheries Management, 27(4):1208–1213, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wertheimer:2007:ESF

[Wer07] Robert H. Wertheimer. Evaluation of a surface flow bypass system for steelhead kelt passage at Bonneville Dam, Washington. North American Journal of Fisheries Management, 27(1):21–29, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:2009:EFC

[WESW09] Curtis P. Wagner, Lisa M. Einfalt, Adam B. Scimone, and David H. Wahl. Effects of fin-clipping on the foraging behavior and growth of age-0 muskellunge. North American Journal of Fisheries Management, 29(6):1644–1652, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weber:2004:ASD

[WF04] Edward D. Weber and Kurt D. Fausch. Abundance and size distribution of ocean-type juvenile Chinook salmon in the Upper Sacramento River margin before and after hatchery releases. North American Journal of Fisheries Management, 24(4):1447–1455, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walker:2007:EED

[WFS07] Jordan R. Walker, Lee Foote, and Michael G. Sullivan. Effectiveness of enforcement to deter illegal angling harvest of northern pike in Alberta. North American Journal of Fisheries Management, 27(4):1369–1377, November 2007. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walsh:2002:CES

[WFW02] Maureen G. Walsh, Daniel B. Fenner, and Dana L. Winkelman. Comparison of an electric seine and prepositioned area electrofishers for sampling stream fish communities. North American Journal of Fisheries Management, 22(1):77–85, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Watkinson:2003:SDW

[WG03] Douglas A. Watkinson and Darren M. Gillis. Stock differentiation of walleye based on the Fourier approximation of averaged scale outline signals. North American Journal of Fisheries Management, 23(1):91–99, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wallmo:2008:CRF

[WG08] Kristy Wallmo and Brad Gentner. Catch-and-release fishing: a comparison of intended and actual behavior of marine anglers. North American Journal of Fisheries Management, 28(5):1459–1471, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wirgin:2007:GER

[WGSW07] Isaac Wirgin, Cheryl Grunwald, Joseph Stabile, and John Waldman. Genetic evidence for relict Atlantic sturgeon stocks along the mid-Atlantic Coast of the USA. North American Journal of Fisheries Management, 27(4):1214–1229, November 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:2004:EHE

[WH04] David L. Ward and Kara D. Hilwig. Effects of holding environment and exercise conditioning on swimming performance of southwestern native fishes. North American Journal of Fisheries Management, 24(3):1083–1087, August 2004. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weeks:2009:WMM

[WH09] Jordan G. Weeks and Michael J. Hansen. Walleye and muskellunge movement in the manitowish chain of lakes, Vilas County, Wisconsin. North American Journal of Fisheries Management, 29(3):791–804, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilberg:2003:HMA

[WHB03] Michael J. Wilberg, Michael J. Hansen, and Charles R. Bronte. Historic and modern abundance of wild lean lake trout in Michigan waters of Lake Superior: Implications for restoration goals. North American Journal of Fisheries Management, 23(1):100–108, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:2009:ECC

[WHC09] Abby Williams, Mark Herrmann, and Keith R. Criddle. The effects of Chilean Coho salmon and rainbow trout aquaculture on markets for Alaskan sockeye salmon. North American Journal of Fisheries Management, 29(6):1777–1796, December 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:2002:TGC

[WHCL02] Daniel J. Willis, Mark V. Hoyer, Daniel E. Canfield, Jr., and William J. Lindberg. Training grass carp to respond to sound for potential lake management uses. North American Journal of Fisheries Management, 22(1):208–212, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Welsh:2001:DJC

[WHHR01] Hartwell H. Welsh, Jr., Garth R. Hodgson, Bret C. Harvey, and Maureen F. Roche. Distribution of juvenile Coho salmon in relation to water temperatures in tributaries of the Mattole River, California. North American Journal of Fisheries Management, 21(3):464–470, August 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:2009:PCC

[WHNC09] Jack E. Williams, Amy L. Haak, Helen M. Neville, and Warren T. Colyer. Potential consequences of climate change to persistence of cutthroat trout populations. North American Journal of Fisheries Management, 29(3):533–548, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Welsh:2008:GAL

[WHQ^{+}08] Amy Welsh, Tracy Hill, Henry Quinlan, Charmaine Robinson, and Bernie May. Genetic assessment of lake sturgeon population structure in the Laurentian Great Lakes. North American Journal of Fisheries Management, 28(2):572–591, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:2002:EUD

[Wil02] Erik H. Williams. The effects of unaccounted discards and misspecified natural mortality on harvest policies based on estimates of spawners per recruit. North American Journal of Fisheries Management, 22(2):311–325, February 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wills:2006:CAS

[Wil06] Todd C. Wills. Comparative abundance, survival, and growth of one wild and two domestic brown trout strains stocked in Michigan rivers. North American Journal of Fisheries Management, 26(3):535–544, August 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winston:2004:SFS

[Win04] Matthew R. Winston. Sampling fish in streams with seines: Costs of recording data by haul. North American Journal of Fisheries Management, 24(2):711–714, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:2007:SGT

[WJKW07] Curtis P. Wagner, Martin J. Jennings, Jeffrey M. Kampa, and David H. Wahl. Survival, growth, and tag retention in age-0 muskellunge implanted with passive integrated transponders. North American Journal of Fisheries Management, 27 (3):873–877, August 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waters:2004:EST

[WKAP04] D. Scott Waters, Thomas J. Kwak, Joshua B. Arnott, and William E. Pine III. Evaluation of stomach tubes and gastric lavage for sampling diets from blue catfish and flathead catfish. North American Journal of Fisheries Management, 24 (1):258–261, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wildhaber:2003:CMR

[WLG03] Mark L. Wildhaber, Peter J. Lamberson, and David L. Galat. A comparison of measures of riverbed form for evaluating distributions of benthic fishes. North American Journal of Fisheries Management, 23(2):543–557, May 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wayman:2008:CSE

[WLHT08] William R. Wayman, Gregory L. Looney, Robert J. Holm, and Terrence R. Tiersch. Cryopreservation of sperm from endangered pallid sturgeon. North American Journal of Fisheries Management, 28(3):740–744, June 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williamson:2005:HFR

[WM05] Kevin S. Williamson and Bernie May. Homogenization of fall-run Chinook salmon gene pools in the Central Valley of California, USA. North American Journal of Fisheries Management, 25(3):993–1009, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilde:2000:BTE

[WMB^{+}00] Gene R. Wilde, Maurice I. Muoneke, Phillip W. Bettoli, Kent L. Nelson, and Bruce T. Hysmith. Bait and temperature effects on striped bass hooking mortality in freshwater. North American Journal of Fisheries Management, 20(3):810–815, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weeder:2005:APG

[WME05] Julie A. Weeder, Anne R. Marshall, and John M. Epifanio. An assessment of population genetic variation in Chinook salmon from seven Michigan rivers 30 years after introduction. North American Journal of Fisheries Management, 25 (3):861–875, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Warner:2004:EBR

[WMJBC04] Daniel A. Warner, Anne L. McMillen-Jackson, Theresa M. Bert, and Charles R. Crawford. The efficiency of a bycatch reduction device used in skimmer trawls in the Florida shrimp fishery. North American Journal of Fisheries Management, 24(3):853–864, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walsh:2008:CRS

[WMO08] Maureen G. Walsh, Aaron P. Maloy, and Timothy P. O’Brien. Comparison of rainbow smelt age estimates from fin rays and otoliths. North American Journal of Fisheries Management, 28(1):42–49, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wang:2009:UGA

[WMSG09] Shizhen Wang, Gary Morishima, Rishi Sharma, and Larry Gilbertson. The use of generalized additive models for forecasting the abundance of Queets River Coho salmon. North American Journal of Fisheries Management, 29(2):423–433, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waters:2004:SSN

[WN04] D. Scott Waters and Richard L. Noble. Spawning season and nest fidelity of largemouth bass in a tropical reservoir. North American Journal of Fisheries Management, 24(4): 1240–1251, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilcox:2003:DCF

[WP03] Chris Wilcox and Caroline Pomeroy. Do commercial fishers aggregate around marine reserves? Evidence from Big Creek Marine Ecological Reserve, Central California. North American Journal of Fisheries Management, 23(2):241–250, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilde:2004:APC

[WP04a] Gene R. Wilde and Kevin L. Pope. Anglers’ probabilities of catching record-size fish. North American Journal of Fisheries Management, 24(3):1046–1049, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilde:2004:RBL

[WP04b] Gene R. Wilde and Kevin L. Pope. Relationship between lake-record weights of fishes and reservoir area and growing season. North American Journal of Fisheries Management, 24(3):1025–1030, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whittier:2008:SAE

[WP08] Thomas R. Whittier and David V. Peck. Surface area estimates of streams and rivers occupied by nonnative fish and amphibians in the Western USA. North American Journal of Fisheries Management, 28(6):1887–1893, December 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Willis:2002:BWB

[WPB02] David W. Willis, Craig P. Paukert, and Brian G. Blackwell. Biology of white bass in Eastern South Dakota glacial lakes.

North American Journal of Fisheries Management, 22(2): 627–636, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilde:2003:EFT

[WPS03] Gene R. Wilde, Kevin L. Pope, and Richard E. Strauss. Estimation of fishing tournament mortality and its sampling variance. North American Journal of Fisheries Management, 23(3):779–786, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winans:2004:GSI

[WPV^{+}04] Gary A. Winans, Melanie M. Paquin, Donald M. Van Doornik, Bruce M. Baker, Perry Thornton, Dan Rawding, Anne Marshall, Paul Moran, and Steven Kalinowski. Genetic stock identification of steelhead in the Columbia River Basin: An evaluation of different molecular markers. North American Journal of Fisheries Management, 24(2):672–685, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Warren:2009:CMB

[WSH09] Melvin L. Warren, Jr., Andrew L. Sheldon, and Wendell R. Haag. Constructed microhabitat bundles for sampling fishes and crayfishes in coastal plain streams. North American Journal of Fisheries Management, 29(2):330–342, April 2009. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:2001:SED

[WSM01] John G. Williams, Steven G. Smith, and William D. Muir. Survival estimates for downstream migrant yearling juvenile salmonids through the Snake and Columbia Rivers hydropower system, 1966–1980 and 1993–1999. North American Journal of Fisheries Management, 21(2):310–317, May 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilson:2008:CGL

[WSM^{+}08] Chris C. Wilson, Wendylee Stott, Loren Miller, Silvia D’Amelio, Martin J. Jennings, and Anne M. Cooper. Conservation genetics of Lake Superior brook trout: Issues, questions, and directions. North American Journal of Fisheries Management, 28(5):1307–1320, August 2008. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilde:2002:IMB

[WSP02] Gene R. Wilde, Calub E. Shavlik, and Kevin L. Pope. Initial mortality of black bass in B.A.S.S. fishing tournaments. North American Journal of Fisheries Management, 22(3): 950–954, August 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilderbuer:2009:SSN

[WT09] Thomas K. Wilderbuer and Benjamin J. Turnock. Sexspecific natural mortality of arrowtooth flounder in Alaska: Implications of a skewed sex ratio on exploitation and management. North American Journal of Fisheries Management, 29(2):306–322, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wollebaek:2008:RSM

[WTH08] Jens Wollebæk, Roy Thue, and Jan Heggenes. Redd site microhabitat utilization and quantitative models for wild large brown trout in three contrasting boreal rivers. North American Journal of Fisheries Management, 28(5):1249–1258, August 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whiteman:2004:AES

[WTW^{+}04] Kasey W. Whiteman, Vincent H. Travnichek, Mark L. Wildhaber, Aaron DeLonay, Diana Papoulias, and Donald Tillett. Age estimation for shovelnose sturgeon: a cautionary note based on annulus formation in pectoral fin rays. North American Journal of Fisheries Management, 24(2):731–734, May 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:2009:EPD

[WVT09] Tyler Wagner, Christopher S. Vandergoot, and Jeff Tyson. Evaluating the power to detect temporal trends in fisheryindependent surveys: a case study based on gill nets set in the Ohio waters of Lake Erie for walleyes. North American Journal of Fisheries Management, 29(3):805–816, June 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weller:2001:SVH

[WW01] Robert R. Weller and Jimmy D. Winter. Seasonal variation in home range size and habitat use of flathead catfish in Buffalo Springs Lake, Texas. North American Journal of Fisheries Management, 21(4):792–800, November 2001. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walsh:2004:AVI

[WW04] Maureen G. Walsh and Dana L. Winkelman. Anchor and visible implant elastomer tag retention by hatchery rainbow trout stocked into an Ozark stream. North American Journal of Fisheries Management, 24(4):1435–1439, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wedemeyer:2008:PRS

[WW08] Gary A. Wedemeyer and Richard S. Wydoski. Physiological response of some economically important freshwater salmonids to catch-and-release fishing. North American Journal of Fisheries Management, 28(5):1587–1596, October 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walsh:2004:EIS

[WWB04] Maureen G. Walsh, Dana L. Winkelman, and Robert J. Bahr. Electrofishing injury and short-term mortality in hatcheryreared rainbow trout stocked into an Ozark stream. North American Journal of Fisheries Management, 24(1):316–321, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yule:2007:UMG

[YASG07] Daniel L. Yule, Jean V. Adams, Jason D. Stockwell, and Owen T. Gorman. Using multiple gears to assess acoustic detectability and biomass of fish species in Lake Superior. North American Journal of Fisheries Management, 27(1):106–126, February 2007. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yule:2008:FAB

[YASG08] Daniel L. Yule, Jean V. Adams, Jason D. Stockwell, and Owen T. Gorman. Factors affecting bottom trawl catches:

Implications for monitoring the fishes of Lake Superior. North American Journal of Fisheries Management, 28(1):109–122, February 2008. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:2009:SSB

[YDJM09] John R. Young, William P. Dey, Steven M. Jinks, and David T. Mosier. Survival of striped bass entrained into the cooling systems of two Hudson River power stations. North American Journal of Fisheries Management, 29(4): 1015–1034, August 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:2004:PCG

[YGG04] Michael K. Young and Paula M. Guenther-Gloss. Population characteristics of greenback cutthroat trout in streams: Their relation to model predictions and recovery criteria. North American Journal of Fisheries Management, 24(1):184–197, February 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:2004:APT

[YH04] Roger G. Young and John W. Hayes. Angling pressure and trout catchability: Behavioral observations of brown trout in two New Zealand backcountry rivers. North American Journal of Fisheries Management, 24(4):1203–1213, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:2004:TSE

[YI04] Shawn P. Young and J. Jeffery Isely. Temporal and spatial estimates of adult striped bass mortality from telemetry and transmitter return data. North American Journal of Fisheries Management, 24(4):1112–1119, November 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:2006:PSD

[YI06] Shawn P. Young and J. Jeffery Isely. Posttournament survival and dispersal of adult striped bass. North American Journal of Fisheries Management, 26(4):1030–1033, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:2003:PPA

[YMR03] Michael K. Young, Bradley K. Micek, and Michael Rathbun. Probable pheromonal attraction of sexually mature brook trout to mature male conspecifics. North American Journal of Fisheries Management, 23(2):276–282, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:2009:CFT

[You09] John R. Young. Comparison of failure time analysis and Abbott’s formula for estimating survival of organisms entrained at power stations. North American Journal of Fisheries Management, 29(2):263–278, April 2009. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yin:2004:BPE

[YS04] Yanshui Yin and David B. Sampson. Bias and precision of estimates from an age-structured stock assessment program in relation to stock and data characteristics. North American Journal of Fisheries Management, 24(3):865–879, August 2004. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yuen:2005:UST

[YS05] Henry Yuen and Rishi Sharma. Using simulation techniques to estimate management parameters on Snake River steelhead: Declines in productivity make rebuilding difficult. North American Journal of Fisheries Management, 25(2): 446–463, May 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yule:2000:CHA

[Yul00] Daniel L. Yule. Comparison of horizontal acoustic and purseseine estimates of salmonid densities and sizes in eleven Wyoming waters. North American Journal of Fisheries Management, 20(3):759–775, August 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yule:2000:USS

[YWM^{+}00] Daniel L. Yule, Roy A. Whaley, Paul H. Mavrakis, Dirk D. Miller, and Steve A. Flickinger. Use of strain, season of stocking, and size at stocking to improve fisheries for rainbow trout in reservoirs with walleyes. North American Journal of Fisheries Management, 20(1):10–18, February 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zoellick:2005:LTC

[ZAF05] Bruce W. Zoellick, Dale B. Allen, and Brian J. Flatter. A long-term comparison of redband trout distribution, density, and size structure in Southwestern Idaho. North American Journal of Fisheries Management, 25(3):1179–1190, August 2005. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zendt:2000:DAA

[ZB00] Joseph S. Zendt and Eric P. Bergersen. Distribution and abundance of the aquatic oligochaete host Tubifex tubifex for the salmonid whirling disease parasite Myxobolus cerebralis in the Upper Colorado River Basin. North American Journal of Fisheries Management, 20(2):502–512, May 2000. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zoellick:2006:ERT

[ZC06] Bruce W. Zoellick and Brian S. Cade. Evaluating redband trout habitat in sagebrush desert basins in Southwestern Idaho. North American Journal of Fisheries Management, 26(2):268–281, May 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zigler:2003:MHU

[ZDK^{+}03] Steven J. Zigler, Michael R. Dewey, Brent C. Knights, Ann L. Runstrom, and Mark T. Steingraeber. Movement and habitat use by radio-tagged paddlefish in the Upper Mississippi River and tributaries. North American Journal of Fisheries Management, 23(1):189–205, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zweifel:2009:HTV

[ZHBB09] Richard D. Zweifel, R. Scott Hale, David B. Bunnell, and Mary T. Bremigan. Hatch timing variations among reservoir gizzard shad populations: Implications for stocked Sander spp. fingerlings. North American Journal of Fisheries Management, 29(2):488–494, April 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhou:2002:UEF

[Zho02] Shijie Zhou. Uncertainties in estimating fishing mortality in unmarked salmon in mark-selective fisheries using doubleindex-tagging methods. North American Journal of Fisheries Management, 22(2):480–493, May 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhou:2003:AAN

[Zho03] Shijie Zhou. Application of artificial neural networks for forecasting salmon escapement. North American Journal of Fisheries Management, 23(1):48–59, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhou:2004:PME

[Zho04] Shijie Zhou. A pipeline model for estimating fishing mortality in salmon mark-selective fisheries. North American Journal of Fisheries Management, 24(3):979–989, August 2004. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zydlewski:2002:RBT

[ZJ02] Gayle B. Zydlewski and Jeffrey R. Johnson. Response of bull trout fry to four types of water diversion screens. North American Journal of Fisheries Management, 22(4):1276– 1282, November 2002. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhou:2000:MCP

[ZK00] Shijie Zhou and Gordon H. Kruse. Modifications of cod pots to reduce Tanner crab bycatch. North American Journal of Fisheries Management, 20(4):897–907, November 2000. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zimmerman:2009:EPR

[ZK09] Mara S. Zimmerman and Charles C. Krueger. An ecosystem perspective on re-establishing native deepwater fishes in the Laurentian Great Lakes. North American Journal of Fisheries Management, 29(5):1352–1371, October 2009. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zymonas:2006:EPF

[ZM06] Nikolas D. Zymonas and Thomas E. McMahon. Effect of pelvic fin ray removal on survival and growth of bull trout. North American Journal of Fisheries Management, 26(4): 953–959, November 2006. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zweifel:2003:BAU

[ZS03] Richard D. Zweifel and John S. Stanovick. Biases associated with using roving angler surveys on the opening day of put-and-take fisheries. North American Journal of Fisheries Management, 23(2):290–296, February 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zimmerman:2003:MPM

[ZSH03] Christian E. Zimmerman, R. Wes Stonecypher, Jr., and Michael C. Hayes. Migration of precocious male hatchery Chinook salmon in the Umatilla River, Oregon. North American Journal of Fisheries Management, 23(4):1006–1014, August 2003. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).